### v1

# Nearly Optimal Sample Complexity of Offline KL-Regularized Contextual Bandits under Single-Policy Concentrability

Abstract
KL-regularized policy optimization has become a workhorse in learning-based decision making, while its theoretical understanding is still very limited. Although recent progress has been made towards settling the sample complexity of KL-regularized contextual bandits, existing sample complexity bounds are either $\widetilde{O}(\epsilon^{-2})$ under single-policy concentrability or $\widetilde{O}(\epsilon^{-1})$ under all-policy concentrability. In this paper, we propose the first algorithm with $\widetilde{O}(\epsilon^{-1})$ sample complexity under single-policy concentrability for offline contextual bandits. Our algorithm is designed for general function approximation and based on the principle of pessimism in the face of uncertainty. The core of our proof leverages the strong convexity of the KL regularization, and the conditional non-negativity of the gap between the true reward and its pessimistic estimator to refine a mean-value-type risk upper bound to its extreme. This in turn leads to a novel covariance-based analysis, effectively bypassing the need for uniform control over the discrepancy between any two functions in the function class. The near-optimality of our algorithm is demonstrated by an $\widetilde{\Omega}(\epsilon^{-1})$ lower bound. Furthermore, we extend our algorithm to contextual dueling bandits and achieve a similar nearly optimal sample complexity.
1 Introduction
Due to the data-hungry and instable nature of reinforcement learning (RL), divergences that are straightforward to estimate via Monte Carlo methods or amenable to constrained optimization stand out from numerous candidates (Rényi, 1961; Müller, 1997; Basseville, 2013) as regularizers; the former family of which is typically $f$-divergence because any of them is an expectation under one policy (thus easy for on-policy estimation) (Levine, 2018; Levine et al., 2020), and the latter are naturally cases of Bregman divergence (Bregman, 1967) induced by strongly convex functions because they are suitable for scalable optimization algorithms over the probability simplex by design (Lattimore and Szepesvári, 2020; Lattimore, 2024). In particular, Kullback-Leibler (KL) divergence is the only one at the intersection of Bregman divergence and $f$-divergence (Jiao et al., 2014, Theorem 5), indicating that it is uniquely favorable among common choices from both computational and statistical aspects. Therefore, the KL-regularized RL objective is arguably the most popular one in practice:
$J(\pi)=\mathbb{E}_{\pi}[r]-\frac{1}{\eta}\mathsf{KL}({\pi}\|{\pi^{\mathsf{ref}}}),$ (1.1)
where $r$ is a reward function, $\pi$ is the policy to be optimized, $\pi^{\mathsf{ref}}$ is a reference policy, $\mathsf{KL}({\pi}\|{\pi^{\mathsf{ref}}})$ is the reverse KL divergence, and $\eta>0$ is the inverse temperature. When $\pi^{\mathsf{ref}}$ is chosen to be the uniform distribution, the objective function reduces to the (Shannon) entropy-regularized RL objective that encourages diverse actions and enhances robustness (Williams, 1992; Ziebart et al., 2008; Ziebart, 2010; Levine and Koltun, 2013; Fox et al., 2015; Levine et al., 2016; Haarnoja et al., 2018; Richemond et al., 2024; Liu et al., 2024). The KL-regularized objective has also been widely used in the RL fine-tuning of large language models (LLMs) (Ouyang et al., 2022; Rafailov et al., 2023, 2024), where $\pi^{\mathsf{ref}}$ is a pretrained LLM.
There has also been a surge of interest in understanding the principle behind KL-regularized RL. Ahmed et al. (2019); Liu et al. (2019) studied by ablation the effect of entropy regularization on the stability and policy improvement in policy optimization, the regret of which has been rigorously settled under the classic online mirror descent framework (Cai et al., 2020; He et al., 2022; Ji et al., 2023). Neu et al. (2017) unified popular KL-regularized policy optimization algorithms under a convex optimization framework, but the interplay of which with data were left untouched. A series of work (Geist et al., 2019; Vieillard et al., 2020; Kozuno et al., 2022) then analyzed the sample complexity of algorithms using KL w.r.t. the previous iteration or/and entropy regularizers with improved dependence on the effective horizon in discounted Markov decision processes (MDPs). However, the metric of suboptimality in all these studies is still with respect to the unregularized reward maximization objective, and its optimal sample complexity is $O(\epsilon^{-2})$ for finding the $\epsilon$-optimal policy.
Several recent studies (Xiong et al., 2024; Xie et al., 2024; Zhao et al., 2024) switched the focus to suboptimality guarantee with respect to the regularized objective in (1.1). In particular, Xiong et al. (2024) proposed an Offline GSHF algorithm via the principle of pessimism in the face of uncertainty, and proved $O(\epsilon^{-2})$ sample complexity under single-policy concentrability (See Section 2.1 for detailed definitions of concentrability.). On the other hand, the sharp analysis in Zhao et al. (2024) yields the optimal sample complexity $O(\epsilon^{-1})$, but requires all-policy concentrability (Zhao et al., 2024, Definition 2.6), i.e., the data-generating policy $\pi^{\mathsf{ref}}$ is required to cover the entire function class for all possible policies. Xie et al. (2024) studied token-level Markov decision processes (MDPs) and proposed a KL-regularized RL algorithm named XPO, which achieves $O(\epsilon^{-2})$ sample complexity under all-policy concentrability. In fact, even for offline contextual bandits (Foster and Rakhlin, 2023, Chapter 3) with $\mathrm{i.i.d.}$ data, which is among the simplest setting for offline decision making, existing sample complexity under KL-regularization are still primitive. Thus, a natural question arises:
Is the $\widetilde{O}(\epsilon^{-1})$ sample complexity w.r.t. 1.1 achievable for offline KL-regularized contextual bandits under single-policy concentrability?
We answer this question affirmatively by proposing a novel yet minimalist approach based on the principle of pessimism (Jin et al., 2021). Our algorithm can deal with general reward function which belongs to a reward function class with finite covering entropy. The optimal sample complexity is also due to our proof technique, which leverages the strong convexity of KL regularization and the conditional non-negativity of the gap between the true reward and its pessimistic estimator to refine a mean-value-type risk upper bound to its extreme. This in turn leads to a novel covariance-based analysis, effectively bypassing the need for uniform control over the discrepancy between any two functions in the function class.
Table 1: Comparison of sample complexity for finding $\epsilon$-optimal policy for offline contextual bandits. $\mathrm{polylog}(\epsilon^{-1})$ and the dependence of other parameters are omitted here. C(D)Bs stands for contextual (dueling) bandits. The last column indicates whether the bounds hold under single-policy concentrability.
Reference [^1] The four existing upper bounds are adapted from the implicit form in Xiong et al. (2024, Theorem 3.1) and Zhao et al. (2024, Theorem 3.3 and Theorem 4.4.). See Section 2.2 for details. CBs CDBs Single?
Xiong et al. (2024) $\widetilde{O}(\epsilon^{-2})$ $\widetilde{O}(\epsilon^{-2})$
Zhao et al. (2024) $\widetilde{O}(\epsilon^{-1})$ $\widetilde{O}(\epsilon^{-1})$
This work $\widetilde{O}(\epsilon^{-1})$ $\widetilde{O}(\epsilon^{-1})$
Lower bounds $\widetilde{\Omega}(\epsilon^{-1})$ $\widetilde{\Omega}(\epsilon^{-1})$  [^2] The lower bounds in terms of $\epsilon$ are the same under both single- and all-policy concentrability, which are detailed in Remarks 2.11 and 3.8 respectively.
1.1 Contributions
  • We propose a new algorithm dubbed KL-PCB and develop a novel proof technique that simultaneously utilizes the curvature of KL-regularized objectives and integrates pessimism with a novel covariance-based argument, for deriving its $\widetilde{O}(\epsilon^{-1})$ sample complexity under single-policy concentrability.
  • To demonstrate the versatility of our techniques beyond absolute reward feedback, we also extend our results to the offline contextual dueling bandit (CDB) setting, and provide a tight sample complexity analysis.
  • The matching lower bounds (up to $\mathrm{polylog}$ factors) in both settings suggest that our sample complexity results are nearly optimal, which renders a complete picture of the sample complexity for KL-regularized offline contextual under single policy concentrability.
For the ease of comparison, we summarize the relevant sample complexity results along with the corresponding coverage assumptions in Table 1.
1.2 Additional Related Work
We review two additional lines of theoretical progress that are closely related to our algorithm design and analysis.
Pessimism in Offline RL
The principle of pessimism has been underpinning offline RL for both the tabular (Rashidinejad et al., 2021) and function approximation (Jin et al., 2021) settings under the name of lower confidence bound (LCB). For contextual bandits, it is behind the adaptively optimal sample complexity analysis (Li et al., 2022). Shi et al. (2022) proposed a LCB-based model-free algorithm for tabular RL with near-optimal guarantee. Jin et al. (2021); Xiong et al. (2022); Di et al. (2023) utilized LCB in conjunction with the classic least-square value iteration paradigm to derive $\widetilde{O}(\epsilon^{-2})$ sample complexity results for model-free RL with function approximation. The line of work from Rashidinejad et al. (2021); Xie et al. (2021b) to Li et al. (2024) settled the sample complexity of tabular model-based RL via pessimistic estimators exploiting the variance information. It is also possible to leverage the idea of pessimism to design model-based algorithms under general function approximation that are at least statistically efficient (Xie et al., 2021a; Uehara and Sun, 2021; Wang et al., 2024).
However, in terms of risk decomposition, to the best of our knowledge, none of these pessimism-based analyses really goes beyond the performance difference lemma (Foster and Rakhlin, 2023, Lemma 13) or simulation lemma (Foster and Rakhlin, 2023, Lemma 23); both of which are not able to capture the strong concavity of KL-regularized objectives even in the bandit setting. The algorithmic idea of using pessimistic least-square estimators under general function approximation in Jin et al. (2021); Di et al. (2023) is similar to ours, but their suboptimality gap is bounded by the sum of bonuses, the direct adaptation of which to our objective cannot lead to the desired sample complexity.
Offline Contextual Dueling Bandits
CDBs (Dudík et al., 2015) is the contextual extension of dueling bandits in classic literature of online learning from pairwise comparisons (Yue et al., 2012; Zoghi et al., 2014). Since the empirical breakthrough of preference-based RL fine-tuning of LLMs (Ouyang et al., 2022), the theory of offline CDBs has received more attention under linear (Zhu et al., 2023; Xiong et al., 2024) and general (Zhan et al., 2022; Zhao et al., 2024; Song et al., 2024; Huang et al., 2024) function approximation. Preference models without stochastic transitivity (Munos et al., 2023; Ye et al., 2024; Wu et al., 2024; Zhang et al., 2024) are beyond the scope of this work, viz., our preference labels are assumed to follow the Bradley-Terry Model (Bradley and Terry, 1952).
Notation
Boldfaced lower case (resp. upper case) letters are reserved for vectors (resp. matrices). The sets ${\mathcal{S}}$ and $\mathcal{A}$ are assumed to be countable throughout this paper to avoid measure-theoretic arguments. For nonnegative sequences $\{x_{n}\}$ and $\{y_{n}\}$, we write $x_{n}=O(y_{n})$ if $\limsup_{n\to\infty}{x_{n}}/{y_{n}}<\infty$, and write $y_{n}=\Omega(x_{n})$ if $x_{n}=O(y_{n})$. We further employ $\widetilde{O}(\cdot)$ and $\widetilde{\Omega}(\cdot)$ to hide $\mathrm{polylog}$ factors. Given a positive definite $\bm{\Sigma}\in\mathbb{R}^{d\times d}$ and $\mathbf{x}\in\mathbb{R}^{d}$, we denote the vector’s Euclidean norm by $\|\mathbf{x}\|_{2}$ and define $\|\mathbf{x}\|_{\bm{\Sigma}}=\sqrt{\mathbf{x}^{\top}\bm{\Sigma}\mathbf{x}}$. For coutable $\mathcal{X}$ and $\mathcal{Y}$, we denote the family of probability kernels from $\mathcal{X}$ to $\mathcal{Y}$ by $\Delta(\mathcal{Y}|\mathcal{X})$. For a pair of probability measures $P\ll Q$ on the same space, the KL divergence from $P$ to $Q$ is ${\mathsf{KL}}\left(P\|Q\right)\coloneqq\int\log({{\,\mathrm{d}}P}/{{\,\mathrm{d}}Q}){\,\mathrm{d}}P$, and their total variation (TV) distance is ${\mathsf{TV}}\left(P\|Q\right)\coloneqq 0.5\int|{\,\mathrm{d}}P-{\,\mathrm{d}}Q|$. For $f:\mathcal{X}\to\mathbb{R}$, its infinity norm is denoted by $\|f\|_{\infty}\coloneqq\sup_{x\in\mathcal{X}}|f(x)|$.
2 KL-Regularized Contextual Bandits
In this section, we propose a pessimism-based algorithm, dubbed PCB-KL, for offline KL-regularized contextual bandits. In the following subsections, we also showcase our key technical novelty which couples the pessimism of our reward estimator to the non-trivial curvature property of KL regularization.
2.1 Problem Setup
We consider contextual bandit, which is denoted by a tuple $({\mathcal{S}},\mathcal{A},r,\pi^{\mathsf{ref}})$. Specifically, ${\mathcal{S}}$ is the context space, $\mathcal{A}$ is the action space and $r:{\mathcal{S}}\times\mathcal{A}\to[0,1]$ is the reward function. In the offline setting, the agent only has access to an $\mathrm{i.i.d.}$ dataset $\mathcal{D}=\{(s_{i},a_{i},r_{i})\}_{i=1}^{n}$. Here $s_{i}\in{\mathcal{S}}$ is the state sampled from a distribution $\rho$, $a_{i}\in\mathcal{A}$ is the action taken from a behavior policy, and $r_{i}$ is the observed reward given by $r_{i}=r(s_{i},a_{i})+\varepsilon_{i}$, where $\varepsilon_{t}$ is $1$-sub-Gaussian (Lattimore and Szepesvári, 2020, Definition 5.2). In this work, we consider the KL-regularized objective
$J(\pi)\coloneqq\mathbb{E}_{(s,a)\sim\rho\times\pi}\bigg{[}r(s,a)-\eta^{-1}\log\frac{\pi(a|s)}{\pi^{\mathsf{ref}}(a|s)}\bigg{]},$ (2.1)
where $\pi^{\mathsf{ref}}$ is a known reference policy and the “inverse temperature” $\eta$ controls the intensity of regularization. For simplicity, we assume that $\pi^{\mathsf{ref}}$ is also the behavior policy that generates the dataset $\mathcal{D}$. This type of “behavior regularization” has been also been studied in Zhan et al. (2022). The unique optimal policy $\pi^{*}\coloneqq\mathop{\mathrm{argmax}}_{\pi\in\Delta(\mathcal{A}|{\mathcal{S}})}J(\pi)$ is in the following closed-form (See, e.g., Zhang 2023, Proposition 7.16)
$\pi^{*}(\cdot|s)\propto\exp\big{(}\eta\cdot r(s,\cdot)\big{)},\forall s\in{\mathcal{S}}.$ (2.2)
A policy $\pi$ is said to be $\epsilon$-optimal if $J(\pi^{*})-J(\pi)\leq\epsilon$ and the goal of the agent is to find one such policy based on $\mathcal{D}$. To ensure $\epsilon$-optimality is achievable, we assume that $r$ lies in a known function class $\mathcal{F}\subset({\mathcal{S}}\times\mathcal{A}\to[0,1])$, from which the agent obtains her estimator $\widehat{f}$.
Assumption 2.1 (General Function Approximation).
There exists a known function class $\mathcal{F}:{\mathcal{S}}\times\mathcal{A}\to[0,1]$ such that $\exists\ f^{*}\in\mathcal{F}$ such that $f^{*}=r$.
We also need a standard condition to control the complexity of $\mathcal{F}$ through the notion of covering number (Wainwright, 2019, Definition 5.1).
Definition 2.2 ($\epsilon$-net and covering number).
Given a function class $\mathcal{G}\subset({\mathcal{S}}\times\mathcal{A}\to\mathbb{R})$, a finite set $\mathcal{G}(\epsilon)\subset\mathcal{G}$ is an $\epsilon$-net of $\mathcal{G}$ w.r.t. $\|\cdot\|_{\infty}$, if for any $g\in\mathcal{G}$, there exists $g^{\prime}\in\mathcal{G}(\epsilon)$ such that $\|g-g^{\prime}\|_{\infty}\leq\epsilon$. The $\epsilon$-covering number is the smallest cardinality $\mathcal{N}_{\mathcal{G}}(\epsilon)$ of such $\mathcal{G}(\epsilon)$.
Assumption 2.3.
For any $\epsilon_{c}>0$, the $\epsilon_{c}$-covering number $\mathcal{N}_{\mathcal{F}}(\epsilon_{c})$ of $\mathcal{F}$ is $\mathrm{poly}(\epsilon_{c}^{-1})$.
Remark 2.4.
When $\mathcal{F}$ is the class of linear functions of dimension $d$ and radius $R$, the covering number of $\mathcal{F}$ is given by $\mathcal{N}_{\mathcal{F}}=O((1+R\epsilon^{-1})^{d})$ (Jin et al., 2020, Lemma D.6), indicating that the Assumption 2.3 on $\mathcal{N}_{\mathcal{F}}$ is mild.
Concentrability
The data quality of $\mathcal{D}$ collected by $\pi^{\mathsf{ref}}$ is typically characterized by concentrability in offline RL (Farahmand et al., 2010; Chen and Jiang, 2019; Jiang and Xie, 2024). In the traditional definition based on density ratio (See, e.g., Zhan et al. 2022, Table 1), a strong concentrability condition intuitively corresponds to a large $\text{supp}(\pi^{\mathsf{ref}})$, which reduces the agent’s difficulty of learning since (with high probability) $\widehat{\pi}$ can hardly underfit on many state-action pairs. The two extreme cases of concentrability coefficients, all- and single-policy concentrability, intuitively correspond to (1) $\text{supp}(\pi^{\mathsf{ref}})$ covers all possible inputs; and (2) $\text{supp}(\pi^{\mathsf{ref}})$ only subsumes $\text{supp}(\pi^{*})$. Our definition of concentrability, however, is different from the density ratio-based one, because we take the function class $\mathcal{F}$ into account. In detail, we first introduce $D^{2}$-divergence as follows.
Definition 2.5.
Given a function class $\mathcal{F}\subset({\mathcal{S}}\times\mathcal{A}\to\mathbb{R})$ and a fixed policy $\pi$, define the $D^{2}$-divergence $D^{2}_{\mathcal{F}}((s,a);\pi)$ as
$\sup_{f,g\in\mathcal{F}}\frac{\big{(}f(s,a)-g(s,a)\big{)}^{2}}{\mathbb{E}_{(s^{\prime},a^{\prime})\sim\rho\times\pi}[(f(s^{\prime},a^{\prime})-g(s^{\prime},a^{\prime}))^{2}]}.$
The “Eluder dimension”-type Definition 2.5 is directly inspired by Di et al. (2023); Zhao et al. (2024), the intuition behind which is that given $(s,a)\in{\mathcal{S}}\times\mathcal{A}$, a small $D^{2}$-divergence indicates that for two functions $f$ and $g$, if they are close under the behavior policy $\pi$, then they will also be close on such pair $(s,a)$. Therefore, the $D^{2}$-divergence quantifies how well the estimation on dataset collected by the behavior policy $\pi$ can be generalized to a specific state-action pair. We are now ready to define the two notions of concentrability conditions.
Assumption 2.6 (All-Policy Concentrability).
Given a reference policy $\pi^{\mathsf{ref}}$, there exists $D>0$ such that $D^{2}=\sup_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}D^{2}_{\mathcal{F}}((s,a);\pi^{\mathsf{ref}})$.
Assumption 2.6 indicates that the errors on any state-action pairs can be bounded by the error on the samples from $\rho\times\pi$ up to a factor $D$.
Assumption 2.7 (Single-Policy Concentrability).
Given a reference policy $\pi^{\mathsf{ref}}$, there exists $D_{\pi^{*}}>0$ such that $D_{\pi^{*}}^{2}=\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}[D^{2}_{\mathcal{F}}((s,a);\pi^{\mathsf{ref}})]$.
Single-policy concentrability indicates that the errors on the distributions of state-action pairs $\rho\times\pi$ can be bounded by the error on the samples from $\rho\times\pi$ up to some constant. The single-policy concentrability assumption is strictly weaker than the all-policy concentrability in Assumption 2.6.
2.2 Review of Existing Results
In this subsection, we discuss the direct adaptation of previous results to our setting and demonstrate that their results cannot imply an $\widetilde{O}(\epsilon^{-1})$ sample complexity without all-policy concentrability. Specifically, Xiong et al. (2024) obtained a performance gap upper bound under linear function approximation
$J(\pi^{*})-J(\pi)\leq\big{\|}\mathbb{E}_{\rho\times\pi^{*}}[\bm{\phi}(s,a)]-\bm{\nu}\big{\|}_{\bm{\Sigma}_{\text{off}}^{-1}}\eqqcolon\text{RHS},$
where $\bm{\nu}$ is the reference vector, $\bm{\phi}(s,a)\in\mathbb{R}^{d}$ is the feature map, and $\Sigma_{\text{off}}=\sum_{i=1}^{n}\bm{\phi}(s_{i},a_{i})\bm{\phi}(s_{i},a_{i})^{\top}$ is the sample covariance matrix. However, we can show that RHS can be bounded from below by
$\big{\|}\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}[\bm{\phi}(s,a)]-\bm{\nu}\big{\|}\sqrt{\lambda_{\min}(\bm{\Sigma}_{\text{off}}^{-1})}$ $=\big{\|}\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}[\bm{\phi}(s,a)]-\bm{\nu}\big{\|}\lambda_{\max}(\bm{\Sigma}_{\text{off}})^{-1/2}$
$\geq\big{\|}\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}[\bm{\phi}(s,a)]-\bm{\nu}\big{\|}\text{tr}(\bm{\Sigma}_{\text{off}})^{-1/2}$
$=\big{\|}\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}[\bm{\phi}(s,a)]-\bm{\nu}\big{\|}\bigg{(}\sum_{i=1}^{n}\|\bm{\phi}(s_{i},a_{i})\|_{2}^{2}\bigg{)}^{-1/2}$
$=\Omega(n^{-1/2}),$
where $\lambda_{\min}$ and $\lambda_{\max}$ is the minimum and maximum eigenvalue of a matrix, the first inequality holds due to the fact that $\mathbf{x}^{\top}\bm{\Sigma}\mathbf{x}\geq\|\mathbf{x}\|_{2}^{2}\lambda_{\min}(\bm{\Sigma})$ and the second inequality holds due to $\lambda_{\max}(\bm{\Sigma})\leq\text{tr}(\bm{\Sigma})$. Zhao et al. (2024) proposed a two-stage learning algorithm and obtained an $\widetilde{O}(\epsilon^{-1})$ sample complexity for online KL-regularized bandits. The algorithm can be adopted to offline learning by removing the second stage [^3] This can be done by setting the $n$ in their paper to 00. and treat the samples from first stage as the offline dataset. An analogous analysis gives a sample complexity of $\widetilde{O}(D^{2}\epsilon^{-1})$, where $D^{2}$ is the all-policy concentrability.
2.3 Algorithm
In this subsection, we present an offline bandit algorithm, KL-PCB, for KL-regularized contextual bandits in Algorithm 1. KL-PCB first leverages least-square estimator to find a function ${\bar{f}}\in\mathcal{F}$ that minimizes its risk on the offline dataset. In Zhao et al. (2024), such ${\bar{f}}$ is directly applied to construct the estimated policy. In contrast, we construct a pessimistic estimator of $f^{*}$ following the well-known pessimism principle in offline RL (Jin et al., 2021).
Specifically, we consider the bonus term defined as
$\Gamma_{n}(s,a)=\beta D_{\mathcal{F}}\big{(}(s,a),\pi^{\mathsf{ref}}\big{)},\forall(s,a)\in{\mathcal{S}}\times\mathcal{A};$ (2.3)
where
$\beta=\sqrt{128\log\big{(}2\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta\big{)}/3n+18\epsilon_{c}}.$ (2.4)
We then obtain our pessimistic estimation ${\widehat{f}}$ by setting ${\widehat{f}}={\bar{f}}-\Gamma_{n}$, which is less than $f^{*}$ with high probability. After obtaining the pessimistic estimation, KL-PCB output the policy $\widehat{\pi}$, which maximizes the estimated objective
$\widehat{J}(\pi)=\mathbb{E}_{(s,a)\sim\rho\times\pi}\bigg{[}{\widehat{f}}(s,a)-\eta^{-1}\log\frac{\pi(a|s)}{\pi^{\mathsf{ref}}(a|s)}\bigg{]},$
the maximizer of which is the counterpart of 2.2, i.e.,
$\widehat{\pi}(a|s)\propto\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta\cdot{\widehat{f}}(s,a)\big{)}.$
Algorithm 1 Offline KL-Regularized Pessimistic Contextual Bandit (KL-PCB)
0:  regularization $\eta$, reference policy $\pi^{\mathsf{ref}}$, function class $\mathcal{F}$, offline dataset $\mathcal{D}=\{(s_{i},a_{i},r_{i})\}_{i=1}^{n}$
1:  Compute the least square estimation of reward function
${\bar{f}}\in\mathop{\mathrm{argmin}}_{f\in\mathcal{F}}\sum_{(s_{i},a_{i},r_{i})\in\mathcal{D}}\big{(}f(s_{i},a_{i})-r_{i}\big{)}^{2}$
2:  Let ${\widehat{f}}\leftarrow{\bar{f}}-\Gamma_{n}$, where $\Gamma_{n}$ is the bonus term in 2.3
2:  $\widehat{\pi}(a|s)\propto\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta\cdot{\widehat{f}}(s,a)\big{)}$
2.4 Theoretical Results
The sample complexity for KL-regularized contextual bandits is settled in this subsection. We first give the upper bound of KL-PCB.
Theorem 2.8.
Under Assumption 2.7, for sufficiently small $\epsilon\in(0,1)$, if we set $\Gamma_{n}$ as in (2.3), then $n=\widetilde{O}(\eta D_{\pi^{*}}^{2}\epsilon^{-1})$ suffices to guarantee the output policy $\widehat{\pi}$ of Algorithm 1 to be $\epsilon$-optimal with probability at least $1-\delta$.
Remark 2.9.
Previously, Zhao et al. (2024) achieved an $\widetilde{O}(\epsilon^{-1})$ sample complexity under Assumption 2.6. As a comparison, KL-PCB achieves the same $\widetilde{O}(\epsilon^{-1})$ sample complexity but only requiring Assumption 2.7, which is weaker than Assumption 2.6
We also provide the sample complexity lower bound of KL-regularized contextual bandits in the following theorem.
Theorem 2.10.
For any $\epsilon\in(0,1),\eta>0$, and algorithm Alg, there is a KL-regularized contextual bandit instance with $O(1)$ single-policy concentrability such that Alg requires at least $\Omega\big{(}\min\{{\eta\log\mathcal{N}_{\mathcal{F}}(\epsilon)}/{\epsilon},{\log\mathcal{N}_{\mathcal{F}}(\epsilon)}/{\epsilon^{2}}\}\big{)}$ samples to return an $\epsilon$-optimal policy.
Remark 2.11.
Theorem 2.10 is concluded by plugging $D^{2}\geq D^{2}_{\pi^{*}}$ to Theorem 3.1 of Zhao et al. (2024) and thus the sample complexity lower bound also holds for the KL-regularized contextual bandit instance with $O(1)$ all-policy concentrability.
Remark 2.12.
Theorem 2.10 shows that when $\epsilon$ is sufficiently small, any algorithm for offline KL-regularized contextual bandits requires at least $\widetilde{\Omega}(\epsilon^{-1})$ samples to output an $\epsilon$-optimal policy. This sample complexity lower bound matches the sample complexity upper bound in Theorem 2.8, indicating that KL-PCB is nearly optimal.
2.5 Proof of Theorem 2.8
In this section, we finish the proof of Theorem 2.8. At a high level, if we consider the regularized objective 1.1 multi-arm bandits, then $P\mapsto{\mathsf{KL}}\left(P\|Q\right)$ is $1$-strongly convex w.r.t. ${\mathsf{TV}}\left(\cdot\|\cdot\right)$ (Polyanskiy and Wu, 2024, Exercise I.37), and thus $J(\pi)$ is strongly concave. Therefore, $J(\pi^{*})-J(\widehat{\pi})$ is possible to be of the order $[{\mathsf{TV}}\left(\pi^{*}\|\widehat{\pi}\right)]^{2}\approx\widetilde{O}(n^{-1})$, pretending that $\pi^{*}$ is the unconstrained maximizer. This intuition guides our analysis for contextual bandits.
We begin the proof with the definition of the event $\mathcal{E}(\delta)$ given $\delta>0$ as
$\mathcal{E}(\delta)\coloneqq\bigg{\{}\sup_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}\Big{[}\big{|}{\bar{f}}-f^{*}\big{|}-\Gamma_{n}\Big{]}(s,a)\leq 0\bigg{\}},$ (2.5)
where $\Gamma_{n}$ is defined in (2.3). Event $\mathcal{E}(\delta)$ holds indicates that the least square estimation ${\bar{f}}$ obtained in Line 1 of Algorithm 1 does not deviate too much from the true function $f^{*}$. More specifically, we have the following lemma, whose proof are deferred to Please refer to Appendix A.1.
Lemma 2.13.
For all $\delta>0$, $\mathcal{E}(\delta)$ holds with probability at least $1-\delta$.
The following covariance-type observation is the first pivot.
Lemma 2.14.
If a bounded random variable $X\leq 0$ almost surely, then $\mathbb{E}[X^{3}]-\mathbb{E}[X^{2}]\mathbb{E}[X]\leq 0$.
Remark 2.15.
While this lemma is elementary, to the best of our knowledge, we are the first to isolate this structure from our non-standard analysis of offline RL, from which the sharp upper bound is derived.
Proof of Lemma 2.14.
We define $Y=-X$. Then it suffices to show that the covariance between $Y$ and $Y^{2}$ is
$\operatorname{\rm Cov}(Y,Y^{2})$ $=\mathbb{E}[Y^{3}]-\mathbb{E}[Y^{2}]\mathbb{E}[Y]$
$\geq\big{(}\mathbb{E}[Y^{2}]\big{)}^{3/2}-\mathbb{E}[Y^{2}]\mathbb{E}[Y]$
$=\big{(}\mathbb{E}[Y^{2}]\big{)}\big{(}\sqrt{\mathbb{E}[Y^{2}]}-\mathbb{E}[Y]\big{)}$
$\geq 0,$
where both inequalities follow from Jensen’s inequality. ∎
We further define the following quantities. For all $\gamma\in[0,1]$, we define
$f_{\gamma}\coloneqq\gamma{\widehat{f}}+(1-\gamma)f^{*},$
and based on which we further denote
$\pi_{\gamma}(\cdot|s)\propto\pi^{\mathsf{ref}}(\cdot|s)\exp\big{(}\eta f_{\gamma}(s,\cdot)\big{)},\forall s\in{\mathcal{S}};$
$G(\gamma)\coloneqq\mathbb{E}_{\rho\times\pi_{\gamma}}\Big{[}\big{(}{\widehat{f}}-{f^{*}}\big{)}^{2}(s,a)\Big{]}.$
The key to our analysis is the monotonicity of the function $G(\gamma)$ in $\gamma$, which is formally stated in the following lemma.
Lemma 2.16.
On event $\mathcal{E}$, $0\in\mathop{\mathrm{argmax}}_{\gamma\in[0,1]}G(\gamma)$.
Proof.
For simplicity, we use $\triangle(s,a)$ to denote $\big{(}{\widehat{f}}-{f^{*}})(s,a)$ in this proof. Then we know that $\triangle(s,a)\leq 0$ for all $(s,a)\in{\mathcal{S}}\times\mathcal{A}$ on event $\mathcal{E}$. The most direct way to prove is to take derivative of $G$ with respect to $\gamma$, which corresponds to the policy gradient (Sutton et al., 1999) of $\pi_{\gamma}$ and thus implying a favorable structure. A direct calculation yields that
$G^{\prime}(\gamma)$ $=\mathbb{E}_{\rho\times\pi_{\gamma}}\big{[}\nabla_{\gamma}\log\pi_{\gamma}(a|s)\triangle(s,a)^{2}\big{]}$
$=\eta\mathbb{E}_{\rho}\mathbb{E}_{a\sim\pi_{\gamma}}\big{[}\triangle^{2}(s,a)\big{(}\triangle(s,a)-\mathbb{E}_{a^{\prime}\sim\pi_{\gamma}}[\triangle(s,a^{\prime})]\big{)}\big{]}$
$=\eta\mathbb{E}_{\rho}\Big{[}\mathbb{E}_{\pi_{\gamma}}\big{[}\triangle^{3}(s,a)\big{]}-\mathbb{E}_{\pi_{\gamma}}\big{[}\triangle^{2}(s,a)\big{]}\mathbb{E}_{\pi_{\gamma}}\big{[}\triangle(s,a)\big{]}\Big{]}$
$\leq 0,$
where $\mathbb{E}_{\rho}$ is the shorthand of $\mathbb{E}_{s\sim\rho}$, $\mathbb{E}_{\pi_{\gamma}}$ is the shorthand of $\mathbb{E}_{a\sim\pi_{\gamma}}$, the first equation is derived from standard policy gradient and the inequality holds conditioned on the event $\mathcal{E}(\delta)$ due to Lemma 2.14. ∎
Finally, we need the following lemma to bound the performance difference between two policy w.r.t. the KL-regularized objective $J$. For any given $f:{\mathcal{S}}\times\mathcal{A}\to\mathbb{R}$, we define $\pi_{f}(\cdot|s)\propto\exp\big{(}\eta\cdot f(s,\cdot)\big{)},\forall s\in{\mathcal{S}}$ to facilitate presentation; then $\pi^{*}=\pi_{f^{*}}$ by definition. Please refer to Appendix A.2 for the proof of the lemma.
Lemma 2.17.
Given $f:{\mathcal{S}}\times\mathcal{A}\to\mathbb{R}$, there exist $\gamma\in[0,1]$ such that for $f_{\gamma}=\gamma f+(1-\gamma)f^{*}$ and $\pi_{\gamma}\coloneqq\pi_{f_{\gamma}}$,
$J(\pi^{*})-J(\pi_{f})\leq\eta\mathbb{E}_{(s,a)\sim\rho\times\pi_{\gamma}}\big{[}\big{(}f^{*}-f\big{)}^{2}(s,a)\big{]}.$
Now we are ready to prove Theorem 2.8.
Proof of Theorem 2.8.
Following the proof of Zhao et al. (2024, Theorem 3.3), we know that there exists $\bar{\gamma}\in[0,1]$ such that
$J(\pi^{*})-J(\widehat{\pi})\leq\eta G(\bar{\gamma})\leq\eta G(0),$ (2.6)
where the first inequality holds due to Lemma 2.17 and the second inequality holds due to the event $\mathcal{E}$ and Lemma 2.16. The term $G(0)$ can be further bounded by
$G(0)$ $=\eta\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}\Big{[}\big{(}{\widehat{f}}-{f^{*}}\big{)}^{2}(s,a)\Big{]}$
$\leq 4\eta\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}[\Gamma_{n}^{2}(s,a)]$
$=4\eta\beta^{2}\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}\big{[}D^{2}_{\mathcal{F}}((s,a);\pi^{\mathsf{ref}})\big{]}$
$=\widetilde{O}(\eta D_{\pi^{*}}^{2}n^{-1}),$ (2.7)
where the second inequality holds conditioned on $\mathcal{E}(\delta)$ because of Lemma 2.16, and the last inequality follows from the definition of $\mathcal{E}(\delta)$ together with Line 2. By Lemma 2.13, we know that event $\mathcal{E}$ holds with probability at least $1-\delta$, which finishes the proof. ∎
3 KL-Regularized Contextual Dueling Bandits
In this section, we extend our algorithm to KL-regularized dueling bandit setting, where the learner receives preference feedback instead of absolute signals. We formally introduce setting and extended algorithm. Except the objective, the core setting follows Zhu et al. (2023); Zhan et al. (2023). Our objective follows Xiong et al. (2024); Zhao et al. (2024).
3.1 Problem Setup
We still [^4] We overload some notations in Section 2 by their dueling counterparts for notational simplicity. consider contextual bandits $({\mathcal{S}},\mathcal{A},r,\pi^{\mathsf{ref}})$ where ${\mathcal{S}}$ is the state space, $\mathcal{A}$ is the action space and $r:{\mathcal{S}}\times\mathcal{A}\to[0,1]$ is the reward function. But only relative preference feedback is available, viz., we have an $\mathrm{i.i.d.}$ offline dataset $\mathcal{D}=\{(s_{i},a^{1}_{i},a^{2}_{i},y_{i})\}_{i=1}^{n}$, where $s_{i}\in{\mathcal{S}}$ is generated from distribution $\rho$ and $a^{1}_{i},a^{2}_{i}\sim\pi^{\mathsf{ref}}$. The binary preference label $y_{i}=1$ indicates $a_{i}^{1}$ is preferred over $a_{i}^{2}$ and 00 for the opposite. Given a state $s$ and two corresponding action $a^{1}$ and $a^{2}$, let $y$ be the preference label such that $y=1$ indicates $a^{1}\succ a^{2}$ and $y=0$ for $a^{2}\succ a^{1}$. In this work we consider the Bradley-Terry Model, where $\mathbb{P}[y=1|s,a^{1},a^{2}]=\sigma(r(s_{i},a^{1}_{i})-r(s_{i},a^{2}_{i}))$, where $\sigma(x)=(1+e^{-x})^{-1}$ is the link function. The objective here identical to (2.1) and thus the optimal policy $\pi^{*}$ also follows the closed form in (2.2). Our goal is still to find an $\epsilon$-optimal policy. To control the complexity of the function class $\mathcal{F}$, we assume that Assumption 2.1 still holds here.
Concentrability
Analogous to Section 2, we need our estimation from offline dataset generalizable to the state-action pairs visited by our obtained policy, and therefore requires similar $D^{2}$-divergence and corresponding policy coverage conditions. However, in dueling bandit, we are unable to observe and estimate the absolute value of rewards. Therefore, intuitively, the best estimation $f$ we can achieve is that for any state $s$ and actions $a^{1},a^{2}$, our estimated $f(s,a^{1})-f(s,a^{2})\approx r(s,a^{1})-r(s,a^{2})$. This implies that there exists some mapping $b:{\mathcal{S}}\to[-1,1]$ such that $f(s,a)-b(s)\approx r(s,a)$ on the offline data. Consequently, it is natural to define the the following modified $D^{2}$-divergence for characterizing how the error $f(s,a)-b(s)-r(s,a)$ generalize to unseen samples.
Definition 3.1.
Given a class of functions $\mathcal{F}\subset({\mathcal{S}}\times\mathcal{A}\to\mathbb{R})$ and some policy $\pi$, let $\mathcal{B}=({\mathcal{S}}\to[-1,1])$ be the function class, define the $D^{2}$-divergence $D^{2}_{\mathcal{F}}((s,a);\pi)$ as
$\sup_{f,g\in\mathcal{F}}\inf_{b\in\mathcal{B}}\frac{\big{(}f(s,a)-g(s,a)-b(s)\big{)}^{2}}{\mathbb{E}_{s\sim\rho}\operatorname{{\rm Var}}_{a^{\prime}\sim\pi(\cdot|s^{\prime})}[f(s^{\prime},a^{\prime})-g(s^{\prime},a^{\prime})]}.$
A similar definition has been introduced in Zhao et al. (2024, Definition 2.6), which underpins the following two assumptions characterizing the coverage ability of $\pi^{\mathsf{ref}}$.
Equipped with the overloaded $D^{2}$-divergence definition, we are now able to define all-policy concentrability and single concentrability similar as in Section 2.
Assumption 3.2 (All-Policy Concentrability).
Given a reference policy $\pi^{\mathsf{ref}}$, there exists $D>0$ such that $D^{2}=\sup_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}D^{2}_{\mathcal{F}}((s,a);\pi^{\mathsf{ref}})$.
Assumption 3.3 (Single-Policy Concentrability).
Given a reference policy $\pi^{\mathsf{ref}}$, there exists $D_{\pi^{*}}>0$ such that $D_{\pi^{*}}^{2}=\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}[D^{2}_{\mathcal{F}}((s,a);\pi^{\mathsf{ref}})]$.
Remark 3.4.
Similar single-policy concentrability assumptions have appeared in previous work in offline contextual dueling bandits. Specifically, Huang et al. (2024) defined an $L_{\infty}$-concentrability coefficient $\mathcal{C}_{\infty}^{\pi^{*}}=\sup_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}\big{(}\frac{\pi^{*}(a|s)}{\pi^{\mathsf{ref}}(a|s)}\big{)}$. Song et al. (2024) introduced a stronger assumption that $\mathcal{C}_{\infty}^{\pi}$ is bounded for all the policy $\pi$ in the KL-ball surrounding $\pi^{*}$. In our Definition 3.1, instead of simply considering the ratio of probability measure, we introduce a more refined eluder-type uncertainty term which also relies on the structure of the model class. As a result, $D_{\pi^{*}}$ can be small even if $L_{\infty}$-concentrability coefficient of $\pi^{*}$ is extremely large as long as the reward model generalizes well after training on the offline dataset. There has also been similar notions of partial coverage for the analysis of model-based RL (Uehara and Sun, 2021; Wang et al., 2024), which fused the information of the transition kernel class to the density-ratio-based formulation.
3.2 Algorithm
We elucidate KL-PCDB for offline KL-regularized contextual dueling bandits, whose pseudocode is summarized in Algorithm 2. KL-PCDB first estimate the ground truth function $f^{*}$ on offline dataset with maximum likelihood estimator (MLE) to estimate a function ${\bar{f}}\in\mathcal{F}$. After that, analogous to Algorithm 1, we adopt the principle of pessimism in the face of uncertainty. Specifically, we define the penalty term
$\Gamma_{n}(s,a)=\beta\sqrt{D_{\mathcal{F}}^{2}((s,a),\pi^{\mathsf{ref}})},$ (3.1)
where
$\beta^{2}=128\log(2\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)/3n+18\epsilon_{c}=\widetilde{O}(n^{-1})$ (3.2)
and then subtract it from the MLE ${\bar{f}}$ to obtain a pessimistic estimator ${\widehat{f}}$. KL-PCB then output the policy $\widehat{\pi}$, maximizing the estimated objective
$\widehat{J}(\pi)=\mathbb{E}_{(s,a)\sim\rho\times\pi}\bigg{[}{\widehat{f}}(s,a)-\eta^{-1}\log\frac{\pi(a|s)}{\pi^{\mathsf{ref}}(a|s)}\bigg{]},$
the maximizer of which is in closed form as the counterpart of 2.2.
$\widehat{\pi}(a|s)\propto\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta\cdot{\widehat{f}}(s,a)\big{)}.$
Algorithm 2 Offline KL-Regularized Pessimistic Contextual Dueling Bandit (KL-PCDB)
0:  regularization $\eta$, reference policy $\pi^{\mathsf{ref}}$, function class $\mathcal{F}$, offline dataset $\mathcal{D}=\{(s_{i},a^{1}_{i},a^{2}_{i},y_{i})\}_{i=1}^{n}$
1:  Compute the maximum likelihood estimator of the reward function
${\bar{f}}=\mathop{\mathrm{argmin}}_{f\in\mathcal{F}}\sum_{i=1}^{n}\Big{[}y_{i}\log\sigma\Big{(}\big{[}f(s_{i},a^{1}_{i})-f(s_{i},a^{2}_{i})\big{]}\Big{)}+(1-y_{i})\log\sigma\Big{(}\big{[}f(s_{i},a^{2}_{i})-f(s_{i},a^{1}_{i})\big{]}\Big{)}\Big{]}$
2:  Let ${\widehat{f}}(s,a)={\bar{f}}(s,a)-\Gamma_{n}(s,a)$, where $\Gamma_{n}(s,a)$ is the bonus term in 3.1
2:  $\widehat{\pi}(a|s)\propto\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta\cdot{\widehat{f}}(s,a)\big{)}$
3.3 Main Results
In this subsection, we provide our main results for KL-regularized contextual dueling bandits. The sample complexity upper bound of KL-PCDB is presented in the following theorem.
Theorem 3.5.
Under Assumption 3.3, if we set $\Gamma_{n}$ according to Lemma 3.1, then for sufficiently small $\epsilon\in(0,1)$, with probability at least $1-\delta$, $n=\widetilde{O}(\eta D_{\pi^{*}}^{2}\epsilon^{-1})$ is sufficient to guarantee the output policy $\widehat{\pi}$ of Algorithm 2 to be $\epsilon$-optimal.
Remark 3.6.
Zhao et al. (2024) achieved an $\widetilde{O}(\epsilon^{-1})$ sample complexity under Assumption 3.2. Comparing to Zhao et al. (2024), KL-PCDB achieves the same $\widetilde{O}(\epsilon^{-1})$ sample complexity but only requiring Assumption 3.3, which is weaker than Assumption 3.2.
The following theorem provides the sample complexity lower bound for KL-regularized dueling contextual bandits.
Theorem 3.7.
For any $\epsilon\in(0,1)$, $\eta>0$, and offline RL algorithm Alg, there is a KL-regularized contextual dueling bandit instance with $O(1)$ single-policy concentrability such that Alg requires at least $\Omega\big{(}\min\{{\eta\log\mathcal{N}_{\mathcal{F}}(\epsilon)}/{\epsilon},{\log\mathcal{N}_{\mathcal{F}}(\epsilon)}/{\epsilon^{2}}\}\big{)}$ samples to return an $\epsilon$-optimal policy.
Proof.
See Appendix A.3. ∎
Remark 3.8.
Theorem 3.7 can be concluded by plugging $D^{2}\geq D^{2}_{\pi^{*}}$ into Zhao et al. (2024, Theorem 4.3). This implies that the sample complexity lower bound in Theorem 3.7 also holds for KL-regularized dueling contextual bandit with all-policy concentrability.
Remark 3.9.
Theorem 3.7 shows that when $\epsilon$ is sufficiently small, any algorithm for offline KL-regularized contextual dueling bandit requires at least $\widetilde{\Omega}(\epsilon^{-1})$ samples to output a $\epsilon$-optimal policy. This sample complexity lower bound matches the sample complexity upper bound in Theorem 3.5, indicating that KL-PCDB is nearly optimal.
3.4 Proof of Theorem 3.5
The proof follows the proof in Section 2. At the beginning, we first define the event $\mathcal{E}(\delta)$ given $\delta>0$ as
$\mathcal{E}(\delta)\coloneqq\Big{\{}\exists\ b:{\mathcal{S}}\to[-1,1],\forall(s,a)\in{\mathcal{S}}\times\mathcal{A},\big{|}{\bar{f}}(s,a)-b(s)-f^{*}(s,a)\big{|}\leq\Gamma_{n}(s,a)\Big{\}}.$ (3.3)
Here, $\Gamma_{n}$ is defined in (3.1). We abuse the notation and define $b(\cdot)$ as
$b=\mathop{\mathrm{argmin}}_{\mathcal{B}}\sup_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}\Phi_{b}(s,a)-\Gamma_{n}(s,a),$ (3.4)
where $\Phi_{b}(s,a)=\big{|}{\bar{f}}(s,a)-b(s)-f^{*}(s,a)\big{|}$ and when $\mathcal{E}$ holds, for all $(s,a)\in{\mathcal{S}}\times\mathcal{A}$, we have $\Phi_{b}(s,a)\leq\Gamma_{n}(s,a)$ This indicates that the least square estimation ${\bar{f}}$ obtained in Line 1 of Algorithm 2, after adjusted by some bias function $b$, is close to the true function $f^{*}$. The following lemma shows that this event holds with high probability. Please refer to Appendix B.1 for the proof.
Lemma 3.10.
For any $\delta>0$, $\mathbb{P}(\mathcal{E}(\delta))\geq 1-\delta$.
We overload the following quantities. For any $\gamma\in[0,1]$ and $(s,a)\in{\mathcal{S}}\times\mathcal{A}$, we define
$f_{\gamma}(s,a)\coloneqq\gamma({\widehat{f}}(s,a)-b(s))+(1-\gamma)f^{*}(s,a).$
Furthermore, we introduce the following quantities
$\pi_{\gamma}(\cdot|\cdot)=\pi_{f_{\gamma}}(\cdot|\cdot)\propto\pi^{\mathsf{ref}}(\cdot|\cdot)\exp\big{(}\eta f_{\gamma}(\cdot,\cdot)\big{)},$
$G(\gamma)\coloneqq\mathbb{E}_{\rho\times\pi_{\gamma}}\big{[}\big{(}{\widehat{f}}(s,a)-b(s)-{f^{*}}(s,a)\big{)}^{2}\big{]},$
where $b(\cdot)$ is defined in (3.4). We still have the monotonicity of the function $G(\gamma)$, which is characterized by the following lemma.
Lemma 3.11.
On event $\mathcal{E}(\delta)$, $0\in\mathop{\mathrm{argmax}}_{\gamma\in[0,1]}G(\gamma)$.
Proof.
For simplicity, we use $\triangle(s,a)$ to denote ${\widehat{f}}(s,a)-b(s)-{f^{*}}(s,a)$ in this proof. Then on event $\mathcal{E}(\delta)$, we know that $\triangle(s,a)\leq 0$ for all $(s,a)\in{\mathcal{S}}\times\mathcal{A}$. Taking derivatives of $G$ w.r.t., $\gamma$ directly, we conclude that for all $\gamma\in[0,1]$,
$G^{\prime}(\gamma)$ $=\eta\mathbb{E}_{\rho}\mathbb{E}_{a\sim\pi_{\gamma}}\big{[}\triangle^{2}(s,a)\big{(}\triangle(s,a)-\mathbb{E}_{a^{\prime}\sim\pi_{\gamma}}[\triangle(s,a^{\prime})]\big{)}\big{]}$
$=\eta\mathbb{E}_{\rho}\Big{[}\mathbb{E}_{\pi_{\gamma}}\big{[}\triangle^{3}(s,a)\big{]}-\mathbb{E}_{\pi_{\gamma}}\big{[}\triangle^{2}(s,a)\big{]}\mathbb{E}_{\pi_{\gamma}}\big{[}\triangle(s,a)\big{]}\Big{]}$
$\leq 0,$
where $\mathbb{E}_{\rho}$ is the shorthand of $\mathbb{E}_{s\sim\rho}$, $\mathbb{E}_{\pi_{\gamma}}$ is the shorthand of $\mathbb{E}_{a\sim\pi_{\gamma}}$ and the inequality holds conditioned on the event $\mathcal{E}(\delta)$ due to Lemma 2.14. ∎
Finally, we have the proposition that adding some bias term $b:{\mathcal{S}}\to\mathbb{R}$ does not affect the resulting policy.
Proposition 3.12.
Let $b:{\mathcal{S}}\to\mathbb{R}$ be some bias function, then for all $f\in\mathcal{F}$ we have $J(\pi_{f})=J(\pi_{f-b})$, where $(f-b)(s,a)=f(s,a)-b(s)$.
Please see Appendix B.2 for the proof of Proposition 3.12. Now we are ready to prove Theorem 3.5.
Proof of Theorem 3.5.
We proceed the proof under the event $\mathcal{E}(\delta)$. By Proposition 3.12, we know that
$J(\pi^{*})-J(\widehat{\pi})$ $=J(\pi^{*})-J(\pi_{{\widehat{f}}})$
$=J(\pi^{*})-J(\pi_{{\widehat{f}}-b}).$
Consequently, there exist some $\gamma\in[0,1]$ and $b:{\mathcal{S}}\to[-1,1]$ such that
$J(\pi^{*})-J(\widehat{\pi})$ $=J(\pi^{*})-J(\pi_{{\widehat{f}}-b})$
$\leq\eta\mathbb{E}_{\rho\times\pi_{\gamma}}\big{[}\big{(}{\widehat{f}}(s,a)-b(s)-{f^{*}}(s,a)\big{)}^{2}\big{]}$
$=\eta G(\gamma),$ (3.5)
where the inequality holds due to Lemma 2.17. Under event $\mathcal{E}(\delta)$, we know that ${\widehat{f}}(s,a)-b(s)\leq{f^{*}}(s,a)$. Together with Lemma 3.11, we obtain $G(\gamma)\leq G(0)$. Therefore, we know that
$J(\pi^{*})-J(\widehat{\pi})$ $\leq G(0)$ (3.6)
$=\eta\mathbb{E}_{\rho\times\pi^{*}}\Big{[}\big{(}{\widehat{f}}(s,a)-b(s)-{f^{*}}(s,a)\big{)}^{2}\Big{]}$
$\leq 4\eta\mathbb{E}_{\rho\times\pi^{*}}\Gamma_{n}^{2}(s,a)$
$=4\eta\beta^{2}\mathbb{E}_{\rho\times\pi^{*}}\big{[}D^{2}_{\mathcal{F}}((s,a);\pi^{\mathsf{ref}})\big{]}$
$=\widetilde{O}(\eta D_{\pi^{*}}^{2}n^{-1}),$ (3.7)
where the inequality holds due to the definition of $\mathcal{E}(\delta)$. Plugging (3.7) into (3.5), we know that $J(\pi^{*})-J(\widehat{\pi})$ has upper bound $\widetilde{O}(D_{\pi^{*}}^{2}n^{-1})$. By Lemma 3.10, event $\mathcal{E}$ with probability at least $1-\delta$, which concludes the proof. ∎
4 Conclusions and Future Work
In this work, we settled the sample complexity of KL-regularized offline contextual bandits. We first propose KL-PCB, which follows the principle of pessimism for reward estimation and yields the optimal policy accordingly. Our analysis leverages the curvature of KL-regularized objectives and integrates pessimism with a newly identified covariance-based observation, which in turn enables a neat refinement of a mean-value-type argument the its extreme point, showing that KL-PCB achieves a nearly optimal sample complexity under single-policy concentrability. The pair of novel techniques are decoupled from tricky algorithmic tweaks, and thus might be of independent interest. We then extended KL-PCB to offline KL-regularized contextual dueling bandit problem and introduced KL-PCDB, the analysis of which manifests the near-optimality of KL-PCDB under single-policy concentration. Currently, our focus is solely on the reverse-KL-regularized objective. It would be an interesting future direction to study more general regularization techniques like $f$-divergence (Csiszár, 1967) regularization, which includes both forward and reverse KL regularizations.
Appendix A Missing Proofs from Section 2
A.1 Proof of Lemma 2.13
We first provide the following lemmas.
Lemma A.1 (Zhao et al. 2024, Lemma C.1).
For any policy $\pi$ and state-action pairs $\{(s_{i},a_{i})\}_{i=1}^{m}$ generated i.i.d. from $\rho\times\pi$, and $\epsilon_{c}<1$, with probability at least $1-\delta$, for any $f_{1}$ and $f_{2}$ we have
$\mathbb{E}_{\rho\times\pi}\big{[}\big{(}f_{1}(s,a)-f_{2}(s,a)\big{)}^{2}\big{]}\leq\frac{2}{n}\sum_{i=1}^{n}\big{(}f_{1}(s_{i},a_{i})-f_{2}(s_{i},a_{i})\big{)}^{2}+\frac{32}{3n}\log(2\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+10\epsilon_{c},$
where $\mathcal{N}_{\mathcal{F}}(\epsilon_{c})$ is the $\epsilon_{c}$-covering number of $\mathcal{F}$.
Lemma A.2 (Zhao et al. 2024, Lemma C.2).
For arbitrary policy $\pi$ and dataset $\{(s_{i},a_{i},r_{i})\}_{i=1}^{m}$ generated i.i.d., from the product of $\pi$, $\rho$ and the Bradley-Terry Model; let ${\bar{f}}$ be the least square estimator of $f^{*}$, then for any $10<\epsilon_{c}<$ 1 and $\delta>0$, with probability at least $1-\delta$ we have
$\sum_{i=1}^{n}\big{(}{\bar{f}}(s_{i},a_{i})-f^{*}(s_{i},a_{i})\big{)}^{2}\leq 16\log(a\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+4n\epsilon_{c}.$
Proof of Lemma 2.13.
We have the following inequality
$\big{(}{\bar{f}}(s,a)-f^{*}(s,a)\big{)}^{2}$ $=\frac{\big{(}{\bar{f}}(s,a)-f^{*}(s,a)\big{)}^{2}}{\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{f}}(s,a)-f^{*}(s,a)\big{)}^{2}\big{]}}\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{f}}(s,a)-f^{*}(s,a)\big{)}^{2}\big{]}$
$\leq\sup_{f_{1},f_{2}\in\mathcal{F}}\frac{\big{(}f_{1}(s,a)-f_{2}(s,a)\big{)}^{2}}{\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}f_{1}(s,a)-f_{2}(s,a)\big{)}^{2}\big{]}}\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{f}}(s,a)-f^{*}(s,a)\big{)}^{2}\big{]}$
$=D_{\mathcal{F}}^{2}((s,a),\pi^{\mathsf{ref}})\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{f}}(s,a)-f^{*}(s,a)\big{)}^{2}\big{]},$ (A.1)
where the inequality holds by taking supremum to $f_{1},f_{2}\in\mathcal{F}$. Now we have
$\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{f}}(s,a)-f^{*}(s,a)\big{)}^{2}\big{]}$ $\leq\frac{2}{n}\sum_{i=1}^{n}\big{(}{\bar{f}}(s_{i},a_{i})-f^{*}(s_{i},a_{i})\big{)}^{2}+\frac{32}{3n}\log(2\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+10\epsilon_{c}$
$\leq\frac{2}{n}\big{[}16\log(\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+4n\epsilon_{c}\big{]}+\frac{32}{3n}\log(2\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+10\epsilon_{c}$
$=\frac{128}{3n}\log(2\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+18\epsilon_{c},$ (A.2)
where the first inequality holds due to Lemma A.1 and second holds due to Lemma A.2. Plugging (A.1) into (A.1) and setting $\epsilon_{c}=O(n^{-1})$ complete the proof. ∎
A.2 Proof of Lemma 2.17
This proof is extracted from the proof of Zhao et al. (2024, Theorem 3.3) and we present it here for completeness. By definition of our objective in (2.1), we have
$J(\pi^{*})-J(\pi_{f})$
$\quad=\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}\bigg{[}f^{*}(s,a)-\eta^{-1}\log\frac{\pi^{*}(a|s)}{\pi^{\mathsf{ref}}(a|s)}\bigg{]}-\mathbb{E}_{(s,a)\sim\rho\times\pi_{f}}\bigg{[}f^{*}(s,a)-\frac{1}{\eta}\log\frac{\pi_{f}(a|s)}{\pi^{\mathsf{ref}}(a|s)}\bigg{]}$
$\quad=\frac{1}{\eta}\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}\bigg{[}\log\frac{\pi^{\mathsf{ref}}(a|s)\cdot\exp\bigl{(}\eta f^{*}(s,a)\bigr{)}}{\pi^{*}(a|s)}\bigg{]}-\frac{1}{\eta}\mathbb{E}_{(s,a)\sim\rho\times\pi_{f}}\bigg{[}\log\frac{\pi^{\mathsf{ref}}(a|s)\cdot\exp\bigl{(}\eta f^{*}(s,a)\bigr{)}}{\pi_{f}(a|s)}\bigg{]}$
$\quad=\frac{1}{\eta}\mathbb{E}_{s\sim\rho}\big{[}\log Z_{f^{*}}(s)\big{]}-\frac{1}{\eta}\mathbb{E}_{s\sim\rho}\big{[}\log Z_{f}(s)\big{]}-\mathbb{E}_{s\sim\rho}\bigg{[}\sum_{a\in\mathcal{A}}\pi_{f}(a|s)\cdot\big{(}f^{*}(s,a)-f(s,a)\big{)}\bigg{]},$
where for all $f\in\mathcal{F}$ we define $Z_{f}(\cdot)$ as follows,
$Z_{f}(\cdot)\coloneqq\sum_{a\in\mathcal{A}}\pi^{\mathsf{ref}}(a|\cdot)\exp\big{(}\eta f(\cdot,a)\big{)}.$
We further denote $\Delta(s,a)=f(s,a)-f^{*}(s,a)$ and $H_{s}(f)=\log Z_{f}(s)-\eta\sum_{a\in\mathcal{A}}\pi_{f}(a|s)\cdot\Delta(s,a)$. It worth noticing that $\eta^{-1}\mathbb{E}_{s\sim\rho}[H_{s}(f^{*})-H_{s}(f)]=J(\pi^{*})-J(\pi_{f})$. Now we take derivative of $H$ with respect to $\Delta(s,a)$,
$\frac{\partial H_{s}(f)}{\partial\Delta(s,a)}$ $=\frac{\partial}{\partial\Delta(s,a)}\bigg{[}\log Z_{f}(s)-\eta\sum_{a\in\mathcal{A}}\pi_{f}(a|s)\cdot\Delta(s,a)\bigg{]}$
$=\frac{1}{Z_{f}(s)}\cdot\pi^{\mathsf{ref}}(a|s)\exp\bigl{(}\eta\cdot f(s,a)\bigr{)}\cdot\eta-\eta\cdot\pi_{f}(a|s)$
$\quad-\eta^{2}\cdot\Delta(s,a)\cdot\frac{\pi^{\mathsf{ref}}(a|s)\cdot\exp\bigl{(}\eta\cdot f(s,a)\bigr{)}}{Z_{f}(s)}+\eta^{2}\cdot\Delta(s,a)\cdot\frac{\bigl{[}\pi^{\mathsf{ref}}(a|s)\cdot\exp\bigl{(}\eta\cdot f(s,a)\bigr{)}\bigr{]}^{2}}{[Z_{f}(s)]^{2}}$
$\quad+\eta\sum_{a^{\prime}\in\mathcal{A}\backslash\{a\}}\frac{\pi^{\mathsf{ref}}(a^{\prime}|x)\cdot\exp\bigl{(}\eta\cdot f(s,a^{\prime})\bigr{)}}{Z_{f}(s)}\cdot\eta\cdot\Delta(s,a^{\prime})\cdot\frac{\pi^{\mathsf{ref}}(a|s)\cdot\exp\bigl{(}\eta\cdot f(s,a)\bigr{)}}{Z_{f}(s)}$
$=-\eta^{2}\pi_{f}(a|s)\Delta(s,a)+\eta^{2}[\pi_{f}(a|s)]^{2}\cdot\Delta(s,a)+\eta^{2}\sum_{a^{\prime}\in\mathcal{A}\backslash\{a\}}\pi_{f}(a^{\prime}|x)\pi_{f}(a|s)\Delta(s,a^{\prime}).$
Therefore, by mean value theorem, there exists $\gamma\in[0,1]$ and $f_{\gamma}=\gamma f+(1-\gamma)f^{*}$ such that
$H_{s}(f)-H_{s}(f^{*})$ $=-\eta^{2}\gamma\sum_{a\in\mathcal{A}}\pi_{f_{\gamma}}(a|s)\Delta(s,a)^{2}+\gamma\eta^{2}\sum_{a_{1}\in\mathcal{A}}\sum_{a_{2}\in\mathcal{A}}\pi_{f_{\gamma}}(a_{1}|x)\pi_{f_{\gamma}}(a_{2}|x)\Delta(s,a_{1})\Delta(s,a_{2})$
$=-\eta^{2}\gamma\mathbb{E}_{a\sim\pi_{f_{\gamma}}}\big{[}\big{(}f^{*}(s,a)-f(s,a)\big{)}^{2}\big{]}+\gamma\eta^{2}\Big{(}\mathbb{E}_{a\sim\pi_{f_{\gamma}}}\big{[}\big{(}f^{*}(s,a)-f(s,a)\big{)}\big{]}\Big{)}^{2}$
$\geq-\eta^{2}\mathbb{E}_{a\sim\pi_{f_{\gamma}}}\big{[}\big{(}f^{*}(s,a)-f(s,a)\big{)}^{2}\big{]},$
where the inequality holds by omitting the second term and $\gamma\leq 1$. Now taking expectation over $\rho$, we have
$J(\pi^{*})-J(\pi_{f})$ $=\eta^{-1}\mathbb{E}_{s\sim\rho}[H_{s}(f^{*})-H_{s}(f)]$
$\leq\eta\mathbb{E}_{(s,a)\sim\rho\pi_{f_{\gamma}}}\big{[}\big{(}f^{*}(s,a)-f(s,a)\big{)}^{2}\big{]},$
which concludes the proof.
A.3 Proof of Theorem 2.10
Proof of Theorem 2.10.
We consider the bandit instance $\mathcal{K}$ and corresponding function class $\mathcal{F}$ constructed in the proof of Zhao et al. (2024, Theorem 3.1). Since the constructed $\pi^{\mathsf{ref}}(\cdot|s)$ in $\mathcal{K}$ is a uniform distribution for any $s\in{\mathcal{S}}$, the all-policy concentrability coefficient $D^{2}$ of $\mathcal{K}$ defined in Assumption 2.6 satisfies $D^{2}=O(1)$, which indicates that the single-policy concentrability coefficient $D_{\pi^{*}}^{2}$ defined in Assumption 2.7 satisfies $D_{\pi^{*}}^{2}\leq D^{2}=O(1)$. Thus, the proof of Zhao et al. (2024, Theorem 3.1) implies that any Alg requires at least $\Omega\big{(}\min\{{\eta\log\mathcal{N}_{\mathcal{F}}(\epsilon)}/{\epsilon},{\log\mathcal{N}_{\mathcal{F}}(\epsilon)}/{\epsilon^{2}}\}\big{)}$ to obtain an $\epsilon$-optimal policy. ∎
Appendix B Missing Proofs from Section 3
B.1 Proof of Lemma 3.10
The following on-poicy concentration bound is the prerequisite.
Lemma B.1 (Zhao et al. 2024, Lemma D.4).
Consider a offline dataset $\{(s_{i},a^{1}_{i},a^{2}_{i},y_{i})\}_{i=1}^{n}$ generated from the product of the context distribution $\rho\in\Delta({\mathcal{S}})$, policy $\pi\in\Delta(\mathcal{A}|{\mathcal{S}})$, and the Bradley-Terry Model defined in Section 3.1. Suppose ${\bar{f}}$ is the result of MLE estimation of Algorithm 2, and we further define $b(s)=\mathbb{E}_{a\sim\pi(\cdot|s)}\big{[}{\bar{f}}(s,a)-f^{*}(s,a)\big{]}$, then with probability at least $1-2\delta$, we have
$\mathbb{E}_{s,a\sim\rho\times\pi}\big{[}\big{(}{\bar{f}}(s,a)-f^{*}(s,a)-b(s)\big{)}^{2}\big{]}\leq O\bigg{(}\frac{1}{n}\log(\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+\epsilon_{c}\bigg{)}.$
Proof of Lemma 3.10.
From Lemma B.1, we have that with probability at least $1-\delta$, it holds that
$\mathbb{E}_{s^{\prime}\sim\rho}\operatorname{{\rm Var}}_{a^{\prime}\sim\pi^{\mathsf{ref}}(\cdot|s^{\prime})}\big{[}{\bar{f}}(s^{\prime},a^{\prime})-f^{*}(s^{\prime},a^{\prime})\big{]}\leq O\bigg{(}\frac{1}{n}\log(\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+\epsilon_{c}\bigg{)}.$ (B.1)
It further holds true that for some $b:{\mathcal{S}}\to\mathbb{R}$
$D_{\mathcal{F}}^{2}((s,a),\pi^{\mathsf{ref}}))\cdot\mathbb{E}_{s\sim\rho}\operatorname{{\rm Var}}_{a\sim\pi^{\mathsf{ref}}(\cdot|s)}\big{[}{\bar{f}}(s,a)-f^{*}(s,a)\big{]}\geq\big{(}{\bar{f}}(s,a)-b(s)-f^{*}(s,a)\big{)}^{2}.$ (B.2)
Substituting (B.1) into (B.2), we have
$\inf_{b}\big{(}{\bar{f}}(s,a)-b(s)-f^{*}(s,a)\big{)}^{2}$
$\quad=\inf_{b}\frac{\big{(}{\bar{f}}(s,a)-b(s)-f^{*}(s,a)\big{)}^{2}}{\mathbb{E}_{s^{\prime}\sim\rho}\operatorname{{\rm Var}}_{a^{\prime}\sim\pi^{\mathsf{ref}}(\cdot|s^{\prime})}\big{[}{\bar{f}}(s^{\prime},a^{\prime})-f^{*}(s^{\prime},a^{\prime})\big{]}}\mathbb{E}_{s^{\prime}\sim\rho}\operatorname{{\rm Var}}_{a^{\prime}\sim\pi^{\mathsf{ref}}(\cdot|s^{\prime})}\big{[}{\bar{f}}(s^{\prime},a^{\prime})-f^{*}(s^{\prime},a^{\prime})\big{]}$
$\quad\leq D_{\mathcal{F}}^{2}((s,a),\pi^{\mathsf{ref}})\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{f}}(s,a)-b(s)-f^{*}(s,a)\big{)}^{2}\big{]}$
$\quad\leq D_{\mathcal{F}}^{2}((s,a),\pi^{\mathsf{ref}})O\bigg{(}\frac{1}{n}\log(\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+\epsilon_{c}\bigg{)},$
where the first inequality holds due to the definition of $D_{\mathcal{F}}^{2}((s,a),\pi^{\mathsf{ref}})$ and the last inequality holds due to Lemma B.1. ∎
B.2 Proof of Proposition 3.12
Proof of Proposition 3.12.
For any fixed state $s\in{\mathcal{S}}$, we have for any $a\in\mathcal{A}$ that,
$\pi_{f}(a|s)$ $=\frac{\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta f(s,a)\big{)}}{\sum_{a^{\prime}\in\mathcal{A}}\pi^{\mathsf{ref}}(a^{\prime}|s)\exp\big{(}\eta f(s,a^{\prime})\big{)}}$
[…]
$=\frac{\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta[f(s,a)-b(s)]\big{)}}{\sum_{a^{\prime}\in\mathcal{A}}\pi^{\mathsf{ref}}(a^{\prime}|s)\exp\big{(}\eta[f(s,a^{\prime})-b(s)]\big{)}}$
$=\pi_{f-b}(a|s),$
which indicates that $\pi_{f}=\pi_{f-b}$. This immediately leads to $J(\pi_{f})=J(\pi_{f-b})$. ∎
B.3 Proof of Theorem 3.7
Proof.
Notice that Assumptions 3.2 and 3.3 imply $D^{2}\geq D^{2}_{\pi^{*}}$, which enables us to adapt Zhao et al. (2024, Theorem 4.3) to our scenario. In particular, we consider the bandit instance $\mathcal{K}$ and the corresponding function class $\mathcal{F}$ constructed in Zhao et al. (2024, Theorem 4.3). By Theorem 4.3 of Zhao et al. (2024), we know that the all-policy concentrability coefficient $D^{2}$ of $\mathcal{K}$ satisify $D^{2}=O(1)$, which indicates that its single-policy concentrability coefficient $D_{\pi^{*}}=O(1)$. Theorem 4.3 of Zhao et al. (2024) further indicates that Alg requires at least $\Omega\big{(}\min\{{\eta\log\mathcal{N}_{\mathcal{F}}(\epsilon)}/{\epsilon},{\log\mathcal{N}_{\mathcal{F}}(\epsilon)}/{\epsilon^{2}}\}\big{)}$ to obtain an $\epsilon$-optimal policy, which finishes the proof. ∎

### v2

@@
-Nearly Optimal Sample Complexity of Offline KL-Regularized Contextual Bandits under Single-Policy Concentrability
+Towards a Sharp Analysis of Offline Policy Learning for $f$-Divergence-Regularized Contextual Bandits
 Abstract
-KL-regularized policy optimization has become a workhorse in learning-based decision making, while its theoretical understanding is still very limited.
-Although recent progress has been made towards settling the sample complexity of KL-regularized contextual bandits, existing sample complexity bounds are either $\widetilde{O}(\epsilon^{-2})$ under single-policy concentrability or $\widetilde{O}(\epsilon^{-1})$ under all-policy concentrability.
-In this paper, we propose the first algorithm with $\widetilde{O}(\epsilon^{-1})$ sample complexity under single-policy concentrability for offline contextual bandits.
-Our algorithm is designed for general function approximation and based on the principle of pessimism in the face of uncertainty.
-The core of our proof leverages the strong convexity of the KL regularization, and the conditional non-negativity of the gap between the true reward and its pessimistic estimator to refine a mean-value-type risk upper bound to its extreme.
-This in turn leads to a novel covariance-based analysis, effectively bypassing the need for uniform control over the discrepancy between any two functions in the function class.
-The near-optimality of our algorithm is demonstrated by an $\widetilde{\Omega}(\epsilon^{-1})$ lower bound.
-Furthermore, we extend our algorithm to contextual dueling bandits and achieve a similar nearly optimal sample complexity.
+Although many popular reinforcement learning algorithms are underpinned by $f$-divergence regularization, their sample complexity with respect to the regularized objective still lacks a tight characterization.
+In this paper, we analyze $f$-divergence-regularized offline policy learning.
+For reverse Kullback–Leibler (KL) divergence, arguably the most commonly used one, we give the first $\widetilde{O}(\epsilon^{-1})$ sample complexity under single-policy concentrability for contextual bandits, surpassing existing $\widetilde{O}(\epsilon^{-1})$ bound under all-policy concentrability and $\widetilde{O}(\epsilon^{-2})$ bound under single-policy concentrability.
+Our analysis for general function approximation leverages the principle of pessimism in the face of uncertainty to refine a mean-value-type argument to its extreme.
+This in turn leads to a novel moment-based technique, effectively bypassing the need for uniform control over the discrepancy between any two functions in the function class.
+We further propose a lower bound, demonstrating that a multiplicative dependency on single-policy concentrability is necessary to maximally exploit the strong convexity of reverse KL.
+In addition, for $f$-divergences with strongly convex $f$, to which reverse KL does not belong, we show that the sharp sample complexity $\widetilde{\Theta}(\epsilon^{-1})$ is achievable even without single-policy concentrability.
+In this case, the algorithm design can get rid of pessimistic estimators.
+We further extend our analysis to dueling bandits, and we believe these results take a significant step toward a comprehensive understanding of $f$-divergence-regularized policy learning.
 1 Introduction
-Due to the data-hungry and instable nature of reinforcement learning (RL), divergences that are straightforward to estimate via Monte Carlo methods or amenable to constrained optimization stand out from numerous candidates (Rényi, 1961; Müller, 1997; Basseville, 2013) as regularizers; the former family of which is typically $f$-divergence because any of them is an expectation under one policy (thus easy for on-policy estimation) (Levine, 2018; Levine et al., 2020), and the latter are naturally cases of Bregman divergence (Bregman, 1967) induced by strongly convex functions because they are suitable for scalable optimization algorithms over the probability simplex by design (Lattimore and Szepesvári, 2020; Lattimore, 2024).
-In particular, Kullback-Leibler (KL) divergence is the only one at the intersection of Bregman divergence and $f$-divergence (Jiao et al., 2014, Theorem 5), indicating that it is uniquely favorable among common choices from both computational and statistical aspects.
-Therefore, the KL-regularized RL objective is arguably the most popular one in practice:
-$J(\pi)=\mathbb{E}_{\pi}[r]-\frac{1}{\eta}\mathsf{KL}({\pi}\|{\pi^{\mathsf{ref}}}),$ (1.1)
-where $r$ is a reward function, $\pi$ is the policy to be optimized, $\pi^{\mathsf{ref}}$ is a reference policy, $\mathsf{KL}({\pi}\|{\pi^{\mathsf{ref}}})$ is the reverse KL divergence, and $\eta>0$ is the inverse temperature.
-When $\pi^{\mathsf{ref}}$ is chosen to be the uniform distribution, the objective function reduces to the (Shannon) entropy-regularized RL objective that encourages diverse actions and enhances robustness (Williams, 1992; Ziebart et al., 2008; Ziebart, 2010; Levine and Koltun, 2013; Fox et al., 2015; Levine et al., 2016; Haarnoja et al., 2018; Richemond et al., 2024; Liu et al., 2024).
-The KL-regularized objective has also been widely used in the RL fine-tuning of large language models (LLMs) (Ouyang et al., 2022; Rafailov et al., 2023, 2024), where $\pi^{\mathsf{ref}}$ is a pretrained LLM.
-There has also been a surge of interest in understanding the principle behind KL-regularized RL.
-Ahmed et al.
-(2019); Liu et al.
-(2019) studied by ablation the effect of entropy regularization on the stability and policy improvement in policy optimization, the regret of which has been rigorously settled under the classic online mirror descent framework (Cai et al., 2020; He et al., 2022; Ji et al., 2023).
-Neu et al.
-(2017) unified popular KL-regularized policy optimization algorithms under a convex optimization framework, but the interplay of which with data were left untouched.
-A series of work (Geist et al., 2019; Vieillard et al., 2020; Kozuno et al., 2022) then analyzed the sample complexity of algorithms using KL w.r.t.
-the previous iteration or/and entropy regularizers with improved dependence on the effective horizon in discounted Markov decision processes (MDPs).
-However, the metric of suboptimality in all these studies is still with respect to the unregularized reward maximization objective, and its optimal sample complexity is $O(\epsilon^{-2})$ for finding the $\epsilon$-optimal policy.
-Several recent studies (Xiong et al., 2024; Xie et al., 2024; Zhao et al., 2024) switched the focus to suboptimality guarantee with respect to the regularized objective in (1.1).
-In particular, Xiong et al.
-(2024) proposed an Offline GSHF algorithm via the principle of pessimism in the face of uncertainty, and proved $O(\epsilon^{-2})$ sample complexity under single-policy concentrability (See Section 2.1 for detailed definitions of concentrability.).
-On the other hand, the sharp analysis in Zhao et al.
-(2024) yields the optimal sample complexity $O(\epsilon^{-1})$, but requires all-policy concentrability (Zhao et al., 2024, Definition 2.6), i.e., the data-generating policy $\pi^{\mathsf{ref}}$ is required to cover the entire function class for all possible policies.
-Xie et al.
-(2024) studied token-level Markov decision processes (MDPs) and proposed a KL-regularized RL algorithm named XPO, which achieves $O(\epsilon^{-2})$ sample complexity under all-policy concentrability.
-In fact, even for offline contextual bandits (Foster and Rakhlin, 2023, Chapter 3) with $\mathrm{i.i.d.}$ data, which is among the simplest setting for offline decision making, existing sample complexity under KL-regularization are still primitive.
-Thus, a natural question arises:
-Is the $\widetilde{O}(\epsilon^{-1})$ sample complexity w.r.t. 1.1 achievable for offline KL-regularized contextual bandits under single-policy concentrability?
-We answer this question affirmatively by proposing a novel yet minimalist approach based on the principle of pessimism (Jin et al., 2021).
-Our algorithm can deal with general reward function which belongs to a reward function class with finite covering entropy.
-The optimal sample complexity is also due to our proof technique, which leverages the strong convexity of KL regularization and the conditional non-negativity of the gap between the true reward and its pessimistic estimator to refine a mean-value-type risk upper bound to its extreme.
-This in turn leads to a novel covariance-based analysis, effectively bypassing the need for uniform control over the discrepancy between any two functions in the function class.
-Table 1: Comparison of sample complexity for finding $\epsilon$-optimal policy for offline contextual bandits.
-$\mathrm{polylog}(\epsilon^{-1})$ and the dependence of other parameters are omitted here.
-C(D)Bs stands for contextual (dueling) bandits.
-The last column indicates whether the bounds hold under single-policy concentrability.
-Reference [^1] The four existing upper bounds are adapted from the implicit form in Xiong et al.
+Due to the data-hungry and instable nature of reinforcement learning (RL), divergences that are straightforward to estimate via Monte Carlo or amenable to constrained optimization stand out from numerous candidates (Rényi, 1961; Csiszár, 1967; Müller, 1997; Basseville, 2013) as regularizers; the former family is typically $f$-divergence (Rényi, 1961) because any of them is an expectation, for which empirical average is a good proxy (Levine, 2018; Levine et al., 2020); and the latter class subsumes those with nice positive curvatures (e.g., Bregman divergence (Bregman, 1967) induced by strongly convex functions).
+In particular, Kullback-Leibler (KL) divergence is the only one at the intersection of $f$-divergence and Bregman divergence (Jiao et al., 2014, Theorem 5), indicating its theoretical advantage among common choices from both computational and statistical aspects.
+Also, the KL-regularized RL objective is arguably the most popular one in practice:
+$J(\pi)=\mathbb{E}_{\pi}[r]-\eta^{-1}\mathsf{KL}({\pi}\|{\pi^{\mathsf{ref}}}),$ (1.1)
+where $r$ is the reward, $\pi^{\mathsf{ref}}$ is a reference policy, $\mathsf{KL}({\pi}\|{\pi^{\mathsf{ref}}})$ is the reverse KL divergence, and $\eta>0$ is the inverse temperature.
+When $\pi^{\mathsf{ref}}$ is uniform, 1.1 reduces to the entropy-regularized objective that encourages diverse actions and enhances robustness (Williams, 1992; Ziebart et al., 2008; Levine and Koltun, 2013; Levine et al., 2016; Haarnoja et al., 2018; Richemond et al., 2024; Liu et al., 2024).
+KL regularization has also been widely used in the RL fine-tuning of large language models (Ouyang et al., 2022; Rafailov et al., 2023), where $\pi^{\mathsf{ref}}$ is the base model.
+Given its widespread use, there has been a surge of interest in understanding the role of KL regularization in RL by both empirical studies (Ahmed et al., 2019; Liu et al., 2019) and theoretical analysis (Geist et al., 2019; Vieillard et al., 2020; Kozuno et al., 2022).
+There are also lines of research on KL regularization in online learning (Cai et al., 2020; He et al., 2022; Ji et al., 2023) and convex optimization (Neu et al., 2017).
+However, the performance metric in most of these studies is still the unregularized reward maximization objective, under which the sample complexity is at least $\Omega(\epsilon^{-2})$.
+(See Section A.1 for detailed discussions)
+Table 1: Comparison of sample complexity bounds for finding $\epsilon$-optimal policy for offline contextual bandits with KL- and (strongly convex) $f$-divergence regularization.
+Constants and $\mathrm{polylog}$ factors are omitted here except the metric entropy $\log\mathcal{N}$.
+“Reverse-KL” stands for KL-regularized contextual bandits and “$f$-divergence w/ s.c., $f$” for the counterpart with an $\alpha$-strongly convex $f$.
+The two existing upper bounds are adapted from the implicit form in Xiong et al.
 (2024, Theorem 3.1) and Zhao et al.
-(2024, Theorem 3.3 and Theorem 4.4.).
-See Section 2.2 for details.
-CBs CDBs Single?
-Xiong et al. (2024) $\widetilde{O}(\epsilon^{-2})$ $\widetilde{O}(\epsilon^{-2})$
-Zhao et al. (2024) $\widetilde{O}(\epsilon^{-1})$ $\widetilde{O}(\epsilon^{-1})$
-This work $\widetilde{O}(\epsilon^{-1})$ $\widetilde{O}(\epsilon^{-1})$
-Lower bounds $\widetilde{\Omega}(\epsilon^{-1})$ $\widetilde{\Omega}(\epsilon^{-1})$  [^2] The lower bounds in terms of $\epsilon$ are the same under both single- and all-policy concentrability, which are detailed in Remarks 2.11 and 3.8 respectively.
+(2024, Theorem 3.3 and Theorem 4.4.), of which the detailed adaptions are deferred to Appendix A.
+Also see Remark 2.8 for discussion on the relationship between $D^{2}_{\pi^{*}}$ and $C^{\pi^{*}}$
+Regularizer Xiong et al. (2024) Zhao et al. (2024) This work
+Reverse KL Upper $d\epsilon^{-2}$ $\eta D^{2}\epsilon^{-1}\log\mathcal{N}$ $\eta D^{2}_{\pi^{*}}\epsilon^{-1}\log\mathcal{N}$
+Lower - $\eta\epsilon^{-1}\log\mathcal{N}$ $\eta C^{\pi^{*}}\epsilon^{-1}\log\mathcal{N}$
+$f$-divergence w/ s.c. $f$ Upper - - $\alpha^{-1}\eta\epsilon^{-1}\log\mathcal{N}$
+Lower - - $\alpha^{-1}\eta\epsilon^{-1}\log\mathcal{N}$
+Several recent papers (Xiong et al., 2024; Xie et al., 2024; Zhao et al., 2024; Foster et al., 2025) switched the focus to analyzing the suboptimality defined via the regularized objective (1.1), under which an $\Omega(\epsilon^{-1})$ sample complexity is possible (Zhao et al., 2024, Theorem 3.6).
+However, even restricted to the pure $\mathrm{i.i.d.}$ setting, existing analyses in this vein either result in still $\widetilde{O}(\epsilon^{-2})$ bounds (Xiong et al., 2024; Xie et al., 2024) or has a stringent all-policy concentrabiity dependency in their upper bounds (Zhao et al., 2024) (See Section 2.1 for details on coverage conditions).
+Thus, there are by far no perfectly matching bounds down to both the dependency of $\epsilon^{-1}$ and concrete coverage conditions for KL-regularized offline decision making.
+Moreover, these theoretical analyses set KL as the right target by default; but reverse KL is the $f$-divergence with $f(x)=x\log x$, which is merely convex.
+Therefore, it is also unknown whether those choices with even nicer (e.g., strongly convex) $f$, whose performance against the reward maximization objective can be provably promising (Zhan et al., 2022; Gabbianelli et al., 2024; Huang et al., 2024), can enjoy a better coverage dependency in the sample complexity under the corresponding regularized objectives.
+These perspectives motivate an open problem:
+What is the weakest coverage condition for offline policy learning to be minimax optimal with respect to $f$-divergence-regularized objectives?
+We are the first to concretely investigate this question by proposing sharp sample complexity analyses for two representative ad subclasses of $f$-divergence.
+In particular, for contextual bandits with KL regularization, we present sample complexity guarantee with linearly dependence on single-policy coverage ratio.
+Our lower bound further indicates that a multiplicative dependency on single-policy concentrability is necessary.
+Surprisingly, for $f$-divergence with $\alpha$-strongly-convex $f$, we prove nearly matching sample complexity bounds of $\widetilde{\Theta}(\alpha^{-1}\eta\epsilon^{-1})$ for the first time, settling the dependence on coverage completely.
+For the ease of comparison, we adapt previous sample complexity bounds to our setting and notation, and summarize them in Table 1.
 1.1 Contributions
-  • We propose a new algorithm dubbed KL-PCB and develop a novel proof technique that simultaneously utilizes the curvature of KL-regularized objectives and integrates pessimism with a novel covariance-based argument, for deriving its $\widetilde{O}(\epsilon^{-1})$ sample complexity under single-policy concentrability.
-  • To demonstrate the versatility of our techniques beyond absolute reward feedback, we also extend our results to the offline contextual dueling bandit (CDB) setting, and provide a tight sample complexity analysis.
-  • The matching lower bounds (up to $\mathrm{polylog}$ factors) in both settings suggest that our sample complexity results are nearly optimal, which renders a complete picture of the sample complexity for KL-regularized offline contextual under single policy concentrability.
-For the ease of comparison, we summarize the relevant sample complexity results along with the corresponding coverage assumptions in Table 1.
+  • For arguably the most commonly used KL regularization, we are the first to employ the principle of pessimism (Jin et al., 2021) to achieve a nearly tight sample complexity under single-policy concentrability.
+Our analysis roadmap leverages the strong convexity of KL and pessimism of the reward estimator, to refine a mean-value-type risk upper bound to its extreme which in turn leads to a novel moment-based analysis, effectively bypassing the need for uniform control over the discrepancy between any two functions in the function class.
+  • Under KL-regularization, we prove a lower bound growing linearly the density-ratio-based single-policy concentrability, which also strictly improved upon previous works (Zhao et al., 2024; Foster et al., 2025) in offline setting.
+This lower bound manifests the necessity of a multiplicative dependency on single-policy concentrability and firstly closes the gap between upper and lower bounds under a technically nontrivial case.
+  • For any $f$-divergence with strongly convex $f$, we are able to design a truly lightweight algorithm free of pessimism-based gadgets and still obtain the $\widetilde{\Theta}(\epsilon^{-1})$ sample complexity certified by a matching lower bound without coverage conditions.
+  • We extend all algorithmic ideas and hard instance constructions above to $f$-divergence-regularized contextual dueling bandits, achieving $\widetilde{\Theta}(\epsilon^{-1})$ sample complexity bounds.
+Moreover, all algorithms work for general reward function classes with finite metric entropy.
 1.2 Additional Related Work
 We review two additional lines of theoretical progress that are closely related to our algorithm design and analysis.
-Pessimism in Offline RL
+Pessimism in offline RL.
 The principle of pessimism has been underpinning offline RL for both the tabular (Rashidinejad et al., 2021) and function approximation (Jin et al., 2021) settings under the name of lower confidence bound (LCB).
@@
 (2021); Di et al.
-(2023) is similar to ours, but their suboptimality gap is bounded by the sum of bonuses, the direct adaptation of which to our objective cannot lead to the desired sample complexity.
-Offline Contextual Dueling Bandits
+(2023) is similar to ours, but their suboptimality gap is bounded by the sum of bonuses, which cannot lead to the desired sample complexity of our objective by direct adaption.
+Offline contextual dueling bandits.
 CDBs (Dudík et al., 2015) is the contextual extension of dueling bandits in classic literature of online learning from pairwise comparisons (Yue et al., 2012; Zoghi et al., 2014).
@@
 Preference models without stochastic transitivity (Munos et al., 2023; Ye et al., 2024; Wu et al., 2024; Zhang et al., 2024) are beyond the scope of this work, viz., our preference labels are assumed to follow the Bradley-Terry Model (Bradley and Terry, 1952).
-Notation
-Boldfaced lower case (resp.
-upper case) letters are reserved for vectors (resp.
-matrices).
-The sets ${\mathcal{S}}$ and $\mathcal{A}$ are assumed to be countable throughout this paper to avoid measure-theoretic arguments.
-For nonnegative sequences $\{x_{n}\}$ and $\{y_{n}\}$, we write $x_{n}=O(y_{n})$ if $\limsup_{n\to\infty}{x_{n}}/{y_{n}}<\infty$, and write $y_{n}=\Omega(x_{n})$ if $x_{n}=O(y_{n})$.
-We further employ $\widetilde{O}(\cdot)$ and $\widetilde{\Omega}(\cdot)$ to hide $\mathrm{polylog}$ factors.
-Given a positive definite $\bm{\Sigma}\in\mathbb{R}^{d\times d}$ and $\mathbf{x}\in\mathbb{R}^{d}$, we denote the vector’s Euclidean norm by $\|\mathbf{x}\|_{2}$ and define $\|\mathbf{x}\|_{\bm{\Sigma}}=\sqrt{\mathbf{x}^{\top}\bm{\Sigma}\mathbf{x}}$.
-For coutable $\mathcal{X}$ and $\mathcal{Y}$, we denote the family of probability kernels from $\mathcal{X}$ to $\mathcal{Y}$ by $\Delta(\mathcal{Y}|\mathcal{X})$.
-For a pair of probability measures $P\ll Q$ on the same space, the KL divergence from $P$ to $Q$ is ${\mathsf{KL}}\left(P\|Q\right)\coloneqq\int\log({{\,\mathrm{d}}P}/{{\,\mathrm{d}}Q}){\,\mathrm{d}}P$, and their total variation (TV) distance is ${\mathsf{TV}}\left(P\|Q\right)\coloneqq 0.5\int|{\,\mathrm{d}}P-{\,\mathrm{d}}Q|$.
-For $f:\mathcal{X}\to\mathbb{R}$, its infinity norm is denoted by $\|f\|_{\infty}\coloneqq\sup_{x\in\mathcal{X}}|f(x)|$.
+Notation.
+The sets ${\mathcal{S}}$ and $\mathcal{A}$ are assumed to be countable throughout the paper.
+For nonnegative sequences $\{x_{n}\}$ and $\{y_{n}\}$, we write $x_{n}=O(y_{n})$ if $\limsup_{n\to\infty}{x_{n}}/{y_{n}}<\infty$, $y_{n}=\Omega(x_{n})$ if $x_{n}=O(y_{n})$, and $y_{n}=\Theta(x_{n})$ if $x_{n}=O(y_{n})$ and $x_{n}=\Omega(y_{n})$.
+We further employ $\widetilde{O}(\cdot),\widetilde{\Omega}(\cdot)$, and $\widetilde{\Theta}$ to hide $\mathrm{polylog}$ factors.
+For countable $\mathcal{X}$ and $\mathcal{Y}$, we denote the family of probability kernels from $\mathcal{X}$ to $\mathcal{Y}$ by $\Delta(\mathcal{Y}|\mathcal{X})$.
+For $g:\mathcal{X}\to\mathbb{R}$, its infinity norm is denoted by $\|g\|_{\infty}\coloneqq\sup_{x\in\mathcal{X}}|g(x)|$.
+For a pair of probability measures $P\ll Q$ on the same space and function $f:\mathbb{R}_{+}\to\mathbb{R}$, their $f$-divergence is $D_{f}(P\|Q)\coloneqq\int f({\,\mathrm{d}}P/{\,\mathrm{d}}Q){\,\mathrm{d}}Q$.
+Specifically, when $f(x)=x\log x$, $f$-divergence becomes KL divergence denoted as ${\mathsf{KL}}\left(P\|Q\right)\coloneqq\int\log({{\,\mathrm{d}}P}/{{\,\mathrm{d}}Q}){\,\mathrm{d}}P$, and when $f(x)=|x-1|/2$, it becomes the total variation (TV) distance, which is denoted as ${\mathsf{TV}}\left(P\|Q\right)\coloneqq 0.5\int|{\,\mathrm{d}}P-{\,\mathrm{d}}Q|$.
 2 KL-regularized Contextual Bandits
@@
 In the offline setting, the agent only has access to an $\mathrm{i.i.d.}$ dataset $\mathcal{D}=\{(s_{i},a_{i},r_{i})\}_{i=1}^{n}$.
-Here $s_{i}\in{\mathcal{S}}$ is the state sampled from a distribution $\rho$, $a_{i}\in\mathcal{A}$ is the action taken from a behavior policy, and $r_{i}$ is the observed reward given by $r_{i}=r(s_{i},a_{i})+\varepsilon_{i}$, where $\varepsilon_{t}$ is $1$-sub-Gaussian (Lattimore and Szepesvári, 2020, Definition 5.2).
+Here $s_{i}^{\prime}s$ are states sampled from $\rho\in\Delta({\mathcal{S}})$, $a_{i}\in\mathcal{A}$ is the action taken from a behavior policy, and $r_{i}$ is the observed reward given by $r_{i}=r(s_{i},a_{i})+\varepsilon_{i}$, where $\varepsilon_{t}$ is $1$-sub-Gaussian (Lattimore and Szepesvári, 2020, Definition 5.2).
 In this work, we consider the KL-regularized objective
@@
 (2022).
-The unique optimal policy $\pi^{*}\coloneqq\mathop{\mathrm{argmax}}_{\pi\in\Delta(\mathcal{A}|{\mathcal{S}})}J(\pi)$ is in the following closed-form (See, e.g., Zhang 2023, Proposition 7.16)
-$\pi^{*}(\cdot|s)\propto\exp\big{(}\eta\cdot r(s,\cdot)\big{)},\forall s\in{\mathcal{S}}.$ (2.2)
+The unique optimal policy $\pi^{*}\coloneqq\mathop{\mathrm{argmax}}_{\pi\in\Delta(\mathcal{A}|{\mathcal{S}})}J(\pi)$ is given by (See, e.g., Zhang 2023, Proposition 7.16)
+$\pi^{*}(\cdot|s)\propto\pi^{\mathsf{ref}}(\cdot|s)\exp\big{(}\eta\cdot r(s,\cdot)\big{)},\forall s\in{\mathcal{S}}.$ (2.2)
 A policy $\pi$ is said to be $\epsilon$-optimal if $J(\pi^{*})-J(\pi)\leq\epsilon$ and the goal of the agent is to find one such policy based on $\mathcal{D}$.
-To ensure $\epsilon$-optimality is achievable, we assume that $r$ lies in a known function class $\mathcal{F}\subset({\mathcal{S}}\times\mathcal{A}\to[0,1])$, from which the agent obtains her estimator $\widehat{f}$.
-Assumption 2.1 (General Function Approximation).
-There exists a known function class $\mathcal{F}:{\mathcal{S}}\times\mathcal{A}\to[0,1]$ such that $\exists\ f^{*}\in\mathcal{F}$ such that $f^{*}=r$.
-We also need a standard condition to control the complexity of $\mathcal{F}$ through the notion of covering number (Wainwright, 2019, Definition 5.1).
+To ensure $\epsilon$-optimality is achievable, we assume that $r$ lies in a known function class ${\mathcal{G}}\subset({\mathcal{S}}\times\mathcal{A}\to[0,1])$, from which the agent obtains the estimator $\widehat{r}$.
+More specifically, we work with general function approximation under realizability, which is as follows.
+Assumption 2.1.
+For this known function class ${\mathcal{G}}\subset({\mathcal{S}}\times\mathcal{A}\to[0,1])$, $\exists{g^{*}}\in{\mathcal{G}}$ with ${g^{*}}=r$.
+We also need a standard condition to control the complexity of ${\mathcal{G}}$ through the notion of covering number (Wainwright, 2019, Definition 5.1).
 Definition 2.2 ($\epsilon$-net and covering number).
@@
 Assumption 2.3.
-For any $\epsilon_{c}>0$, the $\epsilon_{c}$-covering number $\mathcal{N}_{\mathcal{F}}(\epsilon_{c})$ of $\mathcal{F}$ is $\mathrm{poly}(\epsilon_{c}^{-1})$.
-Remark 2.4.
-When $\mathcal{F}$ is the class of linear functions of dimension $d$ and radius $R$, the covering number of $\mathcal{F}$ is given by $\mathcal{N}_{\mathcal{F}}=O((1+R\epsilon^{-1})^{d})$ (Jin et al., 2020, Lemma D.6), indicating that the Assumption 2.3 on $\mathcal{N}_{\mathcal{F}}$ is mild.
-Concentrability
-The data quality of $\mathcal{D}$ collected by $\pi^{\mathsf{ref}}$ is typically characterized by concentrability in offline RL (Farahmand et al., 2010; Chen and Jiang, 2019; Jiang and Xie, 2024).
-In the traditional definition based on density ratio (See, e.g., Zhan et al.
-2022, Table 1), a strong concentrability condition intuitively corresponds to a large $\text{supp}(\pi^{\mathsf{ref}})$, which reduces the agent’s difficulty of learning since (with high probability) $\widehat{\pi}$ can hardly underfit on many state-action pairs.
-The two extreme cases of concentrability coefficients, all- and single-policy concentrability, intuitively correspond to (1) $\text{supp}(\pi^{\mathsf{ref}})$ covers all possible inputs; and (2) $\text{supp}(\pi^{\mathsf{ref}})$ only subsumes $\text{supp}(\pi^{*})$.
-Our definition of concentrability, however, is different from the density ratio-based one, because we take the function class $\mathcal{F}$ into account.
+For any $\epsilon_{c}>0$, the $\epsilon_{c}$-covering number $\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})$ of ${\mathcal{G}}$ is $\mathrm{poly}(\epsilon_{c}^{-1})$.
+When ${\mathcal{G}}$ is the class of linear functions of dimension $d$ and radius $R$, the covering number of ${\mathcal{G}}$ is given by $\mathcal{N}_{{\mathcal{G}}}=O((1+R\epsilon^{-1})^{d})$ (Jin et al., 2020, Lemma D.6), indicating that the Assumption 2.3 on $\mathcal{N}_{{\mathcal{G}}}$ is mild.
+Concentrability.
+The data quality of $\mathcal{D}$ collected by $\pi^{\mathsf{ref}}$ is typically characterized by concentrability in offline RL (Farahmand et al., 2010; Chen and Jiang, 2019; Jiang and Xie, 2024), which quantifies the ability of the behavioral policy to generate diverse actions.
+We first define the density-ratio-based concentrability as follows.
+Definition 2.4 (Density-ratio-based concentrability).
+For policy class $\Pi$, reference policy $\pi^{\mathsf{ref}}$, the density-ratio-based all-policy concentrability $C^{\Pi}$ is $C^{\Pi}\coloneqq\sup_{\pi\in\Pi,s\in{\mathcal{S}},a\in\mathcal{A}}{\pi(a|s)}/{\pi^{\mathsf{ref}}(a|s)}$, whose single-policy counterpart under the optimal policy $\pi^{*}$ is $C^{\pi^{*}}\coloneqq\sup_{s\in{\mathcal{S}},a\in\mathcal{A}}{\pi^{*}(a|s)}/{\pi^{\mathsf{ref}}(a|s)}$.
+In the definition above, small all-policy concentrability intuitively corresponds to $\text{supp}(\pi^{\mathsf{ref}})$ covering all possible inputs.
+On the other hand, small single-policy concentrability means that $\text{supp}(\pi^{\mathsf{ref}})$ only subsumes $\text{supp}(\pi^{*})$.
+In this paper, in addition to density-ratio-based concentrability, we also adopt the following $D^{2}$-based concentrabilites to better capturing the nature of function class ${\mathcal{G}}$.
 In detail, we first introduce $D^{2}$-divergence as follows.
 Definition 2.5.
-Given a function class $\mathcal{F}\subset({\mathcal{S}}\times\mathcal{A}\to\mathbb{R})$ and a fixed policy $\pi$, define the $D^{2}$-divergence $D^{2}_{\mathcal{F}}((s,a);\pi)$ as
-$\sup_{f,g\in\mathcal{F}}\frac{\big{(}f(s,a)-g(s,a)\big{)}^{2}}{\mathbb{E}_{(s^{\prime},a^{\prime})\sim\rho\times\pi}[(f(s^{\prime},a^{\prime})-g(s^{\prime},a^{\prime}))^{2}]}.$
+Given a function class ${\mathcal{G}}\subset({\mathcal{S}}\times\mathcal{A}\to\mathbb{R})$ and a fixed policy $\pi$, define the $D^{2}$-divergence $D^{2}_{{\mathcal{G}}}((s,a);\pi)$ as
+$\sup_{g,h\in{\mathcal{G}}}\frac{\big{(}g(s,a)-h(s,a)\big{)}^{2}}{\mathbb{E}_{(s^{\prime},a^{\prime})\sim\rho\times\pi}[(g(s^{\prime},a^{\prime})-h(s^{\prime},a^{\prime}))^{2}]}.$
 The “eluder dimension”-type Definition 2.5 is directly inspired by Di et al.
 (2023); Zhao et al.
-(2024), the intuition behind which is that given $(s,a)\in{\mathcal{S}}\times\mathcal{A}$, a small $D^{2}$-divergence indicates that for two functions $f$ and $g$, if they are close under the behavior policy $\pi$, then they will also be close on such pair $(s,a)$.
+(2024), the intuition behind which is that given $(s,a)\in{\mathcal{S}}\times\mathcal{A}$, a small $D^{2}$-divergence indicates that for two functions $g$ and $h$, if they are close under the behavior policy $\pi$, then they will also be close on such pair $(s,a)$.
 Therefore, the $D^{2}$-divergence quantifies how well the estimation on dataset collected by the behavior policy $\pi$ can be generalized to a specific state-action pair.
@@
 Assumption 2.6 (All-policy concentrability).
-Given a reference policy $\pi^{\mathsf{ref}}$, there exists $D>0$ such that $D^{2}=\sup_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}D^{2}_{\mathcal{F}}((s,a);\pi^{\mathsf{ref}})$.
-Assumption 2.6 indicates that the errors on any state-action pairs can be bounded by the error on the samples from $\rho\times\pi$ up to a factor $D$.
+Given a reference policy $\pi^{\mathsf{ref}}$, there exists $D<\infty$ such that $D^{2}=\sup_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}D^{2}_{{\mathcal{G}}}((s,a);\pi^{\mathsf{ref}})$.
+Assumption 2.6 indicates that the errors on any state-action pairs can be bounded by the error on the samples from $\rho\times\pi$ up to a factor $D$, whose relaxed counterpart under the same $\pi^{\mathsf{ref}}$ is as follows.
 Assumption 2.7 (Single-policy concentrability).
-Given a reference policy $\pi^{\mathsf{ref}}$, there exists $D_{\pi^{*}}>0$ such that $D_{\pi^{*}}^{2}=\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}[D^{2}_{\mathcal{F}}((s,a);\pi^{\mathsf{ref}})]$.
-Single-policy concentrability indicates that the errors on the distributions of state-action pairs $\rho\times\pi$ can be bounded by the error on the samples from $\rho\times\pi$ up to some constant.
-The single-policy concentrability assumption is strictly weaker than the all-policy concentrability in Assumption 2.6.
-2.2 Review of Existing Results
-In this subsection, we discuss the direct adaptation of previous results to our setting and demonstrate that their results cannot imply an $\widetilde{O}(\epsilon^{-1})$ sample complexity without all-policy concentrability.
+$D_{\pi^{*}}^{2}\coloneqq\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}D^{2}_{{\mathcal{G}}}((s,a);\pi^{\mathsf{ref}})<\infty$.
+Assumption 2.7 indicates that the errors on the distributions of state-action pairs $\rho\times\pi^{*}$ can be bounded by the error on the samples from $\rho\times\pi^{\mathsf{ref}}$ up to some constant.
+For both types, the single-policy concentrability assumption is strictly weaker than the all-policy concentrability assumption.
+However, in general, the two quantities characterizing single-policy concentrability $C^{\pi^{*}}$ and $D^{2}_{\pi^{*}}$ cannot be bounded by each other up to constant factors.
+Remark 2.8.
+In particular, we have $D^{2}_{\pi^{*}}\leq|{\mathcal{S}}||\mathcal{A}|C^{\pi^{*}}$, indicating that $C^{\pi^{*}}$ subsumes $D^{2}_{\pi^{*}}$ when $|{\mathcal{S}}|$ and $|\mathcal{A}|$ can be seen as constant.
+2.2 Algorithm
+In this subsection, we present an offline bandit algorithm, KL-PCB, for KL-regularized contextual bandits in Algorithm 1.
+KL-PCB first leverages least-square estimator to find a function ${\bar{g}}\in{\mathcal{G}}$ that minimizes its risk on the offline dataset.
+In Zhao et al.
+(2024), such ${\bar{g}}$ is directly applied to construct the estimated policy.
+In contrast, we construct a pessimistic estimator of ${g^{*}}$ following the well-known pessimism principle in offline RL (Jin et al., 2021).
+Specifically, we define the bonus term $\Gamma_{n}$ through the confidence radius $\beta=\sqrt{128\log\big{(}2\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta\big{)}/3n+18\epsilon_{c}}$ as
+$\Gamma_{n}(s,a)=\beta D_{{\mathcal{G}}}\big{(}(s,a),\pi^{\mathsf{ref}}\big{)},\forall(s,a)\in{\mathcal{S}}\times\mathcal{A}.$ (2.3)
+We then obtain our pessimistic estimation ${\widehat{g}}$ by setting ${\widehat{g}}={\bar{g}}-\Gamma_{n}$, which is less than ${g^{*}}$ with high probability.
+After obtaining the pessimistic estimation, KL-PCB output the policy $\widehat{\pi}$, which maximizes the estimated objective
+$\widehat{J}(\pi)=\mathbb{E}_{(s,a)\sim\rho\times\pi}\bigg{[}{\widehat{g}}(s,a)-\eta^{-1}\log\frac{\pi(a|s)}{\pi^{\mathsf{ref}}(a|s)}\bigg{]},$
+the maximizer of which is the counterpart of 2.2, i.e.,
+$\widehat{\pi}(a|s)\propto\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta\cdot{\widehat{g}}(s,a)\big{)}.$
+Algorithm 1 Offline KL-Regularized Pessimistic Contextual Bandits (KL-PCB)
+0:  regularization $\eta$, reference policy $\pi^{\mathsf{ref}}$, offline dataset $\mathcal{D}$, function class ${\mathcal{G}}$
+1:  Compute the least square estimation of reward function ${\bar{g}}\in\mathop{\mathrm{argmin}}_{g\in{\mathcal{G}}}\sum_{(s_{i},a_{i},r_{i})\in\mathcal{D}}\big{(}g(s_{i},a_{i})-r_{i}\big{)}^{2}$
+${\bar{g}}\in\mathop{\mathrm{argmin}}_{g\in{\mathcal{G}}}\sum_{(s_{i},a_{i},r_{i})\in\mathcal{D}}\big{(}g(s_{i},a_{i})-r_{i}\big{)}^{2}$
+2:  Let ${\widehat{g}}\leftarrow{\bar{g}}-\Gamma_{n}$, where $\Gamma_{n}$ is the bonus term in 2.3
+2:  $\widehat{\pi}(a|s)\propto\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta\cdot{\widehat{g}}(s,a)\big{)}$
+2.3 Theoretical Results
+The sample complexity for KL-regularized contextual bandits is settled in this subsection. We first give the upper bound of KL-PCB.
+Theorem 2.9.
+Under Assumption 2.7, for sufficiently small $\epsilon\in(0,1)$, if we set $\Gamma_{n}$ as in (2.3), then $n=\widetilde{O}\big{(}\eta D^{2}_{\pi^{*}}\epsilon^{-1}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon)\big{)}$ suffices to guarantee the output policy $\widehat{\pi}$ of Algorithm 1 to be $\epsilon$-optimal with probability at least $1-\delta$.
+Previously, Zhao et al.
+(2024) achieved an $\widetilde{O}(\epsilon^{-1})$ sample complexity under Assumption 2.6.
+As a comparison, KL-PCB achieves the same $\widetilde{O}(\epsilon^{-1})$ sample complexity but only requiring Assumption 2.7, which is weaker than Assumption 2.6.
+We also provide the sample complexity lower bound of KL-regularized contextual bandits in the following theorem.
+Theorem 2.10.
+For $\forall S\geq 1$, $\eta>4\log 2$, $C^{*}\in(2,\exp(\eta/4)]$, and any algorithm Alg, there is a KL-regularized contextual bandit with $C^{\pi^{*}}\leq C^{*}$ such that Alg requires $\Omega\big{(}\min\{\eta\epsilon^{-1},\epsilon^{-2}\}C^{*}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon)\big{)}$ samples to find an $\epsilon$-optimal policy for sufficiently small $\epsilon$.
+Previously, Zhao et al.
+(2024) provided a sample complexity lower bound of $\Omega(\eta\log\mathcal{N}_{{\mathcal{G}}}(\epsilon)/{\epsilon})$ under KL-regularized objective.
+Foster et al.
+(2025) also provided a lower bound of $\Omega(C^{\pi^{*}})$ for KL-regularized objective to show the necessity of coverage.
+Compared to their results, our result shows that the multiplicative dependency on $C^{\pi^{*}}$ is necessary for the first time.
+Remark 2.11.
+Theorem 2.10 shows that when $\epsilon$ is sufficiently small, any algorithm for offline KL-regularized contextual bandits requires at least $\Omega(\eta C^{\pi^{*}})\epsilon^{-1}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon))$ samples to output an $\epsilon$-optimal policy.
+The presence of $\exp(\mathrm{poly}(\eta))$ in the range of $C^{*}$ is inevitable, since we always have $C^{\pi^{*}}\leq\exp(\eta)$ in reverse KL regularized bandits with bounded rewards.
+As discussed in Remark 2.8, when $|{\mathcal{S}}|$ and $|\mathcal{A}|$ are constant, $D^{2}_{\pi^{*}}\lesssim C^{\pi^{*}}$, indicating that KL-PCB is near-optimal under this special case.
+2.4 Proof of Theorem 2.9
+In this section, we finish the proof of Theorem 2.9.
+At a high level, if we consider the regularized objective 1.1 multi-arm bandits, then $P\mapsto{\mathsf{KL}}\left(P\|Q\right)$ is $1$-strongly convex w.r.t.
+${\mathsf{TV}}\left(\cdot\|\cdot\right)$ (Polyanskiy and Wu, 2025, Exercise I.37), and thus $J(\pi)$ is strongly concave.
+Therefore, $J(\pi^{*})-J(\widehat{\pi})$ is possible to be of the order $[{\mathsf{TV}}\left(\pi^{*}\|\widehat{\pi}\right)]^{2}\approx\widetilde{O}(n^{-1})$, pretending that $\pi^{*}$ is the unconstrained maximizer.
+This intuition guides our analysis for contextual bandits.
+We begin the proof with the definition of the event $\mathcal{E}(\delta)$ given $\delta>0$ as
+$\mathcal{E}(\delta)\coloneqq\Big{\{}\sup\nolimits_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}\Big{[}\big{|}{\bar{g}}-{g^{*}}\big{|}-\Gamma_{n}\Big{]}(s,a)\leq 0\Big{\}},$ (2.4)
+where $\Gamma_{n}$ is defined in (2.3).
+Event $\mathcal{E}(\delta)$ holds indicates that the least square estimation ${\bar{g}}$ obtained in Line 1 of Algorithm 1 does not deviate too much from the true function ${g^{*}}$.
+More specifically, we have the following lemma, whose proof are deferred to Appendix B.1.
+Lemma 2.12.
+For all $\delta>0$, $\mathcal{E}(\delta)$ holds with probability at least $1-\delta$.
+The following covariance-type observation is the first pivot.
+Lemma 2.13.
+If a bounded random variable $X\leq 0$ almost surely, then $\mathbb{E}[X^{3}]-\mathbb{E}[X^{2}]\mathbb{E}[X]\leq 0$.
+Remark 2.14.
+While this lemma is elementary, to the best of our knowledge, we are the first to isolate this structure from our non-standard analysis of offline RL, from which the sharp upper bound is derived.
+The intuition behind Lemma 2.13 is very natural: $X$ and $X^{2}$ cannot be positively correlated.
+We further define the following quantities. For all $\gamma\in[0,1]$, we define $g_{\gamma}\coloneqq\gamma{\widehat{g}}+(1-\gamma){g^{*}}$ and further denote
+$\pi_{\gamma}(\cdot|s)\propto\pi^{\mathsf{ref}}(\cdot|s)\exp\big{(}\eta g_{\gamma}(s,\cdot)\big{)},\forall s\in{\mathcal{S}};$
+$G(\gamma)\coloneqq\mathbb{E}_{\rho\times\pi_{\gamma}}\Big{[}\big{(}{\widehat{g}}-{{g^{*}}}\big{)}^{2}(s,a)\Big{]}.$
+The key to our analysis is the monotonicity of the function $G(\gamma)$ in $\gamma$, which is formally stated in the following lemma.
+Lemma 2.15.
+On event $\mathcal{E}$, $0\in\mathop{\mathrm{argmax}}_{\gamma\in[0,1]}G(\gamma)$.
+Proof.
+For simplicity, we use $\triangle(s,a)$ to denote $\big{(}{\widehat{g}}-{{g^{*}}})(s,a)$ in this proof.
+Then we know that $\triangle(s,a)\leq 0$ for all $(s,a)\in{\mathcal{S}}\times\mathcal{A}$ on event $\mathcal{E}$.
+The most direct way to prove is to take derivative of $G$ with respect to $\gamma$, which corresponds to the policy gradient (Sutton et al., 1999) of $\pi_{\gamma}$ and thus implying a favorable structure.
+A direct calculation yields that
+$\ =\mathbb{E}_{\rho\times\pi_{\gamma}}\big{[}\nabla_{\gamma}\log\pi_{\gamma}(a|s)\triangle(s,a)^{2}\big{]}$
+$\ =\eta\mathbb{E}_{\rho}\mathbb{E}_{a\sim\pi_{\gamma}}\big{[}\triangle^{2}(s,a)\big{(}\triangle(s,a)-\mathbb{E}_{a^{\prime}\sim\pi_{\gamma}}[\triangle(s,a^{\prime})]\big{)}\big{]}$
+$\ =\eta\mathbb{E}_{\rho}\Big{[}\mathbb{E}_{\pi_{\gamma}}\big{[}\triangle^{3}(s,a)\big{]}-\mathbb{E}_{\pi_{\gamma}}\big{[}\triangle^{2}(s,a)\big{]}\mathbb{E}_{\pi_{\gamma}}\big{[}\triangle(s,a)\big{]}\Big{]}$
+$\ \leq 0,$
+where $\mathbb{E}_{\rho}$ is the shorthand of $\mathbb{E}_{s\sim\rho}$, $\mathbb{E}_{\pi_{\gamma}}$ is the shorthand of $\mathbb{E}_{a\sim\pi_{\gamma}}$, the first equation is derived from standard policy gradient and the inequality holds conditioned on the event $\mathcal{E}(\delta)$ due to Lemma 2.13.
+∎
+Finally, we need the following lemma to bound the performance difference between two policy w.r.t.
+the KL-regularized objective $J$.
+For any given $g:{\mathcal{S}}\times\mathcal{A}\to\mathbb{R}$, we define $\pi_{g}(\cdot|s)\propto\exp\big{(}\eta\cdot g(s,\cdot)\big{)},\forall s\in{\mathcal{S}}$ to facilitate presentation; then $\pi^{*}=\pi_{{g^{*}}}$ by definition.
+Please refer to Appendix B.3 for the proof of the lemma.
+Lemma 2.16.
+Given $g:{\mathcal{S}}\times\mathcal{A}\to\mathbb{R}$, $\exists\gamma\in[0,1]$ such that for $g_{\gamma}=\gamma g+(1-\gamma)g^{*}$ and $\pi_{\gamma}\coloneqq\pi_{g_{\gamma}}$,
+$J(\pi^{*})-J(\pi_{g})\leq\eta\mathbb{E}_{(s,a)\sim\rho\times\pi_{\gamma}}\big{[}\big{(}{g^{*}}-g\big{)}^{2}(s,a)\big{]}.$
+Now we are ready to prove Theorem 2.9.
+Proof of Theorem 2.9.
+Following the proof of Zhao et al. (2024, Theorem 3.3), we know that there exists $\bar{\gamma}\in[0,1]$ such that
+$J(\pi^{*})-J(\widehat{\pi})\leq\eta G(\bar{\gamma})\leq\eta G(0),$ (2.5)
+where the first inequality holds due to Lemma 2.16 and the second inequality holds due to the event $\mathcal{E}$ and Lemma 2.15.
+The term $G(0)$ can be further bounded with the $D^{2}$-based concentrability as follows
+$G(0)$ $=\eta\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}\Big{[}\big{(}{\widehat{g}}-{{g^{*}}}\big{)}^{2}(s,a)\Big{]}$
+$\leq 4\eta\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}[\Gamma_{n}^{2}(s,a)]$
+$=4\eta\beta^{2}\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}\big{[}D^{2}_{\mathcal{F}}((s,a);\pi^{\mathsf{ref}})\big{]}$
+$=\widetilde{O}(\eta D_{\pi^{*}}^{2}n^{-1}\log_{{\mathcal{G}}}(\epsilon_{c})),$ (2.6)
+where the second inequality holds conditioned on $\mathcal{E}(\delta)$ because of Lemma 2.15, and the last inequality follows from the definition of $\mathcal{E}(\delta)$ together with Line 2.
+By Lemma 2.12, we know that event $\mathcal{E}$ holds with probability at least $1-\delta$, which finishes the proof.
+∎
+3 $f$-divergence-regularized Contextual Bandits
+As we discussed in Section 2, the fast rate is primarily achieved due to the strong convexity of $\pi\mapsto\mathsf{KL}(\cdot\|\pi^{\mathsf{ref}})$.
+However, KL is just an instance of $f$-divergence with $f(x)=x\log x$, which is not strongly convex, viz., only locally strongly convex.
+Motivated by this observation, we examine the use of $f$-divergence regularization with strongly convex $f$, which, in principle, may introduce a more favorable curvature in the performance metric of offline learning.
+3.1 Problem Setup
+We study the same contextual bandit problems which follows the definition in Section 2.1. In this section, we consider the following $f$-divergence regularized objective
+$J(\pi)\coloneqq\mathbb{E}_{(s,a)\sim\rho\times\pi}[r(s,a)]-\eta^{-1}\mathbb{E}_{s\sim\rho}\big{[}D_{f}(\pi||\pi^{\mathsf{ref}})\big{]},$ (3.1)
+where $\eta$ is the inverse temperature regularization intensity and $D_{f}(\pi||\pi^{\mathsf{ref}})$ is given by
+$D_{f}(\pi||\pi^{\mathsf{ref}})=\mathbb{E}_{a\sim\pi^{\mathsf{ref}}}\bigg{[}f\bigg{(}\frac{\pi(a|s)}{\pi^{\mathsf{ref}}(a|s)}\bigg{)}\bigg{]}.$
+We consider those functions $f:(0,+\infty)\to\mathbb{R}$ with nice positive curvature as follows.
+Assumption 3.1.
+$f$ is $\alpha$-strongly convex, twice continuously differentiable, and $f(1)=0$.
+Many elementary functions like quadratic polynomials naturally satisfy Assumption 3.1.
+For instance, the 1-strongly convex $f(x)=(x-1)^{2}/2$ yields $D_{f}(P\|Q)=\chi^{2}(P\|Q)$, which is the popular $\chi^{2}$-divergence recently considered in RL literature (see e.g., Zhan et al.
+(2022); Huang et al.
+(2024); Amortila et al.
+(2024)) and exhibits a promising theoretical potential on relaxing the data coverage requirement for efficient offline policy learning, at least against the unregularized objective.
+3.2 Algorithm and Main Results
+Algorithm 2 Offline $f$-divergence Regularized Contextual Bandits ($f$-CB)
+0:  regularization $\eta$, reference policy $\pi^{\mathsf{ref}}$, function class ${\mathcal{G}}$, offline dataset $\mathcal{D}$
+1:  Compute the least square estimation of reward function
+${\bar{g}}\in\mathop{\mathrm{argmin}}_{g\in{\mathcal{G}}}\sum_{(s_{i},a_{i},r_{i})\in\mathcal{D}}\big{(}g(s_{i},a_{i})-r_{i}\big{)}^{2}$
+2:  Compute the optimal policy under the least-square reward estimator ${\bar{g}}$ for $s\in{\mathcal{S}}$ as
+$\widehat{\pi}(\cdot|s)\leftarrow\mathop{\mathrm{argmax}}_{\pi(\cdot|s)\in\Delta(\mathcal{A})}\langle{\pi(\cdot|s)},{{\bar{g}}(s,\cdot)}\rangle+\eta^{-1}D_{f}\big{(}\pi(\cdot|s)\|\pi^{\mathsf{ref}}(\cdot|s)\big{)}$
+2:  $\widehat{\pi}$
+In this subsection, we present an offline learning algorithm for $f$-divergence regularized bandit, $f$-CB, in Algorithm 2.
+Algorithm 2 first leverages least-square estimator to find a function ${\bar{g}}\in{\mathcal{G}}$ that minimizes its risk on the offline dataset.
+The algorithm then uses the least squares estimation ${\bar{g}}$ to construct the output policy $\widehat{\pi}$.
+Compared to Algorithm 1, $f$-CB does not require any procedure to construct pessimistic reward estimation, whose sample complexity upper bound is given as follows.
+Theorem 3.2.
+Under Assumption 3.1, for sufficiently small $\epsilon\in(0,1)$, with probability at least $1-\delta$, $n=\widetilde{O}(\alpha^{-1}\eta\epsilon^{-1}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon))$ is sufficient to guarantee the output policy $\widehat{\pi}$ of $f$-CB to be $\epsilon$-optimal.
+Remark 3.3.
+Compared to the $D^{2}_{\pi^{*}}$ dependency in Theorem 2.9, Theorem 3.2 shows that the sample complexity of Algorithm 2 gets rid of the dependency on any data coverage conditions when $f$ is strongly convex.
+Intuitively, this is because the $f$-divergence regularization in this case is much stronger, so that both $\pi^{*}$ and $\widehat{\pi}$ are close enough to $\pi^{\mathsf{ref}}$.
+The following hardness result justify the near-optimality of Theorem 3.2 for $f$-divergence-regularized bandits.
+Theorem 3.4.
+For any $\epsilon\in(0,1),\alpha>0,\eta>0$, and algorithm Alg, there is an $\alpha$-strongly-convex function $f$ and an $f$-divergence-regularized contextual bandit instance such that Alg requires at least $\Omega\big{(}\alpha^{-1}\eta\epsilon^{-1}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon)\big{)}$ samples to return an $\epsilon$-optimal policy.
+3.3 Insights Behind the Proof of Theorem 3.2
+We provide an overview of key analysis techniques for proving Theorem 3.2. For the simplicity of presentation, we consider multi-arm bandits here and omit the subscript for context $s$.
+We consider the function $H(\pi)=\eta^{-1}D_{f}(\pi\|\pi^{\mathsf{ref}})$, which is the regularizer in the objective.
+Then its convex conjugate is given by $H^{*}(r)=\sup_{\pi\in\Delta^{d}}\{\left\langle\pi,r\right\rangle-H_{s}(\pi)\}$, which is exactly the expected reward obtained by the optimal policy given reward function $r$.
+One observation is that when $f$ is strongly convex, the induced $f$-divergence, and therefore the function $H$ is also strongly convex.
+Therefore, let $\pi_{r}=\mathop{\mathrm{argmax}}_{\pi}\{\left\langle\pi,r\right\rangle-H_{s}(\pi)\}$ given some reward function $r$, the strong convexity of $H(\pi)$ gives that $\nabla H^{*}(r)=\pi_{r}$.
+This leads to the following regret decomposition, which is one of our key observations.
+$J(\pi^{*})-J(\widehat{\pi})$ $=\mathbb{E}_{a\sim\pi^{*}}[{g^{*}}(a)]-\mathbb{E}_{a\sim\widehat{\pi}}[{g^{*}}(a)]$
+$\quad-\eta^{-1}\big{[}D_{f}(\pi^{*}\|\pi^{\mathsf{ref}})-D_{f}(\widehat{\pi}\|\pi^{\mathsf{ref}})\big{]}$
+$=H^{*}({g^{*}})-H^{*}({\bar{g}})-\left\langle\widehat{\pi},{g^{*}}-{\bar{g}}\right\rangle$
+$=H^{*}({g^{*}})-H^{*}({\bar{g}})-\left\langle\nabla H^{*}({\bar{g}}),{g^{*}}-{\bar{g}}\right\rangle,$
+which is the Bregman divergence of the dual function $H^{*}$ and therefore can be bounded by $({g^{*}}-{\bar{g}})^{\top}H^{*}(\widetilde{g})({g^{*}}-{\bar{g}})$ for some $\widetilde{g}$.
+By Proposition 3.2 in Penot (1994), we can bound $H^{*}(\widetilde{g})$ as follows
+$H^{*}(\widetilde{g})$ $\preceq\big{(}\nabla^{2}H(\pi_{\widetilde{g}})\big{)}^{-1}$
+$\preceq\alpha^{-1}\eta\mathrm{diag}\big{(}\pi^{\mathsf{ref}}(a_{1}),\cdots,\pi^{\mathsf{ref}}(a_{|\mathcal{A}|})\big{)},$
+which enables us to bound $({g^{*}}-{\bar{g}})^{\top}H^{*}(\widetilde{g})({g^{*}}-{\bar{g}})$ by $\alpha^{-1}\eta\mathbb{E}_{\pi^{\mathsf{ref}}}[({g^{*}}-{\widehat{g}})^{2}]$.
+Since $\mathbb{E}_{\pi^{\mathsf{ref}}}[({g^{*}}-{\widehat{g}})^{2}]$ is not related to the optimal $\pi^{*}$, our resulting upper bound does not depend on any notion of concentrability.
+4 Generalization to Contextual Dueling Bandits
+In this section, we extend our algorithm to the problems of regularized contextual dueling bandits, where the learner receives preference comparison instead of absolute signals.
+Our data model follows Zhu et al.
+(2023); Zhan et al.
+(2023) and the notion of suboptimality follows Xiong et al.
+(2024); Zhao et al.
+(2024).
+4.1 Problem Setup
+We still consider contextual bandits $({\mathcal{S}},\mathcal{A},r,\pi^{\mathsf{ref}})$ where ${\mathcal{S}}$ is the state space, $\mathcal{A}$ is the action space and $r:{\mathcal{S}}\times\mathcal{A}\to[0,1]$ is the reward function.
+[^1] We overload some notations in Section 2 by their dueling counterparts for notational simplicity.
+But only relative preference feedback is available, viz., we have an $\mathrm{i.i.d.}$ offline dataset $\mathcal{D}=\{(s_{i},a^{1}_{i},a^{2}_{i},y_{i})\}_{i=1}^{n}$, where $s_{i}\in{\mathcal{S}}$ is generated from distribution $\rho$ and $a^{1}_{i},a^{2}_{i}\sim\pi^{\mathsf{ref}}$.
+The binary preference label $y_{i}=1$ indicates $a_{i}^{1}$ is preferred over $a_{i}^{2}$ (denoted by $a^{1}\succ a^{2}$) and 00 for $a^{2}\succ a^{1}$ given context $s$.
+In this work we consider the Bradley-Terry Model, where $\mathbb{P}[y=1|s,a^{1},a^{2}]=\sigma(r(s_{i},a^{1}_{i})-r(s_{i},a^{2}_{i}))$, where $\sigma(x)=(1+e^{-x})^{-1}$ is the link function.
+The objective here identical to (2.1) for KL-regularization and (3.1) for $f$-divergence regularization.
+Our goal is still to find an $\epsilon$-optimal policy.
+To control the complexity of the function class ${\mathcal{G}}$, we assume that Assumption 2.1 still holds here.
+Concentrability.
+Analogous to Section 2, we need our estimation from offline dataset generalizable to the state-action pairs visited by our obtained policy.
+While density-ratio-based concentrability can be directly adapted to dueling bandit, we need a slightly different notion of $D^{2}$-divergence.
+This is because in dueling bandit, we cannot observe the absolute reward and best estimation $g$ we can achieve is that for any state $s$ and actions $a^{1},a^{2}$, our estimated $g(s,a^{1})-g(s,a^{2})\approx r(s,a^{1})-r(s,a^{2})$.
+This implies that there exists some mapping $b:{\mathcal{S}}\to[-1,1]$ such that $g(s,a)-b(s)\approx r(s,a)$ on the offline data, which leads to the following definition.
+Definition 4.1.
+Given a class of functions ${\mathcal{G}}\subset({\mathcal{S}}\times\mathcal{A}\to\mathbb{R})$ and some policy $\pi$, let $\mathcal{B}=({\mathcal{S}}\to[-1,1])$ be the function class, define the $D^{2}$-divergence $D^{2}_{{\mathcal{G}}}((s,a);\pi)$ as
+$\sup_{g,h\in{\mathcal{G}}}\inf_{b\in\mathcal{B}}\frac{\big{(}g(s,a)-h(s,a)-b(s)\big{)}^{2}}{\mathbb{E}_{s\sim\rho}\operatorname{{\rm Var}}_{a^{\prime}\sim\pi(\cdot|s^{\prime})}[g(s^{\prime},a^{\prime})-h(s^{\prime},a^{\prime})]}.$
+A similar definition has been introduced in Zhao et al.
+(2024, Definition 2.6), which underpins the following two assumptions that characterize the coverage ability of $\pi^{\mathsf{ref}}$ similarly as in Section 2.
+Given a reference policy $\pi^{\mathsf{ref}}$, we define two coverage notions for contextual dueling bandits.
+Assumption 4.2 (All-policy concentrability).
+$D^{2}\coloneqq\sup_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}D^{2}_{{\mathcal{G}}}((s,a);\pi^{\mathsf{ref}})<\infty$.
+Assumption 4.3 (Single-policy concentrability).
+$D_{\pi^{*}}^{2}\coloneqq\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}[D^{2}_{{\mathcal{G}}}((s,a);\pi^{\mathsf{ref}})]<\infty$.
+Similar single-policy concentrability assumptions have appeared in previous work in offline contextual dueling bandits (Huang et al., 2024; Song et al., 2024) and similar notions has also appeared in the analysis of model-based RL (Uehara and Sun, 2021; Wang et al., 2024).
+Although Assumption 4.3 is strictly weaker than Assumption 4.2, $D^{2}_{\pi^{*}}\leq|{\mathcal{S}}||\mathcal{A}|C^{\pi^{*}}$ still holds and, in general, $C^{\pi^{*}}$ and $D^{2}_{\pi^{*}}$ cannot be bounded by each other up to some constant factors.
+4.2 Main Results
+In this subsection, we provide our main results for contextual dueling bandits with regularization.
+We refer the reader to Appendix D for a more detailed discussion.
+We first present the following results for KL-regularized contextual dueling bandit.
+Theorem 4.4 (Informal).
+There exists an algorithm, with high probability, generates $\epsilon$-optimal policy with $n=\widetilde{O}(\eta D_{\pi^{*}}^{2}\epsilon^{-1}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon))$ samples.
+Moreover, for any sufficiently small $\epsilon>0$, and $\eta>0$, and any algorithm Alg, there is a KL-regularized contextual dueling bandit such that Alg requires at least $\Omega\big{(}\min\{{\eta C^{\pi^{*}}\epsilon^{-1}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon)},{(C^{\pi^{*}})^{2}\epsilon^{-2}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon)}\}\big{)}$ samples to return an $\epsilon$-optimal policy.
+The following theorem shows a similar guarantee for $f$-divergence regularized contextual dueling bandits where $f$ is $\alpha$-strongly convex.
+Theorem 4.5 (Informal).
+Under Assumption 3.1, there exists an algorithm generating $\epsilon$-optimal policy with $n=\widetilde{O}(\alpha^{-1}\eta\epsilon^{-1}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon))$ with high probability.
+Moreover, for any sufficiently small $\epsilon>0$, $\eta,\alpha>0$, and any offline RL algorithm Alg, there is an $\alpha$-strongly convex function $f$ such that there exists an $f$-divergence regularized contextual dueling bandit such that Alg requires at least $\Omega\big{(}{\alpha^{-1}\eta\epsilon^{-1}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon)}\big{)}$ samples to return an $\epsilon$-optimal policy.
+Theorems 4.4 and 4.5 are extension of Theorems 2.9 and 3.2 to offline contextual dueling bandits.
+These theorems indicate that, for both KL-regularization and $f$-divergence regularization, our techniques can be adapted to achieve algorithms and achieve similar sample complexities.
+5 Conclusion and Limitation
+In this work, we take the first step towards fully understanding the statistical efficiency with respect to $f$-divergence-regularized objectives of offline policy learning by sharp analyses for two empirically relevant subclasses.
+(1) We are the first to show that single-policy concentrability is nearly the right coverage condition for reverse KL to achieve the fast $\widetilde{\Theta}(\epsilon^{-1})$ sample complexity.
+The novel techniques in algorithm analysis leverages the curvature of KL-regularized objectives and integrates pessimism with a newly identified moment-based observation, enabling a neat refinement of a mean-value-type argument to the extreme; which are decoupled from tricky algorithmic tweaks, and thus might be of independent interest.
+(2) If strong convexity is further imposed on $f$, our fast $\widetilde{\Theta}(\epsilon^{-1})$ sample complexity is provably free of any coverage dependency.
+Unlike those for KL, the upper bound arguments for strongly convex $f$ do not rely on the specific closed-form of the regularized objective minimizer.
+All techniques in this work can be generalized beyond vanilla absolute reward feedback, as certified by contextual dueling bandits.
+However, there is still a gap between the upper bound and lower bound under reverse-KL regularization.
+Also, for general $f$-divergence other than reverse-KL, our results are limited to twice continuously differentiable and strongly convex $f$.
+The fully closing of the gap under reverse-KL regularization and the investigation of general $f$-divergence are left as future directions.
+Appendix A Review of Existing Results
+Additional Notations.
+Besides the notation introduced in Section 1, we will use the following notations in Appendix.
+We denote $[N]\coloneqq\{1,\cdots,N\}$ for any positive integer $N$.
+Boldfaced lower case (resp.
+upper case) letters are reserved for vectors (resp.
+matrices).
+Given a positive definite $\bm{\Sigma}\in\mathbb{R}^{d\times d}$ and $\mathbf{x}\in\mathbb{R}^{d}$, we denote the vector’s Euclidean norm by $\|\mathbf{x}\|_{2}$ and define $\|\mathbf{x}\|_{\bm{\Sigma}}=\sqrt{\mathbf{x}^{\top}\bm{\Sigma}\mathbf{x}}$.
+We use $\mathsf{Bern}(p)$ to denote Bernoulli distribution with expectation $p$ and $\mathsf{Unif}(\mathcal{X})$ for the uniform distribution on finite set $\mathcal{X}$.
+For $x\in\mathbb{R}^{\mathcal{A}}$, we denote $\|x\|_{1}=\sum_{a\in\mathcal{A}}|x_{a}|$.
+We also denote $x_{n}=\Omega(y_{n})$ by $x_{n}\gtrsim y_{n}$ in Appendix.
+We use $d_{H}$ for Hamming distance.
+A.1 Previous Attempts on Understanding KL-regularized RL
+There has been a surge of interest in understanding the principle behind KL-regularized RL.
+Ahmed et al.
+(2019); Liu et al.
+(2019) studied by ablation the effect of entropy regularization on the stability of policy improvement in policy optimization, the regret of which has been rigorously settled under the classic online mirror descent framework (Cai et al., 2020; He et al., 2022; Ji et al., 2023).
+Neu et al.
+(2017) unified popular KL-regularized policy optimization algorithms under a convex optimization framework, but the interplay with the data was left untouched.
+A series of work (Geist et al., 2019; Vieillard et al., 2020; Kozuno et al., 2022) then analyzed the sample complexity of algorithms using KL/entropy-type proximal terms with respect to the previous iteration or/and entropy regularizer with improved dependence on the effective horizon in discounted Markov decision processes.
+However, the performance metric in these studies is still the unregularized reward maximization objective, under which the sample complexity for finding an $\epsilon$-optimal policy is at least equal to the statistical limit $\Omega(\epsilon^{-2})$.
+Convergence under regularized objectives.
+Several recent studies (Xie et al., 2024; Xiong et al., 2024; Zhao et al., 2024, 2025; Foster et al., 2025) switched the focus to analyzing the suboptimality guarantee with respect to the regularized objective (1.1).
+In particular, Xie et al.
+(2024) studied token-level Markov decision processes (MDPs) and proposed a KL-regularized RL algorithm named XPO, which achieves $\widetilde{O}(\epsilon^{-2})$ sample complexity under their notion of all-policy concentrability.
+Xiong et al.
+(2024) proposed an Offline GSHF algorithm via the principle of pessimism in the face of uncertainty, and proved $\widetilde{O}(\epsilon^{-2})$ sample complexity under single-policy concentrability (See Section 2.1 for detailed definitions of concentrability).
+On the other hand, the sharp analysis in Zhao et al.
+(2024) yields the optimal sample complexity $\widetilde{O}(\epsilon^{-1})$, but requires all-policy concentrability (Zhao et al., 2024, Definition 2.6), i.e., the behavior policy $\pi^{\mathsf{ref}}$ is required to cover the entire function class for all possible policies.
+Zhao et al.
+(2025) considered the online episodic MDP setting, which inherently does not need any notion of data coverage and thus their results are not directly adaptable to our offline setting.
+Foster et al.
+(2025) considered an interesting hybrid setting in which the $n$ state-action pairs are still from the offline dataset but $\Omega(n)$ online reward queries and policy switches are allowed; in contrast, in our setting, all reward signals are obtained in a purely offline fashion.
+Here, we briefly discuss the direct adaptation of previous sample complexity results (with respect to KL-regularized objectives) to our setting and demonstrate the reason why theirs cannot imply an $\widetilde{O}(\epsilon^{-1})$ sample complexity without all-policy concentrability.
 Specifically, Xiong et al.
@@
 An analogous analysis gives a sample complexity of $\widetilde{O}(D^{2}\epsilon^{-1})$, where $D^{2}$ is the all-policy concentrability.
-2.3 Algorithm
-In this subsection, we present an offline bandit algorithm, KL-PCB, for KL-regularized contextual bandits in Algorithm 1.
-KL-PCB first leverages least-square estimator to find a function ${\bar{f}}\in\mathcal{F}$ that minimizes its risk on the offline dataset.
-In Zhao et al.
-(2024), such ${\bar{f}}$ is directly applied to construct the estimated policy.
-In contrast, we construct a pessimistic estimator of $f^{*}$ following the well-known pessimism principle in offline RL (Jin et al., 2021).
-Specifically, we consider the bonus term defined as
-$\Gamma_{n}(s,a)=\beta D_{\mathcal{F}}\big{(}(s,a),\pi^{\mathsf{ref}}\big{)},\forall(s,a)\in{\mathcal{S}}\times\mathcal{A};$ (2.3)
-where
-$\beta=\sqrt{128\log\big{(}2\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta\big{)}/3n+18\epsilon_{c}}.$ (2.4)
-We then obtain our pessimistic estimation ${\widehat{f}}$ by setting ${\widehat{f}}={\bar{f}}-\Gamma_{n}$, which is less than $f^{*}$ with high probability.
-After obtaining the pessimistic estimation, KL-PCB output the policy $\widehat{\pi}$, which maximizes the estimated objective
-$\widehat{J}(\pi)=\mathbb{E}_{(s,a)\sim\rho\times\pi}\bigg{[}{\widehat{f}}(s,a)-\eta^{-1}\log\frac{\pi(a|s)}{\pi^{\mathsf{ref}}(a|s)}\bigg{]},$
-the maximizer of which is the counterpart of 2.2, i.e.,
-$\widehat{\pi}(a|s)\propto\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta\cdot{\widehat{f}}(s,a)\big{)}.$
-Algorithm 1 Offline KL-Regularized Pessimistic Contextual Bandit (KL-PCB)
-0:  regularization $\eta$, reference policy $\pi^{\mathsf{ref}}$, function class $\mathcal{F}$, offline dataset $\mathcal{D}=\{(s_{i},a_{i},r_{i})\}_{i=1}^{n}$
-1:  Compute the least square estimation of reward function
-${\bar{f}}\in\mathop{\mathrm{argmin}}_{f\in\mathcal{F}}\sum_{(s_{i},a_{i},r_{i})\in\mathcal{D}}\big{(}f(s_{i},a_{i})-r_{i}\big{)}^{2}$
-2:  Let ${\widehat{f}}\leftarrow{\bar{f}}-\Gamma_{n}$, where $\Gamma_{n}$ is the bonus term in 2.3
-2:  $\widehat{\pi}(a|s)\propto\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta\cdot{\widehat{f}}(s,a)\big{)}$
-2.4 Theoretical Results
-The sample complexity for KL-regularized contextual bandits is settled in this subsection. We first give the upper bound of KL-PCB.
-Theorem 2.8.
-Under Assumption 2.7, for sufficiently small $\epsilon\in(0,1)$, if we set $\Gamma_{n}$ as in (2.3), then $n=\widetilde{O}(\eta D_{\pi^{*}}^{2}\epsilon^{-1})$ suffices to guarantee the output policy $\widehat{\pi}$ of Algorithm 1 to be $\epsilon$-optimal with probability at least $1-\delta$.
-Remark 2.9.
-Previously, Zhao et al.
-(2024) achieved an $\widetilde{O}(\epsilon^{-1})$ sample complexity under Assumption 2.6.
-As a comparison, KL-PCB achieves the same $\widetilde{O}(\epsilon^{-1})$ sample complexity but only requiring Assumption 2.7, which is weaker than Assumption 2.6
-We also provide the sample complexity lower bound of KL-regularized contextual bandits in the following theorem.
-Theorem 2.10.
-For any $\epsilon\in(0,1),\eta>0$, and algorithm Alg, there is a KL-regularized contextual bandit instance with $O(1)$ single-policy concentrability such that Alg requires at least $\Omega\big{(}\min\{{\eta\log\mathcal{N}_{\mathcal{F}}(\epsilon)}/{\epsilon},{\log\mathcal{N}_{\mathcal{F}}(\epsilon)}/{\epsilon^{2}}\}\big{)}$ samples to return an $\epsilon$-optimal policy.
-Remark 2.11.
-Theorem 2.10 is concluded by plugging $D^{2}\geq D^{2}_{\pi^{*}}$ to Theorem 3.1 of Zhao et al.
-(2024) and thus the sample complexity lower bound also holds for the KL-regularized contextual bandit instance with $O(1)$ all-policy concentrability.
-Remark 2.12.
-Theorem 2.10 shows that when $\epsilon$ is sufficiently small, any algorithm for offline KL-regularized contextual bandits requires at least $\widetilde{\Omega}(\epsilon^{-1})$ samples to output an $\epsilon$-optimal policy.
-This sample complexity lower bound matches the sample complexity upper bound in Theorem 2.8, indicating that KL-PCB is nearly optimal.
-2.5 Proof of Theorem 2.8
-In this section, we finish the proof of Theorem 2.8.
-At a high level, if we consider the regularized objective 1.1 multi-arm bandits, then $P\mapsto{\mathsf{KL}}\left(P\|Q\right)$ is $1$-strongly convex w.r.t.
-${\mathsf{TV}}\left(\cdot\|\cdot\right)$ (Polyanskiy and Wu, 2024, Exercise I.37), and thus $J(\pi)$ is strongly concave.
-Therefore, $J(\pi^{*})-J(\widehat{\pi})$ is possible to be of the order $[{\mathsf{TV}}\left(\pi^{*}\|\widehat{\pi}\right)]^{2}\approx\widetilde{O}(n^{-1})$, pretending that $\pi^{*}$ is the unconstrained maximizer.
-This intuition guides our analysis for contextual bandits.
-We begin the proof with the definition of the event $\mathcal{E}(\delta)$ given $\delta>0$ as
-$\mathcal{E}(\delta)\coloneqq\bigg{\{}\sup_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}\Big{[}\big{|}{\bar{f}}-f^{*}\big{|}-\Gamma_{n}\Big{]}(s,a)\leq 0\bigg{\}},$ (2.5)
-where $\Gamma_{n}$ is defined in (2.3).
-Event $\mathcal{E}(\delta)$ holds indicates that the least square estimation ${\bar{f}}$ obtained in Line 1 of Algorithm 1 does not deviate too much from the true function $f^{*}$.
-More specifically, we have the following lemma, whose proof are deferred to Please refer to Appendix A.1.
-Lemma 2.13.
-For all $\delta>0$, $\mathcal{E}(\delta)$ holds with probability at least $1-\delta$.
-The following covariance-type observation is the first pivot.
-Lemma 2.14.
-If a bounded random variable $X\leq 0$ almost surely, then $\mathbb{E}[X^{3}]-\mathbb{E}[X^{2}]\mathbb{E}[X]\leq 0$.
-Remark 2.15.
-While this lemma is elementary, to the best of our knowledge, we are the first to isolate this structure from our non-standard analysis of offline RL, from which the sharp upper bound is derived.
+Appendix B Missing Proofs from Section 2
+B.1 Proof of Lemma 2.12
+We first provide the following lemmas of concentration.
+Lemma B.1 (Zhao et al. 2024, Lemma C.1).
+For any policy $\pi$ and state-action pairs $\{(s_{i},a_{i})\}_{i=1}^{m}$ generated i.i.d.
+from $\rho\times\pi$, and $\epsilon_{c}<1$, with probability at least $1-\delta$, for any $g_{1}$ and $g_{2}$ we have
+$\mathbb{E}_{\rho\times\pi}\big{[}\big{(}g_{1}(s,a)-g_{2}(s,a)\big{)}^{2}\big{]}\leq\frac{2}{n}\sum_{i=1}^{n}\big{(}g_{1}(s_{i},a_{i})-g_{2}(s_{i},a_{i})\big{)}^{2}+\frac{32}{3n}\log(2\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+10\epsilon_{c},$
+where $\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})$ is the $\epsilon_{c}$-covering number of ${\mathcal{G}}$.
+Lemma B.2 (Zhao et al. 2024, Lemma C.2).
+For arbitrary policy $\pi$ and dataset $\{(s_{i},a_{i},r_{i})\}_{i=1}^{m}$ generated i.i.d., from the product of $\pi$, $\rho$ and the Bradley-Terry Model; let ${\bar{g}}$ be the least square estimator of ${g^{*}}$, then for any $10<\epsilon_{c}<$ 1 and $\delta>0$, with probability at least $1-\delta$ we have
+$\sum_{i=1}^{n}\big{(}{\bar{g}}(s_{i},a_{i})-{g^{*}}(s_{i},a_{i})\big{)}^{2}\leq 16\log(a\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+4n\epsilon_{c}.$
+Proof of Lemma 2.12.
+We have the following inequality
+$\big{(}{\bar{g}}(s,a)-{g^{*}}(s,a)\big{)}^{2}$ $=\frac{\big{(}{\bar{g}}(s,a)-{g^{*}}(s,a)\big{)}^{2}}{\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{g}}(s,a)-{g^{*}}(s,a)\big{)}^{2}\big{]}}\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{g}}(s,a)-{g^{*}}(s,a)\big{)}^{2}\big{]}$
+$\leq\sup_{g_{1},g_{2}\in{\mathcal{G}}}\frac{\big{(}g_{1}(s,a)-g_{2}(s,a)\big{)}^{2}}{\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}g_{1}(s,a)-g_{2}(s,a)\big{)}^{2}\big{]}}\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{g}}(s,a)-{g^{*}}(s,a)\big{)}^{2}\big{]}$
+$=D_{{\mathcal{G}}}^{2}((s,a),\pi^{\mathsf{ref}})\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{g}}(s,a)-{g^{*}}(s,a)\big{)}^{2}\big{]},$ (B.1)
+where the inequality holds by taking supremum to $g_{1},g_{2}\in{\mathcal{G}}$. Now we have
+$\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{g}}(s,a)-{g^{*}}(s,a)\big{)}^{2}\big{]}$ $\leq\frac{2}{n}\sum_{i=1}^{n}\big{(}{\bar{g}}(s_{i},a_{i})-{g^{*}}(s_{i},a_{i})\big{)}^{2}+\frac{32}{3n}\log(2\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+10\epsilon_{c}$
+$\leq\frac{2}{n}\big{[}16\log(\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+4n\epsilon_{c}\big{]}+\frac{32}{3n}\log(2\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+10\epsilon_{c}$
+$=\frac{128}{3n}\log(2\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+18\epsilon_{c},$ (B.2)
+where the first inequality holds due to Lemma B.1 and second holds due to Lemma B.2. Plugging (B.1) into (B.1) and setting $\epsilon_{c}=O(n^{-1})$ complete the proof. ∎
+B.2 Proof of Lemma 2.13
 Proof of Lemma 2.13.
@@
 where both inequalities follow from Jensen’s inequality. ∎
-We further define the following quantities. For all $\gamma\in[0,1]$, we define
-$f_{\gamma}\coloneqq\gamma{\widehat{f}}+(1-\gamma)f^{*},$
-and based on which we further denote
-$\pi_{\gamma}(\cdot|s)\propto\pi^{\mathsf{ref}}(\cdot|s)\exp\big{(}\eta f_{\gamma}(s,\cdot)\big{)},\forall s\in{\mathcal{S}};$
-$G(\gamma)\coloneqq\mathbb{E}_{\rho\times\pi_{\gamma}}\Big{[}\big{(}{\widehat{f}}-{f^{*}}\big{)}^{2}(s,a)\Big{]}.$
-The key to our analysis is the monotonicity of the function $G(\gamma)$ in $\gamma$, which is formally stated in the following lemma.
-Lemma 2.16.
-On event $\mathcal{E}$, $0\in\mathop{\mathrm{argmax}}_{\gamma\in[0,1]}G(\gamma)$.
-Proof.
-For simplicity, we use $\triangle(s,a)$ to denote $\big{(}{\widehat{f}}-{f^{*}})(s,a)$ in this proof.
-Then we know that $\triangle(s,a)\leq 0$ for all $(s,a)\in{\mathcal{S}}\times\mathcal{A}$ on event $\mathcal{E}$.
-The most direct way to prove is to take derivative of $G$ with respect to $\gamma$, which corresponds to the policy gradient (Sutton et al., 1999) of $\pi_{\gamma}$ and thus implying a favorable structure.
-A direct calculation yields that
-$G^{\prime}(\gamma)$ $=\mathbb{E}_{\rho\times\pi_{\gamma}}\big{[}\nabla_{\gamma}\log\pi_{\gamma}(a|s)\triangle(s,a)^{2}\big{]}$
-$=\eta\mathbb{E}_{\rho}\mathbb{E}_{a\sim\pi_{\gamma}}\big{[}\triangle^{2}(s,a)\big{(}\triangle(s,a)-\mathbb{E}_{a^{\prime}\sim\pi_{\gamma}}[\triangle(s,a^{\prime})]\big{)}\big{]}$
-$=\eta\mathbb{E}_{\rho}\Big{[}\mathbb{E}_{\pi_{\gamma}}\big{[}\triangle^{3}(s,a)\big{]}-\mathbb{E}_{\pi_{\gamma}}\big{[}\triangle^{2}(s,a)\big{]}\mathbb{E}_{\pi_{\gamma}}\big{[}\triangle(s,a)\big{]}\Big{]}$
-$\leq 0,$
-where $\mathbb{E}_{\rho}$ is the shorthand of $\mathbb{E}_{s\sim\rho}$, $\mathbb{E}_{\pi_{\gamma}}$ is the shorthand of $\mathbb{E}_{a\sim\pi_{\gamma}}$, the first equation is derived from standard policy gradient and the inequality holds conditioned on the event $\mathcal{E}(\delta)$ due to Lemma 2.14.
+B.3 Proof of Lemma 2.16
+This proof is extracted from the proof of Zhao et al. (2024, Theorem 3.3) and we present it here for completeness. By definition of our objective in (2.1), we have
+$J(\pi^{*})-J(\pi_{g})$
+$\quad=\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}\bigg{[}{g^{*}}(s,a)-\eta^{-1}\log\frac{\pi^{*}(a|s)}{\pi^{\mathsf{ref}}(a|s)}\bigg{]}-\mathbb{E}_{(s,a)\sim\rho\times\pi_{g}}\bigg{[}{g^{*}}(s,a)-\frac{1}{\eta}\log\frac{\pi_{g}(a|s)}{\pi^{\mathsf{ref}}(a|s)}\bigg{]}$
+$\quad=\frac{1}{\eta}\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}\bigg{[}\log\frac{\pi^{\mathsf{ref}}(a|s)\cdot\exp\bigl{(}\eta{g^{*}}(s,a)\bigr{)}}{\pi^{*}(a|s)}\bigg{]}-\frac{1}{\eta}\mathbb{E}_{(s,a)\sim\rho\times\pi_{g}}\bigg{[}\log\frac{\pi^{\mathsf{ref}}(a|s)\cdot\exp\bigl{(}\eta{g^{*}}(s,a)\bigr{)}}{\pi_{g}(a|s)}\bigg{]}$
+$\quad=\frac{1}{\eta}\mathbb{E}_{s\sim\rho}\big{[}\log Z_{{g^{*}}}(s)\big{]}-\frac{1}{\eta}\mathbb{E}_{s\sim\rho}\big{[}\log Z_{g}(s)\big{]}-\mathbb{E}_{s\sim\rho}\bigg{[}\sum_{a\in\mathcal{A}}\pi_{g}(a|s)\cdot\big{(}{g^{*}}(s,a)-f(s,a)\big{)}\bigg{]},$
+where for all $g\in{\mathcal{G}}$ we define $Z_{g}(\cdot)$ as follows,
+$Z_{g}(\cdot)\coloneqq\sum_{a\in\mathcal{A}}\pi^{\mathsf{ref}}(a|\cdot)\exp\big{(}\eta g(\cdot,a)\big{)}.$
+We further denote $\Delta(s,a)=g(s,a)-{g^{*}}(s,a)$ and $H_{s}(g)=\log Z_{g}(s)-\eta\sum_{a\in\mathcal{A}}\pi_{g}(a|s)\cdot\Delta(s,a)$.
+It worth noticing that $\eta^{-1}\mathbb{E}_{s\sim\rho}[H_{s}({g^{*}})-H_{s}(g)]=J(\pi^{*})-J(\pi_{g})$.
+Now we take derivative of $H$ with respect to $\Delta(s,a)$,
+$\frac{\partial H_{s}(g)}{\partial\Delta(s,a)}$ $=\frac{\partial}{\partial\Delta(s,a)}\bigg{[}\log Z_{g}(s)-\eta\sum_{a\in\mathcal{A}}\pi_{g}(a|s)\cdot\Delta(s,a)\bigg{]}$
+$=\frac{1}{Z_{g}(s)}\cdot\pi^{\mathsf{ref}}(a|s)\exp\bigl{(}\eta\cdot g(s,a)\bigr{)}\cdot\eta-\eta\cdot\pi_{g}(a|s)$
+$\quad-\eta^{2}\cdot\Delta(s,a)\cdot\frac{\pi^{\mathsf{ref}}(a|s)\cdot\exp\bigl{(}\eta\cdot g(s,a)\bigr{)}}{Z_{g}(s)}+\eta^{2}\cdot\Delta(s,a)\cdot\frac{\bigl{[}\pi^{\mathsf{ref}}(a|s)\cdot\exp\bigl{(}\eta\cdot g(s,a)\bigr{)}\bigr{]}^{2}}{[Z_{g}(s)]^{2}}$
+$\quad+\eta\sum_{a^{\prime}\in\mathcal{A}\backslash\{a\}}\frac{\pi^{\mathsf{ref}}(a^{\prime}|x)\cdot\exp\bigl{(}\eta\cdot g(s,a^{\prime})\bigr{)}}{Z_{g}(s)}\cdot\eta\cdot\Delta(s,a^{\prime})\cdot\frac{\pi^{\mathsf{ref}}(a|s)\cdot\exp\bigl{(}\eta\cdot g(s,a)\bigr{)}}{Z_{g}(s)}$
+$=-\eta^{2}\pi_{g}(a|s)\Delta(s,a)+\eta^{2}[\pi_{g}(a|s)]^{2}\cdot\Delta(s,a)+\eta^{2}\sum_{a^{\prime}\in\mathcal{A}\backslash\{a\}}\pi_{g}(a^{\prime}|x)\pi_{g}(a|s)\Delta(s,a^{\prime}).$
+Therefore, by mean value theorem, there exists $\gamma\in[0,1]$ and $g_{\gamma}=\gamma g+(1-\gamma){g^{*}}$ such that
+$H_{s}(g)-H_{s}({g^{*}})$ $=-\eta^{2}\gamma\sum_{a\in\mathcal{A}}\pi_{g_{\gamma}}(a|s)\Delta(s,a)^{2}+\gamma\eta^{2}\sum_{a_{1}\in\mathcal{A}}\sum_{a_{2}\in\mathcal{A}}\pi_{g_{\gamma}}(a_{1}|x)\pi_{g_{\gamma}}(a_{2}|x)\Delta(s,a_{1})\Delta(s,a_{2})$
+$=-\eta^{2}\gamma\mathbb{E}_{a\sim\pi_{g_{\gamma}}}\big{[}\big{(}{g^{*}}(s,a)-g(s,a)\big{)}^{2}\big{]}+\gamma\eta^{2}\Big{(}\mathbb{E}_{a\sim\pi_{g_{\gamma}}}\big{[}\big{(}{g^{*}}(s,a)-g(s,a)\big{)}\big{]}\Big{)}^{2}$
+$\geq-\eta^{2}\mathbb{E}_{a\sim\pi_{g_{\gamma}}}\big{[}\big{(}{g^{*}}(s,a)-g(s,a)\big{)}^{2}\big{]},$
+where the inequality holds by omitting the second term and $\gamma\leq 1$. Now taking expectation over $\rho$, we have
+$J(\pi^{*})-J(\pi_{g})$ $=\eta^{-1}\mathbb{E}_{s\sim\rho}[H_{s}({g^{*}})-H_{s}(g)]$
+$\leq\eta\mathbb{E}_{(s,a)\sim\rho\pi_{g_{\gamma}}}\big{[}\big{(}{g^{*}}(s,a)-g(s,a)\big{)}^{2}\big{]},$
+which concludes the proof.
+B.4 Proof of Theorem 2.10
+Proof of Theorem 2.10.
+We consider the family of contextual bandits with $S\coloneqq|{\mathcal{S}}|,A\coloneqq|\mathcal{A}|<\infty$ and reward function in some function class $\mathcal{G}$ composed of function ${\mathcal{S}}\times\mathcal{A}\to[0,1]$ as follows.
+$\mathrm{CB}_{\mathcal{G}}\coloneqq\{({\mathcal{S}},\mathcal{A},\rho,r,\pi^{\mathsf{ref}},\eta):r\in\mathcal{G},\rho\in\Delta({\mathcal{S}}),\pi^{\mathsf{ref}}\in\Delta(\mathcal{A}|{\mathcal{S}})\}.$ (B.3)
+Our goal is to prove the following statement.
+Fixing any $S\geq 1$, $\eta>4\log 2$ and $C^{*}\in(2,\exp(\eta/4)]$, then for any estimator $\mathcal{D}\mapsto\widehat{\pi}\in\Delta(\mathcal{A}|{\mathcal{S}})$, for any $n\geq 16SC^{*}$, there exist some function class $\mathcal{G}$, such that $\exists\ \mathrm{inst}=({\mathcal{S}},\mathcal{A},\rho,r,\pi^{\mathsf{ref}},\eta)\in\mathrm{CB}_{\mathcal{G}}$ with single-policy concentrability $C^{\pi^{*}}\leq C^{*}$, regularization coefficient $\eta$, $|{\mathcal{S}}|=S=\Theta(\log|\mathcal{G}|)$, and
+$\mathrm{SubOpt}_{\mathrm{RKL}}(\widehat{\pi};\mathrm{inst})\gtrsim\min\{\eta SC^{*}n^{-1},(SC^{*})^{1/2}n^{-1/2}\}.$ (B.4)
+Since $\log|\mathcal{G}|\geq\log\mathcal{N}_{\mathcal{G}}(\epsilon)$ for any $\epsilon\in(0,1)$, equation (B.4) yields the desired bound.
+[…]
+$\forall s\in{\mathcal{S}},\pi^{\mathsf{ref}}(-1|s)=C^{-1},\pi^{\mathsf{ref}}(+1|s)=1-C^{-1};$
+where $C\geq 1$ is a parameter to be specified later.
+We construct $2^{S}$ Bernoulli reward functions, in particular, $\forall\tau\in\{\pm 1\}^{S}$, the mean function $r_{\tau}$ of the reward (indexed by $\tau$) is defined as
+$r_{\tau}(s,-1)=0.5+\tau_{s}\delta,r_{\tau}(s,+1)=0.5-\alpha$
+for any state $s\in{\mathcal{S}}$, where $\alpha\in(0,1/2)$ and $\delta\in(0,1/4]$ will be specified later.
+We omit the $\mathrm{RKL}$ subscript in the following argument when it is clear in context.
+By 2.2, the optimal policy $\pi_{\tau}^{*}$ under $r_{\tau}$ is
+$\forall s\in{\mathcal{S}},\pi_{\tau}^{*}(-1|s)=\frac{\exp\big{(}\eta(\alpha+\tau_{s}\delta)\big{)}}{\exp\big{(}\eta(\alpha+\tau_{s}\delta)\big{)}+C-1},\pi_{\tau}^{*}(+1|s)=\frac{C-1}{\exp\big{(}\eta(\alpha+\tau_{s}\delta)\big{)}+C-1}.$ (B.5)
+Since $C^{*}\leq\exp(\eta/4)$, we assign $C=C^{*}$ and $\alpha=\eta^{-1}\log(C-1)\Leftrightarrow C-1=\exp(\eta\alpha)$, which gives
+$\forall s\in{\mathcal{S}},$ $\frac{\pi_{\tau}^{*}(-1|s)}{\pi^{\mathsf{ref}}(-1|s)}\leq C\frac{\exp(\eta(\alpha+\tau_{s}\delta))}{C-1+\exp(\eta(\alpha+\tau_{s}\delta))}=C\frac{\exp(\eta\tau_{s}\delta)}{1+\exp(\eta\tau_{s}\delta)}\leq C=C^{*};$
+$\forall s\in{\mathcal{S}},$ $\frac{\pi_{\tau}^{*}(+1|s)}{\pi^{\mathsf{ref}}(+1|s)}=\frac{C}{C-1}\cdot\frac{1}{\exp(\eta\tau_{s}\delta)+1}\leq C=C^{*};$
+where the last inequality is due to the assumption $C^{*}\geq 2$. Therefore, we obtain
+$\max_{\tau\in\{\pm 1\}^{S}}C^{\pi_{\tau}^{*}}\leq C^{*}.$ (B.6)
+We will abuse the notation $\mathrm{SubOpt}(\widehat{\pi};\tau)\coloneqq\mathrm{SubOpt}(\widehat{\pi};r_{\tau})$. Since $\rho=\mathsf{Unif}({\mathcal{S}})$,
+$\mathrm{SubOpt}(\widehat{\pi};\tau)=\frac{1}{S}\sum_{s=1}^{S}\mathrm{SubOpt}_{s}(\widehat{\pi};\tau),$ (B.7)
+where
+$\mathrm{SubOpt}_{s}(\widehat{\pi};\tau)$ $=\langle{\pi_{\tau}^{*}(\cdot|s)},{r_{\tau}(s,\cdot)-\eta^{-1}\log\frac{\pi_{\tau}^{*}(\cdot|s)}{\pi^{\mathsf{ref}}(\cdot|s)}}\rangle-\langle{\widehat{\pi}(\cdot|s)},{r_{\tau}(s,\cdot)-\eta^{-1}\log\frac{\widehat{\pi}(\cdot|s)}{\pi^{\mathsf{ref}}(\cdot|s)}}\rangle$
+$=\frac{1}{\eta}\mathbb{E}_{a\sim\pi^{*}_{\tau}(\cdot|s)}\bigg{[}\log\frac{\pi^{\mathsf{ref}}(a|s)\cdot\exp\big{(}\eta r_{\tau}(s,a)\big{)}}{\pi^{*}_{\tau}(a|s)}\bigg{]}$
+$\quad\ -\frac{1}{\eta}\mathbb{E}_{a\sim\widehat{\pi}(\cdot|s)}\bigg{[}\log\frac{\pi^{\mathsf{ref}}(a|s)\cdot\exp\big{(}\eta r_{\tau}(s,a)\big{)}}{\widehat{\pi}(a|s)}\bigg{]}$
+$=\frac{1}{\eta}\mathbb{E}_{a\sim\pi^{*}_{\tau}(\cdot|s)}\bigg{[}\log\Big{(}\sum_{b\in\mathcal{A}}\pi^{\mathsf{ref}}(b|s)\cdot\exp\big{(}\eta r_{\tau}(s,b)\big{)}\Big{)}\bigg{]}$
+$\quad\ -\frac{1}{\eta}\mathbb{E}_{a\sim\widehat{\pi}(\cdot|s)}\bigg{[}\log\frac{\pi^{\mathsf{ref}}(a|s)\cdot\exp\big{(}\eta r_{\tau}(s,a)\big{)}}{\widehat{\pi}(a|s)}\bigg{]}$
+$=\frac{1}{\eta}\mathbb{E}_{a\sim\widehat{\pi}(\cdot|s)}\bigg{[}\log\frac{\pi^{\mathsf{ref}}(a|s)\cdot\exp\big{(}\eta r_{\tau}(s,a)\big{)}}{\pi^{*}_{\tau}(a|s)}-\log\frac{\pi^{\mathsf{ref}}(a|s)\cdot\exp\big{(}\eta r_{\tau}(s,a)\big{)}}{\widehat{\pi}(a|s)}\bigg{]}$
+$=\eta^{-1}{\mathsf{KL}}\left(\widehat{\pi}\|\pi_{\tau}^{*}\right).$ (B.8)
+We write $\tau\sim_{s}\tau^{\prime}$ if $\tau,\tau^{\prime}\in\{\pm 1\}^{{\mathcal{S}}}$ differ in only the $s$-th coordinate and $\tau\sim\tau^{\prime}$ if $\exists s\in{\mathcal{S}},\tau\sim_{s}\tau^{\prime}$.
+By B.8, $\forall s\in{\mathcal{S}}$, $\forall\tau,\tau^{\prime}\in\{\pm 1\}^{{\mathcal{S}}}$ with $\tau\sim_{s}\tau^{\prime}$,
+$\mathrm{SubOpt}_{s}(\widehat{\pi};\tau)+\mathrm{SubOpt}_{s}(\widehat{\pi};\tau^{\prime})$
+$=\eta^{-1}{\mathsf{KL}}\left(\widehat{\pi}\|\pi_{\tau}^{*}\right)+\eta^{-1}{\mathsf{KL}}\left(\widehat{\pi}\|\pi_{\tau^{\prime}}^{*}\right)$
+$=2\eta^{-1}\sum_{a\in\mathcal{A}}\widehat{\pi}(a|s)\log\frac{\widehat{\pi}(a|s)}{\sqrt{\pi_{\tau}^{*}(a|s)\pi_{\tau^{\prime}}^{*}(a|s)}}$
+$=2\eta^{-1}{\mathsf{KL}}\left(\widehat{\pi}(\cdot|s)\|\bar{\pi}_{\tau,\tau^{\prime}}(\cdot|s)\right)-2\eta^{-1}\mathbb{E}_{a\sim\widehat{\pi}(\cdot|s)}\log\Big{(}\sum_{b\in\mathcal{A}}\sqrt{\pi_{\tau}^{*}(b|s)\pi_{\tau^{\prime}}^{*}(b|s)}\Big{)}$
+$\geq-2\eta^{-1}\log\Big{(}\sum_{b\in\mathcal{A}}\sqrt{\pi_{\tau}^{*}(b|s)\pi_{\tau^{\prime}}^{*}(b|s)}\Big{)}$
+$=\frac{1}{\eta}\log\frac{(\exp(\eta\delta)+1)(\exp(-\eta\delta)+1)}{4},$ (B.9)
+where $\bar{\pi}(\cdot|s)=\sqrt{\pi^{*}_{\tau}(\cdot|s)\pi^{*}_{\tau^{\prime}}(\cdot|s)}/\sum_{b\in\mathcal{A}}\sqrt{\pi^{*}_{\tau}(b|s)\pi^{*}_{\tau^{\prime}}(b|s)}$ for every $s\in{\mathcal{S}}$, the inequality is due to the non-negativity of KL divergence, and the last equality follows from B.5 together with the design choice $C-1=\exp(\eta\alpha)$.
+Case $\eta\delta\leq 2$.
+Recall that $\forall x\in\mathbb{R},(\mathrm{e}^{x}+\mathrm{e}^{-x})/2-1=x^{2}\sum_{k=0}^{\infty}\frac{x^{2k}}{(2k+2)!}\geq x^{2}/2$, which implies
+B.9 $=\frac{1}{\eta}\log\bigg{(}1+\frac{1}{2}\Big{(}\frac{\mathrm{e}^{\eta\delta}+\mathrm{e}^{-\eta\delta}}{2}-1\Big{)}\bigg{)}\geq\frac{1}{\eta}\log\Big{(}1+\frac{\eta^{2}\delta^{2}}{4}\Big{)}\geq\frac{1}{\eta}\cdot\frac{\eta^{2}\delta^{2}/4}{2}=\eta\delta^{2}/8.$ (B.10)
+Here, the last inequality is due to $\eta^{2}\delta^{2}/4\leq 1$ and $\forall x\in[0,1],\log(1+x)\geq x/2$.
+Case $\eta\delta>2$.
+We have $-\eta^{-1}2\log 2\geq-\delta\log 2$, which implies the following bound.
+B.9 $\geq\frac{1}{\eta}\log\frac{\exp(\eta\delta)+1}{4}\geq\frac{\eta\delta-2\log 2}{\eta}=\delta-\eta^{-1}2\log 2\geq(1-\log 2)\delta\geq 3\delta/10.$ (B.11)
+In summary, B.10 and B.11 imply that $\forall s\in{\mathcal{S}}$, $\forall\tau,\tau^{\prime}\in\{\pm 1\}^{{\mathcal{S}}}$ with $\tau\sim_{s}\tau^{\prime}$,
+$\mathrm{SubOpt}_{s}(\widehat{\pi};\tau)+\mathrm{SubOpt}_{s}(\widehat{\pi};\tau^{\prime})\geq\frac{\eta\delta^{2}}{8}\wedge\frac{3\delta}{10}.$ (B.12)
+Let $P_{\tau}$ be the distribution of $(s,a,y)$ where $s\sim\rho,a\sim\pi^{\mathsf{ref}}(\cdot|s)$, and $y\sim\mathsf{Bern}(r_{\tau}(s,a))$.
+Then $\forall x\in{\mathcal{S}}$ $\forall\tau,\tau^{\prime}\in\{\pm 1\}^{{\mathcal{S}}}$ with $\tau\sim_{x}\tau^{\prime}$,
+${\mathsf{KL}}\left(P_{\tau}\|P_{\tau^{\prime}}\right)$ $=\frac{1}{S}\sum_{s,a}\pi^{\mathsf{ref}}(a|s){\mathsf{KL}}\left(\mathsf{Bern}(r_{\tau}(s,a))\|\mathsf{Bern}(r_{\tau^{\prime}}(s,a))\right)$
+$=\frac{1}{S}\cdot C^{-1}{\mathsf{KL}}\left(\mathsf{Bern}(r_{\tau}(x,-1))\|\mathsf{Bern}(r_{\tau^{\prime}}(x,-1))\right)$
+$\leq\frac{4\delta^{2}}{SC(0.25-\delta^{2})}\leq\frac{16\delta^{2}}{3SC},$ (B.13)
+where we use the requirement $\delta\leq 1/4$ and ${\mathsf{KL}}\left(\mathsf{Bern}(p)\|\mathsf{Bern}(q)\right)\leq(p-q)^{2}/\big{(}q(1-q)\big{)}$.
+Then let $P_{\mathcal{D}_{\tau}}$ be the distribution of $\mathcal{D}$ given the mean reward function $r_{\tau}$, we employ B.13 to get
+${\mathsf{KL}}\left(P_{\mathcal{D}_{\tau}}\|P_{\mathcal{D}_{\tau^{\prime}}}\right)=n{\mathsf{KL}}\left(P_{\tau}\|P_{\tau^{\prime}}\right)\leq\frac{16n\delta^{2}}{3SC}.$ (B.14)
+Since $n\geq 16SC^{*}=16SC$ by design, we can set $\delta=\sqrt{SC/n}$ (which ensures $\delta\leq 1/4$) to obtain
+$\sup_{\mathrm{inst}}\mathrm{SubOpt}(\widehat{\pi};\mathrm{inst})$ $\geq\sup_{\tau\in\{\pm 1\}^{{\mathcal{S}}}}\mathrm{SubOpt}(\widehat{\pi};\tau)$
+$\geq\frac{1}{S}\cdot S\cdot\frac{1}{4}\cdot\Big{(}\frac{\eta\delta^{2}}{8}\wedge\frac{3\delta}{10}\Big{)}\min_{\tau\sim\tau^{\prime}}\exp\Big{(}-{\mathsf{KL}}\left(P_{\mathcal{D}_{\tau}}\|P_{\mathcal{D}_{\tau^{\prime}}}\right)\Big{)}$
+$\geq\Big{(}\frac{\eta SC^{*}}{32n}\wedge\frac{3\sqrt{SC^{*}}}{40\sqrt{n}}\Big{)}\exp(-16/3)\gtrsim\frac{\eta SC^{*}}{n}\wedge\sqrt{\frac{SC^{*}}{n}}.$
+where the $S^{-1}$ in the second inequality comes from B.7, the second inequality is by substituting B.12 into Assouad’s Lemma (Lemma F.3), and the last inequality is due to B.14. ∎
+Appendix C Missing Proof from Section 3
+C.1 Proof of Theorem 3.2
+Before coming to the proof, we first introduce some useful properties. The following properties characterize the convexity of $f$-divergence when $f$ is (strongly) convex.
+The strong-convexity of $f$ implies that the corresponding $f$-divergence, $D_{f}(\cdot||\pi^{\mathsf{ref}})$ is also strongly convex with respect to all $\pi:{\mathcal{S}}\rightarrow\Delta(\mathcal{A})$ supported by $\pi^{\mathsf{ref}}$.
+Proposition C.1.
+Given context $s$, $D_{f}(\pi(\cdot|s)||\pi^{\mathsf{ref}}(\cdot|s))$ is strict convex with respect to $\pi$ if $f$ is strictly convex.
+Proposition C.2.
+Given context $s$, $\pi(\cdot|s)\mapsto D_{f}(\pi(\cdot|s)\|\pi^{\mathsf{ref}}(\cdot|s))$ is $4\alpha$-strong convex with respect to the metric $\mathsf{TV}$ if $f$ is $\alpha$-strongly convex.
+Proof of Proposition C.2.
+We first show the gradient of $D_{f}$ with respect to $\pi$.
+$\frac{\partial D_{f}(\pi||\pi^{\mathsf{ref}})}{\pi(a)}=\frac{\partial}{\partial\pi(a)}\sum_{b\in\mathcal{A}}\pi^{\mathsf{ref}}(b)f\bigg{(}\frac{\pi(b)}{\pi^{\mathsf{ref}}(b)}\bigg{)}=f^{\prime}\bigg{(}\frac{\pi(a)}{\pi^{\mathsf{ref}}(a)}\bigg{)}.$
+Now consider $\pi_{1},\pi_{2}\in\Delta(\mathcal{A})$ supported by $\pi^{\mathsf{ref}}$.
+$D_{f}(\pi_{1}||\pi^{\mathsf{ref}})-D_{f}(\pi_{2}||\pi^{\mathsf{ref}})-\left\langle\pi_{1}-\pi_{2},\nabla D_{f}(\pi_{2}||\pi^{\mathsf{ref}})\right\rangle$
+$\quad=\sum_{a\in\mathcal{A}}\pi^{\mathsf{ref}}(a)\Bigg{(}f\bigg{(}\frac{\pi_{1}(a)}{\pi^{\mathsf{ref}}(a)}\bigg{)}-f\bigg{(}\frac{\pi_{2}(a)}{\pi^{\mathsf{ref}}(a)}\bigg{)}\Bigg{)}-\sum_{a\in\mathcal{A}}\big{(}\pi_{1}(a)-\pi_{2}(a)\big{)}f^{\prime}\bigg{(}\frac{\pi_{2}(a)}{\pi^{\mathsf{ref}}(a)}\bigg{)}$
+$\quad=\sum_{a\in\mathcal{A}}\pi^{\mathsf{ref}}(a)\Bigg{(}f\bigg{(}\frac{\pi_{1}(a)}{\pi^{\mathsf{ref}}(a)}\bigg{)}-f\bigg{(}\frac{\pi_{2}(a)}{\pi^{\mathsf{ref}}(a)}\bigg{)}-\bigg{(}\frac{\pi_{1}(a)}{\pi^{\mathsf{ref}}(a)}-\frac{\pi_{2}(a)}{\pi^{\mathsf{ref}}(a)}\bigg{)}f^{\prime}\bigg{(}\frac{\pi_{2}(a)}{\pi^{\mathsf{ref}}(a)}\bigg{)}\Bigg{)}$
+$\quad\geq\frac{\alpha}{2}\sum_{a\in\mathcal{A}}\pi^{\mathsf{ref}}(a)\bigg{(}\frac{\pi_{1}(a)}{\pi^{\mathsf{ref}}(a)}-\frac{\pi_{2}(a)}{\pi^{\mathsf{ref}}(a)}\bigg{)}^{2}$
+$\quad=\frac{\alpha}{2}\sum_{a\in\mathcal{A}}\frac{1}{\pi^{\mathsf{ref}}(a)}\big{(}\pi_{1}(a)-\pi_{2}(a)\big{)}^{2}$
+$\quad\geq\frac{\alpha}{2}\bigg{(}\sum_{a\in\mathcal{A}}\big{|}\pi_{1}(a)-\pi_{2}(a)\big{|}\bigg{)}^{2},$
+where the first inequality holds due to $f$’s strong convexity and the second holds due to Cauchy–Schwarz.
+The proof finishes since $\|\pi_{1}-\pi_{2}\|_{1}=2{\mathsf{TV}}\left(\pi_{1}\|\pi_{2}\right)$.
 ∎
-Finally, we need the following lemma to bound the performance difference between two policy w.r.t.
-the KL-regularized objective $J$.
-For any given $f:{\mathcal{S}}\times\mathcal{A}\to\mathbb{R}$, we define $\pi_{f}(\cdot|s)\propto\exp\big{(}\eta\cdot f(s,\cdot)\big{)},\forall s\in{\mathcal{S}}$ to facilitate presentation; then $\pi^{*}=\pi_{f^{*}}$ by definition.
-Please refer to Appendix A.2 for the proof of the lemma.
-Lemma 2.17.
-Given $f:{\mathcal{S}}\times\mathcal{A}\to\mathbb{R}$, there exist $\gamma\in[0,1]$ such that for $f_{\gamma}=\gamma f+(1-\gamma)f^{*}$ and $\pi_{\gamma}\coloneqq\pi_{f_{\gamma}}$,
-$J(\pi^{*})-J(\pi_{f})\leq\eta\mathbb{E}_{(s,a)\sim\rho\times\pi_{\gamma}}\big{[}\big{(}f^{*}-f\big{)}^{2}(s,a)\big{]}.$
+We first introduce some notation and important properties concerning the convex conjugate of functions.
+Given some context $s$, we denote the regularization term as $H_{s}(\pi)=\eta^{-1}D_{f}(\pi(\cdot|s)\|\pi^{\mathsf{ref}}(\cdot|s))$.
+We use $H^{*}_{s}(r)$ to denote the convex conjugate of $H_{s}$, which is defined as
+$H^{*}_{s}(r)=\sup_{\pi\in{\mathcal{S}}\to\Delta^{|\mathcal{A}|}}\{\left\langle\pi(\cdot|s),r(s,\cdot)\right\rangle-H_{s}(\pi)\}.$
+We have the following properties for the convex conjugate. The first property gives the gradient of convex conjugate (see, e.g., Zhou 2018, Lemma 5).
+Proposition C.3.
+Given context $s$, and convex $f$, let $\pi_{r}\in\mathop{\mathrm{argmax}}_{\pi}\{\left\langle\pi(\cdot|s),r(s,\cdot)\right\rangle-H_{s}(\pi)\}$ for some $r$, then the gradient of $H^{*}_{s}$ is given by $\nabla H^{*}_{s}(r)=\pi_{r}(\cdot|s)$.
+We also need some properties of $\nabla^{2}H^{*}_{s}$, the Hessian matrix of the convex conjugate function. We first give the Hessian matrix of the original function $H_{s}$ as follows.
+$\nabla^{2}H_{s}(\pi)=\eta^{-1}\mathrm{diag}\Bigg{(}\pi^{\mathsf{ref}}(a_{1}|s)^{-1}f^{\prime\prime}\bigg{(}\frac{\pi(a_{1}|s)}{\pi^{\mathsf{ref}}(a_{1}|s)}\bigg{)},\cdots,\pi^{\mathsf{ref}}(a_{|\mathcal{A}|}|s)^{-1}f^{\prime\prime}\bigg{(}\frac{\pi(a_{|\mathcal{A}|}|s)}{\pi^{\mathsf{ref}}(a_{|\mathcal{A}|}|s)}\bigg{)}\Bigg{)}.$ (C.1)
+Furthermore, when $f$ is $\alpha$-strongly convex, we have
+$\nabla^{2}H_{s}(\pi)\succeq\alpha\eta^{-1}\mathrm{diag}\Big{(}\pi^{\mathsf{ref}}(a_{1}|s)^{-1},\cdots,\pi^{\mathsf{ref}}(a_{|\mathcal{A}|}|s)^{-1}\Big{)}.$
+The following lemma, which gives an estimate of $\nabla^{2}H^{*}_{s}$, is the pivot of the proof.
+Lemma C.4.
+For any reward $r:{\mathcal{S}}\times\mathcal{A}\to[0,1]$, we have
+$\nabla^{2}H^{*}_{s}(r)\preceq\alpha^{-1}\eta\mathrm{diag}\big{(}\pi^{\mathsf{ref}}(a_{1}|s),\cdots,\pi^{\mathsf{ref}}(a_{|\mathcal{A}|}|s)\big{)}.$
+Proof of Lemma C.4.
+Given reward function $r:{\mathcal{S}}\times\mathcal{A}\to[0,1]$, we consider
+$\pi_{r}\in\mathop{\mathrm{argmax}}_{\pi\in{\mathcal{S}}\to\Delta^{|\mathcal{A}|}}\{\left\langle\pi(\cdot|s),r(\cdot|s)\right\rangle-H_{s}(\pi)\}.$
+From equation (C.1) we know that $\nabla^{2}H_{s}(\pi_{r})$ is invertible.
+Therefore, by Penot 1994, Proposition 3.2, we have $\nabla^{2}H^{*}_{s}(r)\preceq(\nabla^{2}H_{s}(\pi_{r}))^{-1}$.
+Since $f$ is $\alpha$-strongly convex, we have
+$\nabla^{2}H^{*}_{s}(r)\preceq\alpha^{-1}\eta\mathrm{diag}\big{(}\pi^{\mathsf{ref}}(a_{1}|s),\cdots,\pi^{\mathsf{ref}}(a_{|\mathcal{A}|}|s)\big{)},$
+which finishes the proof. ∎
 Now we are ready to prove Theorem 3.2.
 Proof of Theorem 3.2.
-Following the proof of Zhao et al. (2024, Theorem 3.3), we know that there exists $\bar{\gamma}\in[0,1]$ such that
-$J(\pi^{*})-J(\widehat{\pi})\leq\eta G(\bar{\gamma})\leq\eta G(0),$ (2.6)
-where the first inequality holds due to Lemma 2.17 and the second inequality holds due to the event $\mathcal{E}$ and Lemma 2.16. The term $G(0)$ can be further bounded by
-$G(0)$ $=\eta\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}\Big{[}\big{(}{\widehat{f}}-{f^{*}}\big{)}^{2}(s,a)\Big{]}$
-$\leq 4\eta\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}[\Gamma_{n}^{2}(s,a)]$
-$=4\eta\beta^{2}\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}\big{[}D^{2}_{\mathcal{F}}((s,a);\pi^{\mathsf{ref}})\big{]}$
-$=\widetilde{O}(\eta D_{\pi^{*}}^{2}n^{-1}),$ (2.7)
-where the second inequality holds conditioned on $\mathcal{E}(\delta)$ because of Lemma 2.16, and the last inequality follows from the definition of $\mathcal{E}(\delta)$ together with Line 2.
-By Lemma 2.13, we know that event $\mathcal{E}$ holds with probability at least $1-\delta$, which finishes the proof.
-∎
-3 KL-Regularized Contextual Dueling Bandits
-In this section, we extend our algorithm to KL-regularized dueling bandit setting, where the learner receives preference feedback instead of absolute signals.
-We formally introduce setting and extended algorithm.
-Except the objective, the core setting follows Zhu et al.
-(2023); Zhan et al.
-(2023).
-Our objective follows Xiong et al.
-(2024); Zhao et al.
-(2024).
-3.1 Problem Setup
-We still [^4] We overload some notations in Section 2 by their dueling counterparts for notational simplicity.
-consider contextual bandits $({\mathcal{S}},\mathcal{A},r,\pi^{\mathsf{ref}})$ where ${\mathcal{S}}$ is the state space, $\mathcal{A}$ is the action space and $r:{\mathcal{S}}\times\mathcal{A}\to[0,1]$ is the reward function.
-But only relative preference feedback is available, viz., we have an $\mathrm{i.i.d.}$ offline dataset $\mathcal{D}=\{(s_{i},a^{1}_{i},a^{2}_{i},y_{i})\}_{i=1}^{n}$, where $s_{i}\in{\mathcal{S}}$ is generated from distribution $\rho$ and $a^{1}_{i},a^{2}_{i}\sim\pi^{\mathsf{ref}}$.
-The binary preference label $y_{i}=1$ indicates $a_{i}^{1}$ is preferred over $a_{i}^{2}$ and 00 for the opposite.
-Given a state $s$ and two corresponding action $a^{1}$ and $a^{2}$, let $y$ be the preference label such that $y=1$ indicates $a^{1}\succ a^{2}$ and $y=0$ for $a^{2}\succ a^{1}$.
-In this work we consider the Bradley-Terry Model, where $\mathbb{P}[y=1|s,a^{1},a^{2}]=\sigma(r(s_{i},a^{1}_{i})-r(s_{i},a^{2}_{i}))$, where $\sigma(x)=(1+e^{-x})^{-1}$ is the link function.
-The objective here identical to (2.1) and thus the optimal policy $\pi^{*}$ also follows the closed form in (2.2).
-Our goal is still to find an $\epsilon$-optimal policy.
-To control the complexity of the function class $\mathcal{F}$, we assume that Assumption 2.1 still holds here.
-Concentrability
-Analogous to Section 2, we need our estimation from offline dataset generalizable to the state-action pairs visited by our obtained policy, and therefore requires similar $D^{2}$-divergence and corresponding policy coverage conditions.
-However, in dueling bandit, we are unable to observe and estimate the absolute value of rewards.
-Therefore, intuitively, the best estimation $f$ we can achieve is that for any state $s$ and actions $a^{1},a^{2}$, our estimated $f(s,a^{1})-f(s,a^{2})\approx r(s,a^{1})-r(s,a^{2})$.
-This implies that there exists some mapping $b:{\mathcal{S}}\to[-1,1]$ such that $f(s,a)-b(s)\approx r(s,a)$ on the offline data.
-Consequently, it is natural to define the the following modified $D^{2}$-divergence for characterizing how the error $f(s,a)-b(s)-r(s,a)$ generalize to unseen samples.
-Definition 3.1.
-Given a class of functions $\mathcal{F}\subset({\mathcal{S}}\times\mathcal{A}\to\mathbb{R})$ and some policy $\pi$, let $\mathcal{B}=({\mathcal{S}}\to[-1,1])$ be the function class, define the $D^{2}$-divergence $D^{2}_{\mathcal{F}}((s,a);\pi)$ as
-$\sup_{f,g\in\mathcal{F}}\inf_{b\in\mathcal{B}}\frac{\big{(}f(s,a)-g(s,a)-b(s)\big{)}^{2}}{\mathbb{E}_{s\sim\rho}\operatorname{{\rm Var}}_{a^{\prime}\sim\pi(\cdot|s^{\prime})}[f(s^{\prime},a^{\prime})-g(s^{\prime},a^{\prime})]}.$
-A similar definition has been introduced in Zhao et al. (2024, Definition 2.6), which underpins the following two assumptions characterizing the coverage ability of $\pi^{\mathsf{ref}}$.
-Equipped with the overloaded $D^{2}$-divergence definition, we are now able to define all-policy concentrability and single concentrability similar as in Section 2.
-Assumption 3.2 (All-Policy Concentrability).
-Given a reference policy $\pi^{\mathsf{ref}}$, there exists $D>0$ such that $D^{2}=\sup_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}D^{2}_{\mathcal{F}}((s,a);\pi^{\mathsf{ref}})$.
-Assumption 3.3 (Single-Policy Concentrability).
-Given a reference policy $\pi^{\mathsf{ref}}$, there exists $D_{\pi^{*}}>0$ such that $D_{\pi^{*}}^{2}=\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}[D^{2}_{\mathcal{F}}((s,a);\pi^{\mathsf{ref}})]$.
-Remark 3.4.
-Similar single-policy concentrability assumptions have appeared in previous work in offline contextual dueling bandits.
-Specifically, Huang et al.
-(2024) defined an $L_{\infty}$-concentrability coefficient $\mathcal{C}_{\infty}^{\pi^{*}}=\sup_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}\big{(}\frac{\pi^{*}(a|s)}{\pi^{\mathsf{ref}}(a|s)}\big{)}$.
-Song et al.
-(2024) introduced a stronger assumption that $\mathcal{C}_{\infty}^{\pi}$ is bounded for all the policy $\pi$ in the KL-ball surrounding $\pi^{*}$.
-In our Definition 3.1, instead of simply considering the ratio of probability measure, we introduce a more refined eluder-type uncertainty term which also relies on the structure of the model class.
-As a result, $D_{\pi^{*}}$ can be small even if $L_{\infty}$-concentrability coefficient of $\pi^{*}$ is extremely large as long as the reward model generalizes well after training on the offline dataset.
-There has also been similar notions of partial coverage for the analysis of model-based RL (Uehara and Sun, 2021; Wang et al., 2024), which fused the information of the transition kernel class to the density-ratio-based formulation.
-3.2 Algorithm
+Consider our estimation ${\bar{g}}$ which approximates the ground truth reward function ${g^{*}}$, we know that
+$\widehat{\pi}=\mathop{\mathrm{argmax}}_{\pi\in{\mathcal{S}}\to\Delta(\mathcal{A})}\Big{\{}\mathbb{E}_{(s,a)\sim\rho\times\pi}[{\bar{g}}(s,a)]-\eta^{-1}\mathbb{E}_{s\sim\rho}\big{[}D_{f}(\pi||\pi^{\mathsf{ref}})\big{]}\Big{\}}.$
+We have the following sub-optimality decomposition
+$J(\pi^{*})-J(\widehat{\pi})$ $=\mathbb{E}_{s\sim\rho}\Big{[}\mathbb{E}_{a\sim\pi^{*}}[{g^{*}}(s,a)]-\mathbb{E}_{a\sim\widehat{\pi}}[{g^{*}}(s,a)]-\eta^{-1}\big{[}D_{f}(\pi^{*}\|\pi^{\mathsf{ref}})-D_{f}(\widehat{\pi}\|\pi^{\mathsf{ref}})\big{]}\Big{]}$
+$=\mathbb{E}_{s\sim\rho}\big{[}H^{*}_{s}({g^{*}})-H^{*}_{s}({\bar{g}})-\left\langle\widehat{\pi},{g^{*}}-{\bar{g}}\right\rangle\big{]}$
+$=\mathbb{E}_{s\sim\rho}\big{[}H^{*}_{s}({g^{*}})-H^{*}_{s}({\bar{g}})-\left\langle\nabla H^{*}_{s}({\bar{g}}),{g^{*}}-{\bar{g}}\right\rangle\big{]}$
+$=\mathbb{E}_{s\sim\rho}[({g^{*}}-{\bar{g}})^{\top}\nabla^{2}H^{*}_{s}(\widetilde{g})({g^{*}}-{\bar{g}})],$
+where $\widetilde{g}=\gamma{g^{*}}+(1-\gamma){\bar{g}}$ and $\gamma\in[0,1]$ and the last equation holds due to Taylor’s expansion.
+Now, for any $\delta\in(0,1)$ and $\epsilon_{c}>0$, with probability at least $1-\delta$
+$J(\pi^{*})-J(\widehat{\pi})$ $=\mathbb{E}_{s\sim\rho}[({g^{*}}-{\bar{g}})^{\top}\nabla^{2}H^{*}_{s}(\widetilde{g})({g^{*}}-{\bar{g}})]$
+$\leq\alpha^{-1}\eta\mathbb{E}_{s\sim\rho}\Big{[}({g^{*}}-{\bar{g}})^{\top}\mathrm{diag}\big{(}\pi^{\mathsf{ref}}(a_{1}|s),\cdots,\pi^{\mathsf{ref}}(a_{|\mathcal{A}|}|s)\big{)}({g^{*}}-{\bar{g}})\Big{]}$
+$=\alpha^{-1}\eta\mathbb{E}_{(s,a)\sim\rho\times\pi^{\mathsf{ref}}}\big{[}\big{(}{g^{*}}(s,a)-{\bar{g}}\big{(}s,a))^{2}\big{]}$
+$\leq\alpha^{-1}\eta\bigg{(}\frac{128}{3n}\log(2\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+18\epsilon_{c}\bigg{)},$
+where the first inequality holds due to Lemma C.4 and last inequality holds due to equation (B.1). Setting $\epsilon_{c}=O(n^{-1})$ completes the proof. ∎
+C.2 Proof of Theorem 3.4
+We first provide the following lemma that gives the close form of optimal policy under $\chi^{2}$-divergence regularization.
+Lemma C.5 (Huang et al. (2025, Lemma G.2)).
+Let $\pi^{*}$ be the optimal policy of $\chi^{2}$-divergence regularized objective with reward function $r$, then $\pi^{*}$ has the closed form
+$\pi^{*}(\cdot)=\pi^{\mathsf{ref}}(\cdot)\max\big{\{}0,\eta(r(\cdot)-\lambda)\big{\}},\text{ where }\sum_{a\in\mathcal{A}}\pi^{*}_{f\mathrm{div}}(a)=1.$
+By Proposition C.2, $\pi^{*}_{f\mathrm{div}}=\mathop{\mathrm{argmax}}_{\pi\in\Delta(\mathcal{A})}J_{f\mathrm{div}}(\pi)$ is unique. The sub-optimality gap for $f$-divergence is consequently defined as
+$\mathrm{SubOpt}_{f\mathrm{div}}(\cdot)\coloneqq\mathrm{SubOpt}_{f\mathrm{div}}(\cdot;\mathcal{A},r,\pi^{\mathsf{ref}})=J_{f\mathrm{div}}(\pi^{*}_{f\mathrm{div}})-J_{f\mathrm{div}}(\cdot).$ (C.2)
+Now we are ready to prove Theorem 3.4.
+Proof of Theorem 3.4.
+We still consider the family of contextual bandits $\mathrm{CB}_{\mathcal{G}}$ given by (B.3).
+We, still, aim to prove the following statement.
+Fixing any $S\geq 32\log 2$, $\eta>4\log 2$ and $\alpha$, we set $f(x)\coloneqq\alpha(x-1)^{2}/2$, then for any estimator $\mathcal{D}\mapsto\widehat{\pi}\in\Delta(\mathcal{A}|{\mathcal{S}})$, for any $n$ sufficiently large, there exist some function class $\mathcal{G}$, such that $\exists\ \mathrm{inst}=({\mathcal{S}},\mathcal{A},\rho,r,\pi^{\mathsf{ref}},\eta)\in\mathrm{CB}_{\mathcal{G}}$ with $|{\mathcal{S}}|=S=\Theta(\log|\mathcal{G}|)$, and
+$\mathrm{SubOpt}_{f}(\widehat{\pi};\mathrm{inst})\gtrsim\alpha^{-1}\eta Sn^{-1}.$ (C.3)
+Since $\log|\mathcal{G}|\geq\log\mathcal{N}_{\mathcal{G}}(\epsilon)$ for any $\epsilon\in(0,1)$, equation (C.3) yields the desired bound.
+[…]
+$\mathcal{G}=\{r_{v}(s,-1)=1/2+v_{s}\delta,r_{v^{\prime}}(s,+1)=1/2+v_{s}\delta,\ \forall s\in{\mathcal{S}}|v\in\mathcal{V}\},$
+where $\delta\in(0,\eta^{-1}\alpha]$ is to be specified later.
+Fix some context $s$ and $v_{1}\neq v_{2}$ different at entry $s$ and corresponding reward $r_{1}$ and $r_{2}$.
+Without loss of generality, we assume $r_{1}(s,\cdot)=(1/2+\delta,1/2-\delta)$ and $r_{2}(s,\cdot)=(1/2-\delta,1/2+\delta)$.
+Then direct calculation implies that
+$\pi_{1}^{*}(\cdot|s)=\frac{1}{2}\max\{0,\eta\alpha^{-1}(r_{1}(s,\cdot)-{\lambda})\}=0.5\eta\alpha^{-1}(r_{1}(s,\cdot)-{\lambda}),$
+$\pi_{2}^{*}(\cdot|s)=\frac{1}{2}\max\{0,\eta\alpha^{-1}(r_{2}(s,\cdot)-{\lambda})\}=0.5\eta\alpha^{-1}(r_{2}(s,\cdot)-{\lambda}),$
+where $\lambda=0.5-\eta^{-1}\alpha$. Note that $2{\chi^{2}}(\mu\|\nu)+1=\sum_{a\in\mathcal{A}}[\mu(a)]^{2}/\nu(a)$ and ${\chi^{2}}=D_{f}$, we obtain that $\forall\ \widehat{\pi}$,
+$\mathrm{SubOpt}_{s}(\widehat{\pi}(\cdot|s);r_{1})+\mathrm{SubOpt}_{s}(\widehat{\pi}(\cdot|s);r_{2})$ (C.4)
+$\quad=\langle{r_{1}(s,\cdot)},{\pi^{*}_{1}(\cdot|s)}\rangle+\langle{r_{2}(s,\cdot)},{\pi^{*}_{2}(\cdot|s)}\rangle-\overbrace{\langle{r_{1}(s,\cdot)+r_{2}(s,\cdot)},{\widehat{\pi}(\cdot|s)}\rangle}^{=1}+\overbrace{2\eta^{-1}\alpha{\chi^{2}}(\widehat{\pi}(\cdot|s)\|\pi^{\mathsf{ref}}(\cdot|s))}^{\geq 0}$
+$\quad\quad-\eta^{-1}\alpha\cdot{\chi^{2}}(\pi^{*}_{1}(\cdot|s)\|\pi^{\mathsf{ref}}(\cdot|s))-\eta^{-1}\alpha\cdot{\chi^{2}}(\pi^{*}_{2}(\cdot|s)\|\pi^{\mathsf{ref}}(\cdot|s))$
+$\quad\geq 2\langle{r_{1}(s,\cdot)},{\pi^{*}_{1}(\cdot|s)}\rangle-1-2\eta^{-1}\alpha\cdot{\chi^{2}}(\pi^{*}_{1}(\cdot|s)\|\pi^{\mathsf{ref}}(\cdot|s))$
+$\quad=1+\frac{2\eta\delta^{2}}{\alpha}-1-\frac{\eta\delta^{2}}{\alpha}=\frac{\eta\delta^{2}}{\alpha}.$ (C.5)
+Now we take expectation over all possible contexts and recall that $\|v-v^{\prime}\|_{1}\geq S/2$ for $v\neq v^{\prime}$, we know that for any $r_{1}\neq r_{2}\in\mathcal{G}$
+$\mathrm{SubOpt}(\widehat{\pi};r_{1})+\mathrm{SubOpt}(\widehat{\pi};r_{2})\geq\frac{\eta\delta^{2}}{2\alpha}$
+Given any mean reward function $r\in\mathcal{G}$, let $P_{r}$ be the distribution of $(s,a,\mathtt{r})$ when $s\sim\rho$, $a\sim\pi^{\mathsf{ref}}(\cdot|s)$, and $\mathtt{r}\sim\mathsf{Bern}(r(s,a))$.
+Suppose $P_{\mathcal{D}_{r}}$ is the distribution of the dataset given mean reward function $r$, then ${\mathsf{KL}}\left(P_{\mathcal{D}_{r_{1}}}\|P_{\mathcal{D}_{r_{2}}}\right)=n{\mathsf{KL}}\left(P_{{r_{1}}}\|P_{{r_{2}}}\right)$ for any pair of $r_{1},r_{2}\in\mathcal{G}$.
+Now we invoke Fano’s inequality (Lemma F.2) to obtain
+$\inf_{\pi}\sup_{\mathrm{inst}\in\mathrm{CB}_{\mathcal{G}}}\mathrm{SubOpt}(\widehat{\pi};\mathrm{inst})$ $\geq\frac{\eta\delta^{2}}{4\alpha}\Bigg{(}1-\frac{\max_{r_{1}\neq r_{2}\in\mathcal{G}}{\mathsf{KL}}\left(P_{\mathcal{D}_{r_{1}}}\|P_{\mathcal{D}_{r_{2}}}\right)+\log 2}{\log|\mathcal{G}|}\Bigg{)}$
+$\geq\frac{\eta\delta^{2}}{4\alpha}\Bigg{(}1-\frac{64n\delta^{2}+8\log 2}{S}\Bigg{)},$
+where the second inequality holds due to $\mathsf{KL}(\mathsf{Bern}(p)\|\mathsf{Bern}(q))\leq(p-q)^{2}/[q(1-q)]$. Let $\delta=16\sqrt{\alpha\eta^{-1}n^{-1}}$, then we obtain that for all $\pi$ we have
+$\sup_{\mathrm{inst}\in\mathrm{CB}_{\mathcal{G}}}\mathrm{SubOpt}(\widehat{\pi};\mathrm{inst})\gtrsim\frac{\eta S}{\alpha n},$
+which finishes the proof in that $\log_{2}|\mathcal{G}|=S$. ∎
+Appendix D Missing Details from Section 4
+In this section, we provide additional details that has been omitted in Section 4.
+D.1 Algorithms for KL-regularized Contextual Dueling Bandits
+Algorithm 3 Offline KL-Regularized Pessimistic Contextual Dueling Bandit (KL-PCDB)
+0:  regularization $\eta$, reference policy $\pi^{\mathsf{ref}}$, function class ${\mathcal{G}}$, offline dataset $\mathcal{D}=\{(s_{i},a^{1}_{i},a^{2}_{i},y_{i})\}_{i=1}^{n}$
+1:  Compute the maximum likelihood estimator of the reward function
+${\bar{g}}=\mathop{\mathrm{argmin}}_{g\in{\mathcal{G}}}\sum_{i=1}^{n}\Big{[}y_{i}\log\sigma\Big{(}\big{[}g(s_{i},a^{1}_{i})-g(s_{i},a^{2}_{i})\big{]}\Big{)}+(1-y_{i})\log\sigma\Big{(}\big{[}g(s_{i},a^{2}_{i})-g(s_{i},a^{1}_{i})\big{]}\Big{)}\Big{]}$
+2:  Let ${\widehat{g}}(s,a)={\bar{g}}(s,a)-\Gamma_{n}(s,a)$, where $\Gamma_{n}(s,a)$ is the bonus term in D.1
+2:  $\widehat{\pi}(a|s)\propto\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta\cdot{\widehat{g}}(s,a)\big{)}$
 We elucidate KL-PCDB for offline KL-regularized contextual dueling bandits, whose pseudocode is summarized in Algorithm 3.
-KL-PCDB first estimate the ground truth function $f^{*}$ on offline dataset with maximum likelihood estimator (MLE) to estimate a function ${\bar{f}}\in\mathcal{F}$.
+KL-PCDB first estimate the ground truth function ${g^{*}}$ on offline dataset with maximum likelihood estimator (MLE) to estimate a function ${\bar{g}}\in{\mathcal{G}}$.
 After that, analogous to Algorithm 1, we adopt the principle of pessimism in the face of uncertainty.
 Specifically, we define the penalty term
-$\Gamma_{n}(s,a)=\beta\sqrt{D_{\mathcal{F}}^{2}((s,a),\pi^{\mathsf{ref}})},$ (3.1)
+$\Gamma_{n}(s,a)=\beta\sqrt{D_{{\mathcal{G}}}^{2}((s,a),\pi^{\mathsf{ref}})},$ (D.1)
 where
-$\beta^{2}=128\log(2\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)/3n+18\epsilon_{c}=\widetilde{O}(n^{-1})$ (3.2)
-and then subtract it from the MLE ${\bar{f}}$ to obtain a pessimistic estimator ${\widehat{f}}$. KL-PCB then output the policy $\widehat{\pi}$, maximizing the estimated objective
-$\widehat{J}(\pi)=\mathbb{E}_{(s,a)\sim\rho\times\pi}\bigg{[}{\widehat{f}}(s,a)-\eta^{-1}\log\frac{\pi(a|s)}{\pi^{\mathsf{ref}}(a|s)}\bigg{]},$
+$\beta^{2}=128\log(2\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)/3n+18\epsilon_{c}=\widetilde{O}(n^{-1})$ (D.2)
+and then subtract it from the MLE ${\bar{g}}$ to obtain a pessimistic estimator ${\widehat{g}}$. KL-PCB then output the policy $\widehat{\pi}$, maximizing the estimated objective
+$\widehat{J}(\pi)=\mathbb{E}_{(s,a)\sim\rho\times\pi}\bigg{[}{\widehat{g}}(s,a)-\eta^{-1}\log\frac{\pi(a|s)}{\pi^{\mathsf{ref}}(a|s)}\bigg{]},$
 the maximizer of which is in closed form as the counterpart of 2.2.
-$\widehat{\pi}(a|s)\propto\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta\cdot{\widehat{f}}(s,a)\big{)}.$
-Algorithm 2 Offline KL-Regularized Pessimistic Contextual Dueling Bandit (KL-PCDB)
-0:  regularization $\eta$, reference policy $\pi^{\mathsf{ref}}$, function class $\mathcal{F}$, offline dataset $\mathcal{D}=\{(s_{i},a^{1}_{i},a^{2}_{i},y_{i})\}_{i=1}^{n}$
-1:  Compute the maximum likelihood estimator of the reward function
-${\bar{f}}=\mathop{\mathrm{argmin}}_{f\in\mathcal{F}}\sum_{i=1}^{n}\Big{[}y_{i}\log\sigma\Big{(}\big{[}f(s_{i},a^{1}_{i})-f(s_{i},a^{2}_{i})\big{]}\Big{)}+(1-y_{i})\log\sigma\Big{(}\big{[}f(s_{i},a^{2}_{i})-f(s_{i},a^{1}_{i})\big{]}\Big{)}\Big{]}$
-2:  Let ${\widehat{f}}(s,a)={\bar{f}}(s,a)-\Gamma_{n}(s,a)$, where $\Gamma_{n}(s,a)$ is the bonus term in 3.1
-2:  $\widehat{\pi}(a|s)\propto\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta\cdot{\widehat{f}}(s,a)\big{)}$
-3.3 Main Results
-In this subsection, we provide our main results for KL-regularized contextual dueling bandits. The sample complexity upper bound of KL-PCDB is presented in the following theorem.
-Theorem 3.5.
-Under Assumption 3.3, if we set $\Gamma_{n}$ according to Lemma 3.1, then for sufficiently small $\epsilon\in(0,1)$, with probability at least $1-\delta$, $n=\widetilde{O}(\eta D_{\pi^{*}}^{2}\epsilon^{-1})$ is sufficient to guarantee the output policy $\widehat{\pi}$ of Algorithm 2 to be $\epsilon$-optimal.
-Remark 3.6.
+$\widehat{\pi}(a|s)\propto\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta\cdot{\widehat{g}}(s,a)\big{)}.$
+We provide the following theoretical guarantees for Algorithm 3.
+Theorem D.1.
+Under Assumption 4.3, if we set $\Gamma_{n}$ according to D.1, then for sufficiently small $\epsilon\in(0,1)$, with probability at least $1-\delta$, $n=\widetilde{O}\big{(}\eta D_{\pi^{*}}^{2}\epsilon^{-1}\big{)}$ is sufficient to guarantee the output policy $\widehat{\pi}$ of Algorithm 3 to be $\epsilon$-optimal.
+Remark D.2.
 Zhao et al.
@@
 The following theorem provides the sample complexity lower bound for KL-regularized dueling contextual bandits.
-Theorem 3.7.
-For any $\epsilon\in(0,1)$, $\eta>0$, and offline RL algorithm Alg, there is a KL-regularized contextual dueling bandit instance with $O(1)$ single-policy concentrability such that Alg requires at least $\Omega\big{(}\min\{{\eta\log\mathcal{N}_{\mathcal{F}}(\epsilon)}/{\epsilon},{\log\mathcal{N}_{\mathcal{F}}(\epsilon)}/{\epsilon^{2}}\}\big{)}$ samples to return an $\epsilon$-optimal policy.
+Theorem D.3.
+For any sufficiently small $\epsilon\in(0,1),\eta>0$, $21\leq C^{*}\leq\exp(\eta/2)/$ 2, and any algorithm Alg, there is a KL-regularized contextual dueling bandit instance with single-policy concentrability $D^{2}_{\pi^{*}}\leq 2C^{*}$ such that Alg requires at least $\Omega\big{(}\min\{{\eta C^{*}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})}/{\epsilon},{\log\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})}(C^{*})^{2}/{\epsilon^{2}}\}\big{)}$ samples to return an $\epsilon$-optimal policy.
+Remark D.4.
+Theorem D.3 shows that when $\epsilon$ is sufficiently small, any algorithm for offline KL-regularized contextual dueling bandits requires at least $\Omega(\eta D^{2}_{\pi^{*}}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon)\epsilon^{-1})$ samples to output an $\epsilon$-optimal policy, which matches the sample complexity upper bound in Theorem D.1, indicating that KL-PCB is nearly optimal.
+D.2 Algorithm and Results for $f$-divergence Regularized Contextual Dueling Bandit
+Algorithm 4 Offline $f$-Divergence Regularized Contextual Dueling Bandits ($f$-CDB)
+0:  regularization $\eta$, reference policy $\pi^{\mathsf{ref}}$, function class ${\mathcal{G}}$, offline dataset $\mathcal{D}=\{(s_{i},a^{1}_{i},a^{2}_{i},y_{i})\}_{i=1}^{n}$
+1:  Compute the maximum likelihood estimator of the reward function
+${\bar{g}}=\mathop{\mathrm{argmin}}_{g\in{\mathcal{G}}}\sum_{i=1}^{n}\Big{[}y_{i}\log\sigma\Big{(}\big{[}g(s_{i},a^{1}_{i})-g(s_{i},a^{2}_{i})\big{]}\Big{)}+(1-y_{i})\log\sigma\Big{(}\big{[}g(s_{i},a^{2}_{i})-g(s_{i},a^{1}_{i})\big{]}\Big{)}\Big{]}.$
+2:  Compute the optimal policy with respect to reward ${\bar{g}}$
+$\widehat{\pi}(\cdot|s)\leftarrow\mathop{\mathrm{argmax}}_{\pi(\cdot|s)\in\Delta(\mathcal{A})}\sum_{a\in\mathcal{A}}\pi(a|s){\bar{g}}(s,a)+\eta^{-1}D_{f}\big{(}\pi(\cdot|s)\|\pi^{\mathsf{ref}}(\cdot|s)\big{)}$
+2:  $\widehat{\pi}(a|s)$
+In this subsection, we present an offline learning algorithm for $f$-divergence regularized contextual dueling bandit, $f$-CDB, in Algorithm 4.
+$f$-CDB first leverages maximum likelihood estimator to find a function ${\bar{g}}\in{\mathcal{G}}$ that minimizes its risk on the offline dataset.
+Then the algorithm constructs the output policy $\widehat{\pi}$ that maximizes the f-divergence regularized objective induced by ${\bar{g}}$.
+Similar to Algorithm 2, we do not require any pessimism in $f$-CDB.
+The following theorem provides an upper bound of Algorithm 4.
+Theorem D.5.
+For any sufficiently small $\epsilon\in(0,1)$, and $\eta,\alpha>0$, with probability at least $1-\delta$, $n=\widetilde{O}\big{(}\alpha^{-1}\eta\log{\mathcal{N}}(\epsilon)\epsilon^{-1}\big{)}$ is sufficient to guarantee that the output policy $\widehat{\pi}$ of Algorithm 4 is $\epsilon$-optimal.
+The following theorem provides a lower bound for offline $f$-divergence regularized contextual dueling bandit with strongly convex $f$.
+Theorem D.6.
+For any $\epsilon\in(0,1)$, $\alpha,\eta>0$, and offline RL algorithm Alg, there is an $\alpha$-strongly convex $f$ and $f$-divergence regularized contextual dueling bandit instance such that Alg requires at least $\Omega\big{(}\alpha^{-1}\eta\log{\mathcal{N}}(\epsilon)\epsilon^{-1}\big{)}$ samples to return an $\epsilon$-optimal policy.
+Remark D.7.
+Theorem D.6 indicates that, when $\epsilon$ is sufficiently small, to produce an $\epsilon$-optimal policy, any algorithm for offline $f$-regularized contextual bandits with strongly convex $f$ requires at least $\widetilde{\Omega}(\alpha^{-1}\eta\epsilon^{-1})$ samples.
+This lower bound matches the sample-complexity upper bound in Theorem D.5, indicating that Algorithm 4 is nearly optimal.
+Appendix E Missing Proof from Appendix D
+E.1 Proof of Theorem D.1
+The proof follows the proof in Section 2. At the beginning, we first define the event $\mathcal{E}(\delta)$ given $\delta>0$ as
+$\mathcal{E}(\delta)\coloneqq\Big{\{}\exists\ b:{\mathcal{S}}\to[-1,1],\forall(s,a)\in{\mathcal{S}}\times\mathcal{A},\big{|}{\bar{g}}(s,a)-b(s)-{g^{*}}(s,a)\big{|}\leq\Gamma_{n}(s,a)\Big{\}}.$ (E.1)
+Here, $\Gamma_{n}$ is defined in (D.1). We abuse the notation and define $b(\cdot)$ as
+$b=\mathop{\mathrm{argmin}}_{\mathcal{B}}\sup_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}\Phi_{b}(s,a)-\Gamma_{n}(s,a),$ (E.2)
+where $\Phi_{b}(s,a)=\big{|}{\bar{g}}(s,a)-b(s)-{g^{*}}(s,a)\big{|}$ and when $\mathcal{E}$ holds, for all $(s,a)\in{\mathcal{S}}\times\mathcal{A}$, we have $\Phi_{b}(s,a)\leq\Gamma_{n}(s,a)$ This indicates that the least square estimation ${\bar{g}}$ obtained in Line 1 of Algorithm 3, after adjusted by some bias function $b$, is close to the true function ${g^{*}}$.
+The following lemma shows that this event holds with high probability.
+Lemma E.1.
+For any $\delta>0$, $\mathbb{P}(\mathcal{E}(\delta))\geq 1-\delta$.
 Proof.
-See Appendix A.3. ∎
-Remark 3.8.
-Theorem 3.7 can be concluded by plugging $D^{2}\geq D^{2}_{\pi^{*}}$ into Zhao et al.
-(2024, Theorem 4.3).
-This implies that the sample complexity lower bound in Theorem 3.7 also holds for KL-regularized dueling contextual bandit with all-policy concentrability.
-Remark 3.9.
-Theorem 3.7 shows that when $\epsilon$ is sufficiently small, any algorithm for offline KL-regularized contextual dueling bandit requires at least $\widetilde{\Omega}(\epsilon^{-1})$ samples to output a $\epsilon$-optimal policy.
-This sample complexity lower bound matches the sample complexity upper bound in Theorem 3.5, indicating that KL-PCDB is nearly optimal.
-3.4 Proof of Theorem 3.5
-The proof follows the proof in Section 2. At the beginning, we first define the event $\mathcal{E}(\delta)$ given $\delta>0$ as
-$\mathcal{E}(\delta)\coloneqq\Big{\{}\exists\ b:{\mathcal{S}}\to[-1,1],\forall(s,a)\in{\mathcal{S}}\times\mathcal{A},\big{|}{\bar{f}}(s,a)-b(s)-f^{*}(s,a)\big{|}\leq\Gamma_{n}(s,a)\Big{\}}.$ (3.3)
-Here, $\Gamma_{n}$ is defined in (3.1). We abuse the notation and define $b(\cdot)$ as
-$b=\mathop{\mathrm{argmin}}_{\mathcal{B}}\sup_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}\Phi_{b}(s,a)-\Gamma_{n}(s,a),$ (3.4)
-where $\Phi_{b}(s,a)=\big{|}{\bar{f}}(s,a)-b(s)-f^{*}(s,a)\big{|}$ and when $\mathcal{E}$ holds, for all $(s,a)\in{\mathcal{S}}\times\mathcal{A}$, we have $\Phi_{b}(s,a)\leq\Gamma_{n}(s,a)$ This indicates that the least square estimation ${\bar{f}}$ obtained in Line 1 of Algorithm 2, after adjusted by some bias function $b$, is close to the true function $f^{*}$.
-The following lemma shows that this event holds with high probability.
-Please refer to Appendix B.1 for the proof.
-Lemma 3.10.
-For any $\delta>0$, $\mathbb{P}(\mathcal{E}(\delta))\geq 1-\delta$.
+From Lemma F.1, we have that with probability at least $1-\delta$, it holds that
+$\mathbb{E}_{s^{\prime}\sim\rho}\operatorname{{\rm Var}}_{a^{\prime}\sim\pi^{\mathsf{ref}}(\cdot|s^{\prime})}\big{[}{\bar{g}}(s^{\prime},a^{\prime})-{g^{*}}(s^{\prime},a^{\prime})\big{]}\leq O\bigg{(}\frac{1}{n}\log(\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+\epsilon_{c}\bigg{)}.$ (E.3)
+It further holds true that for some $b:{\mathcal{S}}\to\mathbb{R}$
+$D_{\mathcal{G}}^{2}((s,a),\pi^{\mathsf{ref}}))\cdot\mathbb{E}_{s\sim\rho}\operatorname{{\rm Var}}_{a\sim\pi^{\mathsf{ref}}(\cdot|s)}\big{[}{\bar{g}}(s,a)-{g^{*}}(s,a)\big{]}\geq\big{(}{\bar{g}}(s,a)-b(s)-{g^{*}}(s,a)\big{)}^{2}.$ (E.4)
+Substituting (E.3) into (E.4), we have
+$\inf_{b}\big{(}{\bar{g}}(s,a)-b(s)-{g^{*}}(s,a)\big{)}^{2}$ (E.5)
+$\quad=\inf_{b}\frac{\big{(}{\bar{g}}(s,a)-b(s)-{g^{*}}(s,a)\big{)}^{2}}{\mathbb{E}_{s^{\prime}\sim\rho}\operatorname{{\rm Var}}_{a^{\prime}\sim\pi^{\mathsf{ref}}(\cdot|s^{\prime})}\big{[}{\bar{g}}(s^{\prime},a^{\prime})-{g^{*}}(s^{\prime},a^{\prime})\big{]}}\mathbb{E}_{s^{\prime}\sim\rho}\operatorname{{\rm Var}}_{a^{\prime}\sim\pi^{\mathsf{ref}}(\cdot|s^{\prime})}\big{[}{\bar{g}}(s^{\prime},a^{\prime})-{g^{*}}(s^{\prime},a^{\prime})\big{]}$
+$\quad\leq D_{{\mathcal{G}}}^{2}((s,a),\pi^{\mathsf{ref}})\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{g}}(s,a)-b(s)-{g^{*}}(s,a)\big{)}^{2}\big{]}$ (E.6)
+$\quad\leq D_{{\mathcal{G}}}^{2}((s,a),\pi^{\mathsf{ref}})O\bigg{(}\frac{1}{n}\log(\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+\epsilon_{c}\bigg{)},$ (E.7)
+where the first inequality holds due to the definition of $D_{{\mathcal{G}}}^{2}((s,a),\pi^{\mathsf{ref}})$ and the last inequality holds due to Lemma F.1. ∎
 We overload the following quantities. For any $\gamma\in[0,1]$ and $(s,a)\in{\mathcal{S}}\times\mathcal{A}$, we define
-$f_{\gamma}(s,a)\coloneqq\gamma({\widehat{f}}(s,a)-b(s))+(1-\gamma)f^{*}(s,a).$
+$g_{\gamma}(s,a)\coloneqq\gamma({\widehat{g}}(s,a)-b(s))+(1-\gamma){g^{*}}(s,a).$
 Furthermore, we introduce the following quantities
-$\pi_{\gamma}(\cdot|\cdot)=\pi_{f_{\gamma}}(\cdot|\cdot)\propto\pi^{\mathsf{ref}}(\cdot|\cdot)\exp\big{(}\eta f_{\gamma}(\cdot,\cdot)\big{)},$
-$G(\gamma)\coloneqq\mathbb{E}_{\rho\times\pi_{\gamma}}\big{[}\big{(}{\widehat{f}}(s,a)-b(s)-{f^{*}}(s,a)\big{)}^{2}\big{]},$
-where $b(\cdot)$ is defined in (3.4). We still have the monotonicity of the function $G(\gamma)$, which is characterized by the following lemma.
-Lemma 3.11.
+$\pi_{\gamma}(\cdot|\cdot)=\pi_{g_{\gamma}}(\cdot|\cdot)\propto\pi^{\mathsf{ref}}(\cdot|\cdot)\exp\big{(}\eta g_{\gamma}(\cdot,\cdot)\big{)},$
+$G(\gamma)\coloneqq\mathbb{E}_{\rho\times\pi_{\gamma}}\big{[}\big{(}{\widehat{g}}(s,a)-b(s)-{{g^{*}}}(s,a)\big{)}^{2}\big{]},$
+where $b(\cdot)$ is defined in (E.2). We still have the monotonicity of the function $G(\gamma)$, which is characterized by the following lemma.
+Lemma E.2.
 On event $\mathcal{E}(\delta)$, $0\in\mathop{\mathrm{argmax}}_{\gamma\in[0,1]}G(\gamma)$.
 Proof.
-For simplicity, we use $\triangle(s,a)$ to denote ${\widehat{f}}(s,a)-b(s)-{f^{*}}(s,a)$ in this proof.
+For simplicity, we use $\triangle(s,a)$ to denote ${\widehat{g}}(s,a)-b(s)-{{g^{*}}}(s,a)$ in this proof.
 Then on event $\mathcal{E}(\delta)$, we know that $\triangle(s,a)\leq 0$ for all $(s,a)\in{\mathcal{S}}\times\mathcal{A}$.
@@
 Finally, we have the proposition that adding some bias term $b:{\mathcal{S}}\to\mathbb{R}$ does not affect the resulting policy.
-Proposition 3.12.
-Let $b:{\mathcal{S}}\to\mathbb{R}$ be some bias function, then for all $f\in\mathcal{F}$ we have $J(\pi_{f})=J(\pi_{f-b})$, where $(f-b)(s,a)=f(s,a)-b(s)$.
-Please see Appendix B.2 for the proof of Proposition 3.12. Now we are ready to prove Theorem 3.5.
-Proof of Theorem 3.5.
-We proceed the proof under the event $\mathcal{E}(\delta)$. By Proposition 3.12, we know that
-$J(\pi^{*})-J(\widehat{\pi})$ $=J(\pi^{*})-J(\pi_{{\widehat{f}}})$
-$=J(\pi^{*})-J(\pi_{{\widehat{f}}-b}).$
+Proposition E.3.
+Let $b:{\mathcal{S}}\to\mathbb{R}$ be some bias function, then for all $g\in{\mathcal{G}}$ we have $J(\pi_{g})=J(\pi_{g-b})$, where $(g-b)(s,a)=g(s,a)-b(s)$.
+Proof.
+For any fixed state $s\in{\mathcal{S}}$, we have for any $a\in\mathcal{A}$ that,
+$\pi_{g}(a|s)$ $=\frac{\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta g(s,a)\big{)}}{\sum_{a^{\prime}\in\mathcal{A}}\pi^{\mathsf{ref}}(a^{\prime}|s)\exp\big{(}\eta g(s,a^{\prime})\big{)}}$
+[…]
+$=\frac{\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta[g(s,a)-b(s)]\big{)}}{\sum_{a^{\prime}\in\mathcal{A}}\pi^{\mathsf{ref}}(a^{\prime}|s)\exp\big{(}\eta[g(s,a^{\prime})-b(s)]\big{)}}$
+$=\pi_{g-b}(a|s),$
+which indicates that $\pi_{g}=\pi_{g-b}$. This immediately leads to $J(\pi_{g})=J(\pi_{g-b})$. ∎
+Now we are ready to prove Theorem D.1.
+Proof of Theorem D.1.
+We proceed the proof under the event $\mathcal{E}(\delta)$. By Proposition E.3, we know that
+$J(\pi^{*})-J(\widehat{\pi})$ $=J(\pi^{*})-J(\pi_{{\widehat{g}}})$
+$=J(\pi^{*})-J(\pi_{{\widehat{g}}-b}).$
 Consequently, there exist some $\gamma\in[0,1]$ and $b:{\mathcal{S}}\to[-1,1]$ such that
-$J(\pi^{*})-J(\widehat{\pi})$ $=J(\pi^{*})-J(\pi_{{\widehat{f}}-b})$
-$\leq\eta\mathbb{E}_{\rho\times\pi_{\gamma}}\big{[}\big{(}{\widehat{f}}(s,a)-b(s)-{f^{*}}(s,a)\big{)}^{2}\big{]}$
-$=\eta G(\gamma),$ (3.5)
+$J(\pi^{*})-J(\widehat{\pi})$ $=J(\pi^{*})-J(\pi_{{\widehat{g}}-b})$
+$\leq\eta\mathbb{E}_{\rho\times\pi_{\gamma}}\big{[}\big{(}{\widehat{g}}(s,a)-b(s)-{{g^{*}}}(s,a)\big{)}^{2}\big{]}$
+$=\eta G(\gamma),$ (E.8)
 where the inequality holds due to Lemma 2.16.
-Under event $\mathcal{E}(\delta)$, we know that ${\widehat{f}}(s,a)-b(s)\leq{f^{*}}(s,a)$.
-Together with Lemma 3.11, we obtain $G(\gamma)\leq G(0)$.
+Under event $\mathcal{E}(\delta)$, we know that ${\widehat{g}}(s,a)-b(s)\leq{{g^{*}}}(s,a)$.
+Together with Lemma E.2, we obtain $G(\gamma)\leq G(0)$.
 Therefore, we know that
-$J(\pi^{*})-J(\widehat{\pi})$ $\leq G(0)$ (3.6)
-$=\eta\mathbb{E}_{\rho\times\pi^{*}}\Big{[}\big{(}{\widehat{f}}(s,a)-b(s)-{f^{*}}(s,a)\big{)}^{2}\Big{]}$
-$\leq 4\eta\mathbb{E}_{\rho\times\pi^{*}}\Gamma_{n}^{2}(s,a)$
-$=4\eta\beta^{2}\mathbb{E}_{\rho\times\pi^{*}}\big{[}D^{2}_{\mathcal{F}}((s,a);\pi^{\mathsf{ref}})\big{]}$
-$=\widetilde{O}(\eta D_{\pi^{*}}^{2}n^{-1}),$ (3.7)
+$J(\pi^{*})-J(\widehat{\pi})$ $\leq G(0)$
+$=\eta\mathbb{E}_{\rho\times\pi^{*}}\Big{[}\big{(}{\widehat{g}}(s,a)-b(s)-{{g^{*}}}(s,a)\big{)}^{2}\Big{]}$
+$\leq 4\eta\mathbb{E}_{\rho\times\pi^{*}}\big{[}\Gamma_{n}^{2}(s,a)\big{]}$
+$=4\eta\beta^{2}\mathbb{E}_{\rho\times\pi^{*}}\big{[}D^{2}_{{\mathcal{G}}}((s,a);\pi^{\mathsf{ref}})\big{]}$
+$=\widetilde{O}\big{(}\eta D_{\pi^{*}}^{2}\log{{\mathcal{G}}}(\epsilon_{c})n^{-1}\big{)},$ (E.9)
 where the inequality holds due to the definition of $\mathcal{E}(\delta)$.
-Plugging (3.7) into (3.5), we know that $J(\pi^{*})-J(\widehat{\pi})$ has upper bound $\widetilde{O}(D_{\pi^{*}}^{2}n^{-1})$.
-By Lemma 3.10, event $\mathcal{E}$ with probability at least $1-\delta$, which concludes the proof.
+Plugging (E.9) into (E.8), we know that $J(\pi^{*})-J(\widehat{\pi})$ has upper bound $\widetilde{O}(D_{\pi^{*}}^{2}n^{-1})$.
+By Lemma E.1, event $\mathcal{E}$ with probability at least $1-\delta$, which concludes the proof.
 ∎
-4 Conclusions and Future Work
-In this work, we settled the sample complexity of KL-regularized offline contextual bandits.
-We first propose KL-PCB, which follows the principle of pessimism for reward estimation and yields the optimal policy accordingly.
-Our analysis leverages the curvature of KL-regularized objectives and integrates pessimism with a newly identified covariance-based observation, which in turn enables a neat refinement of a mean-value-type argument the its extreme point, showing that KL-PCB achieves a nearly optimal sample complexity under single-policy concentrability.
-The pair of novel techniques are decoupled from tricky algorithmic tweaks, and thus might be of independent interest.
-We then extended KL-PCB to offline KL-regularized contextual dueling bandit problem and introduced KL-PCDB, the analysis of which manifests the near-optimality of KL-PCDB under single-policy concentration.
-Currently, our focus is solely on the reverse-KL-regularized objective.
-It would be an interesting future direction to study more general regularization techniques like $f$-divergence (Csiszár, 1967) regularization, which includes both forward and reverse KL regularizations.
-Appendix A Missing Proofs from Section 2
-A.1 Proof of Lemma 2.13
-We first provide the following lemmas.
-Lemma A.1 (Zhao et al. 2024, Lemma C.1).
-For any policy $\pi$ and state-action pairs $\{(s_{i},a_{i})\}_{i=1}^{m}$ generated i.i.d.
-from $\rho\times\pi$, and $\epsilon_{c}<1$, with probability at least $1-\delta$, for any $f_{1}$ and $f_{2}$ we have
-$\mathbb{E}_{\rho\times\pi}\big{[}\big{(}f_{1}(s,a)-f_{2}(s,a)\big{)}^{2}\big{]}\leq\frac{2}{n}\sum_{i=1}^{n}\big{(}f_{1}(s_{i},a_{i})-f_{2}(s_{i},a_{i})\big{)}^{2}+\frac{32}{3n}\log(2\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+10\epsilon_{c},$
-where $\mathcal{N}_{\mathcal{F}}(\epsilon_{c})$ is the $\epsilon_{c}$-covering number of $\mathcal{F}$.
-Lemma A.2 (Zhao et al. 2024, Lemma C.2).
-For arbitrary policy $\pi$ and dataset $\{(s_{i},a_{i},r_{i})\}_{i=1}^{m}$ generated i.i.d., from the product of $\pi$, $\rho$ and the Bradley-Terry Model; let ${\bar{f}}$ be the least square estimator of $f^{*}$, then for any $10<\epsilon_{c}<$ 1 and $\delta>0$, with probability at least $1-\delta$ we have
-$\sum_{i=1}^{n}\big{(}{\bar{f}}(s_{i},a_{i})-f^{*}(s_{i},a_{i})\big{)}^{2}\leq 16\log(a\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+4n\epsilon_{c}.$
-Proof of Lemma 2.13.
-We have the following inequality
-$\big{(}{\bar{f}}(s,a)-f^{*}(s,a)\big{)}^{2}$ $=\frac{\big{(}{\bar{f}}(s,a)-f^{*}(s,a)\big{)}^{2}}{\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{f}}(s,a)-f^{*}(s,a)\big{)}^{2}\big{]}}\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{f}}(s,a)-f^{*}(s,a)\big{)}^{2}\big{]}$
-$\leq\sup_{f_{1},f_{2}\in\mathcal{F}}\frac{\big{(}f_{1}(s,a)-f_{2}(s,a)\big{)}^{2}}{\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}f_{1}(s,a)-f_{2}(s,a)\big{)}^{2}\big{]}}\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{f}}(s,a)-f^{*}(s,a)\big{)}^{2}\big{]}$
-$=D_{\mathcal{F}}^{2}((s,a),\pi^{\mathsf{ref}})\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{f}}(s,a)-f^{*}(s,a)\big{)}^{2}\big{]},$ (A.1)
-where the inequality holds by taking supremum to $f_{1},f_{2}\in\mathcal{F}$. Now we have
-$\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{f}}(s,a)-f^{*}(s,a)\big{)}^{2}\big{]}$ $\leq\frac{2}{n}\sum_{i=1}^{n}\big{(}{\bar{f}}(s_{i},a_{i})-f^{*}(s_{i},a_{i})\big{)}^{2}+\frac{32}{3n}\log(2\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+10\epsilon_{c}$
-$\leq\frac{2}{n}\big{[}16\log(\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+4n\epsilon_{c}\big{]}+\frac{32}{3n}\log(2\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+10\epsilon_{c}$
-$=\frac{128}{3n}\log(2\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+18\epsilon_{c},$ (A.2)
-where the first inequality holds due to Lemma A.1 and second holds due to Lemma A.2. Plugging (A.1) into (A.1) and setting $\epsilon_{c}=O(n^{-1})$ complete the proof. ∎
-A.2 Proof of Lemma 2.17
-This proof is extracted from the proof of Zhao et al. (2024, Theorem 3.3) and we present it here for completeness. By definition of our objective in (2.1), we have
-$J(\pi^{*})-J(\pi_{f})$
-$\quad=\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}\bigg{[}f^{*}(s,a)-\eta^{-1}\log\frac{\pi^{*}(a|s)}{\pi^{\mathsf{ref}}(a|s)}\bigg{]}-\mathbb{E}_{(s,a)\sim\rho\times\pi_{f}}\bigg{[}f^{*}(s,a)-\frac{1}{\eta}\log\frac{\pi_{f}(a|s)}{\pi^{\mathsf{ref}}(a|s)}\bigg{]}$
-$\quad=\frac{1}{\eta}\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}\bigg{[}\log\frac{\pi^{\mathsf{ref}}(a|s)\cdot\exp\bigl{(}\eta f^{*}(s,a)\bigr{)}}{\pi^{*}(a|s)}\bigg{]}-\frac{1}{\eta}\mathbb{E}_{(s,a)\sim\rho\times\pi_{f}}\bigg{[}\log\frac{\pi^{\mathsf{ref}}(a|s)\cdot\exp\bigl{(}\eta f^{*}(s,a)\bigr{)}}{\pi_{f}(a|s)}\bigg{]}$
-$\quad=\frac{1}{\eta}\mathbb{E}_{s\sim\rho}\big{[}\log Z_{f^{*}}(s)\big{]}-\frac{1}{\eta}\mathbb{E}_{s\sim\rho}\big{[}\log Z_{f}(s)\big{]}-\mathbb{E}_{s\sim\rho}\bigg{[}\sum_{a\in\mathcal{A}}\pi_{f}(a|s)\cdot\big{(}f^{*}(s,a)-f(s,a)\big{)}\bigg{]},$
-where for all $f\in\mathcal{F}$ we define $Z_{f}(\cdot)$ as follows,
-$Z_{f}(\cdot)\coloneqq\sum_{a\in\mathcal{A}}\pi^{\mathsf{ref}}(a|\cdot)\exp\big{(}\eta f(\cdot,a)\big{)}.$
-We further denote $\Delta(s,a)=f(s,a)-f^{*}(s,a)$ and $H_{s}(f)=\log Z_{f}(s)-\eta\sum_{a\in\mathcal{A}}\pi_{f}(a|s)\cdot\Delta(s,a)$.
-It worth noticing that $\eta^{-1}\mathbb{E}_{s\sim\rho}[H_{s}(f^{*})-H_{s}(f)]=J(\pi^{*})-J(\pi_{f})$.
-Now we take derivative of $H$ with respect to $\Delta(s,a)$,
-$\frac{\partial H_{s}(f)}{\partial\Delta(s,a)}$ $=\frac{\partial}{\partial\Delta(s,a)}\bigg{[}\log Z_{f}(s)-\eta\sum_{a\in\mathcal{A}}\pi_{f}(a|s)\cdot\Delta(s,a)\bigg{]}$
-$=\frac{1}{Z_{f}(s)}\cdot\pi^{\mathsf{ref}}(a|s)\exp\bigl{(}\eta\cdot f(s,a)\bigr{)}\cdot\eta-\eta\cdot\pi_{f}(a|s)$
-$\quad-\eta^{2}\cdot\Delta(s,a)\cdot\frac{\pi^{\mathsf{ref}}(a|s)\cdot\exp\bigl{(}\eta\cdot f(s,a)\bigr{)}}{Z_{f}(s)}+\eta^{2}\cdot\Delta(s,a)\cdot\frac{\bigl{[}\pi^{\mathsf{ref}}(a|s)\cdot\exp\bigl{(}\eta\cdot f(s,a)\bigr{)}\bigr{]}^{2}}{[Z_{f}(s)]^{2}}$
-$\quad+\eta\sum_{a^{\prime}\in\mathcal{A}\backslash\{a\}}\frac{\pi^{\mathsf{ref}}(a^{\prime}|x)\cdot\exp\bigl{(}\eta\cdot f(s,a^{\prime})\bigr{)}}{Z_{f}(s)}\cdot\eta\cdot\Delta(s,a^{\prime})\cdot\frac{\pi^{\mathsf{ref}}(a|s)\cdot\exp\bigl{(}\eta\cdot f(s,a)\bigr{)}}{Z_{f}(s)}$
-$=-\eta^{2}\pi_{f}(a|s)\Delta(s,a)+\eta^{2}[\pi_{f}(a|s)]^{2}\cdot\Delta(s,a)+\eta^{2}\sum_{a^{\prime}\in\mathcal{A}\backslash\{a\}}\pi_{f}(a^{\prime}|x)\pi_{f}(a|s)\Delta(s,a^{\prime}).$
-Therefore, by mean value theorem, there exists $\gamma\in[0,1]$ and $f_{\gamma}=\gamma f+(1-\gamma)f^{*}$ such that
-$H_{s}(f)-H_{s}(f^{*})$ $=-\eta^{2}\gamma\sum_{a\in\mathcal{A}}\pi_{f_{\gamma}}(a|s)\Delta(s,a)^{2}+\gamma\eta^{2}\sum_{a_{1}\in\mathcal{A}}\sum_{a_{2}\in\mathcal{A}}\pi_{f_{\gamma}}(a_{1}|x)\pi_{f_{\gamma}}(a_{2}|x)\Delta(s,a_{1})\Delta(s,a_{2})$
-$=-\eta^{2}\gamma\mathbb{E}_{a\sim\pi_{f_{\gamma}}}\big{[}\big{(}f^{*}(s,a)-f(s,a)\big{)}^{2}\big{]}+\gamma\eta^{2}\Big{(}\mathbb{E}_{a\sim\pi_{f_{\gamma}}}\big{[}\big{(}f^{*}(s,a)-f(s,a)\big{)}\big{]}\Big{)}^{2}$
-$\geq-\eta^{2}\mathbb{E}_{a\sim\pi_{f_{\gamma}}}\big{[}\big{(}f^{*}(s,a)-f(s,a)\big{)}^{2}\big{]},$
-where the inequality holds by omitting the second term and $\gamma\leq 1$. Now taking expectation over $\rho$, we have
-$J(\pi^{*})-J(\pi_{f})$ $=\eta^{-1}\mathbb{E}_{s\sim\rho}[H_{s}(f^{*})-H_{s}(f)]$
-$\leq\eta\mathbb{E}_{(s,a)\sim\rho\pi_{f_{\gamma}}}\big{[}\big{(}f^{*}(s,a)-f(s,a)\big{)}^{2}\big{]},$
-which concludes the proof.
-A.3 Proof of Theorem 2.10
-Proof of Theorem 2.10.
-We consider the bandit instance $\mathcal{K}$ and corresponding function class $\mathcal{F}$ constructed in the proof of Zhao et al.
-(2024, Theorem 3.1).
-Since the constructed $\pi^{\mathsf{ref}}(\cdot|s)$ in $\mathcal{K}$ is a uniform distribution for any $s\in{\mathcal{S}}$, the all-policy concentrability coefficient $D^{2}$ of $\mathcal{K}$ defined in Assumption 2.6 satisfies $D^{2}=O(1)$, which indicates that the single-policy concentrability coefficient $D_{\pi^{*}}^{2}$ defined in Assumption 2.7 satisfies $D_{\pi^{*}}^{2}\leq D^{2}=O(1)$.
-Thus, the proof of Zhao et al.
-(2024, Theorem 3.1) implies that any Alg requires at least $\Omega\big{(}\min\{{\eta\log\mathcal{N}_{\mathcal{F}}(\epsilon)}/{\epsilon},{\log\mathcal{N}_{\mathcal{F}}(\epsilon)}/{\epsilon^{2}}\}\big{)}$ to obtain an $\epsilon$-optimal policy.
-∎
-Appendix B Missing Proofs from Section 3
-B.1 Proof of Lemma 3.10
-The following on-poicy concentration bound is the prerequisite.
-Lemma B.1 (Zhao et al. 2024, Lemma D.4).
+E.2 Proof of Theorem D.3
+Proof of Theorem D.3.
+The proof is similar to the proof of Theorem 2.10.
+Consider the following family of contextual dueling bandit instances with $S\coloneqq|{\mathcal{S}}|,A\coloneqq|\mathcal{A}|<\infty$ and reward in some function class $\mathcal{G}$.
+$\mathrm{CDB}\coloneqq\{({\mathcal{S}},\mathcal{A},\rho,r,\pi^{\mathsf{ref}},\eta):r\in\mathcal{G},\rho\in\Delta({\mathcal{S}}),\pi^{\mathsf{ref}}\in\Delta(\mathcal{A}|{\mathcal{S}})\}.$ (E.10)
+Fixing any $S\geq 1$, $\eta>4\log 2$ and $C^{*}\in(2,\exp(\eta/4)]$, we aim to prove that, for any estimator $\mathcal{D}\mapsto\widehat{\pi}\in\Delta(\mathcal{A}|{\mathcal{S}})$, for any $n\geq 16SC^{*}$, there exist some function class $\mathcal{G}$, such that $\exists\ \mathrm{inst}=({\mathcal{S}},\mathcal{A},\rho,r,\pi^{\mathsf{ref}},\eta)\in\mathrm{CDB}$ with single-policy concentrability $C^{\pi^{*}}\leq 2C^{*}$, regularization coefficient $\eta$, $|{\mathcal{S}}|=S=\Theta(\log|\mathcal{G}|)$, and
+$\inf_{\mathrm{inst}\in\mathrm{CDB}}\mathrm{SubOpt}_{\mathrm{RKL}}(\widehat{\pi};\mathrm{inst})\gtrsim\min\{\eta SC^{*}n^{-1},(SC^{*})^{1/2}n^{-1/2}\}.$ (E.11)
+Since $\log|\mathcal{G}|\geq\log\mathcal{N}_{\mathcal{G}}(\epsilon)$ for any $\epsilon\in(0,1)$, the above bound yields the desired result.
+[…]
+$\forall s\in{\mathcal{S}},\pi^{\mathsf{ref}}(-1|s)=C^{-1},\pi^{\mathsf{ref}}(+1|s)=1-C^{-1};$
+where $C=C^{*}$. Then the total sub-optimality of any $\pi\in\Delta(\mathcal{A}|{\mathcal{S}})$ given any reward function $r:{\mathcal{S}}\times\mathcal{A}\to\mathbb{R}$ is
+$\mathrm{SubOpt}_{f\mathrm{div}}(\pi;r)=\frac{1}{S}\sum_{s=1}^{S}\mathrm{SubOpt}_{f\mathrm{div}}\big{(}\pi(\cdot|s);r(s,\cdot)\big{)}.$ (E.12)
+We further let $\alpha=\eta^{-1}\log(C-1)\Leftrightarrow C-1=\exp(\eta\alpha)$.
+We construct $2^{S}$ Bernoulli reward functions, in particular, $\forall\tau\in\{\pm 1\}^{S}$, the mean function $r_{\tau}$ of the reward (indexed by $\tau$) is defined as
+$r_{\tau}(s,-1)=0.5+\tau_{s}\delta,r_{\tau}(s,+1)=0.5-\alpha.$
+Then, following the derivation of B.10 and B.11, we know that $\max_{\tau\in\{\pm 1\}^{S}}C^{\pi_{\tau}^{*}}\leq C^{*}$, and $\forall s\in{\mathcal{S}}$, $\forall\tau,\tau^{\prime}\in\{\pm 1\}^{{\mathcal{S}}}$ with $\tau\sim_{s}\tau^{\prime}$,
+$\mathrm{SubOpt}_{s}(\widehat{\pi};\tau)+\mathrm{SubOpt}_{s}(\widehat{\pi};\tau^{\prime})\geq\frac{\eta\delta^{2}}{8}\wedge\frac{3\delta}{10}.$ (E.13)
+Let $P_{r}$ be the distribution of $(s,a^{1},a^{2},y)$ for $s\sim\rho$, $a^{1},a^{2}\overset{\scriptsize\mathrm{i.i.d.}}{\sim}\pi^{\mathsf{ref}}(\cdot|s)$ and $y\sim\mathsf{Bern}(\sigma(r(s,a^{1})-r(s,a^{2})))$.
+Now we set $\delta=\sqrt{S/n}$ and conclude that for $\tau\sim\tau^{\prime}$ with $\tau_{s}=-\tau_{s}^{\prime}$,
+${\mathsf{KL}}\left(P_{r_{\tau}}\|P_{r_{\tau^{\prime}}}\right)$
+$\quad=\frac{(C-1)}{SC^{2}}\sum_{s^{\prime},a^{1},a^{2}}{\mathsf{KL}}\left(\mathsf{Bern}(\sigma(r_{\tau}(s^{\prime},a^{1})-r(s^{\prime},a^{2})))\|\mathsf{Bern}(\sigma(r_{\tau^{\prime}}(s^{\prime},a^{1})-r(s^{\prime},a^{2})))\right)$
+$\quad=\frac{2(C-1)}{SC^{2}}\Big{(}{\mathsf{KL}}\left(\mathsf{Bern}(\sigma(\alpha+\delta))\|\mathsf{Bern}(\sigma(\alpha-\delta))\right)\vee{\mathsf{KL}}\left(\mathsf{Bern}(\sigma(\alpha-\delta))\|\mathsf{Bern}(\sigma(\alpha+\delta))\right)\Big{)}.$
+Since $\alpha,\delta\in(0,1/2)$, by the fact $\mathsf{KL}(P\|Q)\leq 2Q_{\min}^{-1}{\mathsf{TV}}\left(P\|Q\right)^{2}$ (see e.g., Polyanskiy and Wu (2025, Section 7.6)), we know that
+${\mathsf{KL}}\left(P_{r_{\tau}}\|P_{r_{\tau^{\prime}}}\right)$ $\leq\frac{2(C-1)}{SC^{2}}\frac{4}{1+\exp(\alpha+\delta)}\bigg{(}\frac{1}{1+\exp(\alpha-\delta)}-\frac{1}{1+\exp(\alpha+\delta)}\bigg{)}^{2}$
+$\leq\frac{4}{3SC}\frac{\exp(2\alpha)(\exp(\delta)-\exp(-\delta))^{2}}{(1+\exp(\alpha-\delta))^{4}}$
+$\leq\frac{4e}{3SC}(\exp(\delta)-\exp(-\delta))^{2}$
+$\leq 36S^{-1}C^{-1}\delta^{2},$ (E.14)
+where the second and third inequality hold due to $\alpha,\delta\leq 1/2$, and last inequality follows from $\exp(x)-\exp(-x)\leq 3x$ for $x\in[0,1/2]$.
+Now we set $\delta=\sqrt{SC/n}\leq 1/4$.
+We substitute E.13 into Assouad’s Lemma (Lemma F.3) and obtain that
+$\inf_{\mathrm{inst}\in\mathrm{CDB}}\mathrm{SubOpt}_{\mathrm{RKL}}(\widehat{\pi};\mathrm{inst})$ $\geq\frac{1}{4}S\cdot\frac{1}{S}\cdot\bigg{(}\frac{\eta\delta^{2}}{8}\wedge\frac{3\delta}{10}\bigg{)}\cdot\min_{\tau\sim\tau^{\prime}}\exp\Big{(}-{\mathsf{KL}}\left(P_{\mathcal{D}_{\tau}}\|P_{\mathcal{D}_{\tau^{\prime}}}\right)\Big{)}$
+$=\frac{1}{4}\bigg{(}\frac{\eta\delta^{2}}{8}\wedge\frac{3\delta}{10}\bigg{)}\exp\Big{(}-n{\mathsf{KL}}\left(P_{r_{\tau}}\|P_{r_{\mathcal{D}_{\tau^{\prime}}}}\right)\Big{)}$
+$\geq\frac{\exp(-36)}{32}\min\{\eta CSn^{-1},S^{2}C^{2}n^{-2}\},$
+where the $1/S$ comes from the denominator of (E.12) and the second inequality follows from E.14. ∎
+E.3 Proof of Theorem D.5
+Proof of Theorem D.5.
+The proof is similar to the proof of Theorem 3.2. Recall that $b(\cdot)$ defined in (E.2), we know that
+$\widehat{\pi}$ $=\mathop{\mathrm{argmax}}_{\pi\in\Delta^{d}}\Big{\{}\mathbb{E}_{(s,a)\sim\rho\times\pi}[{\bar{g}}(s,a)]-\eta^{-1}\mathbb{E}_{s\sim\rho}\big{[}D_{f}(\pi||\pi^{\mathsf{ref}})\big{]}\Big{\}}$
+$=\mathop{\mathrm{argmax}}_{\pi\in\Delta^{d}}\Big{\{}\mathbb{E}_{(s,a)\sim\rho\times\pi}[{\bar{g}}(s,a)-b(s)]-\eta^{-1}\mathbb{E}_{s\sim\rho}\big{[}D_{f}(\pi||\pi^{\mathsf{ref}})\big{]}\Big{\}}.$
+We have the following sub-optimality decomposition
+$J(\pi^{*})-J(\widehat{\pi})$ $=\mathbb{E}_{s\sim\rho}\Big{[}\mathbb{E}_{a\sim\pi^{*}}[{g^{*}}(s,a)]-\mathbb{E}_{a\sim\widehat{\pi}}[{g^{*}}(s,a)]-\eta^{-1}\big{[}D_{f}(\pi^{*}\|\pi^{\mathsf{ref}})-D_{f}(\widehat{\pi}\|\pi^{\mathsf{ref}})\big{]}\Big{]}$
+$=\mathbb{E}_{s\sim\rho}\big{[}H^{*}_{s}({g^{*}})-H^{*}_{s}({\bar{g}}-b)-\left\langle\widehat{\pi},{g^{*}}-{\bar{g}}+b\right\rangle\big{]}$
+$=\mathbb{E}_{s\sim\rho}\big{[}H^{*}_{s}({g^{*}})-H^{*}_{s}({\bar{g}}-b)-\left\langle\nabla H^{*}_{s}({\bar{g}}-b),{g^{*}}-{\bar{g}}+b\right\rangle\big{]}$
+$=\mathbb{E}_{s\sim\rho}[({g^{*}}-{\bar{g}}+b)^{\top}\nabla^{2}H^{*}_{s}(\widetilde{g})({g^{*}}-{\bar{g}}+b)],$
+where $\widetilde{g}=\gamma{g^{*}}+(1-\gamma){\bar{g}}$ and $\gamma\in[0,1]$, $({\bar{g}}-b)(s,a)={\bar{g}}(s,a)-b(s)$ and the last equation holds due to Taylor’s expansion.
+Now, for any $\delta\in(0,1)$ and $\epsilon_{c}>0$, with probability at least $1-\delta$
+$J(\pi^{*})-J(\widehat{\pi})$ $=\mathbb{E}_{s\sim\rho}[({g^{*}}-{\bar{g}}+b)^{\top}\nabla^{2}H^{*}_{s}(\widetilde{g})({g^{*}}-{\bar{g}}+b)]$
+$\leq\alpha^{-1}\eta\mathbb{E}_{s\sim\rho}\Big{[}({g^{*}}-{\bar{g}}+b)^{\top}\mathrm{diag}\big{(}\pi^{\mathsf{ref}}(a_{1}|s),\cdots,\pi^{\mathsf{ref}}(a_{d}|s)\big{)}({g^{*}}-{\bar{g}}+b)\Big{]}$
+$=\alpha^{-1}\eta\mathbb{E}_{(s,a)\sim\rho\times\pi^{\mathsf{ref}}}\big{[}\big{(}{g^{*}}(s,a)-{\bar{g}}\big{(}s,a)+b(s))^{2}\big{]}$
+$\leq\alpha^{-1}\eta\bigg{(}\frac{128}{3n}\log(2\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+18\epsilon_{c}\bigg{)},$
+where the first inequality holds due to Lemma C.4 and last inequality holds due to equation (E.3). Setting $\epsilon_{c}=O(n^{-1})$ completes the proof. ∎
+E.4 Proof of Theorem D.6
+Proof of Theorem D.6.
+We still consider the contextual dueling bandit instance class defined in (E.10).
+We show that given any positive $\alpha,\eta$, for any $n\geq S\cdot\max\{16,\eta^{2}\alpha^{-2}\}$, there exists $f:\mathbb{R}\to\mathbb{R}$ such that $f$ is $\alpha$-strongly convex, $\log|\mathcal{G}|=\Theta(S)$ and
+$\inf_{\widehat{\pi}\in\widehat{\Pi}(\mathcal{D})}\sup_{\mathrm{inst}\in\mathrm{CDB}}\mathrm{SubOpt}_{f\mathrm{div}}(\widehat{\pi};\mathrm{inst})\gtrsim\frac{\eta S}{\alpha n},$ (E.15)
+where $\mathcal{D}=\{(s_{i},a_{i}^{1},a_{i}^{2},y_{i})\}_{i=1}^{n}$ is the offline preference dataset, all (possibly randomized) maps from which to $\Delta(\mathcal{A}|{\mathcal{S}})$ is denoted by $\widehat{\Pi}(\mathcal{D})$.
+Since $S=\Theta(\log|\mathcal{G}|)\gtrsim\log\mathcal{N}_{\mathcal{G}}(\epsilon_{c})$ for all $\epsilon_{c}\in(0,1)$, we can conclude the theorem.
+[…]
+$\mathrm{SubOpt}_{f\mathrm{div}}(\pi;r)=\frac{1}{S}\sum_{s=1}^{S}\mathrm{SubOpt}_{f\mathrm{div}}\big{(}\pi(\cdot|s);r(s,\cdot)\big{)}.$ (E.16)
+We still consider the reward function class $\mathcal{G}$ indexed by {±$\{\pm 1\}^{{\mathcal{S}}}$. For all $\tau\in\{\pm 1\}^{{\mathcal{S}}}$ the reward instance “shaped” by $\tau$ is
+$r_{\tau}(s,a)=\frac{1}{2}+a\tau_{s}\cdot\sqrt{\frac{S}{n}},$ (E.17)
+where $a\tau_{s}=\pm 1$ because $a\in\mathcal{A}=\{\pm 1\}$.
+We thereby refer $\tau\sim\tau^{\prime}$ to any pair in {±$\{\pm 1\}^{{\mathcal{S}}}$ that differs only in one coordinate.
+$\forall\tau,\tau^{\prime}\in\{\pm 1\}^{{\mathcal{S}}}$, if $\tau\sim\tau^{\prime}$, then suppose $\tau_{s}=-\tau_{s}^{\prime}$, we have
+$\mathrm{SubOpt}_{f\mathrm{div}}(\pi(\cdot|s);r_{\tau}(s,\cdot))+\mathrm{SubOpt}_{f\mathrm{div}}(\pi(\cdot|s);r_{\tau^{\prime}}(s,\cdot))\geq\frac{\eta S}{\alpha n},$ (E.18)
+[…]
+${\mathsf{KL}}\left(P_{r_{\tau}}\|P_{r_{\tau^{\prime}}}\right)$ $=\frac{1}{SA^{2}}\sum_{s^{\prime},a^{1},a^{2}}{\mathsf{KL}}\left(\mathsf{Bern}(\sigma(r_{\tau}(s^{\prime},a^{1})-r(s^{\prime},a^{2}))\|\mathsf{Bern}(\sigma(r_{\tau^{\prime}}(s^{\prime},a^{1})-r(s^{\prime},a^{2}))\right)$
+$=\frac{1}{4S}\Big{(}{\mathsf{KL}}\left(\mathsf{Bern}(\sigma(2\delta))\|\mathsf{Bern}(\sigma(-2\delta))\right)+{\mathsf{KL}}\left(\mathsf{Bern}(\sigma(-2\delta))\|\mathsf{Bern}(\sigma(2\delta))\right)\Big{)}$
+$\leq\frac{1}{4S}\Big{(}(\exp(-2\delta)-1)^{2}+(\exp(2\delta)-1)^{2}\Big{)}$
+$\leq\frac{1}{2S}(\exp(2\delta)-1)^{2}\leq\frac{36\delta^{2}}{2S}=\frac{18}{n},$ (E.19)
+where the last inequality follows from $\exp(x)-1\leq 3x$ for $x\in[0,0.5]$ and $\delta=\sqrt{S/n}\leq 0.25$ by assumption. Therefore, we substitute E.18 into Assouad’s Lemma (Lemma F.3) to obtain
+LHS of E.15 $\geq\frac{1}{S}\cdot S\cdot\frac{\eta S}{\alpha n}\cdot\frac{1}{4}\cdot\min_{\tau\sim\tau^{\prime}}\exp\Big{(}-{\mathsf{KL}}\left(P_{\mathcal{D}_{\tau}}\|P_{\mathcal{D}_{\tau^{\prime}}}\right)\Big{)}$
+$=0.25\cdot\frac{\eta S}{\alpha n}\cdot\exp\Big{(}-n{\mathsf{KL}}\left(P_{r_{\tau}}\|P_{r_{\mathcal{D}_{\tau^{\prime}}}}\right)\Big{)}\geq\frac{\eta S}{\alpha n}\cdot\frac{1}{3}\cdot\exp(-18)\gtrsim\frac{\eta S}{\alpha n},$ (E.20)
+where the $1/S$ comes from the denominator of E.16 and the second inequality follows from E.19. ∎
+Appendix F Technical Results and Standard Results
+Lemma F.1 (Zhao et al. 2024, Lemma D.4).
 Consider a offline dataset $\{(s_{i},a^{1}_{i},a^{2}_{i},y_{i})\}_{i=1}^{n}$ generated from the product of the context distribution $\rho\in\Delta({\mathcal{S}})$, policy $\pi\in\Delta(\mathcal{A}|{\mathcal{S}})$, and the Bradley-Terry Model defined in Section 4.1.
-Suppose ${\bar{f}}$ is the result of MLE estimation of Algorithm 2, and we further define $b(s)=\mathbb{E}_{a\sim\pi(\cdot|s)}\big{[}{\bar{f}}(s,a)-f^{*}(s,a)\big{]}$, then with probability at least $1-2\delta$, we have
-$\mathbb{E}_{s,a\sim\rho\times\pi}\big{[}\big{(}{\bar{f}}(s,a)-f^{*}(s,a)-b(s)\big{)}^{2}\big{]}\leq O\bigg{(}\frac{1}{n}\log(\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+\epsilon_{c}\bigg{)}.$
-Proof of Lemma 3.10.
-From Lemma B.1, we have that with probability at least $1-\delta$, it holds that
-$\mathbb{E}_{s^{\prime}\sim\rho}\operatorname{{\rm Var}}_{a^{\prime}\sim\pi^{\mathsf{ref}}(\cdot|s^{\prime})}\big{[}{\bar{f}}(s^{\prime},a^{\prime})-f^{*}(s^{\prime},a^{\prime})\big{]}\leq O\bigg{(}\frac{1}{n}\log(\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+\epsilon_{c}\bigg{)}.$ (B.1)
-It further holds true that for some $b:{\mathcal{S}}\to\mathbb{R}$
-$D_{\mathcal{F}}^{2}((s,a),\pi^{\mathsf{ref}}))\cdot\mathbb{E}_{s\sim\rho}\operatorname{{\rm Var}}_{a\sim\pi^{\mathsf{ref}}(\cdot|s)}\big{[}{\bar{f}}(s,a)-f^{*}(s,a)\big{]}\geq\big{(}{\bar{f}}(s,a)-b(s)-f^{*}(s,a)\big{)}^{2}.$ (B.2)
-Substituting (B.1) into (B.2), we have
-$\inf_{b}\big{(}{\bar{f}}(s,a)-b(s)-f^{*}(s,a)\big{)}^{2}$
-$\quad=\inf_{b}\frac{\big{(}{\bar{f}}(s,a)-b(s)-f^{*}(s,a)\big{)}^{2}}{\mathbb{E}_{s^{\prime}\sim\rho}\operatorname{{\rm Var}}_{a^{\prime}\sim\pi^{\mathsf{ref}}(\cdot|s^{\prime})}\big{[}{\bar{f}}(s^{\prime},a^{\prime})-f^{*}(s^{\prime},a^{\prime})\big{]}}\mathbb{E}_{s^{\prime}\sim\rho}\operatorname{{\rm Var}}_{a^{\prime}\sim\pi^{\mathsf{ref}}(\cdot|s^{\prime})}\big{[}{\bar{f}}(s^{\prime},a^{\prime})-f^{*}(s^{\prime},a^{\prime})\big{]}$
-$\quad\leq D_{\mathcal{F}}^{2}((s,a),\pi^{\mathsf{ref}})\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{f}}(s,a)-b(s)-f^{*}(s,a)\big{)}^{2}\big{]}$
-$\quad\leq D_{\mathcal{F}}^{2}((s,a),\pi^{\mathsf{ref}})O\bigg{(}\frac{1}{n}\log(\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+\epsilon_{c}\bigg{)},$
-where the first inequality holds due to the definition of $D_{\mathcal{F}}^{2}((s,a),\pi^{\mathsf{ref}})$ and the last inequality holds due to Lemma B.1. ∎
-B.2 Proof of Proposition 3.12
-Proof of Proposition 3.12.
-For any fixed state $s\in{\mathcal{S}}$, we have for any $a\in\mathcal{A}$ that,
-$\pi_{f}(a|s)$ $=\frac{\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta f(s,a)\big{)}}{\sum_{a^{\prime}\in\mathcal{A}}\pi^{\mathsf{ref}}(a^{\prime}|s)\exp\big{(}\eta f(s,a^{\prime})\big{)}}$
-[…]
-$=\frac{\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta[f(s,a)-b(s)]\big{)}}{\sum_{a^{\prime}\in\mathcal{A}}\pi^{\mathsf{ref}}(a^{\prime}|s)\exp\big{(}\eta[f(s,a^{\prime})-b(s)]\big{)}}$
-$=\pi_{f-b}(a|s),$
-which indicates that $\pi_{f}=\pi_{f-b}$. This immediately leads to $J(\pi_{f})=J(\pi_{f-b})$. ∎
-B.3 Proof of Theorem 3.7
-Proof.
-Notice that Assumptions 3.2 and 3.3 imply $D^{2}\geq D^{2}_{\pi^{*}}$, which enables us to adapt Zhao et al.
-(2024, Theorem 4.3) to our scenario.
-In particular, we consider the bandit instance $\mathcal{K}$ and the corresponding function class $\mathcal{F}$ constructed in Zhao et al.
-(2024, Theorem 4.3).
-By Theorem 4.3 of Zhao et al.
-(2024), we know that the all-policy concentrability coefficient $D^{2}$ of $\mathcal{K}$ satisify $D^{2}=O(1)$, which indicates that its single-policy concentrability coefficient $D_{\pi^{*}}=O(1)$.
-Theorem 4.3 of Zhao et al.
-(2024) further indicates that Alg requires at least $\Omega\big{(}\min\{{\eta\log\mathcal{N}_{\mathcal{F}}(\epsilon)}/{\epsilon},{\log\mathcal{N}_{\mathcal{F}}(\epsilon)}/{\epsilon^{2}}\}\big{)}$ to obtain an $\epsilon$-optimal policy, which finishes the proof.
-∎
+Suppose ${\bar{g}}$ is the result of MLE estimation of Algorithm 3, and we further define $b(s)=\mathbb{E}_{a\sim\pi(\cdot|s)}\big{[}{\bar{g}}(s,a)-{g^{*}}(s,a)\big{]}$, then with probability at least $1-2\delta$, we have
+$\mathbb{E}_{s,a\sim\rho\times\pi}\big{[}\big{(}{\bar{g}}(s,a)-{g^{*}}(s,a)-b(s)\big{)}^{2}\big{]}\leq O\bigg{(}\frac{1}{n}\log(\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+\epsilon_{c}\bigg{)}.$
+Lemmas F.2 and F.3 are two standard reductions (Le Cam, 1973; Yu, 1997; Polyanskiy and Wu, 2025). See, e.g., Chen et al. (2024, Section 3) for a general proof.
+Lemma F.2 (Fano’s inequality).
+Fix any :-$\mathcal{R}\coloneq\{r_{1},\cdots,r_{S}\}$ and policy class $\Pi$, let $L:\Pi\times\mathcal{R}\to\mathbb{R}_{+}$ be some loss function.
+Suppose there exist some constant $c>0$ such that the following condition holds:
+$\min_{i\neq j}\min_{\pi\in\Pi}L(\pi,r_{i})+L(\pi,r_{j})\geq c.$
+Then we have
+$\inf_{\pi\in\Pi}\sup_{r\in\mathcal{R}}L(\pi,r)\geq\frac{c}{2}\bigg{(}1-\frac{\max_{i\neq j}\mathsf{KL}(P_{r_{i}}\|P_{r_{j}})+\log 2}{\log S}\bigg{)},$
+where $P_{r}$ is the distribution of dataset given model $r\in\mathcal{R}$.
+Lemma F.3 (Assouad’s Lemma).
+Let $\mathcal{R}$ be the set of instances, $\Pi$ be the set of estimators, $\Theta\coloneqq\{\pm 1\}^{S}$ for some $S>0$, and $\{L_{j}\}_{j=1}^{S}$ be $S$ functions from $\Pi\times\mathcal{R}$ to $\mathbb{R}_{+}$.
+Suppose $\{r_{\theta}\}_{\theta\in\Theta}\subset\mathcal{R}$ and the loss function is
+$L(\pi,r)\coloneqq\sum_{j=1}^{S}L_{j}(\pi,r),\forall(\pi,r)\in\Pi\times\mathcal{R}.$
+We denote $\theta\sim_{j}\theta^{\prime}$ if they differ only in the $j$-th coordinate. Further assume that
+$\theta\sim_{j}\theta^{\prime}\Rightarrow\inf_{\pi\in\Pi}L_{j}(\pi,r_{\theta})+L_{j}(\pi,r_{\theta^{\prime}})\geq c$ (F.1)
+for some $c>0$, then
+$\inf_{\pi\in\Pi}\sup_{r\in\mathcal{R}}L(\pi,r)\geq S\cdot\frac{c}{4}\min_{\exists j:\theta\sim_{j}\theta^{\prime}}\exp\Big{(}-{\mathsf{KL}}\left(P_{r_{\theta}}\|P_{r_{\theta^{\prime}}}\right)\Big{)},$
+where $P_{r}$ denotes the distribution of the dataset given $r\in\mathcal{R}$.
+The following Lemma F.4 is due to Gilbert (1952); Varshamov (1957), which is a classical result in coding theory.
+Lemma F.4.
+Suppose $\Sigma$ is a set of characters with $|\Sigma|=q$ where $q\geq 2$ is a prime power and $N>0$ is some natural number.
+Then there exists a subset $\mathcal{V}$ of $\Sigma^{N}$ such that (1) for any $v,v^{\prime}\in\mathcal{V},v\neq v_{j}$ , one has $d_{H}(v,v^{\prime})\geq N/2$ and (2) $\log_{q}|\mathcal{V}|\geq H_{q}(1/2)=\Theta(1)$, where $d_{H}$ is the Hamming distance and the entropy function $H$ is given by
+$H_{q}(x)=x\frac{\log(q-1)}{\log q}-x\frac{\log x}{\log q}-(1-x)\frac{\log(1-x)}{\log q}.$
+For example, when $q=2$, this means that there exists a subset $\mathcal{V}$ of $\{-1,1\}^{S}$ such that (1) $|\mathcal{V}|\geq\exp(S/8)$ and (2) for any $v,v^{\prime}\in\mathcal{V},v\neq v_{j}$ , one has $\|v-v^{\prime}\|_{1}\geq S/2$.

### v3

@@
 Abstract
-Although many popular reinforcement learning algorithms are underpinned by $f$-divergence regularization, their sample complexity with respect to the regularized objective still lacks a tight characterization.
-In this paper, we analyze $f$-divergence-regularized offline policy learning.
-For reverse Kullback–Leibler (KL) divergence, arguably the most commonly used one, we give the first $\widetilde{O}(\epsilon^{-1})$ sample complexity under single-policy concentrability for contextual bandits, surpassing existing $\widetilde{O}(\epsilon^{-1})$ bound under all-policy concentrability and $\widetilde{O}(\epsilon^{-2})$ bound under single-policy concentrability.
-Our analysis for general function approximation leverages the principle of pessimism in the face of uncertainty to refine a mean-value-type argument to its extreme.
-This in turn leads to a novel moment-based technique, effectively bypassing the need for uniform control over the discrepancy between any two functions in the function class.
-We further propose a lower bound, demonstrating that a multiplicative dependency on single-policy concentrability is necessary to maximally exploit the strong convexity of reverse KL.
-In addition, for $f$-divergences with strongly convex $f$, to which reverse KL does not belong, we show that the sharp sample complexity $\widetilde{\Theta}(\epsilon^{-1})$ is achievable even without single-policy concentrability.
-In this case, the algorithm design can get rid of pessimistic estimators.
-We further extend our analysis to dueling bandits, and we believe these results take a significant step toward a comprehensive understanding of $f$-divergence-regularized policy learning.
+Many offline reinforcement learning algorithms are underpinned by $f$-divergence regularization, but their sample complexity defined with respect to regularized objectives still lacks tight analyses, especially in terms of concrete data coverage conditions.
+In this paper, we study the exact concentrability requirements to achieve the $\widetilde{\Theta}(\epsilon^{-1})$ sample complexity for offline $f$-divergence-regularized contextual bandits.
+For reverse Kullback–Leibler (KL) divergence, arguably the most commonly used one, we achieve an $\widetilde{O}(\epsilon^{-1})$ sample complexity under single-policy concentrability for the first time via a novel pessimism-based analysis, surpassing existing $\widetilde{O}(\epsilon^{-1})$ bound under all-policy concentrability and $\widetilde{O}(\epsilon^{-2})$ bound under single-policy concentrability.
+We also propose a near-matching lower bound, demonstrating that a multiplicative dependency on single-policy concentrability is necessary to maximally exploit the curvature property of reverse KL.
+Moreover, for $f$-divergences with strongly convex $f$, to which reverse KL does not belong, we show that the sharp sample complexity $\widetilde{\Theta}(\epsilon^{-1})$ is achievable even without pessimistic estimation or single-policy concentrability.
+We further corroborate our theoretical insights with numerical experiments and extend our analysis to contextual dueling bandits.
+We believe these results take a significant step towards a comprehensive understanding of objectives with $f$-divergence regularization.
 1 Introduction
@@
 There are also lines of research on KL regularization in online learning (Cai et al., 2020; He et al., 2022; Ji et al., 2023) and convex optimization (Neu et al., 2017).
-However, the performance metric in most of these studies is still the unregularized reward maximization objective, under which the sample complexity is at least $\Omega(\epsilon^{-2})$.
-(See Section A.1 for detailed discussions)
+However, most of these works still study the unregularized reward maximization objective, against which the sample complexity is at least $\Omega(\epsilon^{-2})$.
+[^1] See Section A.1 for detailed reasons.
 Table 1: Comparison of sample complexity bounds for finding $\epsilon$-optimal policy for offline contextual bandits with KL- and (strongly convex) $f$-divergence regularization.
 Constants and $\mathrm{polylog}$ factors are omitted here except the metric entropy $\log\mathcal{N}$.
-“Reverse-KL” stands for KL-regularized contextual bandits and “$f$-divergence w/ s.c., $f$” for the counterpart with an $\alpha$-strongly convex $f$.
+“Reverse-KL” stands for KL-regularized contextual bandits and “$f$-divergence w/ s.c., ff” for the counterpart with an $\alpha$-strongly convex $f$.
 The two existing upper bounds are adapted from the implicit form in Xiong et al.
 (2024, Theorem 3.1) and Zhao et al.
-(2024, Theorem 3.3 and Theorem 4.4.), of which the detailed adaptions are deferred to Appendix A.
-Also see Remark 2.8 for discussion on the relationship between $D^{2}_{\pi^{*}}$ and $C^{\pi^{*}}$
+(2024, Theorem 3.3 and Theorem 4.4), of which the detailed adaptions are deferred to Appendix A.
+The relationship between $D^{2}_{\pi^{*}}$ and $C^{\pi^{*}}$ is detailed in Section 2.1.
 Regularizer Xiong et al. (2024) Zhao et al. (2024) This work
@@
 Lower - - $\alpha^{-1}\eta\epsilon^{-1}\log\mathcal{N}$
-Several recent papers (Xiong et al., 2024; Xie et al., 2024; Zhao et al., 2024; Foster et al., 2025) switched the focus to analyzing the suboptimality defined via the regularized objective (1.1), under which an $\Omega(\epsilon^{-1})$ sample complexity is possible (Zhao et al., 2024, Theorem 3.6).
-However, even restricted to the pure $\mathrm{i.i.d.}$ setting, existing analyses in this vein either result in still $\widetilde{O}(\epsilon^{-2})$ bounds (Xiong et al., 2024; Xie et al., 2024) or has a stringent all-policy concentrabiity dependency in their upper bounds (Zhao et al., 2024) (See Section 2.1 for details on coverage conditions).
-Thus, there are by far no perfectly matching bounds down to both the dependency of $\epsilon^{-1}$ and concrete coverage conditions for KL-regularized offline decision making.
-Moreover, these theoretical analyses set KL as the right target by default; but reverse KL is the $f$-divergence with $f(x)=x\log x$, which is merely convex.
-Therefore, it is also unknown whether those choices with even nicer (e.g., strongly convex) $f$, whose performance against the reward maximization objective can be provably promising (Zhan et al., 2022; Gabbianelli et al., 2024; Huang et al., 2024), can enjoy a better coverage dependency in the sample complexity under the corresponding regularized objectives.
-These perspectives motivate an open problem:
-What is the weakest coverage condition for offline policy learning to be minimax optimal with respect to $f$-divergence-regularized objectives?
-We are the first to concretely investigate this question by proposing sharp sample complexity analyses for two representative ad subclasses of $f$-divergence.
-In particular, for contextual bandits with KL regularization, we present sample complexity guarantee with linearly dependence on single-policy coverage ratio.
-Our lower bound further indicates that a multiplicative dependency on single-policy concentrability is necessary.
-Surprisingly, for $f$-divergence with $\alpha$-strongly-convex $f$, we prove nearly matching sample complexity bounds of $\widetilde{\Theta}(\alpha^{-1}\eta\epsilon^{-1})$ for the first time, settling the dependence on coverage completely.
-For the ease of comparison, we adapt previous sample complexity bounds to our setting and notation, and summarize them in Table 1.
+Several recent papers (Xiong et al., 2024; Xie et al., 2024; Zhao et al., 2024; Foster et al., 2025; Aminian et al., 2025) switched the focus to analyzing the sub-optimality defined via the regularized objective (1.1), under which an $\Omega(\epsilon^{-1})$ sample complexity is possible (Zhao et al., 2024, Theorem 3.6).
+However, even restricted to the pure $\mathrm{i.i.d.}$ setting, existing analyses in this vein either result in still $\widetilde{O}(\epsilon^{-2})$ bounds (Xiong et al., 2024; Xie et al., 2024) or has stringent (local) all-policy concentrability dependencies in their upper bounds (Zhao et al., 2024; Aminian et al., 2025).
+[^2] See Section 2.1 for details on coverage conditions.
+Thus, there are by far no tight bounds in terms of both the dependency of $\epsilon^{-1}$ and data coverage conditions for KL-regularized offline decision making.
+In addition, all analyses above set KL as the right target by default; but reverse KL is the $f$-divergence with $f(x)=x\log x$, which is merely convex.
+Therefore, it is also unknown whether $f$-divergence regularizers with even nicer (e.g., strongly convex) $f$, whose performance against the reward maximization objective are provably promising (Zhan et al., 2022; Gabbianelli et al., 2024; Huang et al., 2025b), can enjoy a better coverage dependency in their sample complexity when the corresponding regularized objectives serve as the performance metric.
+Because data coverage (i.e., concentrability) conditions captures the crucial distributional shift issue in offline RL (Levine et al., 2020), the aforementioned perspectives motivate a pivotal open problem:
+What is the weakest coverage condition required for offline learning to be near-optimal with respect to $f$-divergence-regularized objectives?
+We attack this problem by showing near-optimal sample complexity with matching concentrability dependencies for two representative subclasses of $f$-divergence.
+First, for contextual bandits with KL regularization, we achieve a near-optimal sample complexity guarantee with linear dependence on single-policy coverage ratio.
+Our novel lower bound further indicates that this multiplicative dependency on single-policy concentrability is necessary.
+Surprisingly, for $f$-divergence with $\alpha$-strongly-convex $f$, we prove nearly matching sample complexity bounds of $\widetilde{\Theta}(\alpha^{-1}\eta\epsilon^{-1})$, eliminating the dependence on coverage for the first time.
+For the ease of comparison, we adapt existing counterparts under our notation to the offline setting and summarize them in Table 1.
 1.1 Contributions
-  • For arguably the most commonly used KL regularization, we are the first to employ the principle of pessimism (Jin et al., 2021) to achieve a nearly tight sample complexity under single-policy concentrability.
-Our analysis roadmap leverages the strong convexity of KL and pessimism of the reward estimator, to refine a mean-value-type risk upper bound to its extreme which in turn leads to a novel moment-based analysis, effectively bypassing the need for uniform control over the discrepancy between any two functions in the function class.
-  • Under KL-regularization, we prove a lower bound growing linearly the density-ratio-based single-policy concentrability, which also strictly improved upon previous works (Zhao et al., 2024; Foster et al., 2025) in offline setting.
-This lower bound manifests the necessity of a multiplicative dependency on single-policy concentrability and firstly closes the gap between upper and lower bounds under a technically nontrivial case.
-  • For any $f$-divergence with strongly convex $f$, we are able to design a truly lightweight algorithm free of pessimism-based gadgets and still obtain the $\widetilde{\Theta}(\epsilon^{-1})$ sample complexity certified by a matching lower bound without coverage conditions.
-  • We extend all algorithmic ideas and hard instance constructions above to $f$-divergence-regularized contextual dueling bandits, achieving $\widetilde{\Theta}(\epsilon^{-1})$ sample complexity bounds.
-Moreover, all algorithms work for general reward function classes with finite metric entropy.
-1.2 Additional Related Work
-We review two additional lines of theoretical progress that are closely related to our algorithm design and analysis.
+  • For KL regularization, we propose a pessimism-based algorithm achieving the tight sample complexity under single-policy concentrability.
+We also obtain a lower bound that linearly scale with the density-ratio-based single-policy concentrability.
+Both results strictly improves upon previous works (Zhao et al., 2024; Foster et al., 2025) in the offline setting, showing that single-policy concentrability is both sufficient and necessary to achieve the $\widetilde{\Theta}(\epsilon^{-1})$ sample complexity.
+  • Technically speaking, our analysis exploits the strong convexity of KL and pessimism of the reward estimator, to refine a mean-value-type risk upper bound (Lemma 2.14) to its, which in turn leads to a novel moment-based analysis, effectively bypassing the need for uniform control over the discrepancy between any two functions in the function class.
+To the best of our knowledge, this machinery has not been used in the standard analysis of existing offline RL algorithms and may be of independent interest.
+  • For $f$-divergence-regularized objectives with strongly convex $f$, we design a truly lightweight algorithm free of pessimism-based gadgets and still obtain the $\widetilde{\Theta}(\epsilon^{-1})$ sample complexity certified by a matching lower bound without coverage conditions.
+  • We verify the statistical rates above in numerical experiments, and demonstrate the versatility of all algorithmic and constructive proof ideas above by extending them to $f$-divergence-regularized contextual dueling bandits (CDBs), achieving similar $\widetilde{\Theta}(\epsilon^{-1})$ sample complexity bounds.
+Moreover, all algorithms are applicable for reward function classes with small metric entropy.
+1.2 Key Related Work
+We review two key lines of theoretical progress that are relevant to our algorithm design and analysis.
 Pessimism in offline RL.
@@
 It is also possible to leverage the idea of pessimism to design model-based algorithms under general function approximation that are at least statistically efficient (Xie et al., 2021a; Uehara and Sun, 2021; Wang et al., 2024).
+The principle of pessimism has also been applied in counterfactual empirical risk minimization (Swaminathan and Joachims, 2015; London and Sandler, 2019) and offline policy learning (Sakhi et al., 2023; 2024), which are orthogonal to our contributions.
 However, in terms of risk decomposition, to the best of our knowledge, none of these pessimism-based analyses really goes beyond the performance difference lemma (Foster and Rakhlin, 2023, Lemma 13) or simulation lemma (Foster and Rakhlin, 2023, Lemma 23); both of which are not able to capture the strong concavity of KL-regularized objectives even in the bandit setting.
@@
 (2021); Di et al.
-(2023) is similar to ours, but their suboptimality gap is bounded by the sum of bonuses, which cannot lead to the desired sample complexity of our objective by direct adaption.
-Offline contextual dueling bandits.
-CDBs (Dudík et al., 2015) is the contextual extension of dueling bandits in classic literature of online learning from pairwise comparisons (Yue et al., 2012; Zoghi et al., 2014).
-Since the empirical breakthrough of preference-based RL fine-tuning of LLMs (Ouyang et al., 2022), the theory of offline CDBs has received more attention under linear (Zhu et al., 2023; Xiong et al., 2024) and general (Zhan et al., 2022; Zhao et al., 2024; Song et al., 2024; Huang et al., 2024) function approximation.
-Preference models without stochastic transitivity (Munos et al., 2023; Ye et al., 2024; Wu et al., 2024; Zhang et al., 2024) are beyond the scope of this work, viz., our preference labels are assumed to follow the Bradley-Terry Model (Bradley and Terry, 1952).
+(2024) is similar to ours, but their sub-optimality gap is bounded by the sum of bonuses, which cannot directly lead to the desired sample complexity of our objective.
+Offline CDBs.
+CDBs (Dudík et al., 2015) is the contextual extension of dueling bandits in the classic literature of online learning from pairwise comparisons (Yue et al., 2012; Zoghi et al., 2014).
+Since the empirical breakthrough of preference-based RL fine-tuning of LLMs (Ouyang et al., 2022), the theory of offline CDBs has received more attention under linear function approximation (Zhu et al., 2023; Xiong et al., 2024) and general function approximation (Zhan et al., 2022; Zhao et al., 2024; Song et al., 2024; Huang et al., 2025b).
+Preference models without stochastic transitivity (Munos et al., 2023; Ye et al., 2024; Wu et al., 2024; Zhang et al., 2024) are beyond the scope of this work, namely, our preference labels are assumed to follow the Bradley-Terry Model (Bradley and Terry, 1952).
 Notation.
-The sets ${\mathcal{S}}$ and $\mathcal{A}$ are assumed to be countable throughout the paper.
-For nonnegative sequences $\{x_{n}\}$ and $\{y_{n}\}$, we write $x_{n}=O(y_{n})$ if $\limsup_{n\to\infty}{x_{n}}/{y_{n}}<\infty$, $y_{n}=\Omega(x_{n})$ if $x_{n}=O(y_{n})$, and $y_{n}=\Theta(x_{n})$ if $x_{n}=O(y_{n})$ and $x_{n}=\Omega(y_{n})$.
-We further employ $\widetilde{O}(\cdot),\widetilde{\Omega}(\cdot)$, and $\widetilde{\Theta}$ to hide $\mathrm{polylog}$ factors.
-For countable $\mathcal{X}$ and $\mathcal{Y}$, we denote the family of probability kernels from $\mathcal{X}$ to $\mathcal{Y}$ by $\Delta(\mathcal{Y}|\mathcal{X})$.
-For $g:\mathcal{X}\to\mathbb{R}$, its infinity norm is denoted by $\|g\|_{\infty}\coloneqq\sup_{x\in\mathcal{X}}|g(x)|$.
-For a pair of probability measures $P\ll Q$ on the same space and function $f:\mathbb{R}_{+}\to\mathbb{R}$, their $f$-divergence is $D_{f}(P\|Q)\coloneqq\int f({\,\mathrm{d}}P/{\,\mathrm{d}}Q){\,\mathrm{d}}Q$.
-Specifically, when $f(x)=x\log x$, $f$-divergence becomes KL divergence denoted as ${\mathsf{KL}}\left(P\|Q\right)\coloneqq\int\log({{\,\mathrm{d}}P}/{{\,\mathrm{d}}Q}){\,\mathrm{d}}P$, and when $f(x)=|x-1|/2$, it becomes the total variation (TV) distance, which is denoted as ${\mathsf{TV}}\left(P\|Q\right)\coloneqq 0.5\int|{\,\mathrm{d}}P-{\,\mathrm{d}}Q|$.
+[…]
 2 KL-regularized Contextual Bandits
-In this section, we propose a pessimism-based algorithm, dubbed PCB-KL, for offline KL-regularized contextual bandits.
-In the following subsections, we also showcase our key technical novelty which couples the pessimism of our reward estimator to the non-trivial curvature property of KL regularization.
+In this section, we introduce a pessimism-based algorithm, PCB-KL, for offline KL-regularized contextual bandits.
+We then showcase our novel analysis techniques for PCB-KL, which couples the algorithmic pessimism with the curvature property of KL-regularized objectives.
 2.1 Problem Setup
@@
 In this work, we consider the KL-regularized objective
-$J(\pi)\coloneqq\mathbb{E}_{(s,a)\sim\rho\times\pi}\bigg{[}r(s,a)-\eta^{-1}\log\frac{\pi(a|s)}{\pi^{\mathsf{ref}}(a|s)}\bigg{]},$ (2.1)
+$J_{\eta}(\pi)\coloneqq\mathbb{E}_{(s,a)\sim\rho\times\pi}\bigg[r(s,a)-\eta^{-1}\log\frac{\pi(a|s)}{\pi^{\mathsf{ref}}(a|s)}\bigg],$ (2.1)
 where $\pi^{\mathsf{ref}}$ is a known reference policy and the “inverse temperature” $\eta$ controls the intensity of regularization.
-For simplicity, we assume that $\pi^{\mathsf{ref}}$ is also the behavior policy that generates the dataset $\mathcal{D}$.
-This type of “behavior regularization” has been also been studied in Zhan et al.
+For simplicity, we assume that $\pi^{\mathsf{ref}}$ is also the behavior policy that generates the dataset $\mathcal{D}$, which is similar to the type of “behavior regularization” studied in Zhan et al.
 (2022).
-The unique optimal policy $\pi^{*}\coloneqq\mathop{\mathrm{argmax}}_{\pi\in\Delta(\mathcal{A}|{\mathcal{S}})}J(\pi)$ is given by (See, e.g., Zhang 2023, Proposition 7.16)
+The unique optimal policy $\pi^{*}_{\eta}\coloneqq\mathop{\mathrm{argmax}}_{\pi\in\Delta(\mathcal{A}|{\mathcal{S}})}J_{\eta}(\pi)$ is given by (See, e.g., Zhang 2023, Proposition 7.16) [^3] We suppress $J_{\eta}$ into $J$ and $\pi^{*}_{\eta}$ into $\pi^{*}$ when they are clear in context in the following presentation.
 $\pi^{*}(\cdot|s)\propto\pi^{\mathsf{ref}}(\cdot|s)\exp\big(\eta\cdot r(s,\cdot)\big),\forall s\in{\mathcal{S}}.$ (2.2)
-A policy $\pi$ is said to be $\epsilon$-optimal if $J(\pi^{*})-J(\pi)\leq\epsilon$ and the goal of the agent is to find one such policy based on $\mathcal{D}$.
-To ensure $\epsilon$-optimality is achievable, we assume that $r$ lies in a known function class ${\mathcal{G}}\subset({\mathcal{S}}\times\mathcal{A}\to[0,1])$, from which the agent obtains the estimator $\widehat{r}$.
+A policy $\pi$ is said to be $\epsilon$-optimal if $\mathrm{SubOpt}_{\mathrm{RKL}}(\pi)\coloneqq J(\pi^{*})-J(\pi)\leq\epsilon$ and the goal of the agent is to find one such policy using $\mathcal{D}$.
+Note that $\mathrm{SubOpt}_{\mathrm{RKL}}(\cdot)$ is defined through 2.1 and thus depends on $\eta$.
+To ensure that $\epsilon$-optimality is achievable, we assume that $r$ lies in a known function class ${\mathcal{G}}\subset({\mathcal{S}}\times\mathcal{A}\to[0,1])$, from which the agent obtains an estimator $\widehat{r}$.
 More specifically, we work with general function approximation under realizability, which is as follows.
@@
 For this known function class ${\mathcal{G}}\subset({\mathcal{S}}\times\mathcal{A}\to[0,1])$, $\exists{g^{*}}\in{\mathcal{G}}$ with ${g^{*}}=r$.
-We also need a standard condition to control the complexity of ${\mathcal{G}}$ through the notion of covering number (Wainwright, 2019, Definition 5.1).
+We also employ the standard notion of covering number (Wainwright, 2019, Definition 5.1) as the complexity measure of the reward function class $\mathcal{G}$}.
 Definition 2.2 ($\epsilon$-net and covering number).
@@
 For any $\epsilon_{c}>0$, the $\epsilon_{c}$-covering number $\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})$ of $\mathcal{G}$} is $\mathrm{poly}(\epsilon_{c}^{-1})$.
-When ${\mathcal{G}}$ is the class of linear functions of dimension $d$ and radius $R$, the covering number of ${\mathcal{G}}$ is given by $\mathcal{N}_{{\mathcal{G}}}=O((1+R\epsilon^{-1})^{d})$ (Jin et al., 2020, Lemma D.6), indicating that the Assumption 2.3 on $\mathcal{N}_{{\mathcal{G}}}$ is mild.
+Assumption 2.3 allowing $\log\mathcal{N}_{\mathcal{G}}(\epsilon)$ to be roughly negligible is arguably mild.
+For example, when $\mathcal{G}$} is the class of linear functions of dimension $d$ and radius $R$, the covering number is $\mathcal{N}_{{\mathcal{G}}}(\epsilon)=O((1+R\epsilon^{-1})^{d})$ (Jin et al., 2020, Lemma D.6), which satisfies Assumption 2.3.
 Concentrability.
@@
 In this paper, in addition to density-ratio-based concentrability, we also adopt the following $D^{2}$-based concentrabilites to better capturing the nature of function class $\mathcal{G}$}.
-In detail, we first introduce $D^{2}$-divergence as follows.
+In detail, we start with the $D^{2}$-divergence as follows.
 Definition 2.5.
@@
 Therefore, the $D^{2}$-divergence quantifies how well the estimation on dataset collected by the behavior policy $\pi$ can be generalized to a specific state-action pair.
+Remark 2.6.
+For the tabular setting, a direct computation yields $D^{2}(s,a)=(\rho(s)\pi^{\text{ref}}(a|s))^{-1}$, which can be estimated by the visitation frequency empirically.
+Under linear function approximation, it is well known that $D^{2}(s,a)=\|\bm{\phi}(s,a)\|^{2}_{\bm{\Sigma}^{-1}}$ under mild conditions of the parameter space, where $\bm{\Sigma}=\mathbb{E}_{\rho\times\pi^{\text{ref}}}\bm{\phi}(s,a)\bm{\phi}(s,a)^{\top}$ is the covariance matrix, which can be estimated by empirical covariance matrices in practice, potentially with ridge regularization.
+For more general function classes like neural networks, the $D^{2}$ can also be efficiently approximated by heuristics as discussed in Xiong et al.
+(2024); Gupta et al.
+(2024); Xu et al.
+(2025).
 We are now ready to define the two notions of concentrability conditions.
@@
 However, in general, the two quantities characterizing single-policy concentrability $C^{\pi^{*}}$ and $D^{2}_{\pi^{*}}$ cannot be bounded by each other up to constant factors.
-Remark 2.8.
-In particular, we have $D^{2}_{\pi^{*}}\leq|{\mathcal{S}}||\mathcal{A}|C^{\pi^{*}}$, indicating that $C^{\pi^{*}}$ subsumes $D^{2}_{\pi^{*}}$ when $|{\mathcal{S}}|$ and $|\mathcal{A}|$ can be seen as constant.
+In particular, we have $D^{2}_{\pi^{*}}\leq|{\mathcal{S}}||\mathcal{A}|C^{\pi^{*}}$, indicating that $C^{\pi^{*}}$ subsumes $D^{2}_{\pi^{*}}$ when $|{\mathcal{S}}|$ and $|\mathcal{A}|$ can be seen as constants.
+We refer the reader to Appendix B for a further discussion on the relation between $C^{\pi^{*}}$ and $D^{2}_{\pi^{*}}$.
 2.2 Algorithm
@@
 In contrast, we construct a pessimistic estimator of $g^{*}$} following the well-known pessimism principle in offline RL (Jin et al., 2021).
-Specifically, we define the bonus term $\Gamma_{n}$ through the confidence radius $\beta=\sqrt{128\log\big{(}2\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta\big{)}/3n+18\epsilon_{c}}$ as
+Specifically, we define the bonus term $\Gamma_{n}$ through the confidence radius $\beta=\sqrt{128\log\big(2\mathcal{N}_{{\mathcal{G}}}(\epsilon)/\delta\big)/3n+18\epsilon}$ as
 $\Gamma_{n}(s,a)=\beta D_{{\mathcal{G}}}\big((s,a),\pi^{\mathsf{ref}}\big),\forall(s,a)\in{\mathcal{S}}\times\mathcal{A}.$ (2.3)
-We then obtain our pessimistic estimation ${\widehat{g}}$ by setting ${\widehat{g}}={\bar{g}}-\Gamma_{n}$, which is less than ${g^{*}}$ with high probability.
+We then obtain our pessimistic estimation $^{\widehat{g}}$ by setting ${\widehat{g}}={\bar{g}}-\Gamma_{n}$, which is less than $g^{*}$} with high probability.
+Formally, let the event $\mathcal{E}(\delta)$ given $\delta>0$ defined as
+$\mathcal{E}(\delta)\coloneqq\Big\{\sup\nolimits_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}\Big[\big|{\bar{g}}-{g^{*}}\big|-\Gamma_{n}\Big](s,a)\leq 0\Big\},$ (2.4)
+on which the least square estimation $\bar{g}$} obtained in Line 1 of Algorithm 1 does not deviate too much from the true function $g^{*}$} and therefore $^{\widehat{g}}$ is a pessimistic estimation of $g^{*}$}.
+We have the following lemma indicating that this event holds with high probability.
+Lemma 2.9.
+For all $\delta>0$, $\mathcal{E}(\delta)$ holds with probability at least $1-\delta$.
 After obtaining the pessimistic estimation, KL-PCB output the policy $\widehat{\pi}$, which maximizes the estimated objective
@@
 0: regularization $\eta$, reference policy $\pi^{\mathsf{ref}}$, offline dataset $\mathcal{D}$, function class $\mathcal{G}$}
-1:  Compute the least square estimation of reward function ${\bar{g}}\in\mathop{\mathrm{argmin}}_{g\in{\mathcal{G}}}\sum_{(s_{i},a_{i},r_{i})\in\mathcal{D}}\big{(}g(s_{i},a_{i})-r_{i}\big{)}^{2}$
-${\bar{g}}\in\mathop{\mathrm{argmin}}_{g\in{\mathcal{G}}}\sum_{(s_{i},a_{i},r_{i})\in\mathcal{D}}\big{(}g(s_{i},a_{i})-r_{i}\big{)}^{2}$
+1: Least square estimation of reward function ${\bar{g}}\in\mathop{\mathrm{argmin}}_{g\in{\mathcal{G}}}\sum_{(s_{i},a_{i},r_{i})\in\mathcal{D}}\big(g(s_{i},a_{i})-r_{i}\big)^{2}$
 2: Let ${\widehat{g}}\leftarrow{\bar{g}}-\Gamma_{n}$, where $\Gamma_{n}$ is the bonus term in 2.3
-2:  $\widehat{\pi}(a|s)\propto\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta\cdot{\widehat{g}}(s,a)\big{)}$
+[…]
 2.3 Theoretical Results
@@
 As a comparison, KL-PCB achieves the same $\widetilde{O}(\epsilon^{-1})$ sample complexity but only requiring Assumption 2.8, which is weaker than Assumption 2.7.
-We also provide the sample complexity lower bound of KL-regularized contextual bandits in the following theorem.
+We also provide the sample complexity lower bound of KL-regularized contextual bandits in the following theorem, which, together with Theorem 2.10, demonstrates that single-policy concentrability is both necessary and sufficient for near-optimal offline learning evaluated by KL-regularized objectives.
 Theorem 2.11.
@@
 Previously, Zhao et al.
-(2024) provided a sample complexity lower bound of $\Omega(\eta\log\mathcal{N}_{{\mathcal{G}}}(\epsilon)/{\epsilon})$ under KL-regularized objective.
+(2024) provided a sample complexity lower bound of $\Omega(\eta\log\mathcal{N}_{{\mathcal{G}}}(\epsilon)/{\epsilon})$ under KL regularization.
 Foster et al.
@@
 The presence of $\exp(\mathrm{poly}(\eta))$ in the range of $C^{*}$ is inevitable, since we always have $C^{\pi^{*}}\leq\exp(\eta)$ in reverse KL regularized bandits with bounded rewards.
-As discussed in Remark 2.8, when $|{\mathcal{S}}|$ and $|\mathcal{A}|$ are constant, $D^{2}_{\pi^{*}}\lesssim C^{\pi^{*}}$, indicating that KL-PCB is near-optimal under this special case.
-2.4 Proof of Theorem 2.9
-In this section, we finish the proof of Theorem 2.9.
-At a high level, if we consider the regularized objective 1.1 multi-arm bandits, then $P\mapsto{\mathsf{KL}}\left(P\|Q\right)$ is $1$-strongly convex w.r.t.
-${\mathsf{TV}}\left(\cdot\|\cdot\right)$ (Polyanskiy and Wu, 2025, Exercise I.37), and thus $J(\pi)$ is strongly concave.
-Therefore, $J(\pi^{*})-J(\widehat{\pi})$ is possible to be of the order $[{\mathsf{TV}}\left(\pi^{*}\|\widehat{\pi}\right)]^{2}\approx\widetilde{O}(n^{-1})$, pretending that $\pi^{*}$ is the unconstrained maximizer.
-This intuition guides our analysis for contextual bandits.
-We begin the proof with the definition of the event $\mathcal{E}(\delta)$ given $\delta>0$ as
-$\mathcal{E}(\delta)\coloneqq\Big{\{}\sup\nolimits_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}\Big{[}\big{|}{\bar{g}}-{g^{*}}\big{|}-\Gamma_{n}\Big{]}(s,a)\leq 0\Big{\}},$ (2.4)
-where $\Gamma_{n}$ is defined in (2.3).
-Event $\mathcal{E}(\delta)$ holds indicates that the least square estimation ${\bar{g}}$ obtained in Line 1 of Algorithm 1 does not deviate too much from the true function ${g^{*}}$.
-More specifically, we have the following lemma, whose proof are deferred to Appendix B.1.
+Remark 2.13.
+As discussed before, we might have some easy instances with $D^{2}_{\pi^{*}}\leq C^{\pi^{*}}$, where KL-PCB outperforms the lower bound.
+This does not volates Theorem 2.11 since Theorem 2.11 only guarantees that there exist some hard instances that all algorithms require at least $\Omega(\min\{\eta\epsilon^{-1},\epsilon^{-2}\}C^{*}\log\mathcal{N}_{G}(\epsilon))$ samples.
+2.4 Proof Overview of Theorem 2.10
+[…]
 Lemma 2.14.
-For all $\delta>0$, $\mathcal{E}(\delta)$ holds with probability at least $1-\delta$.
-The following covariance-type observation is the first pivot.
+Let $g:{\mathcal{S}}\times\mathcal{A}\to\mathbb{R}$ be any reward function, then there exist some $\gamma\in[0,1]$ such that the sub-optimality gap of $\pi_{g}(\cdot|s)\propto\pi^{\mathsf{ref}}(\cdot|s)\exp\big(\eta g(s,\cdot)\big)$ can be bounded as
+$J(\pi^{*})-J(\pi_{g})\leq\eta\mathbb{E}_{(s,a)\sim\rho\times\pi_{\gamma}}\big[\big({g^{*}}-g\big)^{2}(s,a)\big],$
+where $g_{\gamma}\coloneqq\gamma g+(1-\gamma)g^{*}$ and $\pi_{\gamma}(\cdot|s)\propto\pi^{\mathsf{ref}}(\cdot|s)\exp\big(\eta g_{\gamma}(s,\cdot)\big)$.
+In Zhao et al.
+(2024), because the $g$ in Lemma 2.14 is substituted with only the least-square estimator $\bar{g}$ with no extra structures, the reliance on the “mid-point” policy $\pi_{\gamma}$ can only be controlled all-policy concentrability.
+However, our $g$ is the pessimistic estimator $\widehat{g}$ of $g^{*}$} in Algorithm 1, and thus the presence of $\pi_{\gamma}$ can be eliminated for free: let $G(\gamma)\coloneqq\mathbb{E}_{\rho\times\pi_{\gamma}}\Big[\big({\widehat{g}}-{{g^{*}}}\big)^{2}(s,a)\Big]$ and $\triangle(s,a)\coloneqq\big({\widehat{g}}-{{g^{*}}})(s,a)\leq 0$, then a direct computation (detailed in the proof of Lemma D.3) yields
+$G^{\prime}(\gamma)=\eta\mathbb{E}_{\rho}\Big[\mathbb{E}_{\pi_{\gamma}}\big[\triangle^{3}(s,a)\big]-\mathbb{E}_{\pi_{\gamma}}\big[\triangle^{2}(s,a)\big]\mathbb{E}_{\pi_{\gamma}}\big[\triangle(s,a)\big]\Big]\leq 0.$ (2.5)
+This gives $J(\pi^{*})-J(\widehat{\pi})\leq\eta\mathbb{E}_{\rho\times\pi^{*}}\big[({\widehat{g}}-{{g^{*}}})^{2}(s,a)\big]$, which can be bounded with single-policy concentrability while still achieves the sharp dependency $\epsilon^{-1}$ on $\epsilon$.
+Here, (2.5) holds due to a moment-based machinery in Lemma 2.15.
 Lemma 2.15.
-If a bounded random variable $X\leq 0$ almost surely, then $\mathbb{E}[X^{3}]-\mathbb{E}[X^{2}]\mathbb{E}[X]\leq 0$.
-Remark 2.14.
-While this lemma is elementary, to the best of our knowledge, we are the first to isolate this structure from our non-standard analysis of offline RL, from which the sharp upper bound is derived.
-The intuition behind Lemma 2.13 is very natural: $X$ and $X^{2}$ cannot be positively correlated.
-We further define the following quantities. For all $\gamma\in[0,1]$, we define $g_{\gamma}\coloneqq\gamma{\widehat{g}}+(1-\gamma){g^{*}}$ and further denote
-$\pi_{\gamma}(\cdot|s)\propto\pi^{\mathsf{ref}}(\cdot|s)\exp\big{(}\eta g_{\gamma}(s,\cdot)\big{)},\forall s\in{\mathcal{S}};$
-$G(\gamma)\coloneqq\mathbb{E}_{\rho\times\pi_{\gamma}}\Big{[}\big{(}{\widehat{g}}-{{g^{*}}}\big{)}^{2}(s,a)\Big{]}.$
-The key to our analysis is the monotonicity of the function $G(\gamma)$ in $\gamma$, which is formally stated in the following lemma.
-Lemma 2.15.
-On event $\mathcal{E}$, $0\in\mathop{\mathrm{argmax}}_{\gamma\in[0,1]}G(\gamma)$.
-Proof.
-For simplicity, we use $\triangle(s,a)$ to denote $\big{(}{\widehat{g}}-{{g^{*}}})(s,a)$ in this proof.
-Then we know that $\triangle(s,a)\leq 0$ for all $(s,a)\in{\mathcal{S}}\times\mathcal{A}$ on event $\mathcal{E}$.
-The most direct way to prove is to take derivative of $G$ with respect to $\gamma$, which corresponds to the policy gradient (Sutton et al., 1999) of $\pi_{\gamma}$ and thus implying a favorable structure.
-A direct calculation yields that
-$\ =\mathbb{E}_{\rho\times\pi_{\gamma}}\big{[}\nabla_{\gamma}\log\pi_{\gamma}(a|s)\triangle(s,a)^{2}\big{]}$
-$\ =\eta\mathbb{E}_{\rho}\mathbb{E}_{a\sim\pi_{\gamma}}\big{[}\triangle^{2}(s,a)\big{(}\triangle(s,a)-\mathbb{E}_{a^{\prime}\sim\pi_{\gamma}}[\triangle(s,a^{\prime})]\big{)}\big{]}$
-$\ =\eta\mathbb{E}_{\rho}\Big{[}\mathbb{E}_{\pi_{\gamma}}\big{[}\triangle^{3}(s,a)\big{]}-\mathbb{E}_{\pi_{\gamma}}\big{[}\triangle^{2}(s,a)\big{]}\mathbb{E}_{\pi_{\gamma}}\big{[}\triangle(s,a)\big{]}\Big{]}$
-$\ \leq 0,$
-where $\mathbb{E}_{\rho}$ is the shorthand of $\mathbb{E}_{s\sim\rho}$, $\mathbb{E}_{\pi_{\gamma}}$ is the shorthand of $\mathbb{E}_{a\sim\pi_{\gamma}}$, the first equation is derived from standard policy gradient and the inequality holds conditioned on the event $\mathcal{E}(\delta)$ due to Lemma 2.13.
-∎
-Finally, we need the following lemma to bound the performance difference between two policy w.r.t.
-the KL-regularized objective $J$.
-For any given $g:{\mathcal{S}}\times\mathcal{A}\to\mathbb{R}$, we define $\pi_{g}(\cdot|s)\propto\exp\big{(}\eta\cdot g(s,\cdot)\big{)},\forall s\in{\mathcal{S}}$ to facilitate presentation; then $\pi^{*}=\pi_{{g^{*}}}$ by definition.
-Please refer to Appendix B.3 for the proof of the lemma.
-Lemma 2.16.
-Given $g:{\mathcal{S}}\times\mathcal{A}\to\mathbb{R}$, $\exists\gamma\in[0,1]$ such that for $g_{\gamma}=\gamma g+(1-\gamma)g^{*}$ and $\pi_{\gamma}\coloneqq\pi_{g_{\gamma}}$,
-$J(\pi^{*})-J(\pi_{g})\leq\eta\mathbb{E}_{(s,a)\sim\rho\times\pi_{\gamma}}\big{[}\big{(}{g^{*}}-g\big{)}^{2}(s,a)\big{]}.$
-Now we are ready to prove Theorem 2.9.
-Proof of Theorem 2.9.
-Following the proof of Zhao et al. (2024, Theorem 3.3), we know that there exists $\bar{\gamma}\in[0,1]$ such that
-$J(\pi^{*})-J(\widehat{\pi})\leq\eta G(\bar{\gamma})\leq\eta G(0),$ (2.5)
-where the first inequality holds due to Lemma 2.16 and the second inequality holds due to the event $\mathcal{E}$ and Lemma 2.15.
-The term $G(0)$ can be further bounded with the $D^{2}$-based concentrability as follows
-$G(0)$ $=\eta\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}\Big{[}\big{(}{\widehat{g}}-{{g^{*}}}\big{)}^{2}(s,a)\Big{]}$
-$\leq 4\eta\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}[\Gamma_{n}^{2}(s,a)]$
-$=4\eta\beta^{2}\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}\big{[}D^{2}_{\mathcal{F}}((s,a);\pi^{\mathsf{ref}})\big{]}$
-$=\widetilde{O}(\eta D_{\pi^{*}}^{2}n^{-1}\log_{{\mathcal{G}}}(\epsilon_{c})),$ (2.6)
-where the second inequality holds conditioned on $\mathcal{E}(\delta)$ because of Lemma 2.15, and the last inequality follows from the definition of $\mathcal{E}(\delta)$ together with Line 2.
-By Lemma 2.12, we know that event $\mathcal{E}$ holds with probability at least $1-\delta$, which finishes the proof.
-∎
+If $\mathbb{P}(X\leq 0)=1$ and $\mathbb{E}|X|^{3}<\infty$, then $\mathbb{E}[X^{3}]-\mathbb{E}[X^{2}]\mathbb{E}[X]\leq 0$.
+[…]
 3 $f$-divergence-regularized Contextual Bandits
-As we discussed in Section 2, the fast rate is primarily achieved due to the strong convexity of $\pi\mapsto\mathsf{KL}(\cdot\|\pi^{\mathsf{ref}})$.
-However, KL is just an instance of $f$-divergence with $f(x)=x\log x$, which is not strongly convex, viz., only locally strongly convex.
-Motivated by this observation, we examine the use of $f$-divergence regularization with strongly convex $f$, which, in principle, may introduce a more favorable curvature in the performance metric of offline learning.
+As discussed in Section 2, the fast rate implied by Theorems 2.10 and 2.11 is primarily achieved due to the strong convexity of $\pi\mapsto\mathsf{KL}(\pi\|\pi^{\mathsf{ref}})$.
+However, KL is just an instance of $f$-divergence with $f(x)=x\log x$, which is only locally strongly convex but not strongly convex.
+Motivated by this observation, we further examine $f$-divergence regularization with strongly convex $f$, which may introduce a more favorable curvature in the performance metric of offline learning in principle.
 3.1 Problem Setup
-We study the same contextual bandit problems which follows the definition in Section 2.1. In this section, we consider the following $f$-divergence regularized objective
-$J(\pi)\coloneqq\mathbb{E}_{(s,a)\sim\rho\times\pi}[r(s,a)]-\eta^{-1}\mathbb{E}_{s\sim\rho}\big{[}D_{f}(\pi||\pi^{\mathsf{ref}})\big{]},$ (3.1)
-where $\eta$ is the inverse temperature regularization intensity and $D_{f}(\pi||\pi^{\mathsf{ref}})$ is given by
-$D_{f}(\pi||\pi^{\mathsf{ref}})=\mathbb{E}_{a\sim\pi^{\mathsf{ref}}}\bigg{[}f\bigg{(}\frac{\pi(a|s)}{\pi^{\mathsf{ref}}(a|s)}\bigg{)}\bigg{]}.$
-We consider those functions $f:(0,+\infty)\to\mathbb{R}$ with nice positive curvature as follows.
+We study a contextual bandit setting similar to that in Section 2.1. In this section, we consider the following $f$-divergence regularized objective
+$J_{\eta,D_{f}}(\pi)\coloneqq\mathbb{E}_{(s,a)\sim\rho\times\pi}[r(s,a)]-\eta^{-1}\mathbb{E}_{s\sim\rho}\big[D_{f}\big(\pi(\cdot|s)\|\pi^{\mathsf{ref}}(\cdot|s)\big)\big],$ (3.1)
+where $\eta$ is the regularization intensity and $D_{f}(p\|q)\coloneqq\mathbb{E}_{a\sim q}\Big[f\big({p(a)}/{q(a)}\big)\Big]$ is the $f$-divergence.
+Let the optimal policy be $\pi^{*}_{\eta,D_{f}}\coloneqq\mathop{\mathrm{argmax}}_{\pi\in\Delta(\mathcal{A}|{\mathcal{S}})}J_{\eta,D_{f}}(\pi)$ and we re-define the learning objective as searching for a policy $\pi$ with $\mathrm{SubOpt}_{f\mathrm{div}}(\pi)\coloneqq J(\pi^{*})-J(\pi)\leq\epsilon$.
+[^4] We again suppress $J_{\eta,D_{f}}(\cdot)$ into $J(\cdot)$ and $\pi^{*}_{\eta,D_{f}}$ into $\pi^{*}$ when there is no confusion.
+We consider those functions $f:(0,+\infty)\to\mathbb{R}$ with a nice positive curvature condition in Assumption 3.1.
 Assumption 3.1.
@@
 Many elementary functions like quadratic polynomials naturally satisfy Assumption 3.1.
-For instance, the 1-strongly convex $f(x)=(x-1)^{2}/2$ yields $D_{f}(P\|Q)=\chi^{2}(P\|Q)$, which is the popular $\chi^{2}$-divergence recently considered in RL literature (see e.g., Zhan et al.
+For instance, the 1-strongly convex $f(x)=(x-1)^{2}/2$ induces $D_{f}(P\|Q)=\chi^{2}(P\|Q)$, which is the $\chi^{2}$-divergence recently considered in RL literature (see e.g., Zhan et al.
 (2022); Huang et al.
-(2024); Amortila et al.
-(2024)) and exhibits a promising theoretical potential on relaxing the data coverage requirement for efficient offline policy learning, at least against the unregularized objective.
+(2025b); Amortila et al.
+(2024)).
+This regularization exhibits a promising theoretical potential for relaxing the data coverage requirement for efficient offline policy learning (Huang et al., 2025b) and to be effective in preventing reward hacking (Laidlaw et al., 2025) against unregularized objectives.
+These favorable benefits are primary due to the observation that strongly convex $f$’s impose a stronger penalization on actions out of the coverage of $\pi^{\mathsf{ref}}$.
 3.2 Algorithm and Main Results
@@
 0: regularization $\eta$, reference policy $\pi^{\mathsf{ref}}$, function class $\mathcal{G}$}, offline dataset $\mathcal{D}$
-1:  Compute the least square estimation of reward function
-${\bar{g}}\in\mathop{\mathrm{argmin}}_{g\in{\mathcal{G}}}\sum_{(s_{i},a_{i},r_{i})\in\mathcal{D}}\big{(}g(s_{i},a_{i})-r_{i}\big{)}^{2}$
+1: Least square estimation ${\bar{g}}\in\mathop{\mathrm{argmin}}_{g\in{\mathcal{G}}}\sum_{(s_{i},a_{i},r_{i})\in\mathcal{D}}\big(g(s_{i},a_{i})-r_{i}\big)^{2}$
 2: Compute the optimal policy under the least-square reward estimator $\bar{g}$} for $s\in{\mathcal{S}}$ as
 $\widehat{\pi}(\cdot|s)\leftarrow\mathop{\mathrm{argmax}}_{\pi(\cdot|s)\in\Delta(\mathcal{A})}\langle{\pi(\cdot|s)},{{\bar{g}}(s,\cdot)}\rangle+\eta^{-1}D_{f}\big(\pi(\cdot|s)\|\pi^{\mathsf{ref}}(\cdot|s)\big)$
-2:  $\widehat{\pi}$
+[…]
 In this subsection, we present an offline learning algorithm for $f$-divergence regularized bandit, $f$-CB, in Algorithm 2.
@@
 Intuitively, this is because the $f$-divergence regularization in this case is much stronger, so that both $\pi^{*}$ and $\widehat{\pi}$ are close enough to $\pi^{\mathsf{ref}}$.
-The following hardness result justify the near-optimality of Theorem 3.2 for $f$-divergence-regularized bandits.
+The following hardness result justify the near-optimality of Theorem 3.2 for $f$-divergence-regularized contextual bandits.
 Theorem 3.4.
-For any $\epsilon\in(0,1),\alpha>0,\eta>0$, and algorithm Alg, there is an $\alpha$-strongly-convex function $f$ and an $f$-divergence-regularized contextual bandit instance such that Alg requires at least $\Omega\big{(}\alpha^{-1}\eta\epsilon^{-1}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon)\big{)}$ samples to return an $\epsilon$-optimal policy.
-3.3 Insights Behind the Proof of Theorem 3.2
-We provide an overview of key analysis techniques for proving Theorem 3.2. For the simplicity of presentation, we consider multi-arm bandits here and omit the subscript for context $s$.
+For any $\epsilon\in(0,1),\alpha>0,\eta>0,S>32/3\cdot\log 2$, sufficiently small $\epsilon$, and algorithm Alg, there is an $\alpha$-strongly-convex function $f$ and an $f$-divergence-regularized contextual bandit instance such that Alg requires at least $\Omega\big(\alpha^{-1}\eta\epsilon^{-1}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon)\big)$ samples to return an $\epsilon$-optimal policy.
+3.3 Proof Overview of Theorem 3.2
+We provide an overview of key analysis techniques for proving Theorem 3.2.
+Unlike KL-regularization, the $\pi^{*}$ under $f$-divergence might not have a closed form.
+This means that the proof of Lemma 2.14, which relies on the closed form of $\pi^{*}$, cannot be directly adopted.
+Therefore, we address this from a dual-Bregman perspective, inspired by Abernethy et al.
+(2015).
+For the simplicity of presentation, we consider multi-armed bandits here and omit the subscript for context $s$.
 We consider the function $H(\pi)=\eta^{-1}D_{f}(\pi\|\pi^{\mathsf{ref}})$, which is the regularizer in the objective.
-Then its convex conjugate is given by $H^{*}(r)=\sup_{\pi\in\Delta^{d}}\{\left\langle\pi,r\right\rangle-H_{s}(\pi)\}$, which is exactly the expected reward obtained by the optimal policy given reward function $r$.
-One observation is that when $f$ is strongly convex, the induced $f$-divergence, and therefore the function $H$ is also strongly convex.
+Then its convex conjugate is given by $H^{*}(r)=\sup_{\pi\in\Delta^{d}}\{\left\langle\pi,r\right\rangle-H(\pi)\}$, which is exactly the expected reward obtained by the optimal policy given reward function $r$.
+One observation is that when $f$ is strongly convex, the induced $f$-divergence, and therefore the function $H$ are also strongly convex.
 Therefore, let $\pi_{r}=\mathop{\mathrm{argmax}}_{\pi}\{\left\langle\pi,r\right\rangle-H_{s}(\pi)\}$ given some reward function $r$, the strong convexity of $H(\pi)$ gives that $\nabla H^{*}(r)=\pi_{r}$.
-This leads to the following regret decomposition, which is one of our key observations.
-$J(\pi^{*})-J(\widehat{\pi})$ $=\mathbb{E}_{a\sim\pi^{*}}[{g^{*}}(a)]-\mathbb{E}_{a\sim\widehat{\pi}}[{g^{*}}(a)]$
-$\quad-\eta^{-1}\big{[}D_{f}(\pi^{*}\|\pi^{\mathsf{ref}})-D_{f}(\widehat{\pi}\|\pi^{\mathsf{ref}})\big{]}$
+This leads to the following regret decomposition, which is one of our key observations:
+[…]
 $=H^{*}({g^{*}})-H^{*}({\bar{g}})-\left\langle\widehat{\pi},{g^{*}}-{\bar{g}}\right\rangle$
 $=H^{*}({g^{*}})-H^{*}({\bar{g}})-\left\langle\nabla H^{*}({\bar{g}}),{g^{*}}-{\bar{g}}\right\rangle,$
-which is the Bregman divergence of the dual function $H^{*}$ and therefore can be bounded by $({g^{*}}-{\bar{g}})^{\top}H^{*}(\widetilde{g})({g^{*}}-{\bar{g}})$ for some $\widetilde{g}$.
-By Proposition 3.2 in Penot (1994), we can bound $H^{*}(\widetilde{g})$ as follows
-$H^{*}(\widetilde{g})$ $\preceq\big{(}\nabla^{2}H(\pi_{\widetilde{g}})\big{)}^{-1}$
-$\preceq\alpha^{-1}\eta\mathrm{diag}\big{(}\pi^{\mathsf{ref}}(a_{1}),\cdots,\pi^{\mathsf{ref}}(a_{|\mathcal{A}|})\big{)},$
-which enables us to bound $({g^{*}}-{\bar{g}})^{\top}H^{*}(\widetilde{g})({g^{*}}-{\bar{g}})$ by $\alpha^{-1}\eta\mathbb{E}_{\pi^{\mathsf{ref}}}[({g^{*}}-{\widehat{g}})^{2}]$.
-Since $\mathbb{E}_{\pi^{\mathsf{ref}}}[({g^{*}}-{\widehat{g}})^{2}]$ is not related to the optimal $\pi^{*}$, our resulting upper bound does not depend on any notion of concentrability.
-4 Generalization to Contextual Dueling Bandits
-In this section, we extend our algorithm to the problems of regularized contextual dueling bandits, where the learner receives preference comparison instead of absolute signals.
-Our data model follows Zhu et al.
-(2023); Zhan et al.
-(2023) and the notion of suboptimality follows Xiong et al.
-(2024); Zhao et al.
-(2024).
-4.1 Problem Setup
-We still consider contextual bandits $({\mathcal{S}},\mathcal{A},r,\pi^{\mathsf{ref}})$ where ${\mathcal{S}}$ is the state space, $\mathcal{A}$ is the action space and $r:{\mathcal{S}}\times\mathcal{A}\to[0,1]$ is the reward function.
-[^1] We overload some notations in Section 2 by their dueling counterparts for notational simplicity.
-But only relative preference feedback is available, viz., we have an $\mathrm{i.i.d.}$ offline dataset $\mathcal{D}=\{(s_{i},a^{1}_{i},a^{2}_{i},y_{i})\}_{i=1}^{n}$, where $s_{i}\in{\mathcal{S}}$ is generated from distribution $\rho$ and $a^{1}_{i},a^{2}_{i}\sim\pi^{\mathsf{ref}}$.
-The binary preference label $y_{i}=1$ indicates $a_{i}^{1}$ is preferred over $a_{i}^{2}$ (denoted by $a^{1}\succ a^{2}$) and 00 for $a^{2}\succ a^{1}$ given context $s$.
-In this work we consider the Bradley-Terry Model, where $\mathbb{P}[y=1|s,a^{1},a^{2}]=\sigma(r(s_{i},a^{1}_{i})-r(s_{i},a^{2}_{i}))$, where $\sigma(x)=(1+e^{-x})^{-1}$ is the link function.
-The objective here identical to (2.1) for KL-regularization and (3.1) for $f$-divergence regularization.
-Our goal is still to find an $\epsilon$-optimal policy.
-To control the complexity of the function class ${\mathcal{G}}$, we assume that Assumption 2.1 still holds here.
-Concentrability.
-Analogous to Section 2, we need our estimation from offline dataset generalizable to the state-action pairs visited by our obtained policy.
-While density-ratio-based concentrability can be directly adapted to dueling bandit, we need a slightly different notion of $D^{2}$-divergence.
-This is because in dueling bandit, we cannot observe the absolute reward and best estimation $g$ we can achieve is that for any state $s$ and actions $a^{1},a^{2}$, our estimated $g(s,a^{1})-g(s,a^{2})\approx r(s,a^{1})-r(s,a^{2})$.
-This implies that there exists some mapping $b:{\mathcal{S}}\to[-1,1]$ such that $g(s,a)-b(s)\approx r(s,a)$ on the offline data, which leads to the following definition.
-Definition 4.1.
-Given a class of functions ${\mathcal{G}}\subset({\mathcal{S}}\times\mathcal{A}\to\mathbb{R})$ and some policy $\pi$, let $\mathcal{B}=({\mathcal{S}}\to[-1,1])$ be the function class, define the $D^{2}$-divergence $D^{2}_{{\mathcal{G}}}((s,a);\pi)$ as
-$\sup_{g,h\in{\mathcal{G}}}\inf_{b\in\mathcal{B}}\frac{\big{(}g(s,a)-h(s,a)-b(s)\big{)}^{2}}{\mathbb{E}_{s\sim\rho}\operatorname{{\rm Var}}_{a^{\prime}\sim\pi(\cdot|s^{\prime})}[g(s^{\prime},a^{\prime})-h(s^{\prime},a^{\prime})]}.$
-A similar definition has been introduced in Zhao et al.
-(2024, Definition 2.6), which underpins the following two assumptions that characterize the coverage ability of $\pi^{\mathsf{ref}}$ similarly as in Section 2.
-Given a reference policy $\pi^{\mathsf{ref}}$, we define two coverage notions for contextual dueling bandits.
-Assumption 4.2 (All-policy concentrability).
-$D^{2}\coloneqq\sup_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}D^{2}_{{\mathcal{G}}}((s,a);\pi^{\mathsf{ref}})<\infty$.
-Assumption 4.3 (Single-policy concentrability).
-$D_{\pi^{*}}^{2}\coloneqq\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}[D^{2}_{{\mathcal{G}}}((s,a);\pi^{\mathsf{ref}})]<\infty$.
-Similar single-policy concentrability assumptions have appeared in previous work in offline contextual dueling bandits (Huang et al., 2024; Song et al., 2024) and similar notions has also appeared in the analysis of model-based RL (Uehara and Sun, 2021; Wang et al., 2024).
-Although Assumption 4.3 is strictly weaker than Assumption 4.2, $D^{2}_{\pi^{*}}\leq|{\mathcal{S}}||\mathcal{A}|C^{\pi^{*}}$ still holds and, in general, $C^{\pi^{*}}$ and $D^{2}_{\pi^{*}}$ cannot be bounded by each other up to some constant factors.
-4.2 Main Results
-In this subsection, we provide our main results for contextual dueling bandits with regularization.
-We refer the reader to Appendix D for a more detailed discussion.
-We first present the following results for KL-regularized contextual dueling bandit.
-Theorem 4.4 (Informal).
-There exists an algorithm, with high probability, generates $\epsilon$-optimal policy with $n=\widetilde{O}(\eta D_{\pi^{*}}^{2}\epsilon^{-1}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon))$ samples.
-Moreover, for any sufficiently small $\epsilon>0$, and $\eta>0$, and any algorithm Alg, there is a KL-regularized contextual dueling bandit such that Alg requires at least $\Omega\big{(}\min\{{\eta C^{\pi^{*}}\epsilon^{-1}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon)},{(C^{\pi^{*}})^{2}\epsilon^{-2}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon)}\}\big{)}$ samples to return an $\epsilon$-optimal policy.
-The following theorem shows a similar guarantee for $f$-divergence regularized contextual dueling bandits where $f$ is $\alpha$-strongly convex.
-Theorem 4.5 (Informal).
-Under Assumption 3.1, there exists an algorithm generating $\epsilon$-optimal policy with $n=\widetilde{O}(\alpha^{-1}\eta\epsilon^{-1}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon))$ with high probability.
-Moreover, for any sufficiently small $\epsilon>0$, $\eta,\alpha>0$, and any offline RL algorithm Alg, there is an $\alpha$-strongly convex function $f$ such that there exists an $f$-divergence regularized contextual dueling bandit such that Alg requires at least $\Omega\big{(}{\alpha^{-1}\eta\epsilon^{-1}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon)}\big{)}$ samples to return an $\epsilon$-optimal policy.
-Theorems 4.4 and 4.5 are extension of Theorems 2.9 and 3.2 to offline contextual dueling bandits.
-These theorems indicate that, for both KL-regularization and $f$-divergence regularization, our techniques can be adapted to achieve algorithms and achieve similar sample complexities.
-5 Conclusion and Limitation
+[…]
+[…]
+which enables us to bound $({g^{*}}-{\bar{g}})^{\top}\nabla^{2}H^{*}(\widetilde{g})({g^{*}}-{\bar{g}})$ by $\alpha^{-1}\eta\mathbb{E}_{\pi^{\mathsf{ref}}}[({g^{*}}-{\widehat{g}})^{2}]$.
+Since $\mathbb{E}_{\pi^{\mathsf{ref}}}[({g^{*}}-{\widehat{g}})^{2}]$ is not related to $\pi^{*}$, the upper bound is independent of any notion of concentrability.
+(a) The empirical relation between $\log_{2}n$ and $\log_{2}\mathrm{SubOpt}$.
+The fitted rate means the slope of $\log_{2}n\sim\log_{2}\mathrm{SubOpt}$ estimated via linear regression.
+Here $n$ is the sample size.
+Every point is the average over 100 independent trials.
+4 Experiments
+Simulation on multi-armed bandits.
+We first empirically check the correctness of our matching bounds for KL and $f$-divergence on the simplest testbed: two-armed bandits, i.e., $\mathcal{A}=\{0,1\}$.
+We use one hard instance constructed in the proof of Theorem 2.11 (Section D.5) for the simulation under KL and one hard instance constructed in the proof of Theorem 3.4 (Section E.2) for the simulation under $f$-divergence with $f(x)=\alpha(x-1)^{2}/2$.
+Recall that the dependency on $\epsilon$ in all sample complexity bounds above is $\widetilde{\Theta}(\epsilon^{-1})$, and thus both $\mathrm{SubOpt}_{\mathrm{RKL}}$ and $\mathrm{SubOpt}_{f\mathrm{div}}$ should be roughly proportional to $n^{-1}$ as a function of the sample size $n$, which can be verified from the linear regression between $\log_{2}n$ and $\log_{2}\mathrm{SubOpt}$; i.e., the estimated slope should be approximately 1-1.
+Therefore, the two fitted rates in LABEL:subfig:kl indicates that KL-PCB indeed achieves the near-optimal statistical rate $n^{-1}$ under different $\pi^{\mathsf{ref}}$’s and the counterparts in the LHS of LABEL:subfig:f indicates the near-optimality of $f$-CB empirically.
+The contrast between LABEL:subfig:kl and the LHS of LABEL:subfig:f also corroborates that the sample complexity against the KL-regularized objective positively depend on the concentrability, while that against the $\chi^{2}$-divergence-regularized objective does not vary with the coverage condition of $\pi^{\mathsf{ref}}$.
+Moreover, on top of the hard instance for $f$-divergence, we further set $\alpha=2^{15}/n$ to numerically examine the scaling of $\mathrm{SubOpt}_{f\mathrm{div}}$ w.r.t.
+the strong convexity modulus $\alpha$.
+As shown on the RHS of LABEL:subfig:f, $\mathrm{SubOpt}_{f\mathrm{div}}$ remains stable as $n$ goes up given $n\alpha\equiv 2^{15}$; therefore, LABEL:subfig:f also empirically verified that $\mathrm{SubOpt}_{f\mathrm{div}}$ is inversely proportional to $\alpha$.
+(b) The empirical relation between $\log_{2}n$ and $\log_{2}\mathrm{SubOpt}$ for linear bandits. In the legend, we denote $C^{\pi^{*}}$ (resp. $D^{2}_{\pi^{*}}$) by C (resp. Dˆ2).
+Simulation on linear bandits.
+We then simulate a linear bandit as follows.
+The constructions of the feature map $\bm{\phi}$, ground-truth parameter $\bm{\theta}^{*}$ and the induced reward are detailed in Section C.1.
+The behavior policy is constructed as $\pi^{\mathsf{ref}}=\beta\mathsf{Unif}(\mathcal{A})+(1-\beta)\mathsf{Unif}(\mathcal{A}_{k})$, where $\mathcal{A}_{k}\subset\mathcal{A}$ be the subset such that $\mathcal{A}_{k}$ consists of the $k$ arms with the lowest expected reward.
+We consider three different behavior policies, $(\beta,k)\in\{(1,\cdot),(0.1,4),(0.05,20)\}$, which induces various $C^{\pi^{*}}$ and $D^{2}_{\pi^{*}}$ so as to demonstrate the influence of coverage under different regularization.
+The results are compiled in Figure 2(b).
+Specifically, for the KL-regularized cases depicted in Figure LABEL:subfig:kl-linear, we see that as the coverage coefficients $C^{\pi^{*}}$ and $D^{2}_{\pi^{*}}$ vary, there is a consistent sub-optimality gap margin between these instances.
+In contrast, LABEL:subfig:chi2-linearshows that the sub-optimality gaps under different instances (with distinct coverage coefficient) are very close for sufficiently large sample sizes.
+These results corroborate our theoretical finding that the sample complexity w.r.t.
+KL-regularized objectives is concentrability-dependent but that w.r.t $f$-divergence ones is not (for strongly convex $f$).
+(c) The empirical relation between $\log_{2}n$ and $\log_{2}\mathrm{SubOpt}$ on MNIST dataset.
+Real-world experiments.
+We further verify our theory on a vision dataset, MNIST (LeCun, 1998).
+The construction of the feature map is detailed in Section C.2.
+We consider two reference polices, a uniform policy $\mathsf{Unif}(\mathcal{A})$ and a spiky policy $0.5\mathsf{Unif}(\mathcal{A})+0.5\mathsf{Dirac}(\{0\})$ to obtain instances with different concentrability coefficients.
+Figure 2(c) exhibits the $\mathrm{SubOpt}$ curves, which show that under KL-regularization, when sample size is not large enough, there exists a considerable gap between instances with different behavior policy, but the gap is vanishing as the sample size increases.
+On the other hand, as for $\chi^{2}$-divergence regularization, such a gap vanishes quickly when the sample size becomes moderately large and the sub-optimal gap remains similar for larger sample sizes.
+These results are consistent with the simulation in Section 4 and our theoretical findings.
+5 Conclusion and Future Work
 In this work, we take the first step towards fully understanding the statistical efficiency with respect to $f$-divergence-regularized objectives of offline policy learning by sharp analyses for two empirically relevant subclasses.
@@
 (2) If strong convexity is further imposed on $f$, our fast $\widetilde{\Theta}(\epsilon^{-1})$ sample complexity is provably free of any coverage dependency.
-Unlike those for KL, the upper bound arguments for strongly convex $f$ do not rely on the specific closed-form of the regularized objective minimizer.
-All techniques in this work can be generalized beyond vanilla absolute reward feedback, as certified by contextual dueling bandits.
-However, there is still a gap between the upper bound and lower bound under reverse-KL regularization.
-Also, for general $f$-divergence other than reverse-KL, our results are limited to twice continuously differentiable and strongly convex $f$.
-The fully closing of the gap under reverse-KL regularization and the investigation of general $f$-divergence are left as future directions.
-Appendix A Review of Existing Results
+Unlike those for KL, the upper bound arguments for strongly convex $f$ do not rely on specific closed-form solutions of the regularized objective maximizer.
+All techniques in this work can be generalized beyond vanilla absolute reward feedback, as certified by CDBs, which is detailed in Appendix F under a slightly different notion of $D^{2}$ tailored for pairwise comparison feedback.
+However, for reverse-KL regularization, the $D^{2}_{\pi^{*}}$ in the upper bound and the $C^{\pi^{*}}$ in the lower bound still does not perfectly match.
+Also, for general $f$-divergence other than reverse-KL, our analyses require $f$ to be twice-continuously differentiable and strongly convex.
+Fully closing the gap under reverse-KL regularization and extending the analysis to general $f$-divergences are interesting directions for future work.
+The Use of Large Language Models (LLMs)
+We use LLMs as a tool to refine our writing and correct grammatical errors.
+Acknowledgment
+We thank the anonymous reviewers and area chair for their helpful comments.
+QZ, KJ, HZ and QG are supported in part by the National Science Foundation DMS-2323113 and IIS-2403400.
+HZ is also supported in part by Amazon PhD Fellowship.
+The views and conclusions contained in this paper are those of the authors and should not be interpreted as representing any funding agencies.
+Appendix A Additional Review of Existing Results
 Additional notations.
 Besides the notation introduced in Section 1, we will use the following notations in Appendix.
-We denote $[N]\coloneqq\{1,\cdots,N\}$ for any positive integer $N$.
+We denote $N]\coloneqq\{1,\cdots,N\}$ for any positive integer $N$.
 Boldfaced lower case (resp.
@@
 We use $\mathsf{Bern}(p)$ to denote Bernoulli distribution with expectation $p$ and $\mathsf{Unif}(\mathcal{X})$ for the uniform distribution on finite set $\mathcal{X}$.
-For $x\in\mathbb{R}^{\mathcal{A}}$, we denote $\|x\|_{1}=\sum_{a\in\mathcal{A}}|x_{a}|$.
+For $x\in\mathbb{R}^{|\mathcal{A}|}$, we denote $\|x\|_{1}=\sum_{a\in\mathcal{A}}|x_{a}|$.
 We also denote $x_{n}=\Omega(y_{n})$ by $x_{n}\gtrsim y_{n}$ in Appendix.
@@
 Convergence under regularized objectives.
-Several recent studies (Xie et al., 2024; Xiong et al., 2024; Zhao et al., 2024, 2025; Foster et al., 2025) switched the focus to analyzing the suboptimality guarantee with respect to the regularized objective (1.1).
+Several recent studies (Xie et al., 2024; Xiong et al., 2024; Zhao et al., 2024; 2025; Foster et al., 2025) switched the focus to analyzing the sub-optimality guarantee with respect to the regularized objective (1.1).
 In particular, Xie et al.
@@
 Foster et al.
-(2025) considered an interesting hybrid setting in which the $n$ state-action pairs are still from the offline dataset but $\Omega(n)$ online reward queries and policy switches are allowed; in contrast, in our setting, all reward signals are obtained in a purely offline fashion.
-Here, we briefly discuss the direct adaptation of previous sample complexity results (with respect to KL-regularized objectives) to our setting and demonstrate the reason why theirs cannot imply an $\widetilde{O}(\epsilon^{-1})$ sample complexity without all-policy concentrability.
-Specifically, Xiong et al.
-(2024) obtained a performance gap upper bound under linear function approximation
+(2025) considered an interesting hybrid setting in which the $n$ state-action pairs are still from the offline dataset but $\Omega(n)$ online reward queries and policy switches are allowed; in contrast, in our setting, all reward signals are obtained in a purely offline manner.
+Previous analyses and results in detail.
+Here, we briefly discuss the direct adaptation of previous sample complexity analysis and results (with respect to KL-regularized objectives) to our setting and demonstrate the reason why theirs cannot imply an $\widetilde{O}(\epsilon^{-1})$ sample complexity without all-policy concentrability.
+In previous analysis of pessimism for unregularized objectives (Jin et al., 2021; Xiong et al., 2022), the sub-optimality gap is decomposed via the performance difference lemma as follows
+[…]
+$\leq\mathbb{E}_{a\sim\pi^{*}}[{g^{*}}(a)]-\mathbb{E}_{a\sim\widehat{\pi}}[{\widehat{g}}(a)]-\eta^{-1}\mathsf{KL}(\pi^{*}\|\pi^{\mathsf{ref}})+\eta^{-1}\mathsf{KL}(\widehat{\pi}\|\pi^{\mathsf{ref}})$
+$\leq\mathbb{E}_{a\sim\pi^{*}}[{g^{*}}(a)]-\mathbb{E}_{a\sim\pi^{*}}[{\widehat{g}}(a)]-\eta^{-1}\mathsf{KL}(\pi^{*}\|\pi^{\mathsf{ref}})+\eta^{-1}\mathsf{KL}(\pi^{*}\|\pi^{\mathsf{ref}})$
+$=\mathbb{E}_{a\sim\pi^{*}}[{g^{*}}(a)-{\widehat{g}}(a)],$
+where the first inequality holds due to pessimism and last inequality holds due to $\widehat{\pi}$ is optimal for $^{\widehat{g}}$.
+Notably, the KL-regularization term is canceled out in the analysis, leading to a loose sample complexity $\widetilde{O}(\epsilon^{-2})$ since the curvature of KL-divergence is not exploited.
+Specifically, under linear function approximation, this performance gap, obtained by Xiong et al.
+(2024) becomes
 $J(\pi^{*})-J(\pi)\leq\big\|\mathbb{E}_{\rho\times\pi^{*}}[\bm{\phi}(s,a)]-\bm{\nu}\big\|_{\bm{\Sigma}_{\text{off}}^{-1}}\eqqcolon\text{RHS},$
@@
 (2024) proposed a two-stage learning algorithm and obtained an $\widetilde{O}(\epsilon^{-1})$ sample complexity for online KL-regularized bandits.
-The algorithm can be adopted to offline learning by removing the second stage [^2] This can be done by setting the $n$ in their paper to 00.
+The algorithm can be adopted to offline learning by removing the second stage [^5] This can be done by setting the $n$ in their paper to 0.
 and treat the samples from first stage as the offline dataset.
 An analogous analysis gives a sample complexity of $\widetilde{O}(D^{2}\epsilon^{-1})$, where $D^{2}$ is the all-policy concentrability.
-Appendix B Missing Proofs from Section 2
-B.1 Proof of Lemma 2.12
+Appendix B Additional Discussion of Relation between Coverage Measures
+In this section, we provide more illustrations on the relation between two coverage measures, $D^{2}_{\pi^{*}}$ and $C^{\pi^{*}}$.
+In particular, we provide two cases under linear function approximation, on one of which $D^{2}_{\pi^{*}}=\Theta(dC^{\pi^{*}})$ and on the other we have $D^{2}_{\pi^{*}}\ll C^{\pi^{*}}$, where $d$ is the dimension of the function class.
+We summarized them as two propositions.
+Proposition B.1.
+There exist a KL-regularized linear bandit instance, such that $D^{2}_{\pi^{*}}=\Theta(dC^{\pi^{*}})$.
+Proof.
+We construct the instance as follows.
+Let $d=2A+1$ be some odd number and consider an $2A+1$-armed bandit, such that the feature vector of the $i$-th arm, $\bm{\phi}(a_{i})=\mathbf{e}_{i}\in\mathbb{R}^{d}$, which has $1$ on its $i$-th entry and 0 on all other entries.
+The reference policy $\pi^{\mathsf{ref}}(a_{i})=(2AC)^{-1}$ for $i\in[2A]$ and $\pi^{\mathsf{ref}}(a_{2A+1})=(C-1)/C$, where $2C-1=e^{\eta}$.
+The ground truth reward function $\bm{\theta}^{*}=\sum_{i\leq A}\mathbf{e}_{i}$ and the function class is given by all $\|\bm{\theta}\|_{\infty}\leq 1$.
+By construction, we know that $\pi^{*}(a_{i})\geq\pi^{\mathsf{ref}}(a_{i})$ if and only if $i\in[A]$ and its closed form is given by
+$\pi^{*}(a_{i})=\frac{1}{A}\frac{e^{\eta}}{e^{\eta}+2C-1}=\frac{1}{2A},$
+which gives $C^{\pi^{*}}=C$. Now we compute the $D^{2}_{\pi^{*}}$ of this instance. For all $i\in[A]$, we know that
+$D^{2}(a_{i})=\sup_{\|\bm{\theta}\|_{\infty}\leq 2}\frac{\left\langle\bm{\theta},\mathbf{e}_{i}\right\rangle^{2}}{\mathbb{E}_{\pi^{\mathsf{ref}}}\left\langle\bm{\theta},\mathbf{e}_{j}\right\rangle^{2}}=2CA=\Theta(Cd),$
+where the second equation holds with $\bm{\theta}=\mathbf{e}_{i}$. Taking expectation over $\pi^{*}$, we have
+$D^{2}_{\pi^{*}}\geq\sum_{i\in[A]}D^{2}(a_{i})=\Theta(C^{\pi^{*}}d),$
+which concludes the proof. ∎
+The following proposition provides another instance on which $D^{2}_{\pi^{*}}\ll C^{\pi^{*}}$.
+Proposition B.2.
+For any $C\geq 2$, there exists a KL-regularized linear bandit instance, such that $C^{\pi^{*}}=C/2$ and $D^{2}_{\pi^{*}}=\Theta(1)$.
+Proof.
+We consider the function class of $\bm{\theta}\in\mathbb{R}^{2}$ and $\|\bm{\theta}\|\leq\sqrt{2}$.
+The instance consists of three arms, where $\bm{\phi}(a_{1})=(1,0)$, $\bm{\phi}(a_{2})=(0,1)$, and $\bm{\phi}(a_{3})=(1,1)$.
+The ground truth parameter $\bm{\theta}^{*}=(1,1)$.
+The reference policy is given by $\pi^{\mathsf{ref}}(a_{1})=\pi^{\mathsf{ref}}(a_{2})=1/2-1/2C$ and $\pi^{\mathsf{ref}}(a_{1})=1/C$, where $C-1=e^{\eta}$.
+A direct computation yields that
+$\pi^{*}(a_{3})=\frac{e^{\eta}}{e^{\eta}+C-1},\quad\Rightarrow\quad C^{\pi^{*}}=C\frac{e^{\eta}}{e^{\eta}+C-1}=\frac{C}{2}.$
+On the other hand, we know that for $i=1,2$, we have $D^{2}(a_{i})\leq\pi^{\mathsf{ref}}(a_{i})^{-1}\leq 4$.
+As for $a_{3}$, since we have $\left\langle\bm{\theta},\phi(a_{3})\right\rangle^{2}=\left\langle\bm{\theta},\phi(a_{1})+\phi(a_{2})\right\rangle^{2}\leq 2\left\langle\bm{\theta},\phi(a_{1})\right\rangle^{2}+2\left\langle\bm{\theta},\phi(a_{2})\right\rangle$, which gives that $D^{2}(a_{3})\leq 2D^{2}(a_{1})+2D^{2}(a_{2})\leq 16$.
+Therefore, taking expectation over $\pi^{*}$, we know that $D^{2}_{\pi^{*}}\leq 12$ which is a constant.
+∎
+Appendix C Experimental Details
+C.1 Linear Bandits
+The linear bandit instance used for Figure 2(b) has $d=20$ and $|\mathcal{A}|=100$.
+For each arm $a\in\mathcal{A}$, we randomly generate its feature vector $\bm{\phi}(a)\in\mathbb{R}^{d}$ such that $\|\bm{\phi}(a)\|=1$.
+We then randomly sample the model parameter $\bm{\theta}^{*}\in\mathbb{R}^{d}$ such that $\|\bm{\theta}^{*}\|=1$ and the expected reward is obtained via $r(a)=\left\langle\bm{\theta}^{*},\bm{\phi}(a)\right\rangle$.
+C.2 Real-Wold Experiments
+MNIST consists of 60000 figures, each of which is of $28\times 28$ pixels and consists of a handwritten digit in $\{0,\cdots,9\}$.
+Here, we consider each image as a context and $\mathcal{A}=\{0,\cdots,9\}$ for each context.
+To obtain the feature $\bm{\phi}(s,a)$, we first use the hidden representation of a classifier to embed each image as a vector in $\mathbb{R}^{10}$.
+We then follow the approach in Zhou et al.
+(2020) to obtain the feature of each context-action pair by having $\bm{\phi}(s,a)=\mathbf{x}\otimes\mathbf{e}_{a+1}\in\mathbb{R}^{100}$, where $\mathbf{x}$ is the output of image encoder and $\otimes$ stands for tensor product.
+Appendix D Missing Proofs from Section 2
+D.1 Proof of Lemma 2.9
 We first provide the following lemmas of concentration.
-Lemma B.1 (Zhao et al. 2024, Lemma C.1).
+Lemma D.1 (Zhao et al. 2024, Lemma C.1).
 For any policy $\pi$ and state-action pairs $\{(s_{i},a_{i})\}_{i=1}^{m}$ generated i.i.d.
@@
 where $\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})$ is the $\epsilon_{c}$-covering number of $\mathcal{G}$}.
-Lemma B.2 (Zhao et al. 2024, Lemma C.2).
-For arbitrary policy $\pi$ and dataset $\{(s_{i},a_{i},r_{i})\}_{i=1}^{m}$ generated i.i.d., from the product of $\pi$, $\rho$ and the Bradley-Terry Model; let ${\bar{g}}$ be the least square estimator of ${g^{*}}$, then for any $10<\epsilon_{c}<$ 1 and $\delta>0$, with probability at least $1-\delta$ we have
+Lemma D.2 (Zhao et al. 2024, Lemma C.2).
+For arbitrary policy $\pi$ and dataset $\{(s_{i},a_{i},r_{i})\}_{i=1}^{m}$ generated i.i.d., from the product of $\pi$, $\rho$ and the Bradley-Terry Model; let $\bar{g}$} be the least square estimator of $g^{*}$}, then for any $0<\epsilon_{c}<1$ and $\delta>0$, with probability at least $1-\delta$ we have
 $\sum_{i=1}^{n}\big({\bar{g}}(s_{i},a_{i})-{g^{*}}(s_{i},a_{i})\big)^{2}\leq 16\log(a\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+4n\epsilon_{c}.$
+Now we are ready to prove Lemma 2.9.
 Proof of Lemma 2.9.
@@
 $\leq\sup_{g_{1},g_{2}\in{\mathcal{G}}}\frac{\big(g_{1}(s,a)-g_{2}(s,a)\big)^{2}}{\mathbb{E}_{\pi^{\mathsf{ref}}}\big[\big(g_{1}(s,a)-g_{2}(s,a)\big)^{2}\big]}\mathbb{E}_{\pi^{\mathsf{ref}}}\big[\big({\bar{g}}(s,a)-{g^{*}}(s,a)\big)^{2}\big]$
-$=D_{{\mathcal{G}}}^{2}((s,a),\pi^{\mathsf{ref}})\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{g}}(s,a)-{g^{*}}(s,a)\big{)}^{2}\big{]},$ (B.1)
+$=D_{{\mathcal{G}}}^{2}((s,a),\pi^{\mathsf{ref}})\mathbb{E}_{\pi^{\mathsf{ref}}}\big[\big({\bar{g}}(s,a)-{g^{*}}(s,a)\big)^{2}\big],$ (D.1)
 where the inequality holds by taking supremum to $g_{1},g_{2}\in{\mathcal{G}}$. Now we have
@@
 $\leq\frac{2}{n}\big[16\log(\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+4n\epsilon_{c}\big]+\frac{32}{3n}\log(2\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+10\epsilon_{c}$
-$=\frac{128}{3n}\log(2\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+18\epsilon_{c},$ (B.2)
-where the first inequality holds due to Lemma B.1 and second holds due to Lemma B.2. Plugging (B.1) into (B.1) and setting $\epsilon_{c}=O(n^{-1})$ complete the proof. ∎
-B.2 Proof of Lemma 2.13
-Proof of Lemma 2.13.
-We define $Y=-X$. Then it suffices to show that the covariance between $Y$ and $Y^{2}$ is
-$\operatorname{\rm Cov}(Y,Y^{2})$ $=\mathbb{E}[Y^{3}]-\mathbb{E}[Y^{2}]\mathbb{E}[Y]$
-$\geq\big{(}\mathbb{E}[Y^{2}]\big{)}^{3/2}-\mathbb{E}[Y^{2}]\mathbb{E}[Y]$
-$=\big{(}\mathbb{E}[Y^{2}]\big{)}\big{(}\sqrt{\mathbb{E}[Y^{2}]}-\mathbb{E}[Y]\big{)}$
-$\geq 0,$
-where both inequalities follow from Jensen’s inequality. ∎
-B.3 Proof of Lemma 2.16
+$=\frac{128}{3n}\log(2\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+18\epsilon_{c},$ (D.2)
+where the first inequality holds due to Lemma D.1 and second holds due to Lemma D.2. Plugging (D.1) into (D.1) and setting $\epsilon_{c}=O(n^{-1})$ complete the proof. ∎
+D.2 Proof of Lemma 2.14
 This proof is extracted from the proof of Zhao et al. (2024, Theorem 3.3) and we present it here for completeness. By definition of our objective in (2.1), we have
@@
 $\quad-\eta^{2}\cdot\Delta(s,a)\cdot\frac{\pi^{\mathsf{ref}}(a|s)\cdot\exp\bigl(\eta\cdot g(s,a)\bigr)}{Z_{g}(s)}+\eta^{2}\cdot\Delta(s,a)\cdot\frac{\bigl[\pi^{\mathsf{ref}}(a|s)\cdot\exp\bigl(\eta\cdot g(s,a)\bigr)\bigr]^{2}}{[Z_{g}(s)]^{2}}$
-$\quad+\eta\sum_{a^{\prime}\in\mathcal{A}\backslash\{a\}}\frac{\pi^{\mathsf{ref}}(a^{\prime}|x)\cdot\exp\bigl{(}\eta\cdot g(s,a^{\prime})\bigr{)}}{Z_{g}(s)}\cdot\eta\cdot\Delta(s,a^{\prime})\cdot\frac{\pi^{\mathsf{ref}}(a|s)\cdot\exp\bigl{(}\eta\cdot g(s,a)\bigr{)}}{Z_{g}(s)}$
-$=-\eta^{2}\pi_{g}(a|s)\Delta(s,a)+\eta^{2}[\pi_{g}(a|s)]^{2}\cdot\Delta(s,a)+\eta^{2}\sum_{a^{\prime}\in\mathcal{A}\backslash\{a\}}\pi_{g}(a^{\prime}|x)\pi_{g}(a|s)\Delta(s,a^{\prime}).$
+[…]
+[…]
 Therefore, by mean value theorem, there exists $\gamma\in[0,1]$ and $g_{\gamma}=\gamma g+(1-\gamma){g^{*}}$ such that
@@
 which concludes the proof.
-B.4 Proof of Theorem 2.10
+D.3 Proof of Lemma 2.15
+Proof of Lemma 2.15.
+We define Y=XY=-X. Then it suffices to show that the covariance between $Y$ and $Y^{2}$ is
+$\operatorname{\rm Cov}(Y,Y^{2})$ $=\mathbb{E}[Y^{3}]-\mathbb{E}[Y^{2}]\mathbb{E}[Y]$
+$\geq\big(\mathbb{E}[Y^{2}]\big)^{3/2}-\mathbb{E}[Y^{2}]\mathbb{E}[Y]$
+$=\big(\mathbb{E}[Y^{2}]\big)\big(\sqrt{\mathbb{E}[Y^{2}]}-\mathbb{E}[Y]\big)$
+$\geq 0,$
+where both inequalities follow from Jensen’s inequality. ∎
+D.4 Proof of Theorem 2.10
+To start with, we first define the following quantities. For all $\gamma\in[0,1]$, we define $g_{\gamma}\coloneqq\gamma{\widehat{g}}+(1-\gamma){g^{*}}$ and denote
+$\pi_{\gamma}(\cdot|s)\propto\pi^{\mathsf{ref}}(\cdot|s)\exp\big(\eta g_{\gamma}(s,\cdot)\big),\forall s\in{\mathcal{S}};$
+$G(\gamma)\coloneqq\mathbb{E}_{\rho\times\pi_{\gamma}}\Big[\big({\widehat{g}}-{{g^{*}}}\big)^{2}(s,a)\Big].$
+The key to our analysis is the monotonicity of the function $G(\gamma)$ in $\gamma$, which is formally stated in the following lemma.
+Lemma D.3.
+On event $\mathcal{E}$, $0\in\mathop{\mathrm{argmax}}_{\gamma\in[0,1]}G(\gamma)$.
+Proof.
+For simplicity, we use $\triangle(s,a)$ to denote $\big({\widehat{g}}-{{g^{*}}})(s,a)$ in this proof.
+Then we know that $\triangle(s,a)\leq 0$ for all $(s,a)\in{\mathcal{S}}\times\mathcal{A}$ on event $\mathcal{E}$.
+The most direct way to prove is to take derivative of $G$ with respect to $\gamma$, which corresponds to the policy gradient (Sutton et al., 1999) of $\pi_{\gamma}$ and thus implying a favorable structure.
+A direct calculation yields that
+$\ =\mathbb{E}_{\rho\times\pi_{\gamma}}\big[\nabla_{\gamma}\log\pi_{\gamma}(a|s)\triangle(s,a)^{2}\big]$
+$\ =\eta\mathbb{E}_{\rho}\mathbb{E}_{a\sim\pi_{\gamma}}\big[\triangle^{2}(s,a)\big(\triangle(s,a)-\mathbb{E}_{a^{\prime}\sim\pi_{\gamma}}[\triangle(s,a^{\prime})]\big)\big]$
+$\ =\eta\mathbb{E}_{\rho}\Big[\mathbb{E}_{\pi_{\gamma}}\big[\triangle^{3}(s,a)\big]-\mathbb{E}_{\pi_{\gamma}}\big[\triangle^{2}(s,a)\big]\mathbb{E}_{\pi_{\gamma}}\big[\triangle(s,a)\big]\Big]$
+$\ \leq 0,$
+where $\mathbb{E}_{\rho}$ is the shorthand of $\mathbb{E}_{s\sim\rho}$, $\mathbb{E}_{\pi_{\gamma}}$ is the shorthand of $\mathbb{E}_{a\sim\pi_{\gamma}}$, the first equation is derived from standard policy gradient and the inequality holds conditioned on the event $\mathcal{E}(\delta)$ due to Lemma 2.15 and Lemma 2.15.
+∎
+Now we are ready to prove Theorem 2.10.
+Proof of Theorem 2.10.
+Following the proof of Zhao et al. (2024, Theorem 3.3), we know that there exists $\bar{\gamma}\in[0,1]$ such that
+$J(\pi^{*})-J(\widehat{\pi})\leq\eta G(\bar{\gamma})\leq\eta G(0),$ (D.3)
+where the first inequality holds due to Lemma 2.14 and the second inequality holds due to the event $\mathcal{E}$ and Lemma D.3.
+The term $G(0)$ can be further bounded with the $D^{2}$-based concentrability as follows
+$G(0)$ $=\eta\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}\Big[\big({\widehat{g}}-{{g^{*}}}\big)^{2}(s,a)\Big]$
+$\leq 4\eta\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}[\Gamma_{n}^{2}(s,a)]$
+$=4\eta\beta^{2}\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}\big[D^{2}_{\mathcal{F}}((s,a);\pi^{\mathsf{ref}})\big]$
+$=\widetilde{O}(\eta D_{\pi^{*}}^{2}n^{-1}\log_{{\mathcal{G}}}(\epsilon_{c})),$ (D.4)
+where the second inequality holds conditioned on $\mathcal{E}(\delta)$ because of Lemma D.3, and the last inequality follows from the definition of $\mathcal{E}(\delta)$ together with Line 2.
+By Lemma 2.9, we know that event $\mathcal{E}$ holds with probability at least $1-\delta$, which finishes the proof.
+∎
+D.5 Proof of Theorem 2.11
 Proof of Theorem 2.11.
 We consider the family of contextual bandits with $S\coloneqq|{\mathcal{S}}|,A\coloneqq|\mathcal{A}|<\infty$ and reward function in some function class $\mathcal{G}$ composed of function ${\mathcal{S}}\times\mathcal{A}\to[0,1]$ as follows.
-$\mathrm{CB}_{\mathcal{G}}\coloneqq\{({\mathcal{S}},\mathcal{A},\rho,r,\pi^{\mathsf{ref}},\eta):r\in\mathcal{G},\rho\in\Delta({\mathcal{S}}),\pi^{\mathsf{ref}}\in\Delta(\mathcal{A}|{\mathcal{S}})\}.$ (B.3)
+$\mathrm{CB}_{\mathcal{G}}\coloneqq\{({\mathcal{S}},\mathcal{A},\rho,r,\pi^{\mathsf{ref}},\eta):r\in\mathcal{G},\rho\in\Delta({\mathcal{S}}),\pi^{\mathsf{ref}}\in\Delta(\mathcal{A}|{\mathcal{S}})\}.$ (D.5)
 Our goal is to prove the following statement.
 Fixing any $S\geq 1$, $\eta>4\log 2$ and $C^{*}\in(2,\exp(\eta/4)]$, then for any estimator $\mathcal{D}\mapsto\widehat{\pi}\in\Delta(\mathcal{A}|{\mathcal{S}})$, for any $n\geq 16SC^{*}$, there exist some function class $\mathcal{G}$, such that $\exists\ \mathrm{inst}=({\mathcal{S}},\mathcal{A},\rho,r,\pi^{\mathsf{ref}},\eta)\in\mathrm{CB}_{\mathcal{G}}$ with single-policy concentrability $C^{\pi^{*}}\leq C^{*}$, regularization coefficient $\eta$, $|{\mathcal{S}}|=S=\Theta(\log|\mathcal{G}|)$, and
-$\mathrm{SubOpt}_{\mathrm{RKL}}(\widehat{\pi};\mathrm{inst})\gtrsim\min\{\eta SC^{*}n^{-1},(SC^{*})^{1/2}n^{-1/2}\}.$ (B.4)
-Since $\log|\mathcal{G}|\geq\log\mathcal{N}_{\mathcal{G}}(\epsilon)$ for any $\epsilon\in(0,1)$, equation (B.4) yields the desired bound.
-[…]
+$\mathrm{SubOpt}_{\mathrm{RKL}}(\widehat{\pi};\mathrm{inst})\gtrsim\min\{\eta SC^{*}n^{-1},(SC^{*})^{1/2}n^{-1/2}\}.$ (D.6)
+Since $\log|\mathcal{G}|\geq\log\mathcal{N}_{\mathcal{G}}(\epsilon)$ for any $\epsilon\in(0,1)$, equation (D.6) yields the desired bound.
+We set ${\mathcal{S}}=[S]$, $\mathcal{A}=\{\pm 1\}$, $\rho=\mathsf{Unif}({\mathcal{S}})$, and the reference policy to be
 $\forall s\in{\mathcal{S}},\pi^{\mathsf{ref}}(-1|s)=C^{-1},\pi^{\mathsf{ref}}(+1|s)=1-C^{-1};$
@@
 By 2.2, the optimal policy $\pi_{\tau}^{*}$ under $r_{\tau}$ is
-$\forall s\in{\mathcal{S}},\pi_{\tau}^{*}(-1|s)=\frac{\exp\big{(}\eta(\alpha+\tau_{s}\delta)\big{)}}{\exp\big{(}\eta(\alpha+\tau_{s}\delta)\big{)}+C-1},\pi_{\tau}^{*}(+1|s)=\frac{C-1}{\exp\big{(}\eta(\alpha+\tau_{s}\delta)\big{)}+C-1}.$ (B.5)
+$\forall s\in{\mathcal{S}},\pi_{\tau}^{*}(-1|s)=\frac{\exp\big(\eta(\alpha+\tau_{s}\delta)\big)}{\exp\big(\eta(\alpha+\tau_{s}\delta)\big)+C-1},\pi_{\tau}^{*}(+1|s)=\frac{C-1}{\exp\big(\eta(\alpha+\tau_{s}\delta)\big)+C-1}.$ (D.7)
 Since $C^{*}\leq\exp(\eta/4)$, we assign $C=C^{*}$ and $\alpha=\eta^{-1}\log(C-1)\Leftrightarrow C-1=\exp(\eta\alpha)$, which gives
@@
 where the last inequality is due to the assumption $C^{*}\geq 2$. Therefore, we obtain
-$\max_{\tau\in\{\pm 1\}^{S}}C^{\pi_{\tau}^{*}}\leq C^{*}.$ (B.6)
+$\max_{\tau\in\{\pm 1\}^{S}}C^{\pi_{\tau}^{*}}\leq C^{*}.$ (D.8)
 We will abuse the notation $\mathrm{SubOpt}(\widehat{\pi};\tau)\coloneqq\mathrm{SubOpt}(\widehat{\pi};r_{\tau})$. Since $\rho=\mathsf{Unif}({\mathcal{S}})$,
-$\mathrm{SubOpt}(\widehat{\pi};\tau)=\frac{1}{S}\sum_{s=1}^{S}\mathrm{SubOpt}_{s}(\widehat{\pi};\tau),$ (B.7)
+$\mathrm{SubOpt}(\widehat{\pi};\tau)=\frac{1}{S}\sum_{s=1}^{S}\mathrm{SubOpt}_{s}(\widehat{\pi};\tau),$ (D.9)
 where
@@
 $=\frac{1}{\eta}\mathbb{E}_{a\sim\widehat{\pi}(\cdot|s)}\bigg[\log\frac{\pi^{\mathsf{ref}}(a|s)\cdot\exp\big(\eta r_{\tau}(s,a)\big)}{\pi^{*}_{\tau}(a|s)}-\log\frac{\pi^{\mathsf{ref}}(a|s)\cdot\exp\big(\eta r_{\tau}(s,a)\big)}{\widehat{\pi}(a|s)}\bigg]$
-$=\eta^{-1}{\mathsf{KL}}\left(\widehat{\pi}\|\pi_{\tau}^{*}\right).$ (B.8)
+$=\eta^{-1}{\mathsf{KL}}\left(\widehat{\pi}\|\pi_{\tau}^{*}\right).$ (D.10)
 We write $\tau\sim_{s}\tau^{\prime}$ if $\tau,\tau^{\prime}\in\{\pm 1\}^{{\mathcal{S}}}$ differ in only the $s$-th coordinate and $\tau\sim\tau^{\prime}$ if $\exists s\in{\mathcal{S}},\tau\sim_{s}\tau^{\prime}$.
-By B.8, $\forall s\in{\mathcal{S}}$, $\forall\tau,\tau^{\prime}\in\{\pm 1\}^{{\mathcal{S}}}$ with $\tau\sim_{s}\tau^{\prime}$,
+By D.10, $\forall s\in{\mathcal{S}}$, $\forall\tau,\tau^{\prime}\in\{\pm 1\}^{{\mathcal{S}}}$ with $\tau\sim_{s}\tau^{\prime}$,
 $\mathrm{SubOpt}_{s}(\widehat{\pi};\tau)+\mathrm{SubOpt}_{s}(\widehat{\pi};\tau^{\prime})$
@@
 $\geq-2\eta^{-1}\log\Big(\sum_{b\in\mathcal{A}}\sqrt{\pi_{\tau}^{*}(b|s)\pi_{\tau^{\prime}}^{*}(b|s)}\Big)$
-$=\frac{1}{\eta}\log\frac{(\exp(\eta\delta)+1)(\exp(-\eta\delta)+1)}{4},$ (B.9)
-where $\bar{\pi}(\cdot|s)=\sqrt{\pi^{*}_{\tau}(\cdot|s)\pi^{*}_{\tau^{\prime}}(\cdot|s)}/\sum_{b\in\mathcal{A}}\sqrt{\pi^{*}_{\tau}(b|s)\pi^{*}_{\tau^{\prime}}(b|s)}$ for every $s\in{\mathcal{S}}$, the inequality is due to the non-negativity of KL divergence, and the last equality follows from B.5 together with the design choice $C-1=\exp(\eta\alpha)$.
+$=\frac{1}{\eta}\log\frac{(\exp(\eta\delta)+1)(\exp(-\eta\delta)+1)}{4},$ (D.11)
+where $\bar{\pi}(\cdot|s)=\sqrt{\pi^{*}_{\tau}(\cdot|s)\pi^{*}_{\tau^{\prime}}(\cdot|s)}/\sum_{b\in\mathcal{A}}\sqrt{\pi^{*}_{\tau}(b|s)\pi^{*}_{\tau^{\prime}}(b|s)}$ for every $s\in{\mathcal{S}}$, the inequality is due to the non-negativity of KL divergence, and the last equality follows from D.7 together with the design choice $C-1=\exp(\eta\alpha)$.
 Case $\eta\delta\leq 2$.
 Recall that $\forall x\in\mathbb{R},(\mathrm{e}^{x}+\mathrm{e}^{-x})/2-1=x^{2}\sum_{k=0}^{\infty}\frac{x^{2k}}{(2k+2)!}\geq x^{2}/2$, which implies
-B.9 $=\frac{1}{\eta}\log\bigg{(}1+\frac{1}{2}\Big{(}\frac{\mathrm{e}^{\eta\delta}+\mathrm{e}^{-\eta\delta}}{2}-1\Big{)}\bigg{)}\geq\frac{1}{\eta}\log\Big{(}1+\frac{\eta^{2}\delta^{2}}{4}\Big{)}\geq\frac{1}{\eta}\cdot\frac{\eta^{2}\delta^{2}/4}{2}=\eta\delta^{2}/8.$ (B.10)
+D.11 $=\frac{1}{\eta}\log\bigg(1+\frac{1}{2}\Big(\frac{\mathrm{e}^{\eta\delta}+\mathrm{e}^{-\eta\delta}}{2}-1\Big)\bigg)\geq\frac{1}{\eta}\log\Big(1+\frac{\eta^{2}\delta^{2}}{4}\Big)\geq\frac{1}{\eta}\cdot\frac{\eta^{2}\delta^{2}/4}{2}=\eta\delta^{2}/8.$ (D.12)
 Here, the last inequality is due to $\eta^{2}\delta^{2}/4\leq 1$ and $\forall x\in[0,1],\log(1+x)\geq x/2$.
 Case $\eta\delta>2$.
-We have $-\eta^{-1}2\log 2\geq-\delta\log 2$, which implies the following bound.
-B.9 $\geq\frac{1}{\eta}\log\frac{\exp(\eta\delta)+1}{4}\geq\frac{\eta\delta-2\log 2}{\eta}=\delta-\eta^{-1}2\log 2\geq(1-\log 2)\delta\geq 3\delta/10.$ (B.11)
-In summary, B.10 and B.11 imply that $\forall s\in{\mathcal{S}}$, $\forall\tau,\tau^{\prime}\in\{\pm 1\}^{{\mathcal{S}}}$ with $\tau\sim_{s}\tau^{\prime}$,
-$\mathrm{SubOpt}_{s}(\widehat{\pi};\tau)+\mathrm{SubOpt}_{s}(\widehat{\pi};\tau^{\prime})\geq\frac{\eta\delta^{2}}{8}\wedge\frac{3\delta}{10}.$ (B.12)
+We have $\eta^{-1}2\log 2\geq-\delta\log 2$, which implies the following bound.
+D.11 $\geq\frac{1}{\eta}\log\frac{\exp(\eta\delta)+1}{4}\geq\frac{\eta\delta-2\log 2}{\eta}=\delta-\eta^{-1}2\log 2\geq(1-\log 2)\delta\geq 3\delta/10.$ (D.13)
+In summary, D.12 and D.13 imply that $\forall s\in{\mathcal{S}}$, $\forall\tau,\tau^{\prime}\in\{\pm 1\}^{{\mathcal{S}}}$ with $\tau\sim_{s}\tau^{\prime}$,
+$\mathrm{SubOpt}_{s}(\widehat{\pi};\tau)+\mathrm{SubOpt}_{s}(\widehat{\pi};\tau^{\prime})\geq\frac{\eta\delta^{2}}{8}\wedge\frac{3\delta}{10}.$ (D.14)
 Let $P_{\tau}$ be the distribution of $(s,a,y)$ where $s\sim\rho,a\sim\pi^{\mathsf{ref}}(\cdot|s)$, and $y\sim\mathsf{Bern}(r_{\tau}(s,a))$.
@@
 $=\frac{1}{S}\cdot C^{-1}{\mathsf{KL}}\left(\mathsf{Bern}(r_{\tau}(x,-1))\|\mathsf{Bern}(r_{\tau^{\prime}}(x,-1))\right)$
-$\leq\frac{4\delta^{2}}{SC(0.25-\delta^{2})}\leq\frac{16\delta^{2}}{3SC},$ (B.13)
+$\leq\frac{4\delta^{2}}{SC(0.25-\delta^{2})}\leq\frac{16\delta^{2}}{3SC},$ (D.15)
 where we use the requirement $\delta\leq 1/4$ and ${\mathsf{KL}}\left(\mathsf{Bern}(p)\|\mathsf{Bern}(q)\right)\leq(p-q)^{2}/\big(q(1-q)\big)$.
-Then let $P_{\mathcal{D}_{\tau}}$ be the distribution of $\mathcal{D}$ given the mean reward function $r_{\tau}$, we employ B.13 to get
-${\mathsf{KL}}\left(P_{\mathcal{D}_{\tau}}\|P_{\mathcal{D}_{\tau^{\prime}}}\right)=n{\mathsf{KL}}\left(P_{\tau}\|P_{\tau^{\prime}}\right)\leq\frac{16n\delta^{2}}{3SC}.$ (B.14)
+Then let $P_{\mathcal{D}_{\tau}}$ be the distribution of $\mathcal{D}$ given the mean reward function $r_{\tau}$, we employ D.15 to get
+${\mathsf{KL}}\left(P_{\mathcal{D}_{\tau}}\|P_{\mathcal{D}_{\tau^{\prime}}}\right)=n{\mathsf{KL}}\left(P_{\tau}\|P_{\tau^{\prime}}\right)\leq\frac{16n\delta^{2}}{3SC}.$ (D.16)
 Since $n\geq 16SC^{*}=16SC$ by design, we can set $\delta=\sqrt{SC/n}$ (which ensures $\delta\leq 1/4$) to obtain
@@
 $\geq\Big(\frac{\eta SC^{*}}{32n}\wedge\frac{3\sqrt{SC^{*}}}{40\sqrt{n}}\Big)\exp(-16/3)\gtrsim\frac{\eta SC^{*}}{n}\wedge\sqrt{\frac{SC^{*}}{n}}.$
-where the $S^{-1}$ in the second inequality comes from B.7, the second inequality is by substituting B.12 into Assouad’s Lemma (Lemma F.3), and the last inequality is due to B.14. ∎
-Appendix C Missing Proof from Section 3
-C.1 Proof of Theorem 3.2
+where the $S^{-1}$ in the second inequality comes from D.9, the second inequality is by substituting D.14 into Assouad’s Lemma (Lemma H.3), and the last inequality is due to D.16. ∎
+Appendix E Missing Proof from Section 3
+E.1 Proof of Theorem 3.2
 Before coming to the proof, we first introduce some useful properties. The following properties characterize the convexity of $f$-divergence when $f$ is (strongly) convex.
 The strong-convexity of $f$ implies that the corresponding $f$-divergence, $D_{f}(\cdot||\pi^{\mathsf{ref}})$ is also strongly convex with respect to all $\pi:{\mathcal{S}}\rightarrow\Delta(\mathcal{A})$ supported by $\pi^{\mathsf{ref}}$.
-Proposition C.1.
+Proposition E.1.
 Given context $s$, $D_{f}(\pi(\cdot|s)||\pi^{\mathsf{ref}}(\cdot|s))$ is strict convex with respect to $\pi$ if $f$ is strictly convex.
-Proposition C.2.
+Proposition E.2.
 Given context $s$, $\pi(\cdot|s)\mapsto D_{f}(\pi(\cdot|s)\|\pi^{\mathsf{ref}}(\cdot|s))$ is $4\alpha$-strong convex with respect to the metric $\mathsf{TV}$ if $f$ is $\alpha$-strongly convex.
-Proof of Proposition C.2.
+Proof of Proposition E.2.
 We first show the gradient of $D_{f}$ with respect to $\pi$.
@@
 We have the following properties for the convex conjugate. The first property gives the gradient of convex conjugate (see, e.g., Zhou 2018, Lemma 5).
-Proposition C.3.
+Proposition E.3.
 Given context $s$, and convex $f$, let $\pi_{r}\in\mathop{\mathrm{argmax}}_{\pi}\{\left\langle\pi(\cdot|s),r(s,\cdot)\right\rangle-H_{s}(\pi)\}$ for some $r$, then the gradient of $H^{*}_{s}$ is given by $\nabla H^{*}_{s}(r)=\pi_{r}(\cdot|s)$.
 We also need some properties of $\nabla^{2}H^{*}_{s}$, the Hessian matrix of the convex conjugate function. We first give the Hessian matrix of the original function $H_{s}$ as follows.
-$\nabla^{2}H_{s}(\pi)=\eta^{-1}\mathrm{diag}\Bigg{(}\pi^{\mathsf{ref}}(a_{1}|s)^{-1}f^{\prime\prime}\bigg{(}\frac{\pi(a_{1}|s)}{\pi^{\mathsf{ref}}(a_{1}|s)}\bigg{)},\cdots,\pi^{\mathsf{ref}}(a_{|\mathcal{A}|}|s)^{-1}f^{\prime\prime}\bigg{(}\frac{\pi(a_{|\mathcal{A}|}|s)}{\pi^{\mathsf{ref}}(a_{|\mathcal{A}|}|s)}\bigg{)}\Bigg{)}.$ (C.1)
+$\nabla^{2}H_{s}(\pi)=\eta^{-1}\mathrm{diag}\Bigg(\frac{f^{\prime\prime}\Big(\frac{\pi(a_{1}|s)}{\pi^{\mathsf{ref}}(a_{1}|s)}\Big)}{\pi^{\mathsf{ref}}(a_{1}|s)},\cdots,\frac{f^{\prime\prime}\Big(\frac{\pi(a_{|\mathcal{A}|}|s)}{\pi^{\mathsf{ref}}(a_{|\mathcal{A}|}|s)}\Big)}{\pi^{\mathsf{ref}}(a_{|\mathcal{A}|}|s)}\Bigg).$ (E.1)
 Furthermore, when $f$ is $\alpha$-strongly convex, we have
@@
 The following lemma, which gives an estimate of $\nabla^{2}H^{*}_{s}$, is the pivot of the proof.
-Lemma C.4.
+Lemma E.4.
 For any reward $r:{\mathcal{S}}\times\mathcal{A}\to[0,1]$, we have
 $\nabla^{2}H^{*}_{s}(r)\preceq\alpha^{-1}\eta\mathrm{diag}\big(\pi^{\mathsf{ref}}(a_{1}|s),\cdots,\pi^{\mathsf{ref}}(a_{|\mathcal{A}|}|s)\big).$
-Proof of Lemma C.4.
+Proof of Lemma E.4.
 Given reward function $r:{\mathcal{S}}\times\mathcal{A}\to[0,1]$, we consider
 $\pi_{r}\in\mathop{\mathrm{argmax}}_{\pi\in{\mathcal{S}}\to\Delta^{|\mathcal{A}|}}\{\left\langle\pi(\cdot|s),r(\cdot|s)\right\rangle-H_{s}(\pi)\}.$
-From equation (C.1) we know that $\nabla^{2}H_{s}(\pi_{r})$ is invertible.
+From E.1 we know that $\nabla^{2}H_{s}(\pi_{r})$ is invertible.
 Therefore, by Penot 1994, Proposition 3.2, we have $\nabla^{2}H^{*}_{s}(r)\preceq(\nabla^{2}H_{s}(\pi_{r}))^{-1}$.
@@
 We have the following sub-optimality decomposition
-$J(\pi^{*})-J(\widehat{\pi})$ $=\mathbb{E}_{s\sim\rho}\Big{[}\mathbb{E}_{a\sim\pi^{*}}[{g^{*}}(s,a)]-\mathbb{E}_{a\sim\widehat{\pi}}[{g^{*}}(s,a)]-\eta^{-1}\big{[}D_{f}(\pi^{*}\|\pi^{\mathsf{ref}})-D_{f}(\widehat{\pi}\|\pi^{\mathsf{ref}})\big{]}\Big{]}$
+[…]
 $=\mathbb{E}_{s\sim\rho}\big[H^{*}_{s}({g^{*}})-H^{*}_{s}({\bar{g}})-\left\langle\widehat{\pi},{g^{*}}-{\bar{g}}\right\rangle\big]$
@@
 Now, for any $\delta\in(0,1)$ and $\epsilon_{c}>0$, with probability at least $1-\delta$
-$J(\pi^{*})-J(\widehat{\pi})$ $=\mathbb{E}_{s\sim\rho}[({g^{*}}-{\bar{g}})^{\top}\nabla^{2}H^{*}_{s}(\widetilde{g})({g^{*}}-{\bar{g}})]$
+[…]
 $\leq\alpha^{-1}\eta\mathbb{E}_{s\sim\rho}\Big[({g^{*}}-{\bar{g}})^{\top}\mathrm{diag}\big(\pi^{\mathsf{ref}}(a_{1}|s),\cdots,\pi^{\mathsf{ref}}(a_{|\mathcal{A}|}|s)\big)({g^{*}}-{\bar{g}})\Big]$
@@
 $\leq\alpha^{-1}\eta\bigg(\frac{128}{3n}\log(2\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+18\epsilon_{c}\bigg),$
-where the first inequality holds due to Lemma C.4 and last inequality holds due to equation (B.1). Setting $\epsilon_{c}=O(n^{-1})$ completes the proof. ∎
-C.2 Proof of Theorem 3.4
+where the first inequality holds due to Lemma E.4 and last inequality holds due to equation (D.1). Setting $\epsilon_{c}=O(n^{-1})$ completes the proof. ∎
+E.2 Proof of Theorem 3.4
 We first provide the following lemma that gives the close form of optimal policy under $\chi^{2}$-divergence regularization.
-Lemma C.5 (Huang et al. (2025, Lemma G.2)).
+Lemma E.5 (Huang et al. (2025a, Lemma G.2)).
 Let $\pi^{*}$ be the optimal policy of $\chi^{2}$-divergence regularized objective with reward function $r$, then $\pi^{*}$ has the closed form
 $\pi^{*}(\cdot)=\pi^{\mathsf{ref}}(\cdot)\max\big\{0,\eta(r(\cdot)-\lambda)\big\},\text{ where }\sum_{a\in\mathcal{A}}\pi^{*}_{f\mathrm{div}}(a)=1.$
-By Proposition C.2, $\pi^{*}_{f\mathrm{div}}=\mathop{\mathrm{argmax}}_{\pi\in\Delta(\mathcal{A})}J_{f\mathrm{div}}(\pi)$ is unique. The sub-optimality gap for $f$-divergence is consequently defined as
-$\mathrm{SubOpt}_{f\mathrm{div}}(\cdot)\coloneqq\mathrm{SubOpt}_{f\mathrm{div}}(\cdot;\mathcal{A},r,\pi^{\mathsf{ref}})=J_{f\mathrm{div}}(\pi^{*}_{f\mathrm{div}})-J_{f\mathrm{div}}(\cdot).$ (C.2)
+By Proposition E.2, $\pi^{*}_{f\mathrm{div}}=\mathop{\mathrm{argmax}}_{\pi\in\Delta(\mathcal{A})}J_{f\mathrm{div}}(\pi)$ is unique. The sub-optimality gap for $f$-divergence is consequently defined as
+$\mathrm{SubOpt}_{f\mathrm{div}}(\cdot)\coloneqq\mathrm{SubOpt}_{f\mathrm{div}}(\cdot;\mathcal{A},r,\pi^{\mathsf{ref}})=J_{f\mathrm{div}}(\pi^{*}_{f\mathrm{div}})-J_{f\mathrm{div}}(\cdot).$ (E.2)
 Now we are ready to prove Theorem 3.4.
 Proof of Theorem 3.4.
-We still consider the family of contextual bandits $\mathrm{CB}_{\mathcal{G}}$ given by (B.3).
+We still consider the family of contextual bandits $\mathrm{CB}_{\mathcal{G}}$ given by (D.5).
 We, still, aim to prove the following statement.
 Fixing any $S\geq 32\log 2$, $\eta>4\log 2$ and $\alpha$, we set $f(x)\coloneqq\alpha(x-1)^{2}/2$, then for any estimator $\mathcal{D}\mapsto\widehat{\pi}\in\Delta(\mathcal{A}|{\mathcal{S}})$, for any $n$ sufficiently large, there exist some function class $\mathcal{G}$, such that $\exists\ \mathrm{inst}=({\mathcal{S}},\mathcal{A},\rho,r,\pi^{\mathsf{ref}},\eta)\in\mathrm{CB}_{\mathcal{G}}$ with $|{\mathcal{S}}|=S=\Theta(\log|\mathcal{G}|)$, and
-$\mathrm{SubOpt}_{f}(\widehat{\pi};\mathrm{inst})\gtrsim\alpha^{-1}\eta Sn^{-1}.$ (C.3)
-Since $\log|\mathcal{G}|\geq\log\mathcal{N}_{\mathcal{G}}(\epsilon)$ for any $\epsilon\in(0,1)$, equation (C.3) yields the desired bound.
-[…]
-$\mathcal{G}=\{r_{v}(s,-1)=1/2+v_{s}\delta,r_{v^{\prime}}(s,+1)=1/2+v_{s}\delta,\ \forall s\in{\mathcal{S}}|v\in\mathcal{V}\},$
+$\mathrm{SubOpt}_{f\mathrm{div}}(\widehat{\pi};\mathrm{inst})\gtrsim\alpha^{-1}\eta Sn^{-1}.$ (E.3)
+Since $\log|\mathcal{G}|\geq\log\mathcal{N}_{\mathcal{G}}(\epsilon)$ for any $\epsilon\in(0,1)$, equation (E.3) yields the desired bound.
+We again omit subscripts ffdiv when it is clear in context.
+We set ${\mathcal{S}}=[S]$, $\mathcal{A}=\{-1,+1\}$, and $\rho=\mathsf{Unif}({\mathcal{S}})$.
+For all $s\in{\mathcal{S}}$, $\pi^{\mathsf{ref}}=\mathsf{Unif}(\mathcal{A})$.
+We further consider the following reward function class.
+We leverage Lemma H.4 and obtain a set $\mathcal{V}\in\{-1,+1\}^{S}$ such that (1) $|\mathcal{V}|\geq\exp(S/8)$ and (2) for any $v,v^{\prime}\in\mathcal{V},v\neq v^{\prime}$ , one has $\|v-v^{\prime}\|_{1}\geq S/2$.
+We construct the following reward function class where the reward follows Bernoulli distribution and the mean functions are given by the function class
+$\mathcal{G}=\{r_{v}(s,-1)=1/2+v_{s}\delta,r_{v^{\prime}}(s,+1)=1/2+v_{s}^{\prime}\delta,\ \forall s\in{\mathcal{S}}|v\in\mathcal{V}\},$
 where $\delta\in(0,\eta^{-1}\alpha]$ is to be specified later.
@@
 where $\lambda=0.5-\eta^{-1}\alpha$. Note that $2{\chi^{2}}(\mu\|\nu)+1=\sum_{a\in\mathcal{A}}[\mu(a)]^{2}/\nu(a)$ and ${\chi^{2}}=D_{f}$, we obtain that $\forall\ \widehat{\pi}$,
-$\mathrm{SubOpt}_{s}(\widehat{\pi}(\cdot|s);r_{1})+\mathrm{SubOpt}_{s}(\widehat{\pi}(\cdot|s);r_{2})$ (C.4)
+$\mathrm{SubOpt}_{s}(\widehat{\pi}(\cdot|s);r_{1})+\mathrm{SubOpt}_{s}(\widehat{\pi}(\cdot|s);r_{2})$ (E.4)
 $\quad=\langle{r_{1}(s,\cdot)},{\pi^{*}_{1}(\cdot|s)}\rangle+\langle{r_{2}(s,\cdot)},{\pi^{*}_{2}(\cdot|s)}\rangle-\overbrace{\langle{r_{1}(s,\cdot)+r_{2}(s,\cdot)},{\widehat{\pi}(\cdot|s)}\rangle}^{=1}+\overbrace{2\eta^{-1}\alpha{\chi^{2}}(\widehat{\pi}(\cdot|s)\|\pi^{\mathsf{ref}}(\cdot|s))}^{\geq 0}$
@@
 $\quad\geq 2\langle{r_{1}(s,\cdot)},{\pi^{*}_{1}(\cdot|s)}\rangle-1-2\eta^{-1}\alpha\cdot{\chi^{2}}(\pi^{*}_{1}(\cdot|s)\|\pi^{\mathsf{ref}}(\cdot|s))$
-$\quad=1+\frac{2\eta\delta^{2}}{\alpha}-1-\frac{\eta\delta^{2}}{\alpha}=\frac{\eta\delta^{2}}{\alpha}.$ (C.5)
+$\quad=1+\frac{2\eta\delta^{2}}{\alpha}-1-\frac{\eta\delta^{2}}{\alpha}=\frac{\eta\delta^{2}}{\alpha}.$ (E.5)
 Now we take expectation over all possible contexts and recall that $\|v-v^{\prime}\|_{1}\geq S/2$ for $v\neq v^{\prime}$, we know that for any $r_{1}\neq r_{2}\in\mathcal{G}$
@@
 Suppose $P_{\mathcal{D}_{r}}$ is the distribution of the dataset given mean reward function $r$, then ${\mathsf{KL}}\left(P_{\mathcal{D}_{r_{1}}}\|P_{\mathcal{D}_{r_{2}}}\right)=n{\mathsf{KL}}\left(P_{{r_{1}}}\|P_{{r_{2}}}\right)$ for any pair of $r_{1},r_{2}\in\mathcal{G}$.
-Now we invoke Fano’s inequality (Lemma F.2) to obtain
+Now we invoke Fano’s inequality (Lemma H.2) to obtain
 $\inf_{\pi}\sup_{\mathrm{inst}\in\mathrm{CB}_{\mathcal{G}}}\mathrm{SubOpt}(\widehat{\pi};\mathrm{inst})$ $\geq\frac{\eta\delta^{2}}{4\alpha}\Bigg(1-\frac{\max_{r_{1}\neq r_{2}\in\mathcal{G}}{\mathsf{KL}}\left(P_{\mathcal{D}_{r_{1}}}\|P_{\mathcal{D}_{r_{2}}}\right)+\log 2}{\log|\mathcal{G}|}\Bigg)$
 $\geq\frac{\eta\delta^{2}}{4\alpha}\Bigg(1-\frac{64n\delta^{2}+8\log 2}{S}\Bigg),$
-where the second inequality holds due to $\mathsf{KL}(\mathsf{Bern}(p)\|\mathsf{Bern}(q))\leq(p-q)^{2}/[q(1-q)]$. Let $\delta=16\sqrt{\alpha\eta^{-1}n^{-1}}$, then we obtain that for all $\pi$ we have
+where the second inequality holds due to $\mathsf{KL}(\mathsf{Bern}(p)\|\mathsf{Bern}(q))\leq(p-q)^{2}/[q(1-q)]$. Let $\delta=16^{-1}\sqrt{Sn^{-1}}$, then we obtain that for all $\pi$ we have
 $\sup_{\mathrm{inst}\in\mathrm{CB}_{\mathcal{G}}}\mathrm{SubOpt}(\widehat{\pi};\mathrm{inst})\gtrsim\frac{\eta S}{\alpha n},$
 which finishes the proof in that $\log_{2}|\mathcal{G}|=S$. ∎
-Appendix D Missing Details from Section 4
-In this section, we provide additional details that has been omitted in Section 4.
-D.1 Algorithms for KL-regularized Contextual Dueling Bandits
+Appendix F Generalization to contextual dueling bandits
+In this section, we extend our algorithm to the problems of regularized contextual dueling bandits, where the learner receives preference comparison instead of absolute signals.
+Our setup largely follows Zhu et al.
+(2023); Zhan et al.
+(2023) and the notion of sub-optimality follows Xiong et al.
+(2024); Zhao et al.
+(2024).
+F.1 Problem setup
+We still consider contextual bandits $({\mathcal{S}},\mathcal{A},r,\pi^{\mathsf{ref}})$ where $\mathcal{S}$} is the state space, $\mathcal{A}$ is the action space and $r:{\mathcal{S}}\times\mathcal{A}\to[0,1]$ is the reward function.
+[^6] We overload some notations in Section 2 by their dueling counterparts for notational simplicity.
+But only relative preference feedback is available, viz., we have an $\mathrm{i.i.d.}$ offline dataset $\mathcal{D}=\{(s_{i},a^{1}_{i},a^{2}_{i},y_{i})\}_{i=1}^{n}$, where $s_{i}\in{\mathcal{S}}$ is generated from distribution $\rho$ and $a^{1}_{i},a^{2}_{i}\sim\pi^{\mathsf{ref}}$.
+The binary preference label $y_{i}=1$ indicates $a_{i}^{1}$ is preferred over $a_{i}^{2}$ (denoted by $a^{1}\succ a^{2}$) and 0 for $a^{2}\succ a^{1}$ given context $s$.
+In this work we consider the Bradley-Terry Model, where $\mathbb{P}[y=1|s,a^{1},a^{2}]=\sigma(r(s_{i},a^{1}_{i})-r(s_{i},a^{2}_{i}))$, where $\sigma(x)=(1+e^{-x})^{-1}$ is the link function.
+The objective here identical to (2.1) for KL-regularization and (3.1) for $f$-divergence regularization.
+Our goal is still to find an $\epsilon$-optimal policy.
+To control the complexity of the function class $\mathcal{G}$}, we assume that Assumption 2.1 still holds here.
+Concentrability.
+Analogous to Section 2, we need our estimation from offline dataset generalizable to the state-action pairs visited by our obtained policy.
+While density-ratio-based concentrability can be directly adapted to dueling bandit, we need a slightly different notion of $D^{2}$-divergence.
+This is because in dueling bandit, we cannot observe the absolute reward and best estimation $g$ we can achieve is that for any state $s$ and actions $a^{1},a^{2}$, our estimated $g(s,a^{1})-g(s,a^{2})\approx r(s,a^{1})-r(s,a^{2})$.
+This implies that there exists some mapping $b:{\mathcal{S}}\to[-1,1]$ such that $g(s,a)-b(s)\approx r(s,a)$ on the offline data, which leads to the following definition.
+Definition F.1.
+Given a class of functions ${\mathcal{G}}\subset({\mathcal{S}}\times\mathcal{A}\to\mathbb{R})$ and some policy $\pi$, let $\mathcal{B}=({\mathcal{S}}\to[-1,1])$ be the function class, define the $D^{2}$-divergence $D^{2}_{{\mathcal{G}}}((s,a);\pi)$ as
+$\sup_{g,h\in{\mathcal{G}}}\inf_{b\in\mathcal{B}}\frac{\big(g(s,a)-h(s,a)-b(s)\big)^{2}}{\mathbb{E}_{s\sim\rho}\operatorname{{\rm Var}}_{a^{\prime}\sim\pi(\cdot|s^{\prime})}[g(s^{\prime},a^{\prime})-h(s^{\prime},a^{\prime})]}.$
+A similar definition has been introduced in Zhao et al.
+(2024, Definition 2.6), which underpins the following two assumptions that characterize the coverage ability of $\pi^{\mathsf{ref}}$ similarly as in Section 2.
+Given a reference policy $\pi^{\mathsf{ref}}$, we define two coverage notions for contextual dueling bandits.
+Assumption F.2 (All-policy concentrability).
+$D^{2}\coloneqq\sup_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}D^{2}_{{\mathcal{G}}}((s,a);\pi^{\mathsf{ref}})<\infty$.
+Assumption F.3 (Single-policy concentrability).
+$D_{\pi^{*}}^{2}\coloneqq\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}[D^{2}_{{\mathcal{G}}}((s,a);\pi^{\mathsf{ref}})]<\infty$.
+Similar single-policy concentrability assumptions have appeared in previous work in offline contextual dueling bandits (Huang et al., 2025b; Song et al., 2024) and similar notions has also appeared in the analysis of model-based RL (Uehara and Sun, 2021; Wang et al., 2024).
+Still, while Assumption F.3 is strictly weaker than Assumption F.2, in general cases, the two quantities, $C^{\pi^{*}}$ and $D^{2}_{\pi^{*}}$ cannot be bounded by each other.
+F.2 Algorithms and Results
+F.2.1 Algorithms for KL-regularized contextual dueling bandits
 Algorithm 3 Offline KL-Regularized Pessimistic Contextual Dueling Bandit (KL-PCDB)
@@
 ${\bar{g}}=\mathop{\mathrm{argmin}}_{g\in{\mathcal{G}}}\sum_{i=1}^{n}\Big[y_{i}\log\sigma\Big(\big[g(s_{i},a^{1}_{i})-g(s_{i},a^{2}_{i})\big]\Big)+(1-y_{i})\log\sigma\Big(\big[g(s_{i},a^{2}_{i})-g(s_{i},a^{1}_{i})\big]\Big)\Big]$
-2:  Let ${\widehat{g}}(s,a)={\bar{g}}(s,a)-\Gamma_{n}(s,a)$, where $\Gamma_{n}(s,a)$ is the bonus term in D.1
-2:  $\widehat{\pi}(a|s)\propto\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta\cdot{\widehat{g}}(s,a)\big{)}$
+2: Let ${\widehat{g}}(s,a)={\bar{g}}(s,a)-\Gamma_{n}(s,a)$, where $\Gamma_{n}(s,a)$ is the bonus term in F.1
+[…]
 We elucidate KL-PCDB for offline KL-regularized contextual dueling bandits, whose pseudocode is summarized in Algorithm 3.
@@
 Specifically, we define the penalty term
-$\Gamma_{n}(s,a)=\beta\sqrt{D_{{\mathcal{G}}}^{2}((s,a),\pi^{\mathsf{ref}})},$ (D.1)
+$\Gamma_{n}(s,a)=\beta\sqrt{D_{{\mathcal{G}}}^{2}((s,a),\pi^{\mathsf{ref}})},$ (F.1)
 where
-$\beta^{2}=128\log(2\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)/3n+18\epsilon_{c}=\widetilde{O}(n^{-1})$ (D.2)
-and then subtract it from the MLE ${\bar{g}}$ to obtain a pessimistic estimator ${\widehat{g}}$. KL-PCB then output the policy $\widehat{\pi}$, maximizing the estimated objective
+$\beta^{2}=128\log(2\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)/3n+18\epsilon_{c}=\widetilde{O}(n^{-1})$ (F.2)
+and then subtract it from the MLE $\bar{g}$} to obtain a pessimistic estimator $^{\widehat{g}}$. KL-PCB then output the policy $\widehat{\pi}$, maximizing the estimated objective
 $\widehat{J}(\pi)=\mathbb{E}_{(s,a)\sim\rho\times\pi}\bigg[{\widehat{g}}(s,a)-\eta^{-1}\log\frac{\pi(a|s)}{\pi^{\mathsf{ref}}(a|s)}\bigg],$
@@
 We provide the following theoretical guarantees for Algorithm 3.
-Theorem D.1.
-Under Assumption 4.3, if we set $\Gamma_{n}$ according to D.1, then for sufficiently small $\epsilon\in(0,1)$, with probability at least $1-\delta$, $n=\widetilde{O}\big{(}\eta D_{\pi^{*}}^{2}\epsilon^{-1}\big{)}$ is sufficient to guarantee the output policy $\widehat{\pi}$ of Algorithm 3 to be $\epsilon$-optimal.
-Remark D.2.
+Theorem F.4.
+Under Assumption F.3, if we set $\Gamma_{n}$ according to F.1, then for sufficiently small $\epsilon\in(0,1)$, with probability at least $1-\delta$, $n=\widetilde{O}\big(\eta(D_{\pi^{*}}^{2}\wedge C^{\pi^{*}})\epsilon^{-1}\big)$ is sufficient to guarantee the output policy $\widehat{\pi}$ of Algorithm 3 to be $\epsilon$-optimal.
+Remark F.5.
 Zhao et al.
-(2024) achieved an $\widetilde{O}(\epsilon^{-1})$ sample complexity under Assumption 4.2.
+(2024) achieved an $\widetilde{O}(\epsilon^{-1})$ sample complexity under Assumption F.2.
 Comparing to Zhao et al.
-(2024), KL-PCDB achieves the same $\widetilde{O}(\epsilon^{-1})$ sample complexity but only requiring Assumption 4.3, which is weaker than Assumption 4.2.
+(2024), KL-PCDB achieves the same $\widetilde{O}(\epsilon^{-1})$ sample complexity but only requiring Assumption F.3, which is weaker than Assumption F.2.
 The following theorem provides the sample complexity lower bound for KL-regularized dueling contextual bandits.
-Theorem D.3.
-For any sufficiently small $\epsilon\in(0,1),\eta>0$, $21\leq C^{*}\leq\exp(\eta/2)/$ 2, and any algorithm Alg, there is a KL-regularized contextual dueling bandit instance with single-policy concentrability $D^{2}_{\pi^{*}}\leq 2C^{*}$ such that Alg requires at least $\Omega\big{(}\min\{{\eta C^{*}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})}/{\epsilon},{\log\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})}(C^{*})^{2}/{\epsilon^{2}}\}\big{)}$ samples to return an $\epsilon$-optimal policy.
-Remark D.4.
-Theorem D.3 shows that when $\epsilon$ is sufficiently small, any algorithm for offline KL-regularized contextual dueling bandits requires at least $\Omega(\eta D^{2}_{\pi^{*}}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon)\epsilon^{-1})$ samples to output an $\epsilon$-optimal policy, which matches the sample complexity upper bound in Theorem D.1, indicating that KL-PCB is nearly optimal.
-D.2 Algorithm and Results for $f$-divergence Regularized Contextual Dueling Bandit
+Theorem F.6.
+For any sufficiently small $\epsilon\in(0,1),\eta>0$, $1\leq C^{*}\leq\exp(\eta/2)/2$, and any algorithm Alg, there is a KL-regularized contextual dueling bandit instance with single-policy concentrability $C^{\pi^{*}}\leq C^{*}$ such that Alg requires at least $\Omega\big(\min\{{\eta C^{*}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})}/{\epsilon},{\log\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})}(C^{*})^{2}/{\epsilon^{2}}\}\big)$ samples to return an $\epsilon$-optimal policy.
+Remark F.7.
+Theorem F.6 shows that when $\epsilon$ is sufficiently small, any algorithm for offline KL-regularized contextual dueling bandits requires at least $\Omega(\eta C^{\pi^{*}}\log\mathcal{N}_{{\mathcal{G}}}(\epsilon)\epsilon^{-1})$ samples to output an $\epsilon$-optimal policy, which matches the sample complexity upper bound in Theorem F.4, indicating that KL-PCB is nearly optimal.
+F.2.2 Algorithm and Results for $f$-divergence Regularized CDBs
 Algorithm 4 Offline $f$-Divergence Regularized Contextual Dueling Bandits ($f$-CDB)
@@
 $\widehat{\pi}(\cdot|s)\leftarrow\mathop{\mathrm{argmax}}_{\pi(\cdot|s)\in\Delta(\mathcal{A})}\sum_{a\in\mathcal{A}}\pi(a|s){\bar{g}}(s,a)+\eta^{-1}D_{f}\big(\pi(\cdot|s)\|\pi^{\mathsf{ref}}(\cdot|s)\big)$
-2:  $\widehat{\pi}(a|s)$
-In this subsection, we present an offline learning algorithm for $f$-divergence regularized contextual dueling bandit, $f$-CDB, in Algorithm 4.
+[…]
+We present an offline learning algorithm for $f$-divergence regularized contextual dueling bandit, $f$-CDB, in Algorithm 4.
 $f$-CDB first leverages maximum likelihood estimator to find a function ${\bar{g}}\in{\mathcal{G}}$ that minimizes its risk on the offline dataset.
@@
 The following theorem provides an upper bound of Algorithm 4.
-Theorem D.5.
-For any sufficiently small $\epsilon\in(0,1)$, and $\eta,\alpha>0$, with probability at least $1-\delta$, $n=\widetilde{O}\big{(}\alpha^{-1}\eta\log{\mathcal{N}}(\epsilon)\epsilon^{-1}\big{)}$ is sufficient to guarantee that the output policy $\widehat{\pi}$ of Algorithm 4 is $\epsilon$-optimal.
+Theorem F.8.
+For any sufficiently small $\epsilon\in(0,1)$, and $\eta,\alpha>0$, with probability at least $1-\delta$, $n=\widetilde{O}(\alpha^{-1}\eta\log{\mathcal{N}}(\epsilon)\epsilon^{-1})$ is sufficient to guarantee that the output policy $\widehat{\pi}$ of Algorithm 4 is $\epsilon$-optimal.
 The following theorem provides a lower bound for offline $f$-divergence regularized contextual dueling bandit with strongly convex $f$.
-Theorem D.6.
+Theorem F.9.
 For any $\epsilon\in(0,1)$, $\alpha,\eta>0$, and offline RL algorithm Alg, there is an $\alpha$-strongly convex $f$ and $f$-divergence regularized contextual dueling bandit instance such that Alg requires at least $\Omega\big(\alpha^{-1}\eta\log{\mathcal{N}}(\epsilon)\epsilon^{-1}\big)$ samples to return an $\epsilon$-optimal policy.
-Remark D.7.
-Theorem D.6 indicates that, when $\epsilon$ is sufficiently small, to produce an $\epsilon$-optimal policy, any algorithm for offline $f$-regularized contextual bandits with strongly convex $f$ requires at least $\widetilde{\Omega}(\alpha^{-1}\eta\epsilon^{-1})$ samples.
-This lower bound matches the sample-complexity upper bound in Theorem D.5, indicating that Algorithm 4 is nearly optimal.
-Appendix E Missing Proof from Appendix D
-E.1 Proof of Theorem D.1
+Remark F.10.
+Theorem F.9 indicates that, when $\epsilon$ is sufficiently small, to produce an $\epsilon$-optimal policy, any algorithm for offline $f$-regularized contextual bandits with strongly convex $f$ requires at least $\widetilde{\Omega}(\alpha^{-1}\eta\epsilon^{-1})$ samples.
+This lower bound matches the sample-complexity upper bound in Theorem F.8, indicating that Algorithm 4 is nearly optimal.
+Appendix G Missing proof from Appendix F
+G.1 Proof of Theorem F.4
 The proof follows the proof in Section 2. At the beginning, we first define the event $\mathcal{E}(\delta)$ given $\delta>0$ as
-$\mathcal{E}(\delta)\coloneqq\Big{\{}\exists\ b:{\mathcal{S}}\to[-1,1],\forall(s,a)\in{\mathcal{S}}\times\mathcal{A},\big{|}{\bar{g}}(s,a)-b(s)-{g^{*}}(s,a)\big{|}\leq\Gamma_{n}(s,a)\Big{\}}.$ (E.1)
-Here, $\Gamma_{n}$ is defined in (D.1). We abuse the notation and define $b(\cdot)$ as
-$b=\mathop{\mathrm{argmin}}_{\mathcal{B}}\sup_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}\Phi_{b}(s,a)-\Gamma_{n}(s,a),$ (E.2)
+$\mathcal{E}(\delta)\coloneqq\Big\{\exists\ b:{\mathcal{S}}\to[-1,1],\forall(s,a)\in{\mathcal{S}}\times\mathcal{A},\big|{\bar{g}}(s,a)-b(s)-{g^{*}}(s,a)\big|\leq\Gamma_{n}(s,a)\Big\}.$ (G.1)
+Here, $\Gamma_{n}$ is defined in (F.1). We abuse the notation and define $b(\cdot)$ as
+$b=\mathop{\mathrm{argmin}}_{\mathcal{B}}\sup_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}\Phi_{b}(s,a)-\Gamma_{n}(s,a),$ (G.2)
 where $\Phi_{b}(s,a)=\big|{\bar{g}}(s,a)-b(s)-{g^{*}}(s,a)\big|$ and when $\mathcal{E}$ holds, for all $(s,a)\in{\mathcal{S}}\times\mathcal{A}$, we have $\Phi_{b}(s,a)\leq\Gamma_{n}(s,a)$ This indicates that the least square estimation $\bar{g}$} obtained in Line 1 of Algorithm 3, after adjusted by some bias function $b$, is close to the true function $g^{*}$}.
 The following lemma shows that this event holds with high probability.
-Lemma E.1.
+Lemma G.1.
 For any $\delta>0$, $\mathbb{P}(\mathcal{E}(\delta))\geq 1-\delta$.
 Proof.
-From Lemma F.1, we have that with probability at least $1-\delta$, it holds that
-$\mathbb{E}_{s^{\prime}\sim\rho}\operatorname{{\rm Var}}_{a^{\prime}\sim\pi^{\mathsf{ref}}(\cdot|s^{\prime})}\big{[}{\bar{g}}(s^{\prime},a^{\prime})-{g^{*}}(s^{\prime},a^{\prime})\big{]}\leq O\bigg{(}\frac{1}{n}\log(\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+\epsilon_{c}\bigg{)}.$ (E.3)
+From Lemma H.1, we have that with probability at least $1-\delta$, it holds that
+$\mathbb{E}_{s^{\prime}\sim\rho}\operatorname{{\rm Var}}_{a^{\prime}\sim\pi^{\mathsf{ref}}(\cdot|s^{\prime})}\big[{\bar{g}}(s^{\prime},a^{\prime})-{g^{*}}(s^{\prime},a^{\prime})\big]\leq O\bigg(\frac{1}{n}\log(\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+\epsilon_{c}\bigg).$ (G.3)
 It further holds true that for some $b:{\mathcal{S}}\to\mathbb{R}$
-$D_{\mathcal{G}}^{2}((s,a),\pi^{\mathsf{ref}}))\cdot\mathbb{E}_{s\sim\rho}\operatorname{{\rm Var}}_{a\sim\pi^{\mathsf{ref}}(\cdot|s)}\big{[}{\bar{g}}(s,a)-{g^{*}}(s,a)\big{]}\geq\big{(}{\bar{g}}(s,a)-b(s)-{g^{*}}(s,a)\big{)}^{2}.$ (E.4)
-Substituting (E.3) into (E.4), we have
-$\inf_{b}\big{(}{\bar{g}}(s,a)-b(s)-{g^{*}}(s,a)\big{)}^{2}$ (E.5)
+$D_{\mathcal{G}}^{2}((s,a),\pi^{\mathsf{ref}}))\cdot\mathbb{E}_{s\sim\rho}\operatorname{{\rm Var}}_{a\sim\pi^{\mathsf{ref}}(\cdot|s)}\big[{\bar{g}}(s,a)-{g^{*}}(s,a)\big]\geq\big({\bar{g}}(s,a)-b(s)-{g^{*}}(s,a)\big)^{2}.$ (G.4)
+Substituting (G.3) into (G.4), we have
+$\inf_{b}\big({\bar{g}}(s,a)-b(s)-{g^{*}}(s,a)\big)^{2}$ (G.5)
 $\quad=\inf_{b}\frac{\big({\bar{g}}(s,a)-b(s)-{g^{*}}(s,a)\big)^{2}}{\mathbb{E}_{s^{\prime}\sim\rho}\operatorname{{\rm Var}}_{a^{\prime}\sim\pi^{\mathsf{ref}}(\cdot|s^{\prime})}\big[{\bar{g}}(s^{\prime},a^{\prime})-{g^{*}}(s^{\prime},a^{\prime})\big]}\mathbb{E}_{s^{\prime}\sim\rho}\operatorname{{\rm Var}}_{a^{\prime}\sim\pi^{\mathsf{ref}}(\cdot|s^{\prime})}\big[{\bar{g}}(s^{\prime},a^{\prime})-{g^{*}}(s^{\prime},a^{\prime})\big]$
-$\quad\leq D_{{\mathcal{G}}}^{2}((s,a),\pi^{\mathsf{ref}})\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{g}}(s,a)-b(s)-{g^{*}}(s,a)\big{)}^{2}\big{]}$ (E.6)
-$\quad\leq D_{{\mathcal{G}}}^{2}((s,a),\pi^{\mathsf{ref}})O\bigg{(}\frac{1}{n}\log(\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+\epsilon_{c}\bigg{)},$ (E.7)
-where the first inequality holds due to the definition of $D_{{\mathcal{G}}}^{2}((s,a),\pi^{\mathsf{ref}})$ and the last inequality holds due to Lemma F.1. ∎
+$\quad\leq D_{{\mathcal{G}}}^{2}((s,a),\pi^{\mathsf{ref}})\mathbb{E}_{\pi^{\mathsf{ref}}}\big[\big({\bar{g}}(s,a)-b(s)-{g^{*}}(s,a)\big)^{2}\big]$ (G.6)
+$\quad\leq D_{{\mathcal{G}}}^{2}((s,a),\pi^{\mathsf{ref}})O\bigg(\frac{1}{n}\log(\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+\epsilon_{c}\bigg),$ (G.7)
+where the first inequality holds due to the definition of $D_{{\mathcal{G}}}^{2}((s,a),\pi^{\mathsf{ref}})$ and the last inequality holds due to Lemma H.1. ∎
 We overload the following quantities. For any $\gamma\in[0,1]$ and $(s,a)\in{\mathcal{S}}\times\mathcal{A}$, we define
@@
 $G(\gamma)\coloneqq\mathbb{E}_{\rho\times\pi_{\gamma}}\big[\big({\widehat{g}}(s,a)-b(s)-{{g^{*}}}(s,a)\big)^{2}\big],$
-where $b(\cdot)$ is defined in (E.2). We still have the monotonicity of the function $G(\gamma)$, which is characterized by the following lemma.
-Lemma E.2.
+where $b(\cdot)$ is defined in (G.2). We still have the monotonicity of the function $G(\gamma)$, which is characterized by the following lemma.
+Lemma G.2.
 On event $\mathcal{E}(\delta)$, $0\in\mathop{\mathrm{argmax}}_{\gamma\in[0,1]}G(\gamma)$.
@@
 Finally, we have the proposition that adding some bias term $b:{\mathcal{S}}\to\mathbb{R}$ does not affect the resulting policy.
-Proposition E.3.
-Let $b:{\mathcal{S}}\to\mathbb{R}$ be some bias function, then for all $g\in{\mathcal{G}}$ we have $J(\pi_{g})=J(\pi_{g-b})$, where $(g-b)(s,a)=g(s,a)-b(s)$.
+Proposition G.3.
+Let $b:{\mathcal{S}}\to\mathbb{R}$ be some bias function, then for all $g\in{\mathcal{G}}$ we have $J(\pi_{g})=J(\pi_{g-b})$, where (gb)(s,a)=g(s,a)b(s)(g-b)(s,a)=g(s,a)-b(s).
 Proof.
@@
 which indicates that $\pi_{g}=\pi_{g-b}$. This immediately leads to $J(\pi_{g})=J(\pi_{g-b})$. ∎
-Now we are ready to prove Theorem D.1.
-Proof of Theorem D.1.
-We proceed the proof under the event $\mathcal{E}(\delta)$. By Proposition E.3, we know that
-$J(\pi^{*})-J(\widehat{\pi})$ $=J(\pi^{*})-J(\pi_{{\widehat{g}}})$
+Now we are ready to prove Theorem F.4.
+Proof of Theorem F.4.
+We proceed the proof under the event $\mathcal{E}(\delta)$. By Proposition G.3, we know that
+[…]
 $=J(\pi^{*})-J(\pi_{{\widehat{g}}-b}).$
 Consequently, there exist some $\gamma\in[0,1]$ and $b:{\mathcal{S}}\to[-1,1]$ such that
-$J(\pi^{*})-J(\widehat{\pi})$ $=J(\pi^{*})-J(\pi_{{\widehat{g}}-b})$
+[…]
 $\leq\eta\mathbb{E}_{\rho\times\pi_{\gamma}}\big[\big({\widehat{g}}(s,a)-b(s)-{{g^{*}}}(s,a)\big)^{2}\big]$
-$=\eta G(\gamma),$ (E.8)
+$=\eta G(\gamma),$ (G.8)
 where the inequality holds due to Lemma 2.14.
 Under event $\mathcal{E}(\delta)$, we know that ${\widehat{g}}(s,a)-b(s)\leq{{g^{*}}}(s,a)$.
-Together with Lemma E.2, we obtain $G(\gamma)\leq G(0)$.
+Together with Lemma G.2, we obtain $G(\gamma)\leq G(0)$.
 Therefore, we know that
-$J(\pi^{*})-J(\widehat{\pi})$ $\leq G(0)$
+[…]
 $=\eta\mathbb{E}_{\rho\times\pi^{*}}\Big[\big({\widehat{g}}(s,a)-b(s)-{{g^{*}}}(s,a)\big)^{2}\Big]$
-$\leq 4\eta\mathbb{E}_{\rho\times\pi^{*}}\big{[}\Gamma_{n}^{2}(s,a)\big{]}$
-$=4\eta\beta^{2}\mathbb{E}_{\rho\times\pi^{*}}\big{[}D^{2}_{{\mathcal{G}}}((s,a);\pi^{\mathsf{ref}})\big{]}$
-$=\widetilde{O}\big{(}\eta D_{\pi^{*}}^{2}\log{{\mathcal{G}}}(\epsilon_{c})n^{-1}\big{)},$ (E.9)
-where the inequality holds due to the definition of $\mathcal{E}(\delta)$.
-Plugging (E.9) into (E.8), we know that $J(\pi^{*})-J(\widehat{\pi})$ has upper bound $\widetilde{O}(D_{\pi^{*}}^{2}n^{-1})$.
-By Lemma E.1, event $\mathcal{E}$ with probability at least $1-\delta$, which concludes the proof.
-∎
-E.2 Proof of Theorem D.3
-Proof of Theorem D.3.
+$\leq 4\eta\Big(\mathbb{E}_{\rho\times\pi^{*}}\big[\Gamma_{n}^{2}(s,a)\big]\wedge C^{\pi^{*}}\mathbb{E}_{\rho\times\pi^{*}}\big[\big({\widehat{g}}(s,a)-b(s)-{{g^{*}}}(s,a)\big)^{2}\big]\Big)$
+$=4\eta\Big(\beta^{2}\mathbb{E}_{\rho\times\pi^{*}}\big[D^{2}_{{\mathcal{G}}}((s,a);\pi^{\mathsf{ref}})\big]\wedge C^{\pi^{*}}\mathbb{E}_{\rho\times\pi^{*}}\big[\big({\widehat{g}}(s,a)-b(s)-{{g^{*}}}(s,a)\big)^{2}\big]\Big)$
+$=\widetilde{O}\big(\eta D_{\pi^{*}}^{2}\log{{\mathcal{G}}}(\epsilon_{c})n^{-1}\big),$ (G.10)
+[…]
+G.2 Proof of Theorem F.6
+Proof of Theorem F.6.
 The proof is similar to the proof of Theorem 2.11.
 Consider the following family of contextual dueling bandit instances with $S\coloneqq|{\mathcal{S}}|,A\coloneqq|\mathcal{A}|<\infty$ and reward in some function class $\mathcal{G}$.
-$\mathrm{CDB}\coloneqq\{({\mathcal{S}},\mathcal{A},\rho,r,\pi^{\mathsf{ref}},\eta):r\in\mathcal{G},\rho\in\Delta({\mathcal{S}}),\pi^{\mathsf{ref}}\in\Delta(\mathcal{A}|{\mathcal{S}})\}.$ (E.10)
-Fixing any $S\geq 1$, $\eta>4\log 2$ and $C^{*}\in(2,\exp(\eta/4)]$, we aim to prove that, for any estimator $\mathcal{D}\mapsto\widehat{\pi}\in\Delta(\mathcal{A}|{\mathcal{S}})$, for any $n\geq 16SC^{*}$, there exist some function class $\mathcal{G}$, such that $\exists\ \mathrm{inst}=({\mathcal{S}},\mathcal{A},\rho,r,\pi^{\mathsf{ref}},\eta)\in\mathrm{CDB}$ with single-policy concentrability $C^{\pi^{*}}\leq 2C^{*}$, regularization coefficient $\eta$, $|{\mathcal{S}}|=S=\Theta(\log|\mathcal{G}|)$, and
-$\inf_{\mathrm{inst}\in\mathrm{CDB}}\mathrm{SubOpt}_{\mathrm{RKL}}(\widehat{\pi};\mathrm{inst})\gtrsim\min\{\eta SC^{*}n^{-1},(SC^{*})^{1/2}n^{-1/2}\}.$ (E.11)
+$\mathrm{CDB}\coloneqq\{({\mathcal{S}},\mathcal{A},\rho,r,\pi^{\mathsf{ref}},\eta):r\in\mathcal{G},\rho\in\Delta({\mathcal{S}}),\pi^{\mathsf{ref}}\in\Delta(\mathcal{A}|{\mathcal{S}})\}.$ (G.11)
+Fixing any $S\geq 1$, $\eta>4\log 2$ and $C^{*}\in(2,\exp(\eta/4)]$, we aim to prove that, for any estimator $\mathcal{D}\mapsto\widehat{\pi}\in\Delta(\mathcal{A}|{\mathcal{S}})$, for any $n\geq 16SC^{*}$, there exist some function class $\mathcal{G}$, such that $\exists\ \mathrm{inst}=({\mathcal{S}},\mathcal{A},\rho,r,\pi^{\mathsf{ref}},\eta)\in\mathrm{CDB}$ with single-policy concentrability $C^{\pi^{*}}\leq C^{*}$, regularization coefficient $\eta$, $|{\mathcal{S}}|=S=\Theta(\log|\mathcal{G}|)$, and
+$\inf_{\mathrm{inst}\in\mathrm{CDB}}\mathrm{SubOpt}_{\mathrm{RKL}}(\widehat{\pi};\mathrm{inst})\gtrsim\min\{\eta SC^{*}n^{-1},(SC^{*})^{1/2}n^{-1/2}\}.$ (G.12)
 Since $\log|\mathcal{G}|\geq\log\mathcal{N}_{\mathcal{G}}(\epsilon)$ for any $\epsilon\in(0,1)$, the above bound yields the desired result.
-[…]
+We construct the same reward function class as in the proof of Theorem 2.11.
+In particular, we set ${\mathcal{S}}=[S]$, $\mathcal{A}=\{\pm 1\}$, $\rho=\mathsf{Unif}({\mathcal{S}})$, and the reference policy to be
 $\forall s\in{\mathcal{S}},\pi^{\mathsf{ref}}(-1|s)=C^{-1},\pi^{\mathsf{ref}}(+1|s)=1-C^{-1};$
 where $C=C^{*}$. Then the total sub-optimality of any $\pi\in\Delta(\mathcal{A}|{\mathcal{S}})$ given any reward function $r:{\mathcal{S}}\times\mathcal{A}\to\mathbb{R}$ is
-$\mathrm{SubOpt}_{f\mathrm{div}}(\pi;r)=\frac{1}{S}\sum_{s=1}^{S}\mathrm{SubOpt}_{f\mathrm{div}}\big{(}\pi(\cdot|s);r(s,\cdot)\big{)}.$ (E.12)
+$\mathrm{SubOpt}_{f\mathrm{div}}(\pi;r)=\frac{1}{S}\sum_{s=1}^{S}\mathrm{SubOpt}_{f\mathrm{div}}\big(\pi(\cdot|s);r(s,\cdot)\big).$ (G.13)
 We further let $\alpha=\eta^{-1}\log(C-1)\Leftrightarrow C-1=\exp(\eta\alpha)$.
@@
 $r_{\tau}(s,-1)=0.5+\tau_{s}\delta,r_{\tau}(s,+1)=0.5-\alpha.$
-Then, following the derivation of B.10 and B.11, we know that $\max_{\tau\in\{\pm 1\}^{S}}C^{\pi_{\tau}^{*}}\leq C^{*}$, and $\forall s\in{\mathcal{S}}$, $\forall\tau,\tau^{\prime}\in\{\pm 1\}^{{\mathcal{S}}}$ with $\tau\sim_{s}\tau^{\prime}$,
-$\mathrm{SubOpt}_{s}(\widehat{\pi};\tau)+\mathrm{SubOpt}_{s}(\widehat{\pi};\tau^{\prime})\geq\frac{\eta\delta^{2}}{8}\wedge\frac{3\delta}{10}.$ (E.13)
-Let $P_{r}$ be the distribution of $(s,a^{1},a^{2},y)$ for $s\sim\rho$, $a^{1},a^{2}\overset{\scriptsize\mathrm{i.i.d.}}{\sim}\pi^{\mathsf{ref}}(\cdot|s)$ and $y\sim\mathsf{Bern}(\sigma(r(s,a^{1})-r(s,a^{2})))$.
+Then, following the derivation of D.12 and D.13, we know that $\forall s\in{\mathcal{S}}$, $\forall\tau,\tau^{\prime}\in\{\pm 1\}^{{\mathcal{S}}}$ with $\tau\sim_{s}\tau^{\prime}$,
+$\mathrm{SubOpt}_{s}(\widehat{\pi};\tau)+\mathrm{SubOpt}_{s}(\widehat{\pi};\tau^{\prime})\geq\frac{\eta\delta^{2}}{8}\wedge\frac{3\delta}{10}.$ (G.14)
+Let $P_{r}$ be the distribution of $(s,a^{1},a^{2},y)$ for $s\sim\rho$, $a^{1},a^{2}\overset{\mathrm{i.i.d.}}{\sim}\pi^{\mathsf{ref}}(\cdot|s)$ and $y\sim\mathsf{Bern}(\sigma(r(s,a^{1})-r(s,a^{2})))$.
 Now we set $\delta=\sqrt{S/n}$ and conclude that for $\tau\sim\tau^{\prime}$ with $\tau_{s}=-\tau_{s}^{\prime}$,
@@
 $\leq\frac{4e}{3SC}(\exp(\delta)-\exp(-\delta))^{2}$
-$\leq 36S^{-1}C^{-1}\delta^{2},$ (E.14)
+$\leq 36S^{-1}C^{-1}\delta^{2},$ (G.15)
 where the second and third inequality hold due to $\alpha,\delta\leq 1/2$, and last inequality follows from $\exp(x)-\exp(-x)\leq 3x$ for $x\in[0,1/2]$.
 Now we set $\delta=\sqrt{SC/n}\leq 1/4$.
-We substitute E.13 into Assouad’s Lemma (Lemma F.3) and obtain that
+We substitute G.14 into Assouad’s Lemma (Lemma H.3) and obtain that
 $\inf_{\mathrm{inst}\in\mathrm{CDB}}\mathrm{SubOpt}_{\mathrm{RKL}}(\widehat{\pi};\mathrm{inst})$ $\geq\frac{1}{4}S\cdot\frac{1}{S}\cdot\bigg(\frac{\eta\delta^{2}}{8}\wedge\frac{3\delta}{10}\bigg)\cdot\min_{\tau\sim\tau^{\prime}}\exp\Big(-{\mathsf{KL}}\left(P_{\mathcal{D}_{\tau}}\|P_{\mathcal{D}_{\tau^{\prime}}}\right)\Big)$
@@
 $\geq\frac{\exp(-36)}{32}\min\{\eta CSn^{-1},S^{2}C^{2}n^{-2}\},$
-where the $1/S$ comes from the denominator of (E.12) and the second inequality follows from E.14. ∎
-E.3 Proof of Theorem D.5
-Proof of Theorem D.5.
-The proof is similar to the proof of Theorem 3.2. Recall that $b(\cdot)$ defined in (E.2), we know that
+where the $1/S$ comes from the denominator of (G.13) and the second inequality follows from G.15. ∎
+G.3 Proof of Theorem F.8
+Proof of Theorem F.8.
+The proof is similar to the proof of Theorem 3.2. Recall that $b(\cdot)$ defined in (G.2), we know that
 $\widehat{\pi}$ $=\mathop{\mathrm{argmax}}_{\pi\in\Delta^{d}}\Big\{\mathbb{E}_{(s,a)\sim\rho\times\pi}[{\bar{g}}(s,a)]-\eta^{-1}\mathbb{E}_{s\sim\rho}\big[D_{f}(\pi||\pi^{\mathsf{ref}})\big]\Big\}$
@@
 We have the following sub-optimality decomposition
-$J(\pi^{*})-J(\widehat{\pi})$ $=\mathbb{E}_{s\sim\rho}\Big{[}\mathbb{E}_{a\sim\pi^{*}}[{g^{*}}(s,a)]-\mathbb{E}_{a\sim\widehat{\pi}}[{g^{*}}(s,a)]-\eta^{-1}\big{[}D_{f}(\pi^{*}\|\pi^{\mathsf{ref}})-D_{f}(\widehat{\pi}\|\pi^{\mathsf{ref}})\big{]}\Big{]}$
+[…]
 $=\mathbb{E}_{s\sim\rho}\big[H^{*}_{s}({g^{*}})-H^{*}_{s}({\bar{g}}-b)-\left\langle\widehat{\pi},{g^{*}}-{\bar{g}}+b\right\rangle\big]$
@@
 Now, for any $\delta\in(0,1)$ and $\epsilon_{c}>0$, with probability at least $1-\delta$
-$J(\pi^{*})-J(\widehat{\pi})$ $=\mathbb{E}_{s\sim\rho}[({g^{*}}-{\bar{g}}+b)^{\top}\nabla^{2}H^{*}_{s}(\widetilde{g})({g^{*}}-{\bar{g}}+b)]$
+[…]
 $\leq\alpha^{-1}\eta\mathbb{E}_{s\sim\rho}\Big[({g^{*}}-{\bar{g}}+b)^{\top}\mathrm{diag}\big(\pi^{\mathsf{ref}}(a_{1}|s),\cdots,\pi^{\mathsf{ref}}(a_{d}|s)\big)({g^{*}}-{\bar{g}}+b)\Big]$
@@
 $\leq\alpha^{-1}\eta\bigg(\frac{128}{3n}\log(2\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+18\epsilon_{c}\bigg),$
-where the first inequality holds due to Lemma C.4 and last inequality holds due to equation (E.3). Setting $\epsilon_{c}=O(n^{-1})$ completes the proof. ∎
-E.4 Proof of Theorem D.6
-Proof of Theorem D.6.
-We still consider the contextual dueling bandit instance class defined in (E.10).
+where the first inequality holds due to Lemma E.4 and last inequality holds due to equation (G.3). Setting $\epsilon_{c}=O(n^{-1})$ completes the proof. ∎
+G.4 Proof of Theorem F.9
+Proof of Theorem F.9.
+We still consider the contextual dueling bandit instance class defined in (G.11).
 We show that given any positive $\alpha,\eta$, for any $n\geq S\cdot\max\{16,\eta^{2}\alpha^{-2}\}$, there exists $f:\mathbb{R}\to\mathbb{R}$ such that $f$ is $\alpha$-strongly convex, $\log|\mathcal{G}|=\Theta(S)$ and
-$\inf_{\widehat{\pi}\in\widehat{\Pi}(\mathcal{D})}\sup_{\mathrm{inst}\in\mathrm{CDB}}\mathrm{SubOpt}_{f\mathrm{div}}(\widehat{\pi};\mathrm{inst})\gtrsim\frac{\eta S}{\alpha n},$ (E.15)
+$\inf_{\widehat{\pi}\in\widehat{\Pi}(\mathcal{D})}\sup_{\mathrm{inst}\in\mathrm{CDB}}\mathrm{SubOpt}_{f\mathrm{div}}(\widehat{\pi};\mathrm{inst})\gtrsim\frac{\eta S}{\alpha n},$ (G.16)
 where $\mathcal{D}=\{(s_{i},a_{i}^{1},a_{i}^{2},y_{i})\}_{i=1}^{n}$ is the offline preference dataset, all (possibly randomized) maps from which to $\Delta(\mathcal{A}|{\mathcal{S}})$ is denoted by $\widehat{\Pi}(\mathcal{D})$.
 Since $S=\Theta(\log|\mathcal{G}|)\gtrsim\log\mathcal{N}_{\mathcal{G}}(\epsilon_{c})$ for all $\epsilon_{c}\in(0,1)$, we can conclude the theorem.
-[…]
-$\mathrm{SubOpt}_{f\mathrm{div}}(\pi;r)=\frac{1}{S}\sum_{s=1}^{S}\mathrm{SubOpt}_{f\mathrm{div}}\big{(}\pi(\cdot|s);r(s,\cdot)\big{)}.$ (E.16)
-We still consider the reward function class $\mathcal{G}$ indexed by {±$\{\pm 1\}^{{\mathcal{S}}}$. For all $\tau\in\{\pm 1\}^{{\mathcal{S}}}$ the reward instance “shaped” by $\tau$ is
-$r_{\tau}(s,a)=\frac{1}{2}+a\tau_{s}\cdot\sqrt{\frac{S}{n}},$ (E.17)
+Let ${\mathcal{S}}=[S]$, $\mathcal{A}=\{\pm 1\}$, $\rho=\mathsf{Unif}({\mathcal{S}})$ and $\pi^{\mathsf{ref}}(\cdot|s)=\mathsf{Unif}(\mathcal{A})$ for any $s\in{\mathcal{S}}$.
+Then the total sub-optimality of any $\pi\in\Delta(\mathcal{A}|{\mathcal{S}})$ given any reward function $r:{\mathcal{S}}\times\mathcal{A}\to\mathbb{R}$ is
+$\mathrm{SubOpt}_{f\mathrm{div}}(\pi;r)=\frac{1}{S}\sum_{s=1}^{S}\mathrm{SubOpt}_{f\mathrm{div}}\big(\pi(\cdot|s);r(s,\cdot)\big).$ (G.17)
+We still consider the reward function class $\mathcal{G}$ indexed by $\{\pm 1\}^{{\mathcal{S}}}$. For all $\tau\in\{\pm 1\}^{{\mathcal{S}}}$ the reward instance “shaped” by $\tau$ is
+$r_{\tau}(s,a)=\frac{1}{2}+a\tau_{s}\cdot\sqrt{\frac{S}{n}},$ (G.18)
 where $a\tau_{s}=\pm 1$ because $a\in\mathcal{A}=\{\pm 1\}$.
-We thereby refer $\tau\sim\tau^{\prime}$ to any pair in {±$\{\pm 1\}^{{\mathcal{S}}}$ that differs only in one coordinate.
+We thereby refer $\tau\sim\tau^{\prime}$ to any pair in $\{\pm 1\}^{{\mathcal{S}}}$ that differs only in one coordinate.
 $\forall\tau,\tau^{\prime}\in\{\pm 1\}^{{\mathcal{S}}}$, if $\tau\sim\tau^{\prime}$, then suppose $\tau_{s}=-\tau_{s}^{\prime}$, we have
-$\mathrm{SubOpt}_{f\mathrm{div}}(\pi(\cdot|s);r_{\tau}(s,\cdot))+\mathrm{SubOpt}_{f\mathrm{div}}(\pi(\cdot|s);r_{\tau^{\prime}}(s,\cdot))\geq\frac{\eta S}{\alpha n},$ (E.18)
-[…]
+$\mathrm{SubOpt}_{f\mathrm{div}}(\pi(\cdot|s);r_{\tau}(s,\cdot))+\mathrm{SubOpt}_{f\mathrm{div}}(\pi(\cdot|s);r_{\tau^{\prime}}(s,\cdot))\geq\frac{\eta S}{\alpha n},$ (G.19)
+where the inequality follows from exactly the same calculation in equation (E.5) by setting $f(x)=\alpha(x-1)^{2}/2$.
+[^7] Recall that in this case $D_{f}={\chi^{2}}$, where $2{\chi^{2}}(\mu\|\nu)+1=\sum_{a\in\mathcal{A}}[\mu(a)]^{2}/\nu(a)$.
+Let $P_{r}$ be the distribution of $(s,a^{1},a^{2},y)$ for $s\sim\rho$, $a^{1},a^{2}\overset{\mathrm{i.i.d.}}{\sim}\pi^{\mathsf{ref}}(\cdot|s)$ and $y\sim\mathsf{Bern}(\sigma(r(s,a^{1})-r(s,a^{2})))$.
+Then we denote $\delta=\sqrt{S/n}$ and conclude that for $\tau\sim\tau^{\prime}$ with $\tau_{s}=-\tau_{s}^{\prime}$,
 ${\mathsf{KL}}\left(P_{r_{\tau}}\|P_{r_{\tau^{\prime}}}\right)$ $=\frac{1}{SA^{2}}\sum_{s^{\prime},a^{1},a^{2}}{\mathsf{KL}}\left(\mathsf{Bern}(\sigma(r_{\tau}(s^{\prime},a^{1})-r(s^{\prime},a^{2}))\|\mathsf{Bern}(\sigma(r_{\tau^{\prime}}(s^{\prime},a^{1})-r(s^{\prime},a^{2}))\right)$
@@
 $\leq\frac{1}{4S}\Big((\exp(-2\delta)-1)^{2}+(\exp(2\delta)-1)^{2}\Big)$
-$\leq\frac{1}{2S}(\exp(2\delta)-1)^{2}\leq\frac{36\delta^{2}}{2S}=\frac{18}{n},$ (E.19)
-where the last inequality follows from $\exp(x)-1\leq 3x$ for $x\in[0,0.5]$ and $\delta=\sqrt{S/n}\leq 0.25$ by assumption. Therefore, we substitute E.18 into Assouad’s Lemma (Lemma F.3) to obtain
-LHS of E.15 $\geq\frac{1}{S}\cdot S\cdot\frac{\eta S}{\alpha n}\cdot\frac{1}{4}\cdot\min_{\tau\sim\tau^{\prime}}\exp\Big{(}-{\mathsf{KL}}\left(P_{\mathcal{D}_{\tau}}\|P_{\mathcal{D}_{\tau^{\prime}}}\right)\Big{)}$
-$=0.25\cdot\frac{\eta S}{\alpha n}\cdot\exp\Big{(}-n{\mathsf{KL}}\left(P_{r_{\tau}}\|P_{r_{\mathcal{D}_{\tau^{\prime}}}}\right)\Big{)}\geq\frac{\eta S}{\alpha n}\cdot\frac{1}{3}\cdot\exp(-18)\gtrsim\frac{\eta S}{\alpha n},$ (E.20)
-where the $1/S$ comes from the denominator of E.16 and the second inequality follows from E.19. ∎
-Appendix F Technical Results and Standard Results
-Lemma F.1 (Zhao et al. 2024, Lemma D.4).
-Consider a offline dataset $\{(s_{i},a^{1}_{i},a^{2}_{i},y_{i})\}_{i=1}^{n}$ generated from the product of the context distribution $\rho\in\Delta({\mathcal{S}})$, policy $\pi\in\Delta(\mathcal{A}|{\mathcal{S}})$, and the Bradley-Terry Model defined in Section 4.1.
+$\leq\frac{1}{2S}(\exp(2\delta)-1)^{2}\leq\frac{36\delta^{2}}{2S}=\frac{18}{n},$ (G.20)
+where the last inequality follows from $\exp(x)-1\leq 3x$ for $x\in[0,0.5]$ and $\delta=\sqrt{S/n}\leq 0.25$ by assumption. Therefore, we substitute G.19 into Assouad’s Lemma (Lemma H.3) to obtain
+LHS of G.16 $\geq\frac{1}{S}\cdot S\cdot\frac{\eta S}{\alpha n}\cdot\frac{1}{4}\cdot\min_{\tau\sim\tau^{\prime}}\exp\Big(-{\mathsf{KL}}\left(P_{\mathcal{D}_{\tau}}\|P_{\mathcal{D}_{\tau^{\prime}}}\right)\Big)$
+$=0.25\cdot\frac{\eta S}{\alpha n}\cdot\exp\Big(-n{\mathsf{KL}}\left(P_{r_{\tau}}\|P_{r_{\mathcal{D}_{\tau^{\prime}}}}\right)\Big)\geq\frac{\eta S}{\alpha n}\cdot\frac{1}{3}\cdot\exp(-18)\gtrsim\frac{\eta S}{\alpha n},$ (G.21)
+where the $1/S$ comes from the denominator of G.17 and the second inequality follows from G.20. ∎
+Appendix H Auxiliary Lemmas
+Lemma H.1 (Zhao et al. 2024, Lemma D.4).
+Consider a offline dataset $\{(s_{i},a^{1}_{i},a^{2}_{i},y_{i})\}_{i=1}^{n}$ generated from the product of the context distribution $\rho\in\Delta({\mathcal{S}})$, policy $\pi\in\Delta(\mathcal{A}|{\mathcal{S}})$, and the Bradley-Terry Model defined in Section F.1.
 Suppose $\bar{g}$} is the result of MLE estimation of Algorithm 3, and we further define $b(s)=\mathbb{E}_{a\sim\pi(\cdot|s)}\big[{\bar{g}}(s,a)-{g^{*}}(s,a)\big]$, then with probability at least $1-2\delta$, we have
 $\mathbb{E}_{s,a\sim\rho\times\pi}\big[\big({\bar{g}}(s,a)-{g^{*}}(s,a)-b(s)\big)^{2}\big]\leq O\bigg(\frac{1}{n}\log(\mathcal{N}_{{\mathcal{G}}}(\epsilon_{c})/\delta)+\epsilon_{c}\bigg).$
-Lemmas F.2 and F.3 are two standard reductions (Le Cam, 1973; Yu, 1997; Polyanskiy and Wu, 2025). See, e.g., Chen et al. (2024, Section 3) for a general proof.
-Lemma F.2 (Fano’s inequality).
-Fix any :-$\mathcal{R}\coloneq\{r_{1},\cdots,r_{S}\}$ and policy class $\Pi$, let $L:\Pi\times\mathcal{R}\to\mathbb{R}_{+}$ be some loss function.
+Lemmas H.2 and H.3 are two standard reductions (Le Cam, 1973; Yu, 1997; Polyanskiy and Wu, 2025). See, e.g., Chen et al. (2024, Section 3) for a general proof.
+Lemma H.2 (Fano’s inequality).
+Fix any $\mathcal{R}\coloneq\{r_{1},\cdots,r_{S}\}$ and policy class $\Pi$, let $L:\Pi\times\mathcal{R}\to\mathbb{R}_{+}$ be some loss function.
 Suppose there exist some constant $c>0$ such that the following condition holds:
@@
 Then we have
-$\inf_{\pi\in\Pi}\sup_{r\in\mathcal{R}}L(\pi,r)\geq\frac{c}{2}\bigg{(}1-\frac{\max_{i\neq j}\mathsf{KL}(P_{r_{i}}\|P_{r_{j}})+\log 2}{\log S}\bigg{)},$
-where $P_{r}$ is the distribution of dataset given model $r\in\mathcal{R}$.
-Lemma F.3 (Assouad’s Lemma).
+$\inf_{\pi\in\Pi}\sup_{r\in\mathcal{R}}\mathbb{E}_{\mathcal{D}\sim P_{r}}L\big(\pi(\mathcal{D}),r\big)\geq\frac{c}{2}\bigg(1-\frac{\max_{i\neq j}\mathsf{KL}(P_{r_{i}}\|P_{r_{j}})+\log 2}{\log S}\bigg),$
+where the trajectory distribution of $\pi$ interacting with instance $r\in\mathcal{R}$ is denoted by $P_{r}$.
+Lemma H.3 (Assouad’s Lemma).
 Let $\mathcal{R}$ be the set of instances, $\Pi$ be the set of estimators, $\Theta\coloneqq\{\pm 1\}^{S}$ for some $S>0$, and $\{L_{j}\}_{j=1}^{S}$ be $S$ functions from $\Pi\times\mathcal{R}$ to $\mathbb{R}_{+}$.
@@
 We denote $\theta\sim_{j}\theta^{\prime}$ if they differ only in the $j$-th coordinate. Further assume that
-$\theta\sim_{j}\theta^{\prime}\Rightarrow\inf_{\pi\in\Pi}L_{j}(\pi,r_{\theta})+L_{j}(\pi,r_{\theta^{\prime}})\geq c$ (F.1)
+$\theta\sim_{j}\theta^{\prime}\Rightarrow\inf_{\pi\in\Pi}L_{j}(\pi,r_{\theta})+L_{j}(\pi,r_{\theta^{\prime}})\geq c$ (H.1)
 for some $c>0$, then
-$\inf_{\pi\in\Pi}\sup_{r\in\mathcal{R}}L(\pi,r)\geq S\cdot\frac{c}{4}\min_{\exists j:\theta\sim_{j}\theta^{\prime}}\exp\Big{(}-{\mathsf{KL}}\left(P_{r_{\theta}}\|P_{r_{\theta^{\prime}}}\right)\Big{)},$
-where $P_{r}$ denotes the distribution of the dataset given $r\in\mathcal{R}$.
-The following Lemma F.4 is due to Gilbert (1952); Varshamov (1957), which is a classical result in coding theory.
-Lemma F.4.
+$\inf_{\pi\in\Pi}\sup_{r\in\mathcal{R}}\mathbb{E}_{\mathcal{D}\sim P_{r}}L\big(\pi(\mathcal{D}),r\big)\geq S\cdot\frac{c}{4}\min_{\exists j:\theta\sim_{j}\theta^{\prime}}\exp\Big(-{\mathsf{KL}}\left(P_{r_{\theta}}\|P_{r_{\theta^{\prime}}}\right)\Big),$
+where the trajectory distribution of $\pi$ interacting with instance $r\in\mathcal{R}$ is denoted by $P_{r}$.
+The following Lemma H.4 is due to Gilbert (1952); Varshamov (1957), which is a classical result in coding theory.
+Lemma H.4.
 Suppose $\Sigma$ is a set of characters with $|\Sigma|=q$ where $q\geq 2$ is a prime power and $N>0$ is some natural number.